A Complete Bibliography of the Journal of Applied Econometrics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

28 May 2024 Version 1.13

Title word cross-reference

$105.00 [McC02]. $16.95 [Gan88b]. $17.50 [Arb93]. $24.95 [Mil91]. $25.95 [Ano90g]. $28.95 [Car88]. $35 [Mel92]. $39.95 [Son89a]. $44.95 [Cif88]. $45 [Hen93]. $49.50 [Mal88]. $49.95 [McC02, Riz91]. $54.00 [Vis88]. $54.40 [Col92]. $57.95 [McC99a]. $60.75 [Gan88b]. $94.50/DFL [Hen93]. $94.75 [Pis92]. $98.75 [Wee02]. AR(1) [KM91]. D [PMM17]. I(1) [BRiS11]. I(2) [BF05, EH99, BRiS11]. J [KP16]. N [DH16]. q [Sch90]. T [HPS23].

-Test [KP16].

0 [Kuk02]. 0-07-025214-9 [Mad98]. 0-07-032685 [Car88]. 0-13-35582-4 [Oxl96]. 0-19-506011-3 [Smi95]. 0-19-520655-X [Riz91]. 0-19-828367-9 [Arb93]. 0-19-82854-8 [Per91]. 0-19-828759-3 [Bro94]. 0-19-829340-7 [Arb93]. 0-19-877317 [Hen93]. 0-19-877320x [McK94]. 0-19-877453-2 [Leo98]. 0-19-877583-0 [Gar01a]. 0-387-51454-6 [De 92]. 0-387-53194-7 [Bai93]. 0-387-90908-7 [Vis88]. 0-412-01371-1 [Kus91]. 0-412-28830-3 [Ano90e]. 0-412-35340-7 [Bol92]. 0-415-00405-5 [Mel92]. 0-415-00918-9 [Mil92]. 0-444-70529-5 [Pis92]. 0-471-01171-1 [De 93]. 0-471-05856-4 [Son89a]. 0-471-11856-7 [Fie96]. 0-471-13993-9 [McK97]. 0-471-19189-2 [McK97]. 0-471-59951-4 [Szr94]. 0-471-62414-4 [Ano90c]. 0-471-81930-1 [Mal88]. 0-471-82216-7 [Ano90g]. 0-471-85233-3 [Son89b]. 0-471-86187 [Ano90f]. 0-471-86187-1 [Cif88]. 0-471-89530-X [Ano90h]. 0-471-91787-7 [Nüe91]. 0-471-95359-8 [Mad97]. 0-471-95360-1 [Mad97]. 0-471-96688-6 [Tur98]. 0-471-9847-4 [Ano90a]. 0-521-25280-6 [Dav95]. 0-521-2644-8 [Gan92]. 0-521-26616-5 [Klo92]. 0-521-32196-4 [Sab91]. 0-521-33004-1 [Mil89b]. 0-521-34430-1 [Hal89]. 0-521-34552 [Hal89]. 0-521-38043-X [Imb97]. 0-521-41146-7 [McK98b]. 0-521-42308-2 [McK98b]. 0-521-78830-7 [Kuk02]. 0-521-81696-3 [Wee03]. 0-566-05629-1 [Sor89]. 0-582-29723-0 [Mil89a]. 0-674-17544-1 [Gil93]. 0-7923-1287-2 [Gre93]. 0-8133-7489-8 [Mil91]. 0-8157-1 [Ano90b]. 0-86003-084-9 [And93]. 0521-30444X [Col92].

1-538-85013-2 [Wee02]. 1-55786-264-8 [Abe94]. 1-85278-461-x [Cle94]. 1-85278-657-4 [Bar95]. 1.1a [Woh96]. 1019 [Ano90h]. 1024 [Ano90c]. 121 [Mil91]. 139-5 [Ano90b]. 140-9 [Ano90b]. 146 [Ano90a]. 15.95 [Oxl96]. 171 [Gre93]. 176 [Phi89]. 18 [Ano03]. 187 [McK94]. 189 [Gio89]. 18th [Sor89]. 19 [BHLS23, KH22a, PY22]. 19.95 [Szr94]. 191 [Mel92]. 1920s [PMM96]. 1929 [kHcL16]. 1974 [Nei92]. 1974-1987 [Nei92]. 1975-87 [Ras95]. 1980s [BCDJ99]. 1987 [Ano90f, Nei92]. 1987/88 [GS07]. 1990s [BS04]. 1991 [Ano91g]. 1992 [Ano92j]. 1995/96 [GS07].

2-7178-1667-4 [Klo93]. 2-7178-1668-2 [Klo93]. 2007a [AF17]. 2008 [Pes08]. 2020 [LP22]. 205 [Riz91]. 230 [Leo98]. 232 [Mof93]. 240 [Bro94, Per91]. 244 [Gan92]. 259 [Hal89]. 2nd [Ano90c, Ano90f].

3-540-17257-2 [Car92]. 3-540-54054-7 [Mof93]. 3-7908-0356-1 [Gio89].

30. [Nei92]. 30444X [Col92]. 314 [Car92]. 315 [Ano90g]. 32/64 [CNS09]. 326 [Arb93]. 329 [Son89b]. 334 [Wee03]. 338 [Hal93a]. 348 [Mil92]. 349 [Pep05]. 365 [Col92]. 370 [Cle94].

40.00 [Gar01a]. 405 [Gil93]. 406 [DG03]. 427 [Kus91]. 440 [Cif88]. 476 [Hen93]. 487 [Pis92]. 488 [Nüe91].

5 [DG03]. 50-00 [Abe94]. 507 [Mil89b]. 514 [De 93]. 545 [Bai93]. 554 [Sab91]. 558pp [Imb97]. 577 [Fie96]. 587 [Klo93].

669 [Sor89].

7 [AAV07, Kuk02].

’87 [KK96]. 877320x [McK94].

9/11 [SVW08]. 90-267-1264-3 [Hen93].

A$195.95 [Tur98]. A$57.95 [McK97]. A$89.95 [Tur98]. Abiad [Hua09]. Ability [CM17a, HKMR03, KLNT24, Wil19]. abrupt [CGP99]. absence [BKS18, BJ09]. Absenteeism [AJ17]. absolute [DZ18, FVP02]. absorber [CT86b]. abuse [Ter02]. Academic [Gre93, Tho19]. Accelerated [DR93]. acceleration [Ber01]. accompany [McK97]. account [SPS13, vHBB22]. Accounting [EHP12, Sch16, BP13, HH02, JS04, MW20]. accounts [Sto86]. accumulation [BLS10, Sha86]. Accuracy [CCM15, CL11, CM14, CI20, KLS19, LRV12, Mcc97, Vin00, YY07]. ACF [iKLS19]. Achievement [GPW17, BC23, BBDF14, HKMR03, MM89, NR19]. Ackerberg [DM06b]. Acknowledge [Ano95a]. Acknowledgement [Ano89a, Ano92a, Ano92b, Ano93a, Ano94a, Ano05a, Ano07a, Ano08a, Ano09a, Ano09b, Ano09d, Ano10a, Ano12a, Ano13a, Par11a]. acquisition [Tow92]. across [And91, ABCF19, Bra11, Car00, CLM18, PZ06, SP18, Tho19, TL03]. actions [MKP22]. activation [MvdKH20]. Active [FL15, EO14]. activities [CLLY06, MM87]. activity [CCM22a, Djo20, Djo24, GH96, Ham21, HR20, LMST22]. Actual [FFG19, GSS20, SS05]. Adaptive [HHM15, Hod99]. add [KP22, Vin03]. add-on [Vin03]. addicts [CL09b]. additive [VFL99]. Addressing [BC24]. adjusted [PT13]. adjustment [BK86, Bro87, DH90, DH10, EPW07, EH94, GGS08, Hen93, JK90, KR90, LSV23, MMOL08, Nak93, PP06, Sal91]. administrative [KLT20]. Admission [HKS16]. adopt [McC99a]. adults [PS03b]. Advance [KL24]. Advanced [Arb93, Phi89, DFM21, Wee01, Vis88]. advances [CSvD14, Hal89]. advantage [Gan88b]. advertising [HM17, PS07]. advice [KT01]. advisory [HP09]. AER [Sum17]. affairs [LR04]. affect [BBDF14, DST14, FSK06, GKPP07, HKMR03, Tro24, VV13]. Affected [CS15, MQ09]. affiliated [CPV03]. affine [FO18, RH21]. Africa’ [DV13, CLS21, Cua11, MV06, MP08]. after [Fer23b, GK12, Kau20, LP22, Nag22]. aftermath [FIL13]. Against [Pep05, AD06, BD06, DM06a, DM06b, DS91, NH01, Rot92]. Age [Fuk06, Che98, HSWW07, KA07, Kuo12, MRS10, Par10, Par11b, Tre98]. Age-period-cohort [Fuk06]. Agency [GW15]. agent [Hon22]. agents [BJM^{+}08, KRV95]. agglomeration [Whe03]. Aggregate [EPS17, GP22a, HSF05, Koo92a, AV97, CFLP02, EP14, For95, Fuk06, KL24, MS18b, MBR09, PKH12, SD03]. aggregates [Che98, SZ07]. Aggregation [GM09, Alb03, BE92, CP92, MSW13, PP06, Wol93]. Aging [MRGB16]. agricultural [BSS08, Mel92]. agriculture [Coe96, Hoq91]. Aid [CHW17, BGLS21, Brü13, ERS03]. aids [DGM11, CLS21, MB02]. air [BS07]. aircraft [Ken94, Smi12]. airline [GHS05]. airlines [BGV98]. Akom [DG03]. al [Fu19, GVB17, NT22]. Alain [McK98b]. Alchemy [Abe94]. Alcohol [Ter02, KT01, Una06]. alcoholic [MB02]. Aldershot [Bar95, Cle94, Sor89]. algebra [HH95, LR08]. Algorithm [Jam17, HS90, HMRC05]. algorithms [AK07]. Allan [And93]. Allocation [HRV94, And91, BA98, CKT19, Des93, HMSS14, KP22]. Allocations [HKM15]. allocative [AD09, MR88]. allowing [BRRT21]. almost [BS08, Des03, Pep05]. Alternative [CR15, FH12, Lin92, SM95, SC90, TW87, ZR98, Aka12, Cha02, CK13, DS91, DH10, GP06, IRT08, KW13, LO08, LWY00, MT11, Mur12, TP97]. alternatives [EWB09, SdV18]. aluminium [GN00, Gil95]. amenities [Cha02]. America [And03, CGK14, Can05, GLS07]. American [HKW11, Li12, Mor08]. Americans [ABCP02]. AMI [KPST11]. Amman [Gre93]. Among [DJ16, DZ18, GHL04, KR90, Sem18]. amount [VV93]. Amsterdam [Pis92, Wee01]. Analysing [BCT96, MRV95, PS03a]. Analysis [Ano90a, BHKL17, CJ16, CD16, Dah17, DW16, EM17, GKV15, jLjL05, Luc15, MS15, NS17, Sok16, ACS99, AG13, Ano90g, Ash09, AFHI95, BP03, BW11, BHLTM24, BCR01, BS07, BM11, BA98, BC07a, BJJ11, BC07b, BHMW98, CHL14, CS21a, CLSM10, CT92, CJZ22, Che23, CO96, CMM09, DMM05, Dav95, De 92, De 94, DMT06, DG92, DdMPS07, DLR94, Des03, Die89, DS06b, DM03, ERS03, EO14, EP14, FMN05, Fie96, FM14, For95, FSK06, GMH95, GAA^{+}89, GvSZ05, GR08, Haj94, Hey04, HM03, HSF05, HL10, HJC10, JJVW01, JKvdW14, KvdGVW97, KM91, KL01, KMP08, KvD93, KM14b, KT08, Kon04, KS91, KP94, Koo99, Kuk02, Lee04, LWY00, MV06, MSL89, MT13b, MP10, MM89, MP07, MSP13, NS90, ORM13, OSS18, Oxl96, PU88, PSS01, Phi91b, Poi91]. analysis [Pud93, QT23, RS86b, Sen11, Sep00, Shi92, SS98, SS06, SV96, SZ06, TL03, Tri93, Tuf98, VK87, VV14, WVG14, WC04, Win04, Yat00, vDvDF05, BHP16, Hal93a, Imb97, Bai93, Mil89a]. Analytic [FCP96]. Analyze [Van17]. analyzed [Eic09]. Analyzing [GS20, BP10]. Ancestry [SW18b]. Anchoring [AGR17, CHP23]. Anglo [Per91]. Angrist [Nak10]. Angus [Bro94]. Announcement [Ano87a, Ano87b, Ano87c, Ano93c, Ano93b, Ano94b, Ano95b, Ano96a, Ano05b, Gan88a, Mac91, Pes92a, Pes92b, Pes93, Pes94, Pes95, Pes08, Pes09a, Ano93e, DE22]. Announcements [Ano94c, Ano96b, LLN11]. Annual [Ano01d, Ano01e, EHS09, Wat07, Gre93]. anomalies [Col92, RV23]. anomaly [MP01]. answers [DGM11, KV14]. ante [HM03]. Anticipating [CL15]. Anticipation [CDKS17, CT86b]. Anundsen [Fu19]. anything [HL05]. Aoki [Car92]. apartments [FLS22]. Application [FL15, HL17, HHJC16, HMT15, MM17, Sok16, Uys15, XL16, ALR19, And03, ABP21, AHT94, BW11, BRRT21, Bal10, BEK22, Bas14, BC19, BNP11, BL17, Bro87, BBG04, Cam01, Car00, CGP99, CPUBUW24, Cha11, CEG22, CJS23, CP92, DR93, DMT06, Dej93, DM98, Dep18, EPR11, Far12, FL94, FK03, FP02, Fuk06, Giu11, Gur97, HP04a, Hen10, Hod99, Hoq91, HLMvS98, Hus08, Jag91, JKLGS13, JV22, JKvdW14, KY10, Kau10, KL93, MS98, MC90, Mar01, MS88, MT13b, MR90, MS93, OPM19, PW96, Pri02, PSS04, Ras95, RS08b, SS07, SM95, SLKN04, SK07, SV22, Szy19, Tre98, VD98, WFBM01, WPS18, Wat99, WS97, BCMM16, GKWZ16, Kuk02]. Applications [BH15, Don15, RS16, WCC17, YZ15, AL06, ACF22, BP07, CT86a, CJ97, CSvD14, Cha00, CC07, CKKK18, CKL13, DK01, DKB10, FM19, FM97, GA11, HM12, HS00, JM98, KZ24, KPP90, Kom07, LS09b, MH14, Pro14, She93, Sil01, SGK12]. Applied [Ano91g, Ano92j, Ano93c, Ano95i, Ano95c, Ano95d, Ano95o, Ano96c, Ano96d, Ano96e, Ano96f, Ano01d, Ano01e, Ano03, Ano06, Ano07b, Ano08b, Ano10b, Ano11, Ano12b, Ano12c, Bal99, DG03, Pes01a, Pes01b, Pes01c, Pes07a, Pes08, Pes09b, Pes09c, Pes10b, Pes11, SV96, Wat07, Ano90i, Ano13b, Bol92, MM95a, Pes03b, Pes10a, And93, Gar01a, Mil89a]. Applying [JLR14]. appraisal [CG09, Par87]. appraisals [MST12]. Approach [BEK16, BEK15, BPC15, CPA15, CFH16, EHSJ17, FRF15, GVB17, GKV15, Hur17, LM17, MKT16, SW05, VGCBHJC18, YZ15, Abr06, ATW08, AHKM19, ABKK13, AC87, Bai91, BBS97, BJM^{+}08, BV12, Bla96, BH18, BM12, Buc98, CL09a, CMBS15, CTU13, Che08, Che23, CJS07, DHR11, DK12, DLR06, DT97, DIL14, DMVR22, EP04, EHV94, FLHS07, FOS17b, Fuj11, GGS08, GP97, GS20, HGK18, HKO12, HLV02, HCW12, HP04b, HPY19, HKF19, Jal18, JRA21, JO20, Juo22, KMV19, KPS04, KLT20, KO14, Koo92a, Kra12, KN93, LD22, LLP23, LO08, LS99, LM09, LB19, MS18a, MSS06, Mar00, MS88, MM09, Mor94, Nis05, NO21, Nyh03, Par10, Par11b, Ped07, Per91, Pet89, Pot95, QR86, Rid86, Sci23, Sli23, Sut23, TH04, Tur09, VB02, VNZ97, Wat99]. approach [Web95, Wes20, iKLS19, vHBB22, BHP16, Wee02]. approach* [JJVW01]. Approaches [BKL17, Bol92, DD23, pLK20, PSS01, Sto91]. Approval [BEK15]. approximate [Eis19]. Approximating [LMPTR23, GL14b, SHK11]. Approximation [AFR08, Ano90a]. approximations [BD05b, Byr92, She89]. April [Ano91g]. arbitrage [CCM21]. arbitrary [BM14, QT23]. arch [RRB92, DN89, DKB10, Eng02, FM97, FB02, LWY00, RZ93, VFL99]. archive [Ano93c, Ano95i, Ano95c, Ano95d, Ano96c, Ano96d, Ano96e, Ano96f]. ARDL [NT22]. Area [FS17, LSZ17, PMM17, BL06, CPQ10, CV01, DdMPS07, DO14, Eic09, FMMR18, SW05, SW21]. ARFIMA [BCT96]. Aris [McC02]. arithmetic [QT23]. arms [Li12]. Arnold [Fie96]. Art [BH15, MST12]. Arthur [Mil91]. articles [BLM10]. ascending [LP08]. Asian [LT91]. ask [BNP11]. Assessing [Ash09, AK07, CCM20, CM09, DG11, EMW11, HT21, HR96, Orn06, RV22, AP12, AC87, BSS08, CJS23, DK13]. assessments [BC23]. Asset [Fis94, Hur17, AK19, BG94, BSB86, CL09b, CO96, DKB10, Fei87, GMRO19, GA11, HLV02, HMSS14, KP22, MS95, ORW02, Pro14, RS08b, TW17, Wri99]. asset-pricing [BG94, ORW02, RS08b]. assignment [Joc23]. assistance [KvdK22]. associated [DS91]. Assortative [Lee16]. Assumption [HM15, McC94]. assumptions [ABK23, BM12, Giu11, LS09b, MT11]. Asymmetric [Kom07, Sav16, AS06, BV12, CD96, GHOS04, HPY19, JMPW18, Thi20]. asymmetrical [DV92]. asymmetries [RZ93]. Asymmetry [CPV03, PS03c]. Asymptotic [KPP90, Koe88, CKKK18]. Asymptotically [HF00]. asymptotics [CI20]. Asynchronicity [CPA15]. attainment [AJ19, JPPS14, LT06, Pos19]. Attendance [BFM15, KOW21]. Attention [BFM15, LG23]. attitudes [BT23]. Attribute [BFM15, Hen06]. attributes [HMS01]. Attrition [DLR94, BHLTM24, MV06]. Auction [Per10, AS06, DPR06, Fév08, LP08, Sar99]. Auctions [DIM08, HH15, CPV03, HM17, JL08, LP08, Rez08, Xu13]. Augmented [Kul15, BK21a, KUW19, LB19, Tem98, Yam19, YH22]. Australia [GR08]. Australian [BO94, Coe96, SM95]. Austrian [FSK06, Kau10]. authority [Cho99b]. Authors [Pes08, Ano06, Pes01b, Pes10a, Pes11, Pes13]. auto [BGR10]. Autobox [Bew88]. autocorrelation [CL09a, Des93, RS08b]. automated [CJZ22]. automatic [MWWW24]. Automobile [DV92, AHT94, Van07]. Autoregression [LL16, CM24a, Che23, FP02, LP22, MSH19, MW20, YH22]. Autoregressions [CM17a, FM16, AF19, Ano23a, FHHP23, HIT20, HK23, Pee05, Smi93, Vil09, Yam19]. Autoregressive [Bec01, CFH16, ABC16, AAV07, AAQB17, BC19, CCN22, CB17, CG13, CKL13, Des08, GL14b, HHO21, LLL12, LL96, LB19, PL08, Sch07, TA92a, CG23]. availability [AHKM19]. Avebury [Sor89]. Average [CLU17, HLM17, Ver20, ABKK13, FMS19, LOW17, Mor08, PI09, Sak20, Wes20]. average-based [Wes20]. Averages [RW16]. Averaging [CM10, AP12, EPR11, KvD06, LS09b, LM09, MP08, MT16]. averse [BJM^{+}08]. Aversion [Brz15, ATV19, GL12, HLW14, LP08, NSA98]. avoid [BSS11]. Avoidance [CDKS17].

B [DG03, KS91, Sim91, Vis88]. baby [Wol18]. back [GM09]. backcasting [KY10]. bad [CCN22]. bagging [AM11]. balance [GNNS12, SSX22]. Ball [BKL17]. Baltagi [DG03]. BAMLE [DFM21]. band [CEG22, LOW17]. bandpass [CKZ10]. bands [OPM19]. Bangladesh [Hoq91, MSN86]. bank [DDLY18, FSK06, KDK18, LFZC23, Per98, vOZ19, CPBS20]. banking [DK12, EGI05, FS09, NPT18, WW18]. Banks [MKT16, BB22]. Bargaining [MV07, KW89]. Barker [Mil89b, Mil92]. Barriers [GKZ18, Ped07]. Based [KKM16, LRV17, PR16, YZ15, BV12, BL06, CT86a, Che08, CL09b, CJ20, CKZ10, DHR11, DMVR22, DKB10, FP11, Fer17, FO18, HKS14, Hub14, HHLL20, Jag91, KLNT24, pLK20, LS93b, LK21, MM03, Mar00, Mur12, NSS12, PV89, Pee05, Sch00, SS10, TW17, TH04, Wes20, Wil06a, Wol18, dHW22]. Basic [Mad98]. Basil [Phi89]. Bayes [GK17, HET08, KvD06]. Bayesball [GK17]. Bayesian [Fie96, KS91, MT16, Ade04, ABC16, AG13, ABS93, AFR08, AC87, AD09, BW09, BEK22, BGR10, BHKL17, BKKS20, BJM^{+}08, BV12, BVV16, BCRV16, BJJ11, BHH15, Cam01, CKM11, CCM15, CE17, CE18, CEHK20, CJZ22, CJ16, CG23, CFH16, DMT06, Dej93, DWW17, EPR11, EZB17, FHHP23, FK03, FM16, FSPWWE12, Gal20, GGS08, Gew93, GMT98, GMH14, GVS23, HMS16, HK23, JMY19, JS10, KK97, Kau10, KMP08, KW19, KvD93, KM14b, KT08, Koo92a, Koo92b, Koo99, KPT05, Koo13, Kor13, Kra12, LJ10, LLL12, LL16, LS09b, Li99, LZ08, LM09, LMMF21, MS18a, MP08, MW20, MRGB16, MBR09, Phi91a, Phi91b, Poi91, RH21, RT07, SV91, SW05, SGK12, Sto91, vHBB22]. BDS [BFC02]. be [HLW14, Zaf11]. Bear [KvD17, PS03a]. beat [HL05]. Becker [Edw21]. bed [Oku93]. Beenstock [Ano90e]. Behavior [ARST16, BBvdH09, HW22, JMM07, KW13, SK07, Mof93]. behavioral [SZ18]. behaviors [DST14]. behaviour [ABS99, Bek99, BMV00, Cut86, DJ00, DD00, DLR94, DH90, DJ95, GS04, Han93, Hey04, KPS04, Lui00, MAL04, VK87]. Being [LHL17, HMS01]. Belgium [CD05]. belief [BSS11]. beliefs [ASS21, DST14, RM00]. belonging [DDL23b]. Belsley [Gre93, Son89a]. Belts [JZ16]. benchmarking [Mcc97]. Benefit [CTU13, ALV13, Liu09, NS93, VV13]. benefits [AM11, BBC20, DR11, EO14, HL07, HCW12, VVR00, Wol07]. Bergstrom [Arb93]. Berlin [Bai93, Car92, De 92, Vis88]. Berry [Fie96]. best [BS08, HS00]. beta [HLV14, JLR14]. better [CK20]. Between [BEK15, BCRV16, CHW17, EZB17, GVB17, KKS16, AV97, Blo02, Blo92, CCV23, CEF07, Dem89, EKL10, FB02, Gal96, GT09, LU05, LK09, Med02, MS93, PS03b, RLS00, RRB92, SPS13, SdV18, Tan10, Vog92]. beverages [MB02]. Bewley [Hal89]. Beyond [Li06, ALV13, TL03]. Bias [HI16, Pep05, BSBC^{+}04, BC24, CT14, Dev07, FH12, FFHL20, MST12, SE07, Vel92]. Biased [Cha15]. Biases [SHK11, BBvdH09, Coe96, For95, SdV18, Zaf11]. bid [BNP11, LP08, Sar99]. bidders [AS06]. big [PH24]. bilateral [BCO04, BGLS21, DMP21, DE22]. Binary [BV12, CPUBUW24, LfL09, SW18a, Ger96, KV09, Kor06, Sue95]. binding [CBGIZ19]. birth [Abr06, COQD19, FGV18, LO18]. bit [CNS09]. Bivariate [BM91, ABK23, HL19, Hel06, RWM03]. Bjerkholt [Bol92]. black [GMH14, Lui00]. Blackwell [Abe94, Phi89]. Blanchard [Her19]. blast [DS04]. Bliss [Wee01]. blocking [HKO12]. Blomquist [DM06b]. Bloom [CCN22]. blue [KPS04]. blue-collar [KPS04]. BMA [MH14, Pap11]. BMI [KT08]. board [HP09]. body [Smi12]. Bollen [De 93]. Bond [RM17, BL17, CR13, DMP00, FHHP23, GT06]. Bonds [Mad97, EFR11]. bonu [Bil00]. bonuses [VV13]. Book [Abe94, And93, Ano90a, Ano90b, Ano90c, Ano90e, Ano90d, Ano90f, Ano90g, Ano90h, Ano08b, Arb93, Bai93, Bar95, Bol92, Bro94, Car88, Car92, Cif88, Cle94, Col92, Dav95, De 92, De 93, Fie96, Gan88b, Gan92, Gar01a, Gil93, Gio89, Gre93, Hal89, Hal93a, Hen93, Imb97, Klo92, Klo93, Kuk02, Kus91, Leo98, Mad97, Mad98, Mal88, McC99a, McC02, McK94, McK97, McK98b, Mca97, Mel92, Mil89a, Mil89b, Mil91, Mil92, Mof93, Nei92, Nüe91, Oxl96, Per10, Per91, Phi89, Pis92, Riz91, Sab91, Smi95, Son89a, Son89b, Sor89, Sto09, Szr94, Tur98, Vis88, Wee01, Wee02, Wee03, Fie96]. books [Sch05]. boom [CMS24]. boomer [Wol18]. Booms [BCRV16, Sum17]. Boosting [BN09]. Bootstrap [ACF22, MW17, Yam19, FM19, MNW23a]. Bootstrapping [Vea92, DJ00, MZ97]. Borrowing [RTV16]. Boston [Gre93]. both [BP10]. Bottazzi [BCC15]. Boulder [Mil91]. Bound [HI16, CMV18, KW21, ORW02, WY23]. Boundary [HH15]. Bounded [CKP16]. Bounds [HLM17, PSS01, DS06b, Giu11, Gon05, GR08, NT22, New90, Shi92]. Box [MM00, YT06]. boys [LO18]. branches [Bai91]. brand [CEK09, FP09, HKS14, PF00]. brand-level [FP09]. Brazil [Gou22]. Brazilian [CGP99, Ras95]. Break [BKKR15, CP97, Gal06]. breakfast [MT13b]. Breaks [BiS15, AG02, LPU22, MG08, Mar00, RS08a, Son14]. Breastfeeding [Cha15, Ade04]. Brexit [DE22]. Bridging [FB02, LU05]. brief [MM97a]. Brisbane [De 93]. Britain [Nei92, ALM22, AB00, CM01, CLM89, MP96, Pud04]. British [Mil89b, CJR04, Gre92, HP04a, KA07, MS90, PS00]. Britton [Nei92]. broad [TE01]. broiler [HM03]. Brookings [Ano90b]. Bryant [Ano90b]. Bubble [Anu15]. Bubbles [AGHKS16, Fu19, BW09, HPS99, JK21, vN96]. Bull [KvD17, PS03a]. bunching [AJL24]. Bureau [Ano95o]. Burgess [BB22]. burning [JS87]. Business [CS15, EV16, GM17, KL05, ÖO00, RV18, ANT94, BCDJ99, CKZ10, DIW00, FOS17b, Gor97, HL96, HJ93, MM03, MMOL08, NC94, Rav97, ST99, SW05, SS05, Söd94, Sor89, TA92a, vDvDF05, vOZ19, McC99a]. Bust [Sum17, CMS24]. Busts [BCRV16]. Butter [Hen93]. buyers [RLS00].

C [Ano90b, Ano90c, Ano90h, Ano03, Klo93, KS91, McC99a, McK94, McK97, Nei92, Sab91, Sim91, Wee01, Edd96]. calculate [LT18]. calculating [KW14]. Calculus [LT06, BFS98]. Calendar [Ano86a, Ano86b]. Calender [Ano86c, Ano87d, Ano87e, Ano87f, Ano87g]. calibrated [Can94, HPR08]. calibration [Ano92d, Ano93b, DM20, MW11, Pag94]. California [MS07]. Californian [Kuo12]. Call [Ano86e, Ano86d, Ano90i, Ano93d, Ano94d, Ano08b, MM95a, Wat94a, Wat94b, Ano94b, Ano95b]. calls [Ano96b]. Cambridge [Abe94, Col92, Dav95, Gan88b, Gan92, Hal89, Hal93a, Imb97, Klo92, Kuk02, McC02, McK98b, Mil89b, Nei92, Sab91, Wee03]. Can [BG94, Gra96, PR04, Zaf11, BK21b, CS21b, DK23, IMN08, KW14]. Canada [BBDF14, CPBS20, MS88]. Canadian [Ber89, KR90, Mor94]. cancer [Bas14]. Cannabis [vO06, vOW11, Pud04]. cannot [CS21b]. Canonical [APVY16]. capabilities [FPV12]. capacity [BGV98]. capita [Bas09, MM01]. Capital [BLS10, SNVCXDE17, Sha86, CMBS15, CLS21, Edw21, HLV02, MSS06, Ped07, Pol13, Pos19]. CAPM [ER89]. car [EHS09]. carbon [DK11]. Card [Rop11]. care [DT97, DMT06, FJ04, Ho06, Hug88, HMKN94, LD22, Pri02, RWM03, WS97, Win04]. Carlo [Gar01a, Adk11, AKS19, Ble16, BH18, CM02, CS99, HS14, Lin92, Moo89, Non16, ORW02, RS08b, TM94]. Carolina [Bal06]. Carrot [VV13]. Cartel [BBS19]. cartels [Orm14]. Carter [Szr94, Vis88]. Case [MKT16, AD06, AF07, BBGS24, BRiS11, BT14, Bil00, BMW90b, BMW90a, BD06, Cha02, DS04, DM06a, DM06b, DW91, Got90, Jus01, JA17, Lee95b, LP08, RR95, Sea90, SS05, Sol05, Wil06a, ZK05]. cased [Mad97]. cases [TM94]. cash [HV13]. catastrophic [MM09]. catch [SV14]. Catching [AJ19, MS15, Orm14]. Categorical [LOR13, MRY15]. category [FP09]. category-level [FP09]. CATS [Tuf98]. Causal [Per98, Uys15, CKT19, FFHL20, HIT20, Hub14, KZ24]. causal-noncausal [HIT20]. Causality [AAQB17, DW16, DWW17, PRS05]. causation [Hil07]. caused [Pee05]. causes [SW86]. cautionary [KOO21]. CBS [VNZ97]. CCE [KUW19, WPN19]. CCEP [Bal10, SS07]. Censored [FP02, MW23, AS14, AM20, DP11, HN89, Pac19, Vel92]. censoring [JL03, Min03, Sak24, Szy19]. centered [Bas14]. central [And13, LFZC23]. century [BL14b]. certainty [GKPP07]. Certificate [CFK95]. Certificate-of-need [CFK95]. CES [Bai91, MP04, Ste18]. Chain [Non16, BW11, FSPWWE12, MT21]. chains [GL14b]. Chalfin [CM24b]. Challenges [Pis92]. Challoner [Fie96]. Change [BDG20, AB00, BP03, CMS24, CH96, Coe96, DGG13, De 92, DLR06, DLRS21, GKP13, IT94, MS18a, Mee90, Nag22, Raj92, Wol93, ZLKH05, ZK05]. change-point [DLRS21]. Changes [FP16, Kra17, KP17, PY15, BFK19, BS04, BP13, BS95, Buc98, BCDJ99, CH11, FSK06, HPS97, Hon13, IS02, KLLO18, MM05b, PS23, Tsi21, YH22]. Changing [GW15, AGM07, EMW11, FH13, Par10, Par11b]. channel [Per98, SS22]. chaotic [FRSRAF05, LGH92]. Chapman [Ano90e, Bol92, Kus91]. characteristic [BGV98]. Characteristics [BSW15, DMP21, GB89, MV06]. Characterizing [TA92a]. Charemza [Cle94]. Charles [Sto09]. charts [KPST11]. Chatterjee [Ano90g]. check [LN14]. Chemical [BEK16]. Chichester [Ano90f, De 93, Nüe91]. Child [JPPS14, JZ16, EF13, HL10, Hug88, LS21]. childbirth [PPSV21]. Childhood [MM17]. children [CKT19, GvSZ05, Sem18, Tro24]. China [BEK16, Dup21, GvSZ05, kHcL16, HCW12, JR02, KH22a, MZK20, Pud93, She13]. Chinese [Sch98]. Choice [BFM15, BVV16, CKM17, DJ16, Gil88, PR16, BF09, BS07, BP10, BL17, BBP03, CM06, CL11, CEK09, CGS19, DS06a, DD21, DR11, Eis19, GM13, GHS05, Hen06, Ho06, HP04b, KW13, Kuk02, LfL09, Li12, MM87, MBR09, PF00, SF09, SW08, SK07, Sti00, Vel92, Wee03, Zaf11]. Choices [CCM15, AP11, CCLZ18, KPST11]. Choosing [CFM14]. Chow [Hal88]. Christian [McK98b]. Christians [CM17b]. Christopher [Sim91]. cigarette [Jon89]. CIRET [Sor89]. citations [MRS10]. Civil [CHW17]. claiming [HSZ04]. Clarendon [Per91]. clarification [Pes00]. Class [Kri15, CMM09, Fév08, KK03, Sue95, WDBR93]. class-specific [CMM09]. Classical [EZB17]. Classification [CP92, War17]. classifications [Kau20]. classifying [HP05]. clearing [Bro87]. click [HM17]. click-through [HM17]. climacteric [CLM89]. Climbing [BJJ11]. Cliometrics [Cra87]. Clive [HP09]. clock [DPR06]. closer [AW96]. cloth [Ano90a, Ano90c, Ano90d, Bai93, Gan88b, Gan92, Gio89, Kus91, Mal88, Mel92, Nüe91, Per91, Phi89, Pis92, Riz91, Sab91, Son89a, Son89b]. Club [Kra17]. Cluster [MW17, Ver16, ALR19, MNW23a, FM19]. cluster-IV [FM19]. cluster-robust [MNW23a]. clustered [FOS17b]. clustering [CMM09, Cre07, FSPWWE12, HPR08, Kau10, LMPTR23]. Clusters [HMS01, Fer23a]. co [DTV96, Pet89, LT96]. Co-integration [LT96]. co-movement [DTV96]. co-ordination [Pet89]. coal [JS87]. coal-burning [JS87]. cocaine [vO06]. code [Rom04]. Codependence [Sch07]. Coefficient [LHJC17, CFK95, CLH04, Hoq91, LOR13, MSS06, Mur12, Ver20, Web95]. coefficients [ABP21, CJ24, CE17, MR06, MBR09, Tsi02]. coffee [Vog92]. cognitive [HL10, Zaf11]. cohort [Dev07, Fuk06, KA07]. coincident [MM03]. Coint [Ng95a]. Cointegrated [KT15, BM06, CH95, Hod99, HR96, Sch07, WH11]. Cointegration [BiS15, BMM17, HPS97, HT21, KN90, Tim95, Wag15, BF05, BdJ07, BD05b, Bus06, Fac07, GHL04, KL01, Lan00, Mac96, MHM99, MN14, NT22, Phi95, SN06, SV96, Tuf98, Tur09, Wes08, Oxl96]. Cojumps [WBL16, LLN11]. Colin [Bar95]. collaboration [CEF07]. collapsing [HPS99]. collar [KPS04]. Collective [DM18, SV22]. College [Wee02, GSS20]. colleges [ERS03]. collinearity [Son89a]. Colorado [Mil91]. Columbia [HP04a]. Combination [LM17, PP16]. Combinations [PR16, GL14a, GVZ19, MW11, SXY20]. Combining [CD16, FL15, HHO21, JMV10, OvDvdW17, JRA21]. Comment [AD06, Got90, Lea91, Sim91, Str09, VD98, XL16, BMW90a, BD06, Cua11, IS02, KS91, Lil00, Pap11, Poi91, Sto91, Wan06]. Comments [LS09a, Tob97, DW09b, Gra02]. commercial [Smi12]. Commodity [BT19, BSB86, CMR15, GP22b, GL17, BM91, Bec01, CFG22, DL95, DTV96, FS01, HPS18, HM03, Lor91, Mel92, Smi05]. Common [CFG22, Cub99, GUW16, HLP20, HK21, Bal10, Fév08, Juo22, KG96, LK04, Sar99, SS07, WZ14]. common-value [Sar99]. commonality [CCM20]. communication [BDNS02, Koo99, LFZC23]. communities [GB89]. Comovements [Djo20, Eic09, HPY19]. comparative [Gan88b, Rot92, SLR97, Woh96]. Comparing [CLM18, CI20, Des08, Han93, ILOS96, JW05, KW13, MvdKH20, SW05]. Comparison [AP12, GVB17, Aka12, AS06, BKA90, BFC02, BCDJ99, CE18, CJZ22, CR88, DG03, EKL10, FFS01, HL05, Jar10, KR90, KPST11, Kon01, Lin92, Med02, Moo89, Ols98, Par87, Ros22, Rus93, Sar99, SW08, Sil96, Yue94]. Comparisons [WCC17, GR10, HS00, MY07, Zhe04, CT86a]. compatible [EB88]. compensating [BJ09]. Compensatory [HO09]. Competing [BHH15, BC07a, HH90, MP96]. Competition [ABH^{+}17, Cam11, Cor12, FH13, HL07, PS07]. Competitive [GL17]. Complementarities [Van07]. Complementary [Gal20]. completion [GR08, Li06]. Complex [Day92]. complexity [BBS97, KP22]. compliance [BBvdH09, EGP11]. component [Bus06, CCM22b, iSG23, CK20, Joc20, WBAB07, Ano90a]. components [Bas09, FMN05, FMS19, Fos17a, GRW06, HJC10, KL93, LK21, MM89, VD98, Wes20]. Composite [CEHK20]. composition [Bie09, Che98, EMW11].

Compositional [Wol93, Hon13]. comprehensive [CSS12, FHKN22]. Compulsory [CJ16, CR22]. Computation [BP03, CT15, CL18, FCP96, Koo99, Kus91]. Computational [Gre93]. compute [BFC02]. computer [Cho11, HH95, LR08, SZ06]. computers [GAA^{+}89]. Computing [Fai96, Kus91, Swa01, Tey98]. computing-numerical [Kus91]. concave [ABH^{+}07]. concavity [Ter96]. concepts [TP97]. Conceptual [Col92]. concerning [MM05a, Vog92]. Concessions [GLS07]. condition [AM23]. Conditional [BW11, CM17a, HM15, MS15, MS17, TH04, BRRT21, Bec01, BDT18, BFS09, DLR06, FS01, HK21, HS21, HV13, LOW17, Lee24, LJ00, MM19, NTK22, OSS18, PL08, Pro14, Thi20, Tse98]. Conditionally [Dov13, BR99, IP12]. Conditions [GW15, CCV23, Far12, FS09, Han92, Han96, Ter96, Woo05]. conduct [MQ04]. Conducting [Rom99]. Conference [Ano91g, Ano92j, Ano94e, Ano95o, Pes01a, Sor89, Ano91b, Kap93]. Confidence [LMZ16, LOW17, OPM19, PR06, Sam91]. Conflict [CHW17, BGLS21]. Confronting [WKN17b]. Congress [Hal89]. conjugate [HHO21, HK23]. connectedness [DDLY18, MHL22]. consequences [DP11]. conservation [CPUBUW24]. consideration [CEK09, Hen06]. considerations [KPP90]. Consistent [BPC15, Koh89, BK08, Dav87a, Dav87b]. console [CNS09]. constrained [Gew86]. constraint [LT96]. Constraints [RTV16, ACS99, AM95, BSB86, CBGIZ19, HI07, ORD01]. Constructing [FP89, GL14a]. Construction [BPW16, HZ19, TA92b]. consume [CMS24]. Consumer [HL07, LMZ16, Mil91, SF09, Tsi21, AGM90, AHT94, BDP98, Cla06, CS91, DMTvdK11, Kal12, KW13, MGSW13, PS10]. Consumption [BM16, EPS17, LMZ16, AV97, AP11, BBS97, Bru18, CMS21, CO96, DD00, EP14, EP22, For95, GP22a, GAR01b, HPS97, JR02, Jon89, KT01, NV92, ÖO00, Osb88, Pud04, SV22, TW17, Wil06a, Bro94]. consumption-based [TW17, Wil06a]. contagion [DR18, GY21]. Contagious [OPS22, Tho04]. content [AM23, BK21b, Moe12, Nyh03]. context [BCCM22]. contingent [BF09]. Continuous [ABFN10, Cha92, FK03, HHLL20, KV09, Arb93]. Continuous-time [ABFN10, Cha92]. contraceptive [CM06]. contracting [Wol07]. contractionary [NPT18]. Contracts [GW15, FJS21, Kie88]. Contribution [BKKR15]. Contributions [Sor89]. Control [GVB17, Gre93, PP16, PS16, BK17, BEK12, Che20, HS90, HT21, Pep05, Gan92, Hal93a]. Controlling [Wil19]. Conventional [Dah17]. Convergence [BD95, CH05, Kra17, LK04, Wan06, Bia97, CL09a, CK13, CF13, GKZ18, HF00, LPS97, Shi06, Sta20]. converging [LK04]. converting [SM95]. cooperatives [HET08]. Copula [VGCBHJC18, BC19, HM12, KLLO18, YHLL21]. Copulas [HHJC16, LMSM18, OSS18, SGK12]. coresidence [JM14]. corporate [BL17, CR13, EZ23]. Corporation [Ano90d]. corps [EO14]. correct [Tur09]. corrected [CT14, MRW09]. Correcting [SE07]. Correction [Ano22a, Ano23a, GUW16, HH15, KOO22, QR87, DH90, GL89, JKLGS13, KR90, Lee04, MKV98, PSS04, TE01]. corrections [CH96, HK21]. Correlated [HL17, Bas09, BH13, HLP20, HK21, Juo22, Kie11, Sti00, Ver20]. correlates [DFM21]. Correlation [DH16, AH99, BH03, Che08, Joc20, LGH92]. Correlations [APVY16, PS03b]. Correspondence [Ano90a, AV97]. corrupt [KP08]. ‘Corruption [AF17]. Cost [MKT16, San15, EH94, FS09, FP89, GN00, HPH07, Hug88, JLR14, Nak93, Oku93, Ras95, Wei96, Xu13, ZR98]. costly [Kau20]. costs [AP11, BK17, Fis94, FJ04, HET08, MGSW13, RS99, SS06]. Count [BPC15, CJ97, HKS16, LRV17, BNP11, CT86a, Cin97, CD97, Hel06, KTR^{+}22, KT01, Lee04, Mro12, Mul97, RWM03, Rom04, VV93, WDBR93, WS97, Mun22]. count-amount [VV93]. count-data [Rom04]. Counterfactual [CS21a, MM05b, Che23, FFG19, KMP08]. counterfactuals [HZ19]. Countries [DK16, And04, AAV07, DF97, DGM11, EP14, FNR08, GM19, Haj94, IRT08, KMS13, SPS13, dHW22]. Country [SKL17, ATW08, BCCM22, CLM18, DJ95, FLS01, ILL06, Jal18, KL93, LPS97, LS99, VD98]. counts [GC91, Tri97]. county [PS23]. couples [ACG04, BG06, CM06]. course [Gil93]. Covariance [CKM17, CPA15, HKO12, JMY19]. covariances [ER89]. Covariate [SSX22, LSV23, NO21]. Covariates [EGH15, Pac19]. Cover [DDL23a, KH22b, MNW23b, Bro94]. coverage [LP23, Ols98]. COVID [BHLS23, KH22a, PY22]. COVID-19 [BHLS23, KH22a, PY22]. Cox [MM00, TP97, YT06]. CPS [CP92]. Crack [HH02]. crash [HKW11, KK96]. Cravino [Gou22]. creation [EHP12]. credibility [Ash09, FGV18]. Credit [AGHKS16, DK16, FRF15, Gal96, RTV16, SKL17, Sum17, AHKM19, GS20, KL05, MZK17, RV18, Sch05, WVG14]. Crespo [Pap11]. crime [Bal06, BK17]. Crimes [LRV17]. Crises [AGHKS16, Dah17, War17, DM07, Fu19, GQZ23, Haj94, IT94]. Crisis [ACCM17, HHJC16, LO18, OSS18, YHLW06]. criteria [And03, BM86]. criterion [CD18]. critic [Lea91, Sim91]. Critical [Kul15, Tur09]. criticism [Lea89]. criticize [KS91, Lea91, Phi91b, Poi91, Sim91]. critics [KS91, Phi91b, Poi91]. critique [LS10b]. Cross [Ano16, BKP16, BHP16, BiS15, BCC15, DH16, EM17, RW16, Ver16, AH99, BP07, BCDJ99, Byr92, CLM18, DJ95, FLS01, FN23, GRLM08, Hei17, HK21, ILL06, Kap07, LS99, Nis05, Oku93, Pes07b, Rac97, Wes20, MSS16]. cross-country [CLM18, DJ95, FLS01, ILL06, LS99]. cross-Euler [Nis05]. cross-national [BCDJ99]. Cross-Section [BiS15, RW16, AH99, Pes07b, Wes20]. Cross-Sectional [Ano16, BKP16, BHP16, BCC15, DH16, EM17, Ver16, Byr92, GRLM08, Hei17, HK21, Kap07, Oku93]. cross-sectionally [FN23]. crossing [DMM05]. Crowd [TY15]. CRS [HHJC14]. crude [AK10, BP13, CNT22, GM91, KM14a, LN14, Pak18, RS21, SW20]. crude-oil [GM91]. Crystal [BKL17]. Cuaresma [Pap11]. currency [BMV00]. current [SPS13, BMW24]. curse [RS09]. curvature [Moe12]. Curve [HGRSS15, Kri15, LN07, Wag15, AGR17, BÇCV14, CEG22, DIL14, DNRS23, FS23, MS09, NS08, QR86, QR87, Una06, WX20, CKP16]. Curves [BD15, CLM18, DP12, DM98, FFG19, FP89, Kee86]. customers [ANT94, AFHI95]. Cuthbertson [And93, Mad97]. Cycle [EV16, GM17, AM95, AAQB17, BCDJ99, CLH04, CKZ10, DIW00, EJ20, Gor97, HL96, HP05, HJ93, JMM07, LK04, MMOL08, NC94, NV92, ÖO00, Osb88, Pud93, Rav97, Söd94, Sor89, vDvDF05]. Cycles [CS15, CH05, Cub99, FOS17b, KL05, MM03, RV18, ST99, SW05, SS05, TA92a]. Cyclical [ALM22, Söd94, Web95, Ber01, Han93, KvD06, Mor94, PS03c]. Cyclicality [KDK18, vOvGvGB19]. Cygwin [Rac00].

D [Ano90b, Cle94, Gre93, Kuk02, Mel92, Oxl96, Per91, Son89a, Wee01]. D.M [Gio89]. Dahlberg [DM06b]. Daily [BPW16, BMV00, Sok16, ABFN10, Fei87, RTÅ98, Wat99]. dairy [MS95]. Dakota [BNR21]. Damodar [Mad98]. Danger [War17]. Danzer [Tro24]. Data [AG17, Ano95i, Ano95c, Ano95d, Ano96c, Ano96d, Ano96e, Ano96f, Ano16, BiS15, BPC15, DH16, EM17, GVB17, GK17, HKM15, HKS16, KP16, LRS17, MM17, MB16, Ver16, Abr06, AS14, Aka12, AAB09, AW96, AHT94, AP11, BBGS24, BP06, BKA90, Bal06, BP07, BM14, BCM24, BÇCV14, Bil00, BDG20, BC07b, BL06, Bru18, BH21, CT86a, CJ97, CJ24, CGP99, CS14, CL23, CF94, Cin97, CT03a, CG13, CG23, CT03b, CR88, CP92, CD97, CP97, Day92, DLR94, DM13, DG03, EMvS98, EP14, FMN05, FZ12, FL94, FK03, FSK06, Fuk06, GO22, GA19, GPA89, GQ10, HLW14, HT21, HPS23, Hei08, Hel06, Hil06, HLMvS98, HSF05, HCW12, Isa04, Jag91, JK21, JRA21, JL03, JS10, KY10, Kas98, Kau10, KLT20]. data [KW89, KW21, Kon04, KLS19, KP08, KS19, LZ08, LfL09, Liu09, LMSM18, LP08, MM97a, MC90, MSW13, Mro12, Mul97, MSP21, MBR09, NV92, PV89, PR04, Pri02, PH24, RRB92, RWM03, RS92, Rom04, Sak20, Sch90, Sea90, SW13, SW18a, SE07, Sil01, SM95, SLKN04, Son89a, SW21, Sue95, Ver20, WPS18, WVG14, Wea92, WDBR93, WS97, Woo05, WL86, XQ12, Zaf11, AB16, Per10, PP16, EB88, McC02]. data-arch [RRB92]. data-fit [EB88]. database [CPBS20]. datasets [BM14, CKM11, CLM18]. dated [TZ22]. Dating [Kau10, Lee16, BBS19]. David [Abe94, Imb97, Kuk02]. Davidson [Smi95]. day [LS93c]. DC [Ano90b]. Deadman [Cle94]. Deal [dRS10]. Deaton [Bro94]. debate [Brü13]. Debian [Mac99]. debiased [MWWW24]. Debt [TY15, BP18, Dav05, DZ18, Haj94, SZ91]. decade [AB00]. decide [LGH92]. Decision [HMS16, PR16, dRS10, BP10, GRV00, Kap93, KKR95, Pro95, RS99, Van11, Sto09]. Decision-Based [PR16]. decisions [ACS99, BP10, DMM05, DM18, DK13, Hoc03, KV10, MSL89, RT88, VB02, VV14]. Declining [FLS22]. Decomposing [Lak19, RR18]. decomposition [DMP00, Fuk06, LK04, MM05b, Mor08]. decompositions [CMV18, KvD06]. Deep [KSN23]. Default [EPR11, Kie11, KL05]. defaults [Kie11]. deficit [Mar00, TW87]. Deficits [KKS16, SZ91]. deflation [Kau20]. Dehli [Ano90d]. Delineation [BCMM16]. delta [FFG19]. Demand [APP15, CDKS17, CS91, DT97, Mil91, MRGB16, RTV16, SSS16, VGCBHJC18, AHT94, BL91, BdJ07, BS08, BIKY07, Bey98, Bla96, BDP98, BR99, Cha92, CS14, COQD19, CV01, CH11, CNT22, CT86b, DPU97, DU98, Des03, DH90, EH94, FFS01, Fle97, Hel06, Her19, HSF05, HMKN94, IZ03, Kal12, KR90, LD22, Lew03, LTW99, LN14, MM88a, Mil89a, MB02, Ng95b, Nis05, PS10, RH21, RWM03, SS91, SLR97, TE01, Una06, VNZ97, Vea86, WY23, WS97, Wol93]. demands [BSB86, BC07b, Che98, MRS87]. demeaning [PW20]. Democracy [LHJC17, KW19]. democratic [Wri13]. Demographics [EV16]. Denis [HP01, Phi89]. Denmark [Mor23]. densities [JMV10]. Density [AFR17, CG23, PMM17, Wol15, DM20, GL14a, Jal18, KOO21, LL06, MW11, MW23, OvDvdW17]. Dependence [BKP16, BiS15, BCC15, DK16, EM17, Leu22, Ver16, BP07, BBP07, BEH04, Bie09, CT14, CCV23, CD05, DGM03, DKB10, FLS12, Hei17, HK21, HPY19, JMPW18, JM98, Kap07, Pes07b, RvdB13, SGK12, Ano16, BHP16]. Dependence-robust [Leu22]. Dependent [MSS16, CLP22, Dej93, Fac07, FN23, Lee95b, Smi87, Zhe04]. depends [Sak20]. deposit [Lin95]. deposits [SS22]. depreciation [Pol13]. Depression [kHcL16]. Deregulation [HST16, Lin95, RTK14]. derivatives [FCP96, Hen10]. Deriving [DPJF06]. Desai [Phi89]. describe [SS05]. Describing [BL91]. description [MM97a]. Descriptive [Phi01]. Design [MM97b, SSS16, BT19, CCE23, DGM11, Jal18]. designs [BBvdH09, HS21]. despite [BT19]. destabilizing [FLM13]. details [DKT12]. Detecting [AG02, HPS99]. detection [HLST21, Son89b]. determinant [LS09a, Str09]. determinants [AP12, BL17, BO94, DW09a, DW09b, EPR11, FZ12, HGK18, MS93, MT16]. Determination [APVY16, VV98]. Determining [RLS00]. deterministic [Gor97]. Detrending [HJ93, PS03c]. deutsche [MM88b]. devaluations [Gou22]. Developed [ABT15, Djo20, Djo24, FLS22]. Developing [RTV16, DGM11, Haj94, Jal18, KL93, VD98]. Development [MC90, MM17, SNVCXDE17, WKN17b, BG18b, CKT19, HET08, HL10, KÜ24, LS21, MSS06, SW18b, Tan10]. development-The [KÜ24]. developments [KW19]. Devereux [DM06b]. deviations [KZ02]. Diagnostic [PV89, DG92]. diagnostics [ACF22, Son89a]. Dichotomous [BF09]. Dichotomous-choice [BF09]. Dickey [Kul15]. Did [CMS24, Edw21, Nag22, She13, And03]. didn’t [BD05a]. Diebold [KW14]. diet [Che98]. Difference [BG18a, And13, BFS09, Fer23a, Hon13]. Difference-in-differences [BG18a, BFS09, Fer23a, Hon13]. Differences [EZB17, TWW89, Ano03, BFS09, BG18a, Fer23a, GLUW03, Hon13, MS93, Sch98]. Differencing [CJ20]. Different [DD23, EHSJ17, MW17, Cua11, HMKN94, Kon01, PF00, Pat06]. Differential [MSN86, Pet89]. differentials [BJ09, Car00]. differentiated [Kal12]. Diffusion [EM17, ALR19, BN09, BG18b, Bra11, CFK95, MRS10, SP18, SXY20]. Dimension [Ano90a]. Dimensional [HKM15, FGLS18, HKO12, Nib24, NO21, PSW19]. Diplomatic [AF17]. Direct [HHLL20, Che23, MM19]. Directed [DH16]. Directional [MKT16, CH07]. directions [Cle94]. directive [SKG21]. Dirichlet [Cam01, CLM18, HMT15]. Disability [BBG04, BSBC^{+}04, KP08]. Disaggregate [Cla06, DS04, HSF05]. disaggregated [BR99, Kas98]. Disaggregation [Mil92]. disagreement [KK22, LS10a]. Disasters [Brz15, FIL13]. discipline [CML04]. disconnected [Ber14]. Discontinuity [CJ16, Don15, CCE23, CTU13, HS21, Jal18]. discount [Cho99b, FLS22, MP01, Tim95]. discounted [Dav05]. Discounting [GKPP07]. Discrete [GHS05, BL17, Cam01, CGS19, Eis19, HHP92, HP04b, KSD98, KLLO18, Koo94, MT00, Mro12, Sil01, SK07, Sti00, Vel92, Wee03]. discrete-choice [BL17, Eis19, SK07]. discriminate [SdV18]. discrimination [SZ06]. Disease [KÜ24]. diseconomies [Whe03]. Disentangling [LN14]. disequilibrium [Bro87, HHP92, QR86, QR87, Sal91]. disk [McK97]. dispersion [CK14, RRB92, VVR00]. disputandum [Phi91a]. Dissertation [Ano12b, Pes07a, Pes09b, Ano10b, Ano11, Ano13b]. Distance [EP04, MKT16, BG18b, CT02, DK12, HET08, Sci23, TW21]. distance-to-market [HET08]. distant [GKPP07]. Distinguished [Pes08, Pes13, Ano06, Pes01b, Pes10a, Pes11]. distortion [GK12, SS98]. distorts [Wol18]. distributed [HHJC11, SS23a]. distributed-lags [SS23a]. Distribution [BD05b, Brz15, MC90, AHM21, BV12, Buc98, BCDJ99, CLM18, DSW01, FB02, FJ04, GT06, HV18, HD91, Kom07, Mac96, MHM99, MN14, MM00, MM95b, Par10, Par11b, PS23, SSX22, Thi20, YT06, Yat00]. Distributional [FP16, Kra17, ABFN10, BE92, BM12, Che20, EO14, For95, KSN23]. distributions [BKN95, BH21, FNR08, Fer17, HJC10, JLR14, Lee24, LJ00, MM05b]. disturbances [CB17, IP12]. Divergence [Cha02]. diversification [MZK17]. diversify [She13]. diversion [EHP12]. dividends [Lof19, PSS04]. Divorce [KO14]. DM [Car92]. Do [BBvdH09, BBDF14, BB22, HKM15, LPT04, MKP22, NPT18, Sem18, AK10, FSK06, Hen06, ILL06, Kil99, Liu09, Pak18, SS05, SV14, VV98]. doctor [Van11, Win04]. Does [BFS98, GM19, GW15, HKMR03, JM14, KZ20, KP22, MRW09, MM89, TY15, Tro24, ASS21, BMW24, BZ98, GKPP07, GQ10, HL05, JS11, NS93, Par10, Par11b]. doing [Pra10]. Dollar [CI91, GHHP93, LT91, Tse98]. Domains [Sal15]. Domesday [Per91]. Domestic [AGMA22, PMM17, MT21]. dominance [And03, MY07]. Donald [Fie96]. dont [BF09]. Doppelhofer [Str09]. double [GY21, Jon89, LL96]. double-edged [GY21]. double-hurdle [Jon89]. double-threshold [LL96]. Doubly [LOW17, Uys15]. down [KZ20]. downturns [Tsi21]. dramatic [BMW24]. drastic [KH22a]. drawing [KZ20]. Driehuis [Pis92]. drifting [CCM21]. drive [EHS09]. Drives [ABT15]. driving [DT20, Kuo12]. dropouts [LPT04]. dropping [LPT04]. drug [BJJ11]. Drugs [SSS16]. drunk [DT20, Kuo12]. DSGE [And13, CHL14, CFM14, CSvD14, EKL10, FGLL19, FP10, GQ10, HS14, HI16, Hur17, ORM13, QT23, Sal15, Sch00, SW05, WCC17, Wol15]. Dual [BEK12, GH11, KT11]. Duality [Kal12]. dummy [HL19]. durable [CS91]. durables [And91]. Duration [CD05, RvdB13, Van17, WKN17b, Ahn23, Bel01, BE92, Bil00, CTU13, Cam01, Dej93, FPS94, HH90, Jag91, Rom99, Sak24, Sue95, WVG14]. Durbin [fLY16]. During [Dah17, FG16, BHLS23, CFJvdB18, DM07, KH22a, SW08]. Dutch [GPW17, HVW11, MAL04]. Dykstra [Nüe91]. Dynamic [Ble16, BR23, BBG04, CB17, FS19, FGLS18, HM12, Hal93a, HGRSS15, HL10, HP21, Kap07, KvdK22, KLLO18, LLP23, McK98b, MM17, RSU22, Ver16, XL16, Aka12, ACF22, BKKS20, BFS09, BIKY07, BD05a, CMBS15, CM06, CFG08, Cha92, Cha11, DJ00, DR93, DKP22, DLR94, Des03, DH90, EPW07, Eic09, Eis19, FVRR05, Fer17, Fle97, Gal88, HLP20, HH20, HPS23, Hey04, JS10, JKvdW14, KY10, Kap93, KKR95, KPS04, Kau10, KS19, pLK20, LD22, LZ08, MRS87, Mil89b, Nak86, PF00, SW13, Sti00, Tsi06, VV98, Wol93, Woo05, ZLKH05]. Dynamics [Ano91g, APVY16, Fuj11, Gre93, HP04b, Smi12, AV01, ABC21, AG02, BDKK12, BNP11, BMW90b, BMW90a, Buc98, Bur92, CGK14, CLLY06, CJR04, DLOZ24, DK12, Day92, DD19, DN89, DSW01, Eva89, FP11, FRSRAF05, FJ04, FSPWWE12, Got90, GT06, Hil06, Hos12, Kas98, KHS14, Lan06, MT11, MS09, MSP13, PS23, PP92, PZ06, Sen11, SN06, Smi05, Ste07, Sto94, SW20, vDvDF05, GKWZ16].

each [Klo93]. Early [DL22, Gra17, LS21, MM17, Pee05]. Early-life [DL22]. Earnings [CM01, BBGS24, FHKN22, FSPWWE12, HSWW07, JRA21, Min03, Nag22, NT22, PY96, Tre98, Yue94]. east [Jar10, Pro95, BFS09, BHMW98]. East-West [BHMW98]. easy [DP13]. Easyreg [Sep98]. EBA [Shi92]. EC [APVY16]. ECB [ARST16, WBL16]. EconLit [Ros22]. Econometric [Ano91b, Ano92j, Anu15, AFHI95, BD03, BPC15, Bri89, CT86a, Cha00, CN97, Dav95, DW96, FHP16, Hal89, Imb97, Lil00, Mal88, McC99b, Mil92, PW96, Phi89, Tri97, Vis88, Abe94, And93, Ano93b, Arb93, Bai03, BBS97, BDNS02, Bro87, CFP09, Cin97, Cle94, CNJ97, CNZ99, Dav87a, Dav87b, DW88, Din91, EB88, GT06, HHJ86, HS90, Hor92, KvdGVW97, Klo93, Koe88, KZ09, Koo94, Lar90, MP01, McC02, MR09, MT11, MP10, Mrk01, MvdKH20, Oxl96, PU88, Pag94, Pud04, QR86, SC90, Ter88, VK87, Vea92, Win04, WL86, ZLKH05, Car88, Mel92]. Econometrics [Ano91g, Ano01d, Ano01e, Ano06, Ano07b, Ano10b, Ano11, Ano12b, Ano12c, Arb93, BEL07, CS09, DG03, DIM08, Edd00, Per10, Pes01a, Pes01b, Pes01c, Pes07a, Pes09b, Pes09c, Pes10b, Pes11, Rez08, Wat07, Wee02, Ano90i, Ano92d, Ano13b, BD02, Bai09, Bal99, BLM10, BEK12, CGR04, Edd96, Fie96, Hal89, Kap93, Kid92, Klo92, MM95a, MS06, PP92, Pes03b, Pes10a, Phi01, RH02, Rob88, Tri93, Ano92j, Ano95o, McK97, Ano90c, Ano95i, Ano95c, Ano95d, Ano96c, Ano96d, Ano96e, Ano96f, Ano03, Mad98, Ano90h, Gil93, Gre93, McK97, Smi95, Szr94, Wee01, Abe94, Ano93c, Ano08b, Hal88, Pes08]. econométriques [Klo93]. Economic [Ano95o, CPQP15, CF94, CM17b, GMRZ97, GW15, Gre93, Hal93a, KTR^{+}22, KH22a, LO18, Leo98, MSS06, PS09, SHM17, BW11, BF94, Bal06, BK17, CCM22a, CPBS20, CCV23, CSS12, CM09, CT02, Day92, De 92, DRT21, DKT12, FZ12, GH96, GRW06, Gra96, Gre92, HGK18, Ham21, HR20, HS00, HCW12, Isa04, JL24, Kal90, KvD06, LMST22, LT06, LGH92, McK94, MVW92, ORD01, RV22, RR18, Tan10, TZ14, Tsi21, WG23]. Economica [Klo93]. Economics [Col92, Ger88, Mad97, McC99a, BLM10, Gre93, LSV23, RV22, Sor89, Wee01, Bol92]. Economies [ABT15, MZK17, Djo20, DO14, DFM21, FVRR05, LS93a, Whe03, WW18, Yat00, Ano90b]. economists [YL08]. Economy [BEK15, RTV16, ASS21, ABS93, Ano90d, Dav87a, Dav87b, Djo24, DFM21, Hua09, JW05, JJVW01, JP10, LP23, MVW92, Nak86, Nym91, Per91, Sen11, SS23b, ZDE12, Mil89b]. ECU [FB02]. ed [Ano90c, Ano90f, Bar95, De 92]. edged [DMW94, GY21]. Edited [Ano90b, Fie96, Gre93, Hal89, Hen93, Ano90a, Ano90e, Mil89b, Mil92, Phi89, Sor89]. edition [Ano90h, Car88, Col92]. Editorial [Ano91a, Pes86, Pes00]. editors [Pis92, Wee01, vDvDF05]. eds [Bol92]. Eduardo [DW09b]. Education [KS16, BH03, FHKN22, Giu11, KV10]. educational [AJ19, BH20, Fio12, Isa04, JPPS14, MM89, Pos19]. Edward [Bar95, Cle94]. Effect [Lee16, MS15, Abr06, BT23, BKS18, DK11, EF13, GMS11, HL07, HR20, HK21, Hon13, HL10, KT01, Lee04, LS14b, LOW17, LS09b, MV06, MMOL08, MT11, NS93, Ols98, PI09, Sak20, Sam91, Wes08, BLV16]. effective [ALV13]. Effectiveness [AG17]. Effects [APP15, BBE15, FHKN22, HL17, HMT15, HLM17, HST16, Sav16, Uys15, AM20, Ala87, ALR19, ABS93, ABKK13, Bas14, BCO04, BFS09, BDG20, BL17, BSS08, BR23, BH13, CLSM10, CCM20, CT14, CK02, CNS09, CPY18, CPB19, DS04, DD21, DDL23b, DE22, Dov13, DK13, DKB10, DJK18, EPW07, EHP12, ERS03, FG18, For95, FP02, FFHL20, Gal96, GvSZ05, GHHP93, GP00, GHOS04, GQ10, HLP20, HKS14, HPS23, Ho06, HV18, HS21, HL10, HHLL20, HSZ04, IZ03, Jal18, JMM07, JV22, Juo22, KA07, KLT20, Ker19, KPST11, KMP05, KO14, KZ24, Lak19, LP23, jLjL05, LS14b, LLP23, Lee24, LMPTR23, Li12, LLD13, LMZ23, LS93c, LB19, MP96, MR06, MT21, MT13b, MU09, MSP21, NR19, Orn06, PF00, PW20, RvdB13, RWM03, RSU22, Sak20, SS23a]. effects [Sli23, TK14, VVR00, Vea86, Ver20, ZK14]. efficacy [KLNT24, KRS10]. efficiencies [SZ07]. Efficiency [Fei87, FS09, GMH95, IRT08, MS07, Ver16, CL23, DK12, FH12, GS07, GM91, Han93, Kal90, KPP90, KH95, New90, PMM96, MSS16]. Efficient [CE17, CR88, Hei17, Hil07, MSW13, TW21, BKA90, BS95, DG03, FLHS07, IS02, MS18a]. efficiently [HLV02]. EGARCH [CK02]. Egypt [Cha15, Sea90, TA92b]. Egyptian [GS07]. elasticities [BP13, CGPX21, PKH12]. Elasticity [Bai91, CDKS17, GP11, AJL24]. elderly [AHM22, ABC21, DT97, JM14, LK09, MAL04]. electric [JS87]. Electricity [HST16, AG89, ANT94, AFHI95, BHP13, BPPB10, KSN23, Lee95a, MS07, Wol07, Yat00]. Elements [Kus91]. Elgar [Bar95, Cle94]. elicitation [DGM11]. Eliciting [DGM11]. eliminate [BBvdH09]. elliptical [HLV02]. elsewhere [PW20]. elusive [Day92]. Embarrassingly [DP13]. emergency [LD22]. emerging [Djo20, Djo24, YHLW06, ABT15]. emissions [DK11]. Empirical [Ala87, BS16, BBS97, CEF07, Dah17, GVB17, GK17, MR16, OW10, PC15, PZ06, ACS99, And91, ABS99, BL91, BKA90, Blo02, BBvdH09, BLS10, CL09b, Cho11, Cho88, CS99, CR88, CFJvdB18, DMM05, DLR94, DPR06, DG03, DM03, ERS03, EMP14, FPS94, FMN05, FFS01, For95, GMH95, HP01, HM03, JP10, Kie88, KP94, KU02, KMS13, LPS97, LN07, LfL09, LT96, LWY00, LS10b, LB19, MS98, Mar01, MP01, MM19, MS95, Mof93, MM87, MP07, MSP13, OSS18, Pak18, PMM96, Phi01, RS86b, Ros93, She89, SW86, Sor89, Sto94, Ano90b]. empirically [MvdKH20]. Empirics [MB16]. employed [Gil88]. employees [BMW90b, BMW90a, Gil88, Got90]. employing [And03]. Employment [GJSvN23, Kie88, BFS09, Bie09, Bur92, EPW07, EF13, Gil88, IMN08, LS02, MRV95, MSL89, MT11, Nak93, RS86b, RvdB13, Rid86, Sem18, Sli23, Ste07, Ter02, TA92b, VV13, WL86]. empty [FLHS07]. EMS [De 94, KSD98, Miz92]. EMU [SPS13]. Encompassing [EP22, CH10b, HN00]. Endogeneity [FFHL20, HLM17, WS97, WKN17b, DMT06, KT08, LK09]. Endogenous [ABC21, BD15, BCMM16, CT15, HL16, MR16, PY15, Szy19, BFK19, CPUBUW24, CHP23, EWB09, HL19, KT01, KG96, Kuo12, LLD13, Mar00, Sak24, TW87, VV93]. endogenously [DPJF06, FOS17b]. endpoints [vDKvdWW14]. Energy [IT94, BIKY07, BDKK12, Bol92, HH02]. Enforcement [AF17]. Engel [BD15, DP12, DM98, FP89, Kee86, Una06]. engine [Fer23b, Ken94]. England [Sor89]. Enjoying [RTK14]. enrollment [GSS20, LS14b]. enrolment [Fio12]. Ensembles [War17]. entropic [ABH^{+}07]. Entry [AG13, FSPWWE12, Xu13]. environment [Bai09, CNJ97, CNZ99, GS07, KLS19, Tey98]. Environmental [MR16, Wag15, CS91, JS04, MM05a]. environments [CN97, GR10]. epidemic [TP97]. episodic [Raj92]. equal [CM14]. equality [BH21, ORD01]. equalization [KN93]. equation [BH18, Czi04, KPT05, NT22, Nis05, RPSF05, SP04]. Equations [De 93, AAB09, IP12, JV22, KL93, Lee95b, MM88a, VD98]. equilibria [AFR08]. equilibrium [BSS11, Ber89, BG94, BKN95, Des03, FVRR05, Gal88, KL01, KRV95, Raj95, Smi12, VVR00]. Equity [BLV16, LG23, DM11, Fis94, GH11, HLST21, NSA98]. Equity-premium [LG23]. equivalence [Pen05, SS23a, SSW06, Wil06a]. equivalent [GKPP07]. Erasmus [Ano92j]. Erik [Wee01]. Erratum [Ano92c, Ano03, Ano09c, Ano09d, Han96]. erroneous [Kau20]. Error [GUW16, HHM15, AH99, BD05a, Cif88, CGS19, DH90, DK23, GL89, JRA21, JKLGS13, Joc20, JPPS14, KY10, KR90, KL93, KT11, MKV98, MM89, PSS04, RdM18, SC90, TE01, VD98, WBAB07, Wea92, Wee01]. error-component [Joc20]. error-components [KL93, VD98]. error-correction [MKV98, PSS04]. Errors [Don15, ALL24, LK09, Orn06]. escalation [BGLS21]. Essays [Fie96, Abe94]. est [Phi91a]. establishment [KLS23]. establishments [MS90]. ESTAR [PP06]. Estate [NS17, Hon22].

Estimate [PY15, PMM17, CFM14, LP22, LO08]. Estimated [FM16, HI16, DS12, Fos17a]. Estimates [EZB17, HMKN94, SSW06, BH03, BBG04, CKT19, CLM18, DM13, DT20, Fai96, Fer23b, FCP96, GVS23, HET08, Hon13, JL03, KR90, KMMP20, MRS10, NV92, Ols98, PR04, PY96, Pra10, Wol18, HLW14, Wol93]. Estimating [Abr06, AAB09, ALR19, AF07, AHT94, AT06, Bal06, BNS02, Bas14, BK17, BT14, Ble16, CMS21, Cho99b, CMM09, CD97, DHR11, DK11, DP11, DP12, DL95, DDLY18, EH99, FVRR05, GKWZ16, GP11, GA19, GL17, HMS16, HGRSS15, HV18, HMT15, Isa04, Jal18, KY10, LP08, MW20, MRGB16, ORD01, Sal99, Sea90, Smi93, Wei96, WM01, Aka12, BS95, DS91, DJK18, IS02, KLNT24, MRS87, Pro14, Vea86, iKLS19, LLD13]. Estimation [AS14, ABH^{+}17, ABP21, AS06, BKP16, BD15, BR99, BKN95, Cha92, CPA15, EGH15, FFG19, FP16, FLS12, FL15, HL19, HH15, HHJC16, HR12, Koo94, KS19, KP17, pLK20, LRS17, MZK20, MP96, MT13b, Min03, Pat06, PT13, Sak20, SW13, Uys15, VGCBHJC18, Ver16, Wea92, XL16, ABH^{+}07, ABS93, AHKM19, AG96, AIK99, AD09, BdJ07, BM91, BKA90, Bal10, BEK22, BM14, BB19, BSS11, BJM^{+}08, BK08, BDP98, BDG20, BBP03, CHP24, Cam01, CM02, CE17, Cha11, CK14, CEG22, CLH04, CPY18, CR88, Doh13, DS06b, DG03, FT90, Fer17, FO18, Fév08, FM14, FP10, GP06, Ger96, GQ10, Gur97, HM12, Haj94, HH90, HLP20, HST10, HKO12, Hei17, HHJC11, HHJC14, Hod99, HJdP20, HSF05, IP12, JV22, Juo22, KK97, Kal12, KUW19, Kee86]. estimation [KK03, Kie11, KL93, KPP90, Koh89, Kum12, LS93b, Lee95b, LMPTR23, LK21, Lil00, LB95, Mar01, MS93, MGSW13, MBR09, Nak10, Orn06, PHK07, Pen05, PS10, PI09, PS00, RH21, RWM03, RPSF05, Sch98, Sen99, SS07, SZ18, Sil96, TZ14, TW17, TW21, VD98, WZP19, Wil06a, Xu13, YHLL21, vHBB22, LPT04, Dav95, Smi95]. estimator [AKS19, AJL24, BR99, CT14, DG03, Mro12]. estimators [AAB09, BKA90, BD99, CT86a, CLM18, CL18, CR88, FL07, IMN08, KPST11, KG96, KOO21, Lin92, Moo89, RS92, SS06]. Ethic [CM17b]. Ethiopian [HET08]. Ethnic [Pos19, Sch98]. EU [Bal10, CM09, DF97, DRT21, PZ06, SS07]. Euclidean [HM13]. Euler [AAB09, Nis05]. Euro [CR13, FB02, BL06, CPQ10, CV01, DdMPS07, DO14, Eic09, FMMR18, SW05, SW21, CS15, CH05, FS17, LSZ17, PMM17]. euro-sting [CPQ10]. Euro-zone [CH05]. Euromind [PMM17]. Euromind- [PMM17]. Europe [BCM24, CGK14, CF13, GS20, Jar10, MT11]. European [BPPB10, Jus01, LK04, WZ14]. Euroyen [Lin95]. Eurozone [BCRV16, CGPX21]. evaluate [GP06, MvdKH20]. Evaluating [BFS09, CT03a, HSWW07, Kuo12, MW11, ORW02, Wol15, Wri13, Bas14, LS14b]. Evaluation [FL15, GVB17, Bek99, Che20, CH07, CH10a, CJS07, CFJvdB18, CBGIZ19, DHKB22, FFHL20, FHL04, FP10, HCW12, HW10, JMM07, JM10, KvdK22, KLNT24, Li99, MW23, Mof93, Par87, RS86a, Rid86, Sch00, Sto94]. evaluations [SF09]. even [HK22]. event [AH99, Rid86, SS23a, Tri97, WVG14]. Events [Ano86c, Ano86a, Ano86b, Ano87d, Ano87e, Ano87f, Ano87g, GO22, HT21, Tho04]. everything [Pep05].

Evidence [ARST16, AF17, AG17, BP18, BNR21, CF94, DK16, DFM21, EK87, EZ23, GPW17, HMSU23, ILOS96, JL24, KLT20, Lee16, MZK20, MS17, MSP21, RRB92, Sol05, TA92b, Whe03, dHW22, Ala87, AV97, BL14a, Bas09, BÇCV14, Bek99, Bia97, BC23, BLS10, Bru18, CL09a, CGK14, CM02, CS14, CL23, CLS21, Cho88, Cin97, Cla06, CLM89, CEF07, DST14, Dem89, EPW07, EK07, FNR08, FPS94, FS09, FL07, FJS21, Fuj11, Gil88, Gou22, GP00, HPR08, Hod99, Iva92, JL08, KW21, KU02, KRS10, KMS13, Li06, Lui00, MP01, Oku93, PY22, PMM96, PZ06, RS86a, Raj92, Rop11, Scr22, Sha86, SP18, Shi01, SW18b, TL03, YHLW06, ASS21, BP06, BBC20, BBDF14, CP97, DJK18, ILL06, KP22, Nag22, Pee05, RS09, SPS13, Zaf11, Gan88b]. EViews [Lil00, McC99b, MT07, MT12]. evolution [BH20, WW18]. evolving [NS21]. EVT [BDT18]. Ex [She89, HM03]. Ex-post [She89]. Exact [CL18, Gew86, AAB09, RH21, Shi01]. Examination [LHL17, AW96, Ber89, Coo09, DJ00, LT96, Sch90]. Examining [Cha15]. example [CGS19, Vea92]. examples [Alb03]. Excess [BGV98, EPS17, DTV96, JA17, Mul97]. Exchange [BKKS20, GMH14, HMSU23, Kil99, Li99, AL06, BCO04, Ber14, BMV00, CI91, Cho99a, CC07, CH07, CLM96, De 94, DN89, EFR11, FO95, GHHP93, HMW97, Hal87, Hod99, HD91, JA17, KHS14, KL95, LMZ23, Lui00, Miz92, MSP13, PMM96, RS08a, RV23, SN06, SW86, SLKN04, SPS05, Tse98, TH04, WFBM01, vN96, Mad97]. Exciting [MFM17]. exclusion [Kra12, MT13b]. exertion [RLS00]. existence [MR88]. exit [ALV13, CD05, VV13]. Exogeneity [MB16, Smi87]. Exogenous [AF19, Ano23a, BD05b, CPY18, HIT20, Koh89]. expand [NPT18]. Expanding [AHM22, Dav05, ZK14]. expansions [CJ97]. Expectation [LN07, BM11, Sut23]. Expectations [KP17, AGR17, DL95, DGM11, DM11, DNRS23, FT90, Gan92, GH11, GV87, HM03, HKW11, HVW11, Iva92, KR90, KZ22, LfL09, Mad18, SM95, Tsi21, Zaf11]. Expected [Lof19, Cha00, GRLM08, KMG22, WL86]. ExpEnd [Rom04]. Expenditure [And91, AGM90, CGPX21, DMT06, GM13, GvSZ05, JL03, Lea97, PS10, Smi89]. Expenditures [BD15, AGM90, Kee86, MV96]. experience [KH22a, SV14]. experiment [AFHI95, Bil00, DE22, GvdK12, MM95a, MM97b, MM97c, RS09, Col92]. experimental [HV13, LSV23]. Experiments [BFM15, Adk11, AG89, BHLTM24, BD99, Hen06, LKL12, SF09, SW08]. expert [CJS23, Kie11, MST12]. Expertise [HMS16]. experts [FSPWWE12]. Exploiting [SSS16, SdV18, SV05, BK21a]. Exploring [DdMPS07, KW14]. Exponent [BKP16, LGH92, TW21]. exponential [JV22]. exponential-regression [JV22]. exponents [DG92, FRSRAF05]. export [Cut86, LS93a]. exporters [EMP14]. exports [BCO04, MZK20]. exposure [DL22]. extended [BÇCV14]. Extension [JZ16, KP16, Des03, Nak10, QR87, WG23]. external [AF19, Ano23a, Haj94, Lak19]. externalities [EK07, IRT08]. extortionary [Off07]. Extracting [CPQP15, CKZ10]. extraction [Kap07, LK21]. extramarital [LR04]. Extremal [MHL22, HD91]. Extreme [LGR19, SVW08, BDT18, MM09, Shi92].

F [Abe94, Ano03, Cle94, DW09b, Hal89, Hen93, Imb97, Nüe91, Sto09]. facility [CGS19]. Factor [AB16, DW15, KKM16, Luc15, ORM13, PP16, Sch16, Sen11, BKP21, BM14, BK21a, CKKK18, CJ20, Coe96, DKP22, DD23, DN89, Dov13, DMP00, Eic09, FS19, FGLS18, FOS17b, GMRO19, Hei17, HMSU23, JMY19, JKvdW14, Kap07, KUW19, KS17, KLNT24, Lin92, LB19, MRS87, Pro14, RS08b, Tim95, Wol93, Yam19, YH22]. factor-augmented [KUW19, LB19, Yam19, YH22]. Factor-Based [KKM16, CJ20]. Factors [SKL17, BK21a, CR13, CFG22, DD23, Djo20, Djo24, Fos17a, GB89, JS04, KUW19, MWWW24, MP07, SS07, Bal10]. Facts [DW15, BL14b, BHMW98, HJ93, RTÅ98]. fads [KK96]. Failure [PC15, SN06, SW86]. Failures [CF15]. Fair [Wel03]. FAMA [RS08b]. Family [EJ20, BH20, GS04]. famine [DL22]. far [BK21b]. farmers [ILOS96]. Farrell [SZ07]. Farrell-type [SZ07]. Fase [Hen93, Pis92]. Fast [MNW23a]. Fat [TW17, ALL24, FNR08]. fat-tailed [ALL24, FNR08]. fatalities [ALS24]. fatigue [SW08]. FAVAR [LB19, BMM17]. Favero [Gar01a]. FDI [EP04, MS15]. Feasible [AD09, Rac97]. Featured [DDL23a, KH22b, MNW23b]. FECM [BMM17]. Fed [CL23, KOS05]. federal [BS04, BMW90b, BMW90a, CJS07, EKL10, Got90, HP04b, Den06, Ilb12, QW23, RS16]. feedback [Bie09]. feedforward [KL95]. female [Buc98, HSWW07, Kum12, Lee95b, Mar01, MT11]. Fertility [Cha15, MT11, LS14a, LO18]. Fetters [kHcL16]. fever [TP97]. FF [Klo93]. field [DE22, GvdK12]. fields [Tho19]. Fifth [Hal89]. Filter [GNS17, And13, CKZ10, Sab91]. filtered [BÇCV14]. filtering [CF13, TM94, Wat99]. Finance [Ano08b, DIL14]. Financial [Dah17, FM02, LSZ17, Mad97, PEM16, San15, Sep00, War17, AM23, ASS21, AG02, AP11, BMR23, CC23, CCV23, CSS12, CT03a, DF97, DM07, EFR11, ERS03, EP22, GGS08, GP22a, GA11, Hua09, JK21, Kon04, LS14a, Lin95, MS90, MKP22, RM00, Sci23, Thi20, VV14, ZDE12, GKWZ16]. financing [TW87]. Finding [SL15, GMS11]. findings [ZDE12]. Fine [BH15, MST12]. Finite [Aka12, BIKY07, FL07, MR01, ATW08, Blo92, DT97, DKB10, Eis19, GL14b, Mun22, TL03]. finite-horizon [Eis19]. Finite-sample [Aka12, DKB10, TL03]. finite-state [GL14b]. Finland [CEF07]. FinMetrics [Kon04]. Finnish [RT88]. Firm [BEK16, HJC10, KLS23, TK14, Ano03, AD09, BEK22, CF94, Cin97, Gal88, GLUW03, GY21, Hus08, MM00, MZK20, MSL89, ORD01, RT88, Sch90, YT06]. firm- [ORD01]. Firm-Level [BEK16, CF94, MZK20]. firm-size [YT06]. firm-specific [AD09]. Firms [ABH^{+}17, Car00, EZ23, MR88, Pro95, Wol93]. Firpo [BG17]. First [Ano93e, HH15, CPV03, Col92, Xu13]. First-Price [HH15, CPV03]. Fiscal [FRF15, FS17, PPQ23, Pet89, BRiS11, BT14, BT19, CML04, Dav05, LT18, MU09, RS86a, TW87]. Fisher [SP04, Wes08]. fishing [FLHS07]. Fisman [AF17]. fit [EB88]. fitted [BG94]. Fitting [She93].

Fixed [PW20, AM20, Ala87, BL17, CT14, CPY18, CI20, DJK18, EPW07, JV22, KLT20, KL01, KO14, LMPTR23, Sak20, SS23a, WPN19, Wes20, GW15]. fixed- [Wes20]. fixed-effects [CPY18, KLT20]. fixed-smoothing [CI20]. fixed-T [WPN19]. Fixed-Term [GW15]. Flat [KM91]. fleet [DDL23b]. Fletcher [Ano90f]. flexibility [BNR21]. Flexible [HH90, HHJC16, MRGB16, VGCBHJC18, BH13, FFS01, Fle97, Mro12, Nak86, Pri02, Ter96, WFBM01]. floor [Sea90]. floppy [McK97]. flow [BRiS11, Dav87a, Dav87b, PT13]. flows [CGK14, CT19, CFLP02, Fuj11, SD03]. fluctuations [BF94, CF94, DIW00, Koo92a, MVW92, MS18b, SVW08]. focal [KV14]. Focused [LMMF21, CD18, OvDvdW17]. Fomby [Vis88]. Food [CHW17, AV97, BBS97, DPU97, Lea97, SLR97, VNZ97]. force [DM18, Mar01]. forced [KOW21]. Forecast [Ano08b, Bai89, CCM15, CH10a, CH10b, GR10, HW10, KK22, KvD17, Leb92, RS16, BK21b, CM14, CJS23, EKL10, GL14a, GVZ19, GM19, HL05, HN00, HR96, JM10, JMV10, KLS19, LS10a, LPU22, MW11, RRB92, Sam91]. Forecaster [BSW15, Cle22]. Forecasters [ARST16, CJS23, DHKB22, ILL06, PPQ23]. Forecasting [BCM24, BKKR15, Ber14, CKM17, CKM11, CM24a, CR15, CM17a, CH95, CG09, CM16, CFH16, DL17, FGLS18, FM16, FPR15, GO22, GNS17, HHJ86, HGRSS15, Koo13, KL95, LMZ16, LM09, War17, ZHJJ20, vDKvdWW14, Ano86d, AM11, BCCM22, BM98, CL23, CV11, CJ20, CD96, CS99, CG13, CG23, CK20, CLSS22, DGG13, De 92, DW96, ELT22, FVP02, GMH14, GNNS12, GQZ23, KTR^{+}22, KY10, Kau10, KSN23, Kor13, KP22, LLL12, Lar90, LRV12, LT96, Lou19, MG08, SV05, SS10, TZ22, WZP19, WW13, Ano95o, Sab91]. Forecasts [AFR17, FM16, HHM15, RS16, Wol15, CL23, CM09, CM10, CT03a, CJS07, CH10b, DRT21, DM20, EKL10, EMW11, GL14a, Gra96, LL06, MRW09, MM19, MW11, MW23, Miz92, NS21, OvDvdW17, SKG21, Wri13]. Foreign [Mad97, CH07, IRT08, KL93, LLP23, She13, VD98]. forest [Cou24]. Form [Van17, BK21a, DS06a, EKL10, Fle97, MM88b, Mil91]. Formation [HL16, RT88, DU98, DKP22, Lor91, Sha03, Sut23]. formats [AS06]. forms [FFS01, HMKN94, IRT08, Ter96, ZR98]. Formulating [GV87]. Forthcoming [Ano94f, Ano94g, Ano94h, Ano95e, Ano95f, Ano95g, Ano95h, Ano96g, Ano96h, Ano96i, Ano96j, Ano01a, Ano01b, Ano01c]. Fortin [BG17]. Fortran [Whi89]. Forward [Sut23, DH90, FP11, Gia07, HMW97, Hal87, Hod99, JKR22, MP01, OW10, PMM96, SPS05, Wol07]. forward-looking [DH90, FP11, Gia07, OW10]. Forword [Ano90d]. Fostering [Fio12]. Foundation [Ano95o, Kri15]. Foundations [Imb97]. four [Cre07, Lew03]. Fourier [FS09, ILOS96]. fractional [BRRT21, GAR01b, MN14, PW96]. fractionalization [Tan10]. fractionally [BCT96, TH04]. fragility [GY21]. fragmentation [RV22, BLV16]. Framework [KP16, GNNS12, HP05, MR09, PS03a, BCMM16]. France [CLM89, LS02].

Franses [Leo98]. FRED [ELT22]. FRED-MD [ELT22]. free [CH07, McC99a, WZ14]. freedom [YL08]. freight [MV07]. French [AS06, FL94, ILOS96, RS08b]. Frequencies [CD16]. Frequency [HHM15, HKM15, KKS16, MFM17, Sal15, BP06, CCM22a, CL23, CT03a, FG18, FM14, FMS19, GO22, GHHP93, KMMP20, KLLO18, Liu09, NSS12, SS10, Van11]. frictional [KRV95]. frictions [SW05]. Friendship [HL16]. Frontier [MS15, MSS16, AHKM19, CPUBUW24, HW22, Hug88, pLK20, PY96, Sea90, Sen99, Tsi02, Tsi06]. frontiers [CHP24, Eng02]. Frugal [EWB09]. FTP [Ano93c]. fuel [ABS93]. Fueling [BGLS21]. Full [FT90]. Fuller [Cif88, Kul15]. Function [MKT16, AG96, BdJ07, Bai91, CL09a, Cho99b, CD97, DK12, DM13, FS09, HPS97, HST10, HSF05, Hug88, LOW17, LTW99, NV92, PHK07, Ras95, Sch00, Sto94, Wei96, Yue94]. function-based [Sch00]. Functional [Ano90a, BH21, DS06a, FFS01, HMKN94, HJC10, Ter96, ZR98, Mil91]. functions [AHM21, AV97, DPU97, DR13, HM17, Mac96, MHM99, MN14, Orn06, PR06, WM01, Yam19, ZR98, iKLS19]. fund [DG11]. fundamentals [BW09, GRW06, Gil95, GMH14, HMSU23, Kil99]. funds [BS04, HP04b, MHL22]. furnaces [DS04]. Further [BL14a, CD96, Gil88, Mis88, SP18, TH04, YHLW06]. future [DF97, DMTvdK11, DR11, GKPP07, Li06, SS10]. futures [AK10, BM91, BHP13, FS01, HH02, HPS18, LS93c, MKV98, MM88b, SV05, Smi05]. Fuzzy [CJ16, LMPTR23].

G [And93, Ano90b, Ano90c, Ano90h, Ano03, Col92, Hen93, Klo92, McK97, Per91, Pis92, Smi95, Sor89, Szr94, AAV07]. G-7 [AAV07]. Gaim [Har93]. gain [And03]. gains [LPT04, TZ22]. game [CNS09, LLD13, LKL12, Pet89]. Games [Hal93a, AFR08, BDNS02, Koo94, RS99]. gameshow [HLW14]. gamma [Jag91]. gap [Dep18, FB02, GA19, LU05, Mac17, Mor08, MW20, PR04, QW23, Sch98, Wol18]. GARCH [AL06, BM91, BCT96, BLR06, BBP03, CM02, Che08, Cho88, CK20, FCP96, FB02, GP00, HL05, HHS12, HLV14, KvD93, LRV12, Lin92, LB95, RS08a, WFBM01, vdW02]. GARCH-M [GP00]. GARCH-stable [LB95]. GARCH-type [Che08]. gas [SS98]. Gasoline [BKL17, CDKS17, CS14, DK11, DH10, KZ22]. GAUSS [And92, CN97, Rom04, Sep00, Shi92, Vin00, LU05, Rus93]. Gaussian [BHKL17, DR93, JS10, WCC17]. GAUSSX [Pie96]. GCC [Edd96]. GDP [BCM24, Cle22, KMS13, LK04]. Gender [AJ17, Cha15, PS00, AJ19, GvSZ05, HMRC05, KOO21, Mac17]. Gender-Biased [Cha15]. General [BLM10, FHHP23, Jam17, WKN17a, BEH04, JMPW18, JM98, Rac97]. General-interest [BLM10]. Generalised [HKS16]. generalization [HL19, SS23a]. Generalized [CEG22, CKL13, SN06, BC19, CD18, HGK18, HHLL20, Jag91, LPT04, PS00, Ras95, TK14, vdW02]. generated [DMP21]. generating [JS87]. generation [BD03]. generic [Blo92, Day92]. generosity [ZK14]. Genetic [BG18b]. Gensaki [Lin95]. geoadditive [Ade04].

Geographic [JR02, DMP21]. geography [Tan10]. George [Szr94]. German [BL06, BHMW98, LTW99, Nak86]. Germany [BFS09, Bey98, CR22, CEF07, GMRZ97, KS16, Nag22, Tre98, Wil06a]. Gernot [Str09]. Geroski [Ano03]. Geweke [Fie96]. Gibson [CO89]. gifts [HO09]. gilt [DMW94]. gilt-edged [DMW94]. Girls [LO18, AJ19, PS03b]. give [EL91, Ter88]. Glenn [BMW90a]. Global [CC23, CFH16, SKL17, AT06, BEK22, CCM22b, CNT22, DdMPS07, DDLY18, Djo20, Djo24, FS09, FS23, GM19, GY21, Ham21, KM14a, RS21]. globalization [EP22]. GMM [AAB09, BM12, CD18, CG16, CL18, CD97, DJ00, Far12]. GNP [Han92, Han96, Pot95, Shi01]. GNU [BD03, Mac99, Rac00, YL08]. GNU/Linux [Mac99, YL08]. gnuplot [Rac06]. go [HKF19, Van11, vdW02]. God [DKT12]. Godfrey [Klo92]. gold [FLM13]. Goldberger [Gil93, Mil91]. Gone [Sum17]. Gonzélez [Ros22]. good [AM95, CCN22, WZP19]. Goods [BH15, CS91, Cut86, EUW23]. Gotz [BMW90a]. Gourieroux [Klo93, McK98b]. Government [CD16, GLS07, TY15, BP18, FHHP23, KW21, MSP21, SZ91]. Government-led [GLS07]. governmental [CGP99]. governments [DJ00]. Goyal [Ros22]. grade [CPB19]. grades [LT06, Tho19]. gradient [JS11]. Granger [McK94, DW16, DWW17, HP09]. Grants [Pes01a]. graphical [ABC16]. grass [Pud04]. Gravity [Bal10, SS07, BEK12, HM08, JV22, RT07]. Great [FG16, kHcL16, CDG14, KMP08, ALM22]. Greece [BT14]. GRETL [BD03, MS06, YY07, Adk11]. Grier [Sav16]. Griffith [Ano90h]. Griffiths [Ano90c, McK97, Szr94]. Griliches [Pak00]. Gross [PMM17, AGMA22]. ground [And03]. Group [AB16, BBDF14, KW19, LLD13]. grouped [KW19, LMPTR23, Sue95, AB16]. groups [ABC21, Joc23]. growing [BT23]. Growth [CMR15, EK07, EPS17, LPS97, MT16, MB16, Sav16, SHM17, Alb03, ATW08, AM23, AP12, Ano86d, ABKK13, Ber01, BLS10, Brü13, CPQ10, CCV23, CR24, CG09, CG13, Coe96, DF97, DW09a, DW09b, DMVR22, DJ95, DKT12, EPR11, EMP14, EP14, FNR08, FZ12, FLS01, FFG19, FIL13, GT09, Gre92, GNNS12, GP00, GHOS04, Hen10, JR02, JS04, KMMP20, LS09a, LS09b, LS99, MP08, MH14, PS09, ST21, Sta20, Str09, Whe03, dHW22, Ano90d]. growth-rate [FNR08]. Guangdong [BEK22]. guerrilla [Off07]. guidance [JKR22, Sut23]. guide [McK97, PPQ23]. guided [FIM05]. Guidelines [Ano94i, Ano94j, Ano95i, DW16]. guilt [BSS11]. Gujarati [Mad98]. gun [Pep05]. Guns [Pep05]. GVAR [GNNS12].

H [Ano90c, Ano90h, Bai93, Dav95, DG03, Gio89, Gre93, Pis92, Sor89]. habit [CL09b, DU98, Osb88, Sha03]. habit-based [CL09b]. Habits [BC07b]. Hack [De 92]. Hadi [Ano90g]. Half [CPY18, CK13, KZ02]. half-life [CK13, KZ02]. Half-panel [CPY18]. Hall [And93, Ano90e, Bol92, Kus91]. Hallam [Mel92]. Hallett [Gan92]. Han [Per10]. hand [Koh89]. Handbook [Tur98]. Hans [Leo98]. Hansen [ORW02]. hard [BJJ11, Bro94]. hardback [Arb93, Dav95, Gar01a, Gre93, Leo98, McC99a, Tur98, And93]. hardbound [Abe94, Ano90g, De 93, Fie96, Gil93, McK98b, Mil92, Sor89, Bol92, De 92, Smi95]. Hardcover [Ano90h, Col92, Ano90e, Cle94, Hal89, Klo92]. Hargreaves [Bar95]. harmless [AKS19]. Harold [Fer23b]. Harris [Oxl96]. Harry [Per10]. Hartog [Pis92]. Harvard [Gil93, Sto09]. harvest [Pro95]. Harvester [Oxl96]. Harvey [Sab91]. Hashem [Mil92]. Hausman [BEK16, EP04]. Haven [MR16]. hazard [Ala87, ACG04, Rom99, Szy19, VB02]. hazards [McC94]. hb [Bar95, McC02, Wee02]. hb. [Imb97]. heads [MM88a]. Health [BG06, CLH04, LK09, MRGB16, Win04, AHM22, ABC21, CT14, CJR04, CGS19, DMT06, DL22, FJ04, JM14, JPPS14, JS11, Ols98, Pri02, RWM03, WS97, vOW11]. healthcare [CGPX21, JLR14]. heard [Pep05]. HEAVY [NSS12, SS10, KLS23, LJ00]. heavy-tailed [LJ00]. Hedge [BM91, DG11, MHL22]. hedging [BMV00, HH02]. Hedonic [BBE15, AH99, AG96, AC87, BSS08, CMM09, HST10, PHK07, SZ06]. Heidelberg [De 92, Gio89, Vis88]. Helmut [Tur98]. help [GM19, IMN08, KZ20]. Hemel [And93]. Hempstead [And93]. Henderson [Ano90b]. Hendry [Abe94, Imb97, BKK20]. Henry [Sav16]. Hensher [Kuk02]. here [SW21]. Heterogeneity [ATV19, BP07, BHH15, DLOZ24, DJ16, For95, Mul97, San15, AS14, Ahn23, ATW08, Bel01, BKN95, BC07b, BH13, CT14, Cha05, Eic09, EGI05, FPV12, Hos12, JL03, JS11, KW13, KW19, KT04, LZ08, Mac17, MP04, McC94, Mor23, Ped07, RvdB13, Sak20, SZ18, TK14, Woo05]. Heterogeneous [BH15, DD21, EZ23, HMS16, HL17, Alb03, ABC21, ABP21, Bal10, BFK19, BL17, CMBS15, CJ24, HLP20, Jag91, KRV95, LS14b, PKH12, SS07, WDBR93]. Heteroscedastic [PS16, LL96, YT06]. heteroscedasticity [Bec01, KV09, MM88b, Tse98, TH04]. heteroskedastic [CB17, Dov13, IP12, LMSM18]. heteroskedasticity [BRRT21, NTK22, Pro14, RS08b]. Hicksian [Cha02]. Hidden [KW19, YZ15, LD22, RTÅ98, Son14]. Hierarchical [DDL23b, GA11, GGS08, TL03]. High [HHM15, HKM15, Li06, MFM17, PSW19, BP06, BM19, CL23, CT03a, CPB19, DD21, FG18, GO22, GHHP93, HKO12, KLLO18, Liu09, Nib24, NSS12, NO21, SS10, SK07, TL03]. High-Dimensional [HKM15, PSW19, HKO12, Nib24, NO21]. High-Frequency [HHM15, HKM15, MFM17, BP06, CL23, CT03a, FG18, GO22, GHHP93, KLLO18, Liu09]. high-frequency-based [NSS12, SS10]. higher [BEK22, Edw21, KV10, KMMP20, LMM05]. higher-order [BEK22]. highlands [HET08]. Hill [Ano90c, Ano90h, Car88, Mad98, McK97, Szr94, Vis88]. Hindustan [Ano90d]. historical [Kau20, KW21]. histories [Bie09, MRV95]. history [Day92, Rid86, Per91]. HIV [CLS21, DST14]. HIV/AIDS [CLS21]. Holland [Hen93, Pis92, Wee01]. Hollinger [Mal88]. Holly [Gan92]. Holtham [Ano90b]. Home [Van17, ABCP02]. Homogeneity [WPS18, Ng95b]. Hong [De 92, Per10, HCW12]. honor [Fie96]. Hooper [Ano90b]. horizon [CN04, Eis19, Kil99, LN07, RW05, WW13]. horizons [PR06]. Hospital [GKV15, HKS16, Ho06, KPST11, Oku93, WC04]. hospitalization [GMRZ97]. Hours [Bas09, CLSM10, Han93, Ols98]. House [AGHKS16, BPW16, Mil15, ABP21, BL14a]. Household [CJR04, ACS99, Bie09, BPR00, CS14, CMS21, Che98, DM18, EHV94, Fuk06, GM13, Hel06, KK03, MV06, MP10, MM87, VB02]. households [HKW11, HVW11, KOO21, MAL04, SV22]. Housing [BBE15, HHJC16, Luc15, Bra11, CM01, CMS24, DD19, Hon22, Hor92, HPY19, IZ03, RV18, Scr22, SP18, BCMM16]. HPC [Cre07]. HSIAO [GVB17]. HT [Bal10, SS07]. Huang [ZDE12]. Hubei [KH22a]. Hughes [Gan92]. Human [SNVCXDE17, CLS21, Edw21, MSS06, Pol13]. hundred [Sta20]. hunger [DL22]. hurdle [BNP11, Jon89, Gur97]. hurt [MKP22]. Hylleberg [Hen93]. hyperparameters [GVS23]. Hypotheses [Cha15, Vog92, CP18, Sil01]. Hypothesis [MR16, Cho99a, CC07, CT86b, EK87, FO95, GV87, MM88b, Raj95, Sal99, SPS05].

i-ix [Gan92]. IBM [EB88]. IBM-compatible [EB88]. iceberg [Orm14]. IDE [Rac12]. ideal [BS08, Des03]. Identification [ABH^{+}17, BD15, CF15, CKKK18, CT03b, DJ16, DZ18, DKP22, Dep18, DS06b, FP16, KKM16, KT15, KK03, KM14b, fLY16, PP16, SS23a, SV22, WBAB07, WBL16, AF19, Ano23a, BDKK12, CS21a, CCLZ18, Fév08, FM14, FFHL20, Ker19, Lee24, QT23, Sak24, Sto09]. Identification-Robust [KKM16, BDKK12]. identified [ALL24, RV23]. Identify [KvD17, EWB09]. Identifying [ABS99, BM16, BR23, DD23, Djo24, DR18, FLHS07, Hub14, Kas98, KS17, LS14a, LP23, LMZ23, MWWW24, MRS10, NS08, Son89a, YH22, ABK23, DIW00, Pro14]. identity [KOO21]. idiosyncratic [Fos17a]. If [PPSV21, KL12]. ignorance [KM91]. ignore [DK23]. II [Hal89, Dav87b]. iii [Mel92]. illness [KPST11]. illustrations [Phi01]. imbalance [AJ19]. imbalances [SPS13]. Imbens [Nak10]. Immigrant [Yue94]. Impact [Van17, AAZS23, BMW24, BMR23, CLS21, FZ12, GP22a, HT21, KH22a, Kum12, MST12, MKP22, NV92, Par10, Par11b, RV22, XQ12, dHW22]. Impacts [FP16, CK14, HSWW07]. Imperfect [CT15, Ash09, CT03b, EGP11]. imperfections [DM13]. implementation [BS07, OPM19, Swa01]. implications [ABFN10, HP04a, NC94, PP06, Ped07, Ros93]. implied [GV87, Rom99]. import [DU98, Kas98, Mel88, Nis05]. Importance [BHM16, BC23, DR93, Ped07, ST21, Yue94]. important [BP10, DJK18, GB89, LLD13]. Improve [HKM15, BFS98, FM19, Gra96, JM14, PR04]. Improved [IP12]. improvements [BS07, Leb92, MR01]. Improving [LM18]. impulse [PR06, TZ14, WM01, Yam19]. impulses [DIW00]. imputed [CLP22]. Inattention [AAZS23]. Incentive [HMT15, RWM03]. incentives [GMRZ97, LS14a, VV14]. incidence [FFG19, WVG14]. include [BK21a, Pro14]. including [Vin00]. Income [DSW01, LHL17, LHJC17, Pud93, Alb03, ALM22, AJL24, BH20, BCDJ99, CJ04, CH05, CGPX21, CF13, EJ20, GAR01b, HPH07, Hil07, KOO21, MC90, MM95b, Mur12, PS03b, PKH12, PZ06, Raj95, TW21, Tsi21]. incomes [PS23]. Incomplete [DD00, ILOS96, LLD13, MRV95]. inconsistency [KG96]. incorporate [ILL06]. Incorporating [Ter96, DK12, WVG14]. increased [HL07]. Increasing [FGV18]. Independence [HM15]. Independent [BM16, Rac12, Fer23a, Gew93, Her19, Sar99, Van11]. independent-private-values [Sar99]. Indeterminacy [CF15]. Index [GKV15, CK02, FS01, FP89, KV09, Leb92, LMST22, MM03, MKV98, RH21, WG23, YHLL21]. indexes [SXY20]. India [ALR19]. Indian [Ano90d, DS04, DK12]. indicator [AAV07, CPQ10]. Indicators [CPQP15, BMW24, CPQ02, CG09, DF97]. Indices [BH15, BPW16, AM23, BN09]. Indirect [GMR93, RdM18, HHLL20, MRV95, ST21]. Individual [Che98, Cle22, DR11, EGH15, JL03, PKH12, SV22, AV97, Bie09, BDG20, HPS23, Hos12, MRV95, PV89]. individuals [HRV94]. Indonesia [CLU17]. Industrial [BEL07, CP92, Ger88]. industrialized [GM19, KMS13]. industries [RT88]. Industry [BEK16, SKL17, Bai91, FS09, Lee95a, MQ04, RR95, SS98, WC04]. Inefficiency [EGI05, AD09, CHP24, MR88, TK14]. inequalities [CCM18]. Inequality [FP16, And03, Gew86, Gou22, JO20, MZ97, Mor23, ORD01, Pud93, Rom99]. Inference [Ano92j, BKP16, DWW17, Fer23a, LRV17, MW17, MFM17, PP16, Tsi06, ACF22, Ano91b, ABP21, Ash09, BIKY07, BH18, Can94, CGS18, Dav95, Dej93, DHKB22, FFG19, FL94, FM19, FK03, FSPWWE12, Gew86, GRV00, GMT98, GMR93, HM12, KK97, KW19, KZ24, KPT05, Lee24, Leu22, LZ08, MM05a, MNW23a, MRV95, MR01, McC02, MZ97, RT07, Rom99, RdM18, Sak24, ST21, SZ07, Smi95, SGK12, Yam19, Nüe91]. inferences [And04]. Inferring [JK21, KP08, Nyh03]. infidelity [Wel03]. infinite [FW03, FGLS18, Son14]. infinite-dimensional [FGLS18]. infinite-variance [FW03]. inflated [BNP11]. Inflation [CKP16, DW15, DNRS23, ES16, KKS16, Non16, Sav16, Ano90d, BCT96, BCR01, BÇCV14, BM98, BKK20, CCM22b, CHP23, Cla06, CG13, Coo09, CP97, DF97, DMTvdK11, Dem89, DS91, DD19, EUW23, EIT01, GP22b, GM19, Gou22, GNNS12, GP00, GHOS04, Hen01, JJVW01, KZ22, KRS10, LL06, Lan06, MM01, NS21, Nym91, PR04, RM00, Tsi21]. influence [BLM10, Hon22, SS91]. influential [Son89a]. Information [Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano19f, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano24a, Ano24b, CT19, CFJvdB18, DIL14, LS93c, MT16, TZ22, YZ15, Ano95c, Ano95d, CD18, DD00, DV92, FT90, FP10, Hen06, HMSS14, HR12, Ker19, KPST11, Kie11, KM14b, LLD13, Mis88, Nyh03]. Information-Based [YZ15]. informational [BBS97]. informative [Wri13]. informativeness [HJdP20]. informed [PT13]. infrastructure [CMBS15, GLS07]. ing [NS90]. initial [SV14, Woo05]. initiation [Ade04, BJJ11]. Innovation [BCC15, Cor12, CD97, HKF19]. innovations [CFK95, FGLL19]. input [DK12, Han93]. inputs [DK12, KL01]. Insights [MS15, SW21]. Instabilities [RS16, CM10, HMSU23, JMV10]. instability [ZHJJ20]. Institute [Ano92j, HHJ86]. Institution [Ano90b]. Institutional [Raj95]. institutions [FH13, LS02, Tan10]. instruction [BC23]. Instructor [McK97]. instructors [McC99a]. instrument [CGS18, DRT21, FGV18, KÜ24, LMZ23]. Instrumental [FL15, JV22, Kra12, Spi16, AIK99, Ash09, BKA90, Bas14, CL18, CGS18, CR88, DG03, IP12, KZ24, Nak10, SPS05]. Instrumental-variable [JV22]. Instruments [CT15, GP11, AF19, Ano23a, Ash09, BD99, BR23, DMP21, FL07, FFHL20, Lak19]. Insulated [kHcL16]. Insurance [BBG04, HMT15, AHM22, Bel01, BG06, CMS21, CLH04, CGR04, DDL23b, DV92, GP06, KH95, MQ09, Ols98, Pol13, Szy19, ZK14]. integer [BNP11]. integrated [BCT96, iSG23, Med02, TH04]. integration [BRRT21, Bus06, GAR01b, GY21, Hei08, HCW12, Jus01, KL12, LT96, TM94]. intellectual [BLM10]. intensity [FJS21, HH20, LGR19]. intentions [BHMW98]. Inter [CLLY06, HO09]. Inter-state [CLLY06]. interaction [CT03b, MH14, Pap11]. Interactions [HL16, CCLZ18, DM03, HL21, LfL09, SK07]. interactive [BL17, EB88, HPS23, KLT20, KO14, PW20, Rac06, Tey98]. Intercept [CH96]. intercepts [Des03]. Interconnections [BCRV16]. Interdependence [RZH03, BEK22, EK07, LD22, PSW19]. Interdependent [KvdGVW97, Ano90b, ABS99]. Interest [WBL16, AM11, BLM10, Bek99, BL14b, BK19, CCM21, DLM06, Doh13, FMMR18, GL12, HSF05, JKvdW14, KN93, Lan00, Lan06, LT91, LVG19, MP07, SXY20, WY23, WX20, vDKvdWW14]. Interference [KKM16]. Intergenerational [Bru18, FN06, BH03, DFF12, DP11, Min03, Mur12, Pos19]. intermediate [BFS98, LT06]. internal [Fac07]. International [Ano91b, CL09a, CFG08, EM17, JKR22, Kap93, KW21, KRS10, KN93, MY07, Raj92, BD95, BC23, CCM20, DdMPS07, Djo20, DH90, DMP00, DO14, EFR11, ER89, EP22, FM97, FOS17b, Gan88b, GY21, KU02, SV05]. internationalisation [MT21]. interplay [Tan10]. interpret [GH11]. Interpretation [SKG21, BM06, KW89]. Interpreting [DH90, GHK20, DD23, DM11]. interrelated [Ste07]. intersecting [AHM21]. interstate [SS98]. Intertemporal [AP11, DU98, GP11, SZ06, GL12, MS88, Nis05]. Interval [AM20, CT03a, GRLR20, Pac19]. interval-censored [Pac19]. interval-valued [GRLR20]. intervals [PR06, Sam91]. Interventions [FP16, CGP99]. Intifada [AAQB17]. Intra [LS93c, Bal10, Fei87, SS07]. intra-daily [Fei87]. Intra-day [LS93c]. intra-EU [Bal10, SS07]. intraday [KSN23, MKV98]. Intrahousehold [BA98]. intratemporal [AM95]. Intriligator [Wee01]. Introducing [CPQ10, CPBS20, Pes03a]. Introduction [BM11, BMV93, CSvD14, CFP09, DW96, DIM08, DV10, GRV00, HLMvS98, McA89, Pag94, Per10, BP07, FM02, Nag22, PP92, vDvDF05, EB88, Ano90c, Bai93]. Introductory [Wee02]. Invalid [Ker19, DMP21]. invariant [DM20, KS19, Shi01]. invention [CLLY06]. inventories [BM06, BL14b, KM14a, Ros93]. inventory [GL89]. inverse [Hub14]. Investigating [LTW99]. Investigation [GL89, BDNS02, CCN22, CL09b, MM19, ORW02, Pro95, SW86]. Investment [DK13, TY15, CLS21, Cho11, FPS94, GHK20, IRT08, Sch90, vOvGvGB19]. investor [BHLS23]. investors [HMSS14]. Iranian [LP23]. Irregular [MM17]. irrelevant [KS17]. irreversibility [CK02]. irreversible [FPS94]. ISBN [Abe94, And93, Ano90a, Ano90b, Ano90c, Ano90e, Ano90d, Ano90f, Ano90g, Ano90h, Arb93, Bai93, Bar95, Bol92, Bro94, Car88, Car92, Cif88, Cle94, Col92, Dav95, De 92, De 93, Fie96, Gan92, Gar01a, Gil93, Gio89, Gre93, Hal89, Hal93a, Hen93, Imb97, Klo92, Klo93, Kuk02, Kus91, Leo98, Mad97, Mad98, Mal88, McK94, McK97, McK98b, Mel92, Mil89a, Mil89b, Mil91, Mil92, Mof93, Nüe91, Oxl96, Per91, Pis92, Riz91, Sab91, Smi95, Son89a, Son89b, Sor89, Szr94, Tur98, Vis88, Wee03]. Isolating [EGH15]. Isolde [Mof93]. Israel [ANT94, Mel88]. issuance [BL17]. Issue [Ano08b, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano19f, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano24a, Ano24b, Ano86d, Ano92d, Ano93b, CFP09, CGR04, DIM08, DM03, Fio12, HP01, Ano16]. issues [BBP03, Dep18, KM14b]. Italian [Giu11]. Italy [ACS99, CD16, Dep18, Fac07, Jus01, MM01]. item [DP12]. iterated [MM19]. IV [BD99, EWB09, FM19, FL07, HLM17, KKM16]. ix [Gan92, Gio89, Mal88, Oxl96].

J [Ano90a, Bol92, Car88, DW09b, Kuk02, Mal88, McK94, Nei92, Per10, Per91, Riz91, Smi95, Son89b]. Jackknife [AIK99, Nak10, BD99, CPY18, MNW23a]. Jacobson [DJK18]. Jagannathan [ORW02]. James [Tob97]. Japan [HSF05, Lin95, Nak93]. Japanese [Fuk06, GAR01b, HPS97]. Jayachandran [Cha15]. Jeffrey [Wee02]. JIVE’ [AD06, DM06b, BD06, DM06a]. Job [EGP11, GP06, Bel01, BMW90b, BMW90a, CTU13, CFLP02, EO14, GMS11, Got90, HRV94, HSWW07, KvdK22, Kie88, KL24, jLjL05, MP96, MvdKH20, RSU22, Rom99, SD03]. job-training [jLjL05]. Jobs [DJ16, IMN08, vO06]. Joffre [Kuk02]. Johansen [Tur09]. John [Ano90c, Ano90f, Ano90g, Cif88, De 93, Fie96, Mad97, Mal88, McK97, Nüe91, Son89a, Son89b, Szr94, HP01]. Johnson [Vis88]. Johnston [Car88]. Joint [BHKL17, Des93, LSZ17, MRGB16, ACS99, ACG04, BDT18, DLM06, GT06, HHS12]. Jointness [DW09a, LS09a, DW09b, Str09]. Jordan [Kuk02]. Journal [Ano86d, Ano01d, Ano01e, Ano06, Ano07b, Ano10b, Ano11, Ano12b, Ano13b, DG03, Pes01a, Pes01b, Pes03b, Pes07a, Pes09b, Pes10a, Pes11, Wat07, Ano90i, Ano92j, Ano95o, HP09, Ano93c, Ano95i, Ano95c, Ano95d, Ano96c, Ano96d, Ano96e, Ano96f, Ano03, Ano08b, Pes08]. journals [BLM10]. JPE [Ros22]. Jr [Pep05]. Judge [Ano90c, Ano90h, McK97, Szr94]. Jumps [GRLM08, LLN11, MFM17]. June [Ano92j, Gre93].

Kahneman [GH11]. Kalman [Sab91, And13, CPA15, GNS17]. Kalman-em [CPA15]. Kathryn [Fie96]. Keeping [Pud04]. Keith [Mad97]. Kennan [Jag91]. Kenneth [De 93, Wee03]. Kernel [BS16, MS17, DS90, HHJC11]. Keynesian [CEG22, BÇCV14, DIW00, DW88, KRS10, NS08]. Khanti [DG03]. Khanti-Akom [DG03]. Kilian [Che23]. Kingdom [AGMA22, DSW01, EK87, JL24, KMMP20, MB02]. Kink [LHL17, SSS16, BKS18]. Klein [Ano90d]. Klump [Ste18]. Kluwer [Gre93]. know [BF09]. Knowledge [BCC15, ALR19, MRS10, PL08]. Kong [De 92, HCW12]. Korea [Lee16]. Korean [jLjL05]. Krämer [Gio89]. Krishnamurty [Ano90d]. Krueger [Nak10, Rop11]. Kuh [Mal88, Son89a]. Kuziemko [Cha15]. Kuznets [Wag15].

L [Ano90a, Ano90d, Col92, Gre93, Hal93a, Klo92, McK97, Mil89a, Nüe91]. Labor [DJ16, FSPWWE12, BMW24, BK08, CD20, Cha11, CS21b, Dev07, DM13, DM18, DFM21, EO14, Kum12, LS21, PPSV21, RvdB13]. Labour [ACS99, Ano90e, FJS21, FL15, KH95, LS02, ALM22, ABS99, Blo02, BM86, BSB86, DS06a, DLR94, EHV94, EH94, EMvS98, FK03, Ger96, HHP92, Han93, Kie88, Lee95b, Mar01, Mof93, RS09, Sto94, TWW89, XL16, dHW22]. Labour-use [KH95]. Lag [BBE15, Kul15]. lagged [LS93b]. lags [Kuo12, SS23a, Tri93]. Lalonde [DJK18]. LAM [Swa01]. Land [CL09b]. language [Gon05]. Laplace [BV12]. Large [BF94, BGR10, CKM17, DH16, Gou22, KMV19, Koo13, BMM17, BSBC^{+}04, BR99, BL17, Cam11, CKM11, CEHK20, DO14, EZ23, FMN05, HS90, HKF19, JMY19, MW20, ORM13, TWW89]. Large- [DH16]. large-scale [BMM17, HKF19, TWW89]. Lasso [Spi16]. last [Cre07]. Late [CLM89]. Latent [DJ16, De 93, BH13, Czi04, DR93, DKP22, DN89, Djo24, DMP00, FP02, HKF19, LS93b, MQ04, WDBR93, Wee01]. later [DL22]. later-life [DL22]. Latin [Can05, GLS07]. Laurent [Fle97]. Lavy [Tro24]. Law [Brz15, Fer17, KO14]. laws [ALS24, Kuo12]. layoff [KL24]. layoffs [MO04]. Lazarova [Ano03]. lead [CPQ02]. leading [AAV07, CPQ02, CG09]. Leamer [Gan88b, Shi92]. learn [AK10, IMN08, Kil99, Pak18]. learned [HR20]. Learning [BDNS02, KT04, LKL12, MG08, RS99, SW08, BKKS20, BC24, BFS98, Cha05, CLSS22, FP11, KP22, MWWW24, Pra10, Szr94]. learning-by-doing [Pra10]. Least [FPR15, KPP90, Rez08]. least-squares [KPP90]. leave [Tro24]. Lecture [Ano01d, Ano01e, Wat07, Sto86]. Lectures [Phi89]. led [GLS07]. Lee [Ano90c, Ano90h]. Leeuw [Ano90a]. left [Koh89]. left-hand [Koh89]. Legal [AF17]. Leij [Ros22]. Lemieux [BG17]. lending [FSK06, LLD13, Per98]. Length [HKS16]. lengthen [NS93]. lengths [Pat06]. lens [EUW23, HJC10, PH24]. Leroy [Son89b]. Lessons [Col92]. Levchenko [Gou22]. Level [BEK16, Blo92, iSG23, CF94, Cin97, Dem89, FvDF05, FP09, FVP02, GA19, Hus08, MZK20, Moe12, MSP21, PSS01, SV14, VB02]. levels [GT09, HRV94, Isa04]. leverage [Dov13]. Levitt [DT20, JZ16]. Lewbel [RH21]. Lewis [Che23]. Ley [DW09b]. liberalization [GP22a]. Library [Whi89]. life [AM95, CK13, CLH04, DL22, EJ20, JMM07, KZ02, NV92, Osb88, Pud93, RTK14]. life-cycle [AM95, EJ20]. Lifeboat [kHcL16]. lifetime [Pen05]. Likelihood [CBGIZ19, LRV17, WCC17, AJL24, BM14, CM02, CEHK20, DHR11, FVRR05, GVS23, Han92, Han96, pLK20, LB95, MHM99, MS93, Smi87]. Likelihood-Based [LRV17, pLK20]. Limdep [Mel87, Paa94, McC99b, MT03]. Limited [FP10, MT16, AP11, Dej93, DR13, KM14b, Lee95b, Smi87]. Limited-Information [MT16]. limiting [HD91]. limits [BT14]. LIML [BD99]. Linear [BPR00, CT15, WCC17, Alb03, AV01, And13, Ano90g, BM06, BR99, CJ24, CPY18, CLM96, EH94, FVRR05, Gew86, Gew93, Gio89, HM08, KL12, KK22, KS19, NO21, Nym91, Pac19, Pro14, TM94, TE01, Wat99, Wea92, Riz91]. linearities [LS99, ÖO00]. linearity [LTW99]. linearized [LT18]. link [EP22, LS10a]. linkage [Lin95]. Linkages [PEM16, DdMPS07, DM07, DO14, JL24, YHLW06]. linked [KLT20]. Linux [Mac99, YL08]. liquidity [HI07]. LISREL [Czi04]. list [Hor92]. living [FP89, MV06]. load [Hen06]. loadings [HMSU23]. loads [GMRO19]. Loan [GM17, Bro87, WVG14]. loans [KDK18]. Local [HHM15, Sok16, CD20, CT03b, CMM09, DJ00, Fer17, GB89, MSH19, dHW22]. local-time-based [Fer17]. locally [CS21a]. location [CT03b, GMS11]. lockdown [KH22a]. Locus [PP16]. log [Alb03, SHK11]. log-linear [Alb03]. logistic [Ano86d]. Logit [EZB17, HST16, Jam17, CL11, DFM21, EPW07, FPV12, KP94, Nib24, WBAB07]. logit-mixture [WBAB07]. London [Ano90e, Bol92, De 92, Kus91, Mel92, Mil89a, Mil92, Oxl96]. lonely [MP08]. Long [APVY16, BPPB10, CN04, CL15, Doh13, DGM03, Fac07, Her19, HK22, And91, Bar95, EIT01, FW03, FHL04, GP22a, GL03, Hil07, HMSS14, Kil99, KL01, Nak93, NC94, PF00, Pat90, PR06, Raj95, RW05, Ros22, Ros93, SN06, WW13]. long-horizon [CN04, Kil99, RW05]. Long-range [DGM03]. Long-run [APVY16, BPPB10, CN04, Doh13, Fac07, Her19, HK22, FW03, Hil07, KL01, NC94, PF00, Pat90, Raj95, Ros93]. Long-Term [CL15, FHL04, Nak93]. Longer [BPW16]. Longer-run [BPW16]. Longitudinal [GK17, DLR94]. Longman [Mil89a]. Look [Kra17, BLS10, CSS12, SD03, ST99, Ter02]. looking [DH90, FP11, Gia07, Hal87, OW10, ALV13]. Lorenz [CLM18, FFG19]. Lose [KL12]. Loss [HM17, Sch00, CD96, KL24]. losses [HGK18, KDK18]. lost [CL23]. Lott [Pep05]. Louis [RS86a]. Louviere [Kuk02]. Low [KKS16, BM19, CJ04, GO22, HSF05, Sar99, SV14, Ste07]. Low-Frequency [KKS16, GO22]. low-price [Sar99]. low-wage [Ste07]. Lower [HI16, CMV18, KW21]. lowflation [SW21]. loyalty [Sha03]. LQAC [EH99]. Lucas [LS10b]. Lutkepohl [Ano90c, Ano90h, Tur98]. LÜTKEPOHL. [Bai93]. Lyapunov [DG92, FRSRAF05].

M [And93, Ano90e, Fie96, Hal93a, Hen93, Mil92, Pis92, Son89b, Wee01, Wee02, GP00]. M1 [LTW99]. M3 [CV01]. machine [BC24, CLSS22, KP22, MWWW24]. Mackinnon [Smi95]. Macro [PEM16, ACF22, Dav87a, Dav87b, DW88, DIL14, LLN11]. macro-econometric [Dav87a, Dav87b, DW88]. Macro-Financial [PEM16]. Macroeconometric [Par87, Fai96, TWW89, Ano90d]. macroeconometrics [HP01, Gar01a]. Macroeconomic [APP15, BCCM22, BKKR15, Brz15, CR15, DGG13, FPR15, GQZ23, KLS19, Pis92, PS16, AM11, BF94, Ber01, BFS98, CCM20, CG23, CLSS22, DK01, DV10, ELT22, EMW11, FG18, FGS19, GO22, HMSU23, KM14b, Koo92a, KN90, LFZC23, Lou19, PH24, SV96, Tay89, Bar95, Nei92]. Macroeconomics [Ano08b, Ano90b]. macroeconomy [Cou24, DLM06, JKLGS13]. macroprudential [BR23]. Magazine [Wil06b]. magnitude [BS95, DZ18, IS02]. Magnus [Riz91]. mail [SW08]. mainframe [Cho11]. Mainland [HCW12]. Maintaining [PP16]. maintenance [Ken94]. major [Bri89, DFM21, EMP14]. makes [BKS18]. Making [And04, KTR^{+}22, GRV00, Kap93, KKR95, dRS10]. Malawi [DST14]. Malay [Sch98]. Malaysia [Sch98]. management [GHS05]. Managing [Bai03]. Mandatory [ALS24]. Manski [Sto09, GH11]. manual [McC99a]. manufactories [Sea90]. manufactured [Cut86]. manufacturing [ABS93, DK13, FvDF05, GS07, LS93a, MZK20, Mor94, RT88, SM95, Wol93, WL86]. Many [War17, Far12, KMS13, LMZ16]. Maple [HH95]. March [LP22, Pes08]. Marginal [CLU17, KL12, ABKK13, CMS24, EZ23, Gal88, GVS23, HR12]. Marginalized [WCC17]. mark [MM88b, DW09b]. Market [BW09, BH15, CL15, FL15, HMT15, HST16, JP15, Luc15, TA92b, ASS21, AG02, AF07, AHT94, BEK22, BP13, BMW24, Blo02, BSB86, Bro87, CI91, CNS09, CF94, CT19, CNT22, DMW94, DLR94, DM13, DM07, DFM21, EHV94, EO14, FH13, FLS22, FK03, FSPWWE12, Ger96, Gil95, GM91, HHP92, HL96, Ho06, Hod99, HET08, HM03, HKW11, HVW11, Kal12, KM14a, LS02, LT91, Lof19, LN14, MRW09, Mee90, MS07, MV07, PMM96, PPSV21, RLS00, RS21, RvdB13, RS09, RM00, SV05, Scr22, Sen11, SZ91, SZ06, Sto94, SVW08, TWW89, YHLW06, Zho20, dHW22, BLV16, GKWZ16, Ano90e]. marketing [CFP09, CNS09, FP09, PF00]. marketing-mix [CNS09, PF00]. Markets [KvD17, San15, Sok16, AP11, AT06, BMR23, Cam11, CD20, CH07, CMM09, DZ18, EFR11, FM97, GMH14, HH02, Hor92, KU02, Lin95, Lor91, Lui00, MKP22, PS03a, Smi12, Wol07, BCMM16, Mel92]. Markov [Non16, BW11, BCRV16, BH18, Des08, DWW17, FSPWWE12, GA11, GL14b, HPS99, Han92, Han96, HGRSS15, Hur17, LD22, LM18, PSS04, PRS05, RTÅ98, SS05, Son14, SPS05, YZ15]. Markov-Switching [BCRV16, Hur17, BH18, HPS99]. markup [BCR01]. markups [Mor94]. marriage [ABCP02]. Married [XL16, ACG04, BG06, Cha11]. Marsh [McK97]. Mart [HL07]. martingale [FO95, MM88b]. Mary [Imb97]. Masanao [Car92]. Masanjala [Cua11, Pap11]. Massachusetts [Gan88b]. master [Bai89]. Masthead [Ano86f, Ano86g, Ano86h, Ano86i, Ano87h, Ano87i, Ano87j, Ano87k, Ano88a, Ano88b, Ano88c, Ano88d, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano90j, Ano90k, Ano90l, Ano90m, Ano91c, Ano91d, Ano91e, Ano91f, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano94k, Ano94l, Ano94m, Ano94n, Ano94o, Ano95j, Ano95k, Ano95l, Ano95m, Ano95n, Ano96k, Ano96l, Ano96m, Ano96n, Ano96o, Ano96p]. match [CTU13, GQ10, Scr22]. matched [Abr06, BHLTM24]. matched-pair [BHLTM24]. Matching [FL15, PY22, AK07, Bel01, GL14b, IMN08, PI09, Pra10, SS06, Sto94]. Mathematica [Vin03]. Mathematical [Nüe91, McC99a]. mathStatica [Vin03]. Mating [Lee16]. MATLAB [CNJ97, Rus93]. Matrices [CKM17, LRS17, JMY19, RR18, Tur98]. Matrix [Edd96, KP87]. matter [BZ98, BB22, GQ10, MM89, Sem18]. maturity [WVG14]. Maxima [LR08]. maximization [HP04a]. maximizing [ELW10]. Maximum [BM14, CM02, LB95, AJL24, MS93]. maxmin [Cha00]. McAdam [Ste18]. McCrary [CM24b]. McDonald [Per91]. McFadden [Ras95]. McGraw [Car88, Mad98]. McGraw-Hill [Car88, Mad98]. MD [ELT22]. Mean [MP10, BG94, HK21, HP21, KU02, SLKN04]. Mean-variance [MP10]. meaningful [BK21a]. Means [HKS16, BK08, HPH07]. means-testing [HPH07]. measure [CK13, Mur12, Shi06]. measured [Wea92]. Measurement [ARST16, BKP21, BM11, BSW15, Coe96, CGS19, JRA21, MQ04, CH10a, DK23, JPPS14, KY10, Kom07, Mee90, Orn06, Cif88, Wee01]. Measures [AG17, FP16, MFM17, BDT18, Cha02, DS06b, EP22, FH12, HHS12, HLV14, LK09, MRW09, MZ97, Tre98]. Measuring [AHKM19, BSS11, Bie09, Bra11, CNS09, CM17b, DMTvdK11, DK01, DM11, GH96, Ham21, HL07, Hon13, LS10a, LVG19, LMST22, MR06, NS21, OSS18, SP18, VVR00, XQ12, Dej93, EFR11, HCW12, JK90, Kal90, LB19, Sli23, TZ22]. Mechanism [FS17, AGM07]. mechanisms [Hub14, Jus01]. media [ELT22]. median [Fai96]. Medicaid3 [Gur97]. Medical [GM13, DT97, Ho06, HMKN94]. Medigap [HMT15]. Mediterranean [CS15]. Medium [Koo13, ANT94]. medium-size [ANT94]. meeting [Gre93]. meets [GH11]. Meghnad [Phi89]. Meijer [Wee01, JRA21]. Melvyn [Str09]. mem [CEG13]. member [HP09]. membership [BO94, AB16]. memorial [Sto86]. memory [HP01, HP09, SN06]. men [AB00, KA07, VV98]. mental [JPPS14, vOW11]. Merger [Tow92]. Meta [BS07]. Meta-analysis [BS07]. Method [GVB17, Alb03, BS95, Che20, DS91, Gal20, GL14b, IS02, PI09]. methodological [EB88]. methodology [Abe94]. Methods [Ble16, DWW17, FHP16, GK17, Aka12, ABH^{+}07, Ano90f, AD09, BK17, BIKY07, BC24, CSvD14, CEHK20, DJ00, Fer17, FM19, GKP13, HN89, HLMvS98, KK97, Kuk02, LJ10, MNW23a, MvdKH20, Nag22, PW96, RPSF05, SXY20, Wee03, Car88, Vis88]. Mexican [Mor08]. micro [BFS98, HLMvS98, JR02, Kap93, VK87, Shi92]. micro- [BFS98]. micro-data [HLMvS98]. MICRO-EBA [Shi92]. micro-econometric [VK87]. micro-econometrics [Kap93]. microcomputers [Bri89]. microdata [For95]. microeconometric [BM86, FHL04, Hei08]. microeconometrics [KKR95]. Microeconomic [Ble16]. Microfit [McK93, McK98a, MT13a]. Microtsp [HL92, Woh96]. MIDAS [CG09, CK20, AFR17]. Migration [Dup21, BHMW98]. migrations [Fac07]. Miguel [AF17]. miles [EHS09]. military [Smi89]. milk [HET08]. milk-market [HET08]. Millennium [Gra17, Pee05, Kra17]. million [Cre07]. millionaire [HLW14]. minimal [FLHS07]. minimum [EGP11, Jal18, JMM07, Nag22, Sci23, Sli23, TW21]. Mining [MSH19]. Minneapolis [Gre93]. Minnesota [Gre93]. Minute [Fei87]. mirror [SV96]. misclassification [vHBB22]. Mismatch [FG16]. mismeasured [HR12, Pro14]. misreporting [BF09]. Missing [CG16, CPA15, BM14, CHL14, Ker19, LS10a]. Misspecification [Klo92]. mistrust [DV13]. Mittelhammer [McC99a]. mix [CNS09, PF00]. Mixed [CL11, CD16, EZB17, FM14, FMS19, GHL04, HIT20, Jam17, MT00, ACG04, Szy19]. Mixed-frequency [FM14, FMS19]. Mixture [FHP16, HMT15, AL06, ATW08, DT97, DS06b, GA11, GRLM08, HKF19, JRA21, WBAB07]. mixtures [DHR11, FSPWWE12, MW11, Mun22]. mixtures-of-experts [FSPWWE12]. MM [Jam17]. MNL [MT00]. mobility [Bru18, DFF12, DP11, EGP11, FN06, MP96, Min03, Mur12, Nag22, RR18, Tre98]. modality [HPR08]. mode [CK20]. Model [BBE15, BCRV16, BKK20, BHH15, CKP16, CH07, DV10, FLS01, Fos17a, GL17, HGRSS15, HI16, HL16, HMT15, HST16, LHL17, Mil15, MT16, MB16, MH14, MRGB16, Non16, PR16, Van17, WKN17b, YZ15, AM95, Ala87, AP12, ABS93, ALL24, ABS99, BCT96, Bal06, BL14a, BM06, BS07, BS08, BDT18, BNP11, BM86, BL17, BG94, Bur92, BBG04, CM06, CFK95, CE18, CJZ22, CEK09, Cho11, CP18, CKKK18, CD96, CFLP02, CKZ10, CLM96, CF13, Dav87a, Dav87b, De 94, DL95, DMT06, DW88, Des03, DDL23b, DLM06, DN89, DS06b, DH90, DPR06, EKL10, EPW07, EPR11, EHP12, Eic09, EJ20, EH94, EH99, EMP14, FPS94, FvDF05, FPV12, FGLS18, FB02, FK03, FP89, GP06, Ger96, Gew86, Gew93, Gia07, Gil95, Gio89, GRLM08, GHHP93, GP97, GKPP07]. model [GHK20, GT06, HHJ86, Hal87, HHP92, Han92, Han96, HHS12, HLV14, Hel06, HM13, HLV02, Hoq91, Hor92, HW10, Jag91, JR02, JRA21, JKLGS13, Jon89, JP10, KPS04, KLNT24, Ken94, KPST11, KM91, KL93, KvD06, KV09, KT08, KZ24, KSD98, KRV95, KU02, KMS13, KP22, pLK20, LLL12, LPS97, LS09b, LL96, LB95, LM09, MC90, MKV98, MP04, MP08, MP96, Mil89b, MS95, MW11, MR90, MM87, MB02, Mun22, Mur12, Nak86, Nak93, Nib24, Nym91, OW10, ORM13, Osb88, PF00, Pac19, PL08, Par87, PY22, Pra10, Pri02, PS00, Pud04, Rac97, RS86a, RT07, RS08b, RRB92, Ros93, RTÅ98, Sal91, SD03, SW86, Smi05, Smi12, Söd94, SK07, Son14, SPS05, SW21, SXY20, Szy19, Tem98, TK14, VNZ97, VV93, Vea92, VV98, VD98, WFBM01]. model [WDBR93, Wel03, Wol18, WL86, ZHJJ20, Zho20, vdW02]. model-averaging [LM09]. model-based [CKZ10, KLNT24, Wol18]. Model-free [CH07]. modèles [Klo93]. Modeling [AM11, BVV16, BPC15, BPW16, CKM17, Car92, CPB19, FP09, LSZ17, Thi20, BHP13, FHHP23, HKS14, Lou19, MT11, Sci23]. Modelling [Bey98, CJ04, CV11, DMW94, DO14, ES16, FHP16, FS01, FVP02, Gil95, Hen01, HKS16, Hos12, LL06, MSS16, MB02, SV05, SGK12, Son14, YT06, Ade04, AL06, ACG04, Ano90d, Bar95, BBS97, Bol92, Bri89, Car00, Czi04, Die89, Din91, EB88, FB02, GHS05, JS10, Lar90, MM01, McC02, MV96, Mel92, SLR97, Ter88, Tey05, Hen93, Mil92, Pis92, Arb93, Oxl96, Ano90e, McK94, Mof93]. Models [AFR17, APVY16, BHKL17, BKKR15, BP10, Ble16, CT15, CFH16, DW15, DH16, DWW17, EZB17, GC91, HHM15, HKS16, HM15, Jam17, KT15, Kri15, KP17, LRV17, MM17, PY15, PP16, PS16, Sal15, Smi89, Ver16, WCC17, Wol15, XL16, AS14, ABK23, ABC16, Aka12, ATW08, AH99, ABFN10, AAV07, And13, ACF22, ABP21, AAQB17, BP03, BP07, Bal10, BM14, BMM17, Bas09, BÇCV14, BLR06, BSS11, Ber01, BIKY07, BC19, BK08, BH18, BSS08, BBP03, Bus06, Byr92, CHL14, CT86a, Cam01, Can94, CFM14, CSvD14, CJ24, CKM11, CB17, CPUBUW24, CM02, CL11, Cha11, Che08, CL09b, CL18, Che23, CFP09, Cin97, CM14, CS99, CG09, CG13, CG23, CT03b, CMM09, CKL13, CGS19, CBGIZ19, CP97, DLOZ24, DR93, Dej93, DKP22, DM98]. models [Des08, DD23, Dev07, DS12, DH10, Dov13, DLRS21, DKB10, DFM21, EKL10, Eis19, Eng02, FT90, Fai96, FP11, FGLL19, FL94, Fév08, FM19, FS19, FLS12, FM97, FGS19, FM14, FMS19, FP10, GA11, GMT98, GMH14, GL89, Gur97, HM12, HS90, HH90, HL19, HL05, HLP20, HH20, HHO21, HPS23, Hei17, Hei08, HS14, Hod99, HK21, HM03, HN89, HW22, HR12, HKF19, HP21, Hus08, JW05, JMY19, JKLGS13, Joc20, JV22, KK97, KL12, KY10, KUW19, KS17, KW13, KW89, KK03, KL93, KvD93, KM14b, KSN23, Klo93, KP94, KPT05, KLLO18, Koo94, KRS10, KS19, KP22, LLL12, LS93b, LRV12, Li99, LPT04, LZ08, LOR13, LJ00, Lin92, LT18, LM18, LWY00, LS10b, MS98, Mar01, MO04, MM88b, MT00, MR06, Mul97, MBR09, NC94]. models [NSS12, NO21, ORW02, PU88, PV89, Pat06, PS07, Pro14, PSS04, QT23, RZ93, RS08a, Rom99, Rom04, Rot92, Sab91, Sak24, Sam91, Sch07, SS23a, Sch00, SW13, SW18a, SS07, She93, SS10, Sil01, Smi87, Smi93, SS05, Sol05, SC90, Sti00, Sto94, Sue95, TA92a, Tim95, TW17, Tri97, Tsi02, Tsi06, TWW89, Vel92, Ver20, VD98, WBAB07, WPS18, Wat99, Wei96, WS97, Woo05, YH22, YHLL21, Zaf11, ZLKH05, ZK05, vOZ19, AB16, Ano16, fLY16, Mal88, McK98b, Cif88]. models* [FO18]. moderation [KMP08]. Modern [CM17b, Wee02]. Modes [HPR08, HP05, SW08, TW87]. modified [Ras95]. Mody [Hua09]. moment [CCM18, Far12, FO18, GL14b, JO20]. moment-based [FO18]. moment-matching [GL14b]. Moments [EHSJ17, Gal20, HJdP20, LMM05]. Monetary [AG17, CF15, CT86b, Dah17, DD19, Hur17, JJVW01, JP10, Luc15, RV23, Tay89, WZ14, WBL16, AK19, AAZS23, BL06, CCN22, Cho99b, CN04, CCM18, ELW10, FMN05, FMMR18, FSK06, Gia07, Jar10, Jus01, Ker19, Kil99, KOS05, KMG22, Lak19, LMZ23, LB19, MS09, NC94, NPT18, OW10, Pet89, PW12, SW86, TW87]. Money [AW96, BdJ07, Bla96, GKWZ16, MS88, Bey98, CT86b, DH90, EFR11, Fle97, Gal96, Hil07, HSF05, KR90, LTW99, MM88a, PRS05, Raj95, SS91, TE01, WY23]. money-income [Hil07]. money-output [PRS05]. Monfort [Klo93, McK98b]. Monitoring [ZLKH05, CML04, GKP13]. Monographs [Dav95, Hal89]. monopoly [Ger88]. monotone [XQ12]. Monotonicity [BS16, DFF12, Giu11, HS21, Ter96]. Monte [Non16, Adk11, AKS19, Ble16, BH18, CM02, CS99, HS14, Lin92, Moo89, ORW02, RS08b, TM94]. Monthly [PMM17, MM03]. Moody [MS93]. Moraga [Ros22]. Morbidity [AJ17]. Morgan [Imb97]. Mortality [AJ17, KÜ24, MSN86]. mortgage [AHKM19, Mee90]. most [KH22a, Shi01]. Mostly [AKS19]. motherhood [SS06]. motives [Hoc03]. movement [DTV96, Off07]. mover [FK03]. mover-stayer [FK03]. moving [FMS19]. moving-average [FMS19]. MPI [Swa01]. MS [GAA^{+}89]. MS/PC [GAA^{+}89]. MTS [Wat89]. much [BB22, Fer23a, GKPP07, MQ09]. Multi [BFM15, JL08, AM95, BCCM22, DPR06, FvDF05, LL06, LPS97]. Multi-Attribute [BFM15]. multi-country [BCCM22, LPS97]. multi-good [AM95]. multi-level [FvDF05]. multi-period [LL06]. Multi-round [JL08]. multi-unit [DPR06]. multicointegration [GL89]. multicollinearity [GP97]. Multidimensional [BEK15, KLT20]. multifaceted [BMR23]. multilateral [BEK12]. multilevel [Car00, CKKK18]. multimarket [Sal91]. multimodality [Bia97, Hen10]. Multinomial [BVV16, Jam17, EPW07, Nib24, PF00]. Multiple [CG16, DJ95, HS00, LS14b, LSV23, SS91, WVG14, AG02, BP03, Bai93, BH18, iSG23, HN00, JPPS14, KPT05, MS18a, Mar00, NO21, ZK05]. multiple-equation [BH18]. Multiplicative [HHM15, CK20]. multiplier [BP18, KW21]. Multipliers [CD16, LT18]. multiproduct [ILOS96, Nak86]. multisectoral [Mil89b]. multistep [MM19]. Multivalued [Uys15, CCE23]. Multivariate [ACG04, BRRT21, BLR06, BHP13, BPC15, BBP03, CCLZ18, CPA15, DW16, DKB10, GGS08, GKP13, HH20, Miz92, NSS12, AT06, BEK22, BNP11, Bus06, CKM11, CV11, CKZ10, DN89, DM20, DMP00, HLV14, KN93, LRV12, LK04, MS18a, MY07, Pat06, PR06, vdW02]. music [Hon13]. myopic [Fer23b]. myth [TZ14].

N [DS90, Mad98]. N-kernel [DS90]. NAG [Whi89]. Nairu [CKP16]. naive [Kil93]. Napster [Hon13]. narrative [BR23, LFZC23, Sut23]. narratives [BHLS23]. Narrow [AF17, Ano09e, Bal10, CM24b, DV13, GP11, Mil15, Ros22, NO21]. Nash [AFR08]. National [Ano95o, FMMR18, BCDJ99, HHJ86, Ano95o]. Nations [Kra17, And04]. natural [BK19, DE22, EHP12, FIL13, FMMR18, LVG19, Sal99, SS98]. nature [Mor94, PS23]. Near [Lan00]. nearest [Miz92]. nearest-neighbour [Miz92]. NECLM [WBAB07]. need [CFK95, LG23]. Neese [Mal88]. Negative [WX20]. negativity [KP87]. neighbour [Miz92]. Neighbourhood [IZ03]. neighbouring [PS03b]. Nelson [HGRSS15, Kri15]. Nested [BBE15, DM98, HW10, MM89, Sil01, SS91]. nested-error [MM89]. Netherlands [Ano92j, AV97, Blo02, DPU97, Hey04, SLR97, VNZ97]. NETS [BB19]. Network [BB19, BEK22, DLOZ24, DDLY18, DMVR22, GHS05, GY21, GS20, PY22]. network-based [DMVR22]. networks [BNS18, HL21, KL95, Sci23, GKWZ16]. Neuberg [Col92]. Neumark [Rop11]. neural [KL95]. neutral [FO18, LLP23]. neutrality [CN04, Her19, NC94]. new-Keynesian [KRS10]. Newey [Far12]. newmat [Edd96]. News [ELT22, GHHP93, ILL06]. newspaper [KTR^{+}22, LP23]. Newspapers [Sok16, AF07]. next [Sta20]. Ngai [Scr22]. NIG [FB02]. Nigeria [Ade04]. night [Cre07]. NLSY [TL03]. No [Ano03, CCM21, Dav95, DH16, JZ16, Tan10, dRS10, CM17b, CR22]. No-arbitrage [CCM21]. Nobel [Sto86]. noise [iSG23, Hil06, MRW09]. noise-corrected [MRW09]. Noisy [LRS17, AAB09, GVS23]. nominal [GP00]. Non [AG17, And13, Giu11, JS10, KvD93, LS99, TE01, AV01, BHKL17, BÇCV14, Bia97, BBvdH09, Bro87, CLM96, Cub99, DM98, Eic09, FLHS07, FFHL20, GP06, GL89, LLP23, ÖO00, Phi01, Pri02, Pud93, SPS13, Sil01, SS91, TM94, Wat99, WVG14, XQ12]. non-clearing [Bro87]. non-compliance [BBvdH09]. non-empty [FLHS07]. non-filtered [BÇCV14]. Non-Gaussian [JS10, BHKL17]. Non-linear [And13, TE01, AV01, CLM96, TM94, Wat99]. Non-linearities [LS99, ÖO00]. non-monotone [XQ12]. non-nested [DM98, Sil01, SS91]. non-neutral [LLP23]. non-normal [Pri02, TM94]. Non-parametric [Giu11, Bia97, Pud93, SPS13]. non-response [FFHL20]. Non-Standard [AG17]. Non-stationarity [KvD93]. non-stationary [Cub99, Eic09, GP06, Phi01]. non-stochastic [WVG14]. non-symmetric [GL89]. Noncausal [LL16, HIT20, LLL12]. Noncompliance [HLM17]. nondifferencing [CJ20]. nondurable [CS91]. nondurables [And91]. nonfinancial [GS20]. nonignorability [XQ12]. noninvertible [FGS19]. Nonlinear [AAV07, Ano91g, ABKK13, BHKL17, BCO04, Bur92, CPQP15, CT92, DW16, Lan06, MSP21, PP92, AAQB17, BdJ07, DL95, DG11, FVRR05, GT06, HS90, Hei08, HK21, HR12, Lil00, McK94, Pot95, She93, Smi93, Woo05]. Nonlinearities [PH24, DNRS23, MP04, TA92a]. Nonlinearity [KOS05, KMP05, Wag15]. Nonmarket [SL15]. Nonnegativity [AM95]. Nonparametric [ABH^{+}07, Fer17, Fév08, FL15, Gon05, HH15, HW22, Kum12, MS15, PHK07, WC04, Xu13, BC07a, DG92, DS06b, FFHL20, HST10, Hen10, HHJC14, HZ19, HPY19, KMV19, KT08, LLP23, MY07, Mor08, MGSW13, Nag22, NS90, Par10, Par11b, Shi06]. nonpoint [BSS08]. Nonresponse [NV92, BMW24, DP12, KV14, XQ12]. nonstandard [Han92, Han96]. nonstationarity [MP07]. Nonstationary [Phi95, CC07, HP04b, Ng95b, Ped07]. Normal [AL06, ABS22, DHR11, Gew86, GA11, Pri02, TM94, WBAB07]. normality [Fei87, LJ00]. Normalized [Ste18]. Norman [Per91]. Norms [AF17]. North [Hen93, Pis92, Wee01, Bal06, BNR21, CGK14]. North-Holland [Hen93, Pis92, Wee01]. Norway [EK87]. Norwegian [Nym91]. note [FZ12, KW14, LVG19]. notices [KL24]. Nowcasting [AGMA22, CCM22a, FN23, BBGS24, GJSvN23, KMS13]. np [Har08]. npRmpi [HHJC11]. nuclear [RR95]. Number [ABH^{+}17, Win04]. numbers [CD97]. Numerical [KK97, Mac96, MHM99, MN14, AD09, Hei08, Kus91, Mcc97, Vin00, YY07]. nursing [PS00].

O [Ano90h, Bol92]. Object [Edd96, Bai09]. Object-oriented [Edd96, Bai09]. Objective [Koo92b, KS91, Phi91b, Poi91]. observables [KUW19]. observation [EHV94]. observational [McC02]. observations [SV22, VV93]. observed [Ahn23, BG18a, RdM18]. observers [JPPS14]. observing [AGM90]. Obvious [Day92]. occasionally [CBGIZ19]. Occupational [MP96, HMRC05]. Octave [Edd00]. odds [Sar99]. odyssey [MT03]. OECD [CGPX21, EP14, FNR08, IRT08, LS93a, dHW22]. off [BP06, Pud04]. offices [KH95]. Oil [ABT15, APP15, CHP23, JP15, KZ22, SS23b, AK10, ASS21, BP13, CNT22, Djo24, DK13, EMP14, FPS94, FS23, GVZ19, GM91, HR20, KM14a, KZ20, LN14, Pak18, RS21, Zho20]. Okun [Web95]. old [MP01]. Olekalns [Sav16]. OLS [PY15]. Olsen [Bol92]. one [BG18a, CK20, NO21, RS99, Tan10]. one-person [RS99]. Online [ABH^{+}17, Lee16, HM17, SW08]. only [BG18a, KL12, McC99a]. Ontario [KP94]. open [DO14, JJVW01, JP10, LR08, MVW92]. open-economy [JP10]. openness [DMVR22]. operating [Mac99]. operational [Bla96, HGK18]. opinion [DGM03]. Oppenländer [Sor89]. Optimal [BM98, LPU22, PR16, PW12, PS16, RR95, BM91, CCM18, GL14a, Gia07, HS90, Szy19, Gan92]. optimistic [BJM^{+}08]. optimization [GVS23, Liu09, MS88, Ano90f]. option [FO18, GSS20, JK21, MRW09]. options [LMM05, Sen99, TZ22]. Order [DM20, Kul15, BEK22, LMM05, PT13, Nüe91]. Order-invariant [DM20]. ordered [FPV12, KP94, Lee04, Mro12, PS00]. ordering [Per98]. orderings [KW14]. ordination [Pet89]. Organization [BEL07, MM97c]. organized [BK17]. oriented [Bai09, Edd96]. origins [DV13]. Orleans [TP97]. orthogonal [vdW02]. Other [HLM17, Mro12]. Out-of-Sample [LM17]. outcome [LPT04]. outcomes [Abr06, KV09, LO18, Mad18, PPSV21, Tro24]. outlets [HL07]. Outlier [BVV16, KZ24, Son89b]. Outlier-Robust [BVV16].

Outliers [BF94, VFL99]. Output [EIT01, Eva89, GKV15, Sav16, AV01, AAV07, BD95, CL09a, CLSM10, CG09, CG13, CH11, DF97, FNR08, Gal96, GA19, GNNS12, GP00, GHOS04, JK90, KG96, KMMP20, MM01, MSL89, MW20, PR04, PRS05, QW23, Raj92, Web95, Wol18, WL86]. Outputs [MKT16, DK12]. outsourcing [EPW07]. overconfidence [JO20]. overhang [BP18]. overidentifying [GK12]. overlapping [CM14, Smi05]. overview [EB88, Pag94]. overweight [KL12]. Ownership [Van17, LLP23]. Ox [CN97, LU05]. Oxford [Abe94, Arb93, Bro94, Gar01a, Hen93, Leo98, McK94, Per91, Phi89, Riz91, Smi95]. OxGauss [LU05].

P [And93, Ano03, De 92, Mal88, Son89b]. Paarsch [Per10]. package [Die89, EB88, HHJC11, HHJC14, Sep00, Ter88]. packages [Bri89]. pages [Ano03, DG03]. paid [Tro24]. pair [BHLTM24]. pairwise [SS91]. Pakistan [BA98]. Pande [BB22]. pandemic [KH22a]. Pandit [Ano90d]. Panel [AB16, BBGS24, BBP07, BBE15, BiS15, BCRV16, BH13, CJR04, DH16, DK16, DM13, EM17, GVB17, HZ19, KP16, fLY16, LRV17, MM17, MB16, PY96, Ver16, Wes08, XL16, Abr06, AS14, Aka12, AG13, And03, BKA90, Bal06, BP07, BEK22, BDG20, BC07b, CJ24, Cha11, CPY18, Cin97, CR88, CD97, CP97, DLR94, DG03, EMW11, EP14, Fac07, FvDF05, HLP20, HPS23, Hei08, Hil06, HK21, HCW12, ILOS96, JL03, JS10, KL12, Kap07, KS19, LL06, LZ08, MSW13, MP07, NV92, Ped07, Pes07b, RWM03, RS92, Rom04, Sak20, Sea90, SW13, SW18a, SLKN04, Ver20, WPS18, WZP19, Wol93, Woo05, XQ12, Ano16]. Panels [GUW16, Bal10, BFK19, BL17, CC07, CPY18, FN23, FN06, HKS14, SS07, WH11, WKN17a, WPN19, Wes20, MSS16]. Panic [RW16]. Panicca [RW16]. Papageorgiou [Cua11, Pap11]. paper [Ano86d, Gan88b, Mil89a, Mil91]. Paperback [Arb93, Car92, Klo93, Mad97, Mad98, McK94, McK98b, Tur98]. Papers [Ano86e, Ano96g, Ano96h, Ano96i, Ano96j, Ano90i, Ano93d, Ano94b, Ano94d, Ano94f, Ano94g, Ano94h, Ano95b, Ano95e, Ano95f, Ano95g, Ano95h, Ano96b, Ano01a, Ano01b, Ano01c, Ano08b, Gre93, Sor89, Wat94a, Wat94b]. paradigms [Sar99]. Paradox [AJ17, CO89, HSF05]. parallel [DP13, HHJC11, Swa01]. ParallelKnoppix [Cre07]. Parameter [HI16, CE18, EWB09, ELW10, FHHP23, Gia07, HMSS14, HP21, KMV19, MP04, Ped07, Raj95, Sam91, SW21, WZP19, ZHJJ20]. Parameters [BHKL17, BEH04, BDKK12, WBAB07]. Parametric [Ger96, HN89, LMM05, Mar01, MV96, Sok16, ABKK13, Bia97, CK14, CMM09, Giu11, Gur97, HH90, HZ19, Hus08, JMY19, KP87, Lee95b, LWY00, Nag22, Ols98, Pri02, Pud93, Sch98, SPS13, WFBM01, Wil06a]. parent [JM14, LS21]. parental [EF13, Tro24]. Pareto [HGK18, TW21]. Paris [De 92, Klo93, BBE15, GMS11]. Parity [KFHO15, PC15, BF05, EK87, Pat90, PP06, Shi06, Sol05]. Parking [AF17]. part [DMT06, Koh89, Dav87a, Dav87b]. Partial [DH90, FPR15, Lee24, Sak24, And03, CS21a, GGS08, KR90]. Partially [Smi05]. participants [MM95a]. Participation [KV10, AP11, BEK22, DM18, GvdK12, Ger96, HPH07, PW96]. Particle [Non16]. Partitioning [PI09]. partner [EHP12]. Partnership [CS15]. Party [Sha03]. patch [BNR21]. patent [CD97, MRS10]. Patents [Cin97]. Path [JM10, Tan10]. pattern [BM14]. patterns [CT02, GvSZ05, KW19]. Pau [Gre93]. pay [BSS11, BKS18, PS00]. pb [McC02]. Pc [Kil93, EL91, GAA^{+}89, Ter88]. PC-give [EL91, Ter88]. Pc-naive [Kil93]. PCBRAP [Koo92c]. PCgive [Esc95, Hay04, Mar98, Woh96]. PCs [EB88]. PDL [Bai09]. Peak [Kra17, Vea86]. peaks [HH20]. peaks-over-threshold [HH20]. Peer [ERS03, HKMR03, Joc23, LLD13]. peers [BBC20, BBDF14, LKL12]. Peersman [Gra17]. Penalized [CHP24, HL17]. Pendakur [RH21]. Pennsylvania [Bil00]. pensioner [HPH07]. pensions [DS91, GS04, JMM07, MP96]. Perceived [GSS20, Gra96, MST12]. perceptions [Cle22]. perfect [HF00]. Performance [CR15, AK07, CS99, CEF07, DS06b, EKL10, EPR11, HR96, JS87, LS93a, MS90, OW10, RS08b]. Peri [BCC15]. perils [Kau20]. period [BG18a, Fuk06, Koh89, LL06]. periodic [GMT98]. Periodically [Dav05, HPS99]. Periodicity [Leo98]. Perl [Bai03]. PERM [Din91]. Permanent [GRW06, KA07, Rav97, Ano03, BGV98, GLUW03, IMN08, KMP05, RS21, Sen11]. Permutation [BH21]. Persistence [Pro96, BZ98, Cho88, Cla06, Cle22, DRT21, JK90, JA17, KRS10, LS93c, Osb88, Raj92, GKWZ16]. Persistent [SZ91, TK14]. person [Bas14, RS99]. person-centered [Bas14]. personal [MC90, MM95b, SZ06]. perspective [BW09]. Perturbed [GNS17]. pervasive [BJ09]. Pesaran [Mil92, NT22]. pet [Bas14]. Peter [KS91, Sim91]. Peterson [Mil89b]. Petit [Hal93a]. Petroleum [KZ20]. pharmaceutical [AG13]. pharmacies [Oku93]. Philip [And93, Leo98]. Philippines [AHM22]. Phillips [BÇCV14, CKP16, CEG22, DNRS23, KS91, LN07, NS08, QR86, QR87, Sim91]. Physica [Gio89]. physician [KT01]. physicians [FJS21]. PIGLOG [FP89]. pitfalls [KOO21]. plan [PW96]. Planning [Lea89, EJ20]. plans [DS91]. Platform [Rac12, YL08]. Platform-Independent [Rac12]. please [BS08]. plotting [Rac06]. PLUS [Kon04, CN97, Hal93b]. Point [Wol15, BDG20, DLRS21, GL14a, KV14, SKG21]. points [Gor97, Kau10, Yue94]. pointwise [Shi01]. Poisson [WDBR93]. polarization [And04]. Policies [FRF15, FL15, BR23]. Policy [AG17, CF15, Dah17, FS17, Hal93a, Hur17, Luc15, MS18b, WBL16, AK19, AAZS23, BT19, BMR23, CCN22, CHL14, CML04, Che20, Cho99b, CS91, CCM18, Dav05, Den06, DD19, ELW10, FMN05, FM14, FMMR18, FSK06, GS07, Gia07, HSF05, JJVW01, Jar10, JP10, Ker19, KOS05, KMG22, Kuo12, Lak19, LMZ23, LB19, MU09, MS09, Nei92, NPT18, OW10, Pet89, PW12, Pol13, RS86a, RV23, Tay89, TW87, WX20]. Political [Sch16, DGM03, DFM21, HCW12, Hua09, ZDE12]. poll [DGM03]. Pollution [MR16, BSS08]. Polynomial [Byr92, CP18, CMM09]. polynomially [BM06]. pool [KK22, WZP19]. poolability [WH11]. pooled [CJS07]. Pooling [KMS13]. pools [CJS23]. Poor [MS93]. Population [BMW24, MRGB16, MO04]. Populations [HLM17, PKH12]. portable [Rac06]. Porter [DT20]. Portfolio [CKM17, HKM15, PR16, DR11, GM13, Hoc03, Liu09]. portfolios [MP10]. Portugal [Mar01]. Poser [Sor89]. positions [GL03]. possibly [Pat06]. post [She89]. Posterior [BÇCV14, CT15, Sar99]. Posterior-predictive [BÇCV14]. potential [CK14, KOO21]. potentially [ALL24]. Poverty [And03, LRS17, Zhe04, And04, Bie09, JR02]. Power [Brz15, CM16, KFHO15, PC15, AF07, BF05, Blo92, CF94, CP18, EK87, Fer17, Ger88, Kom07, Pat90, PP06, RLS00, Rot92, RR95, Shi06, Sol05]. Power-Law [Brz15]. powerful [Shi01, SLKN04]. powers [MV07]. Pp [Ano90h, Col92, Gan88b, Gan92, Mel92, Sab91, Abe94, Ano90a, Ano90b, Ano90c, Ano90e, Ano90d, Ano90f, Ano90g, Arb93, Bai93, Bar95, Bol92, Bro94, Car88, Car92, Cif88, Cle94, Dav95, De 92, De 93, Gar01a, Gil93, Gio89, Gre93, Hal89, Hal93a, Hen93, Klo92, Klo93, Kus91, Leo98, Mad98, Mal88, McC99a, McC02, McK94, McK97, McK98b, Mil89a, Mil89b, Mil91, Mil92, Mof93, Nei92, Nüe91, Oxl96, Pep05, Per91, Phi89, Pis92, Riz91, Smi95, Son89a, Son89b, Sor89, Szr94, Tur98, Vis88, Wee03]. pp.xxii [Fie96]. pp470 [Mad97]. PPP [CK13, FW03, GM09, KZ02]. Praclical [Ano90f]. Practical [CHL14, Alb03, Hen93, Par87]. Practice [Ano90c, Ano90h, BKP21, Koe88, Cle94]. practicing [Szr94]. practising [BD02]. Precautionary [APP15, Hoc03, CNT22]. precision [BS95, IS02, QT23]. Predictability [LM17, WKN17a, BKKS20, CH07, DK01, EP14, HLST21, HK22, RW05]. predicted [HM17, KÜ24]. Predicting [BKL17, Pak18, SXY20, Ano86d, AT06, BK86, Gal06]. Prediction [GRLR20, LRV17, Sto09, TM94, CSS12, LPT04, LG23, LMMF21]. Predictions [BPW16]. Predictive [CM17a, WCC17, AM23, BÇCV14, BK21b, CI20, EPR11, EMvS98, GP97, Moe12]. Predictor [LR04]. Predictors [LMZ16, MRY15, War17, WKN17a, KMS13]. preface [Ano91a]. Preference [Cha15, San15, Cha02, VV14]. Preferences [HL17, AHT94, Den06, HM13, HL21, Ilb12, KvdGVW97, KP94]. Preliminary [Kie88]. premia [HPS18, KMG22, NSA98]. premium [CI91, Fis94, HLST21, LG23]. Prescription [SSS16]. Presence [BCC15, MRY15, RS16, BHLTM24, BH13, CGP99, CG13, DV92, LMZ23, MZK20, McC94, Mil19, Min03, Pes07b, PW20, Raj92, VFL99]. present [She89, Tim95]. present-value [She89]. presented [Gre93, Sor89]. Presidential [BEK15]. Press [Arb93, Bro94, Col92, Dav95, Gan88b, Gan92, Gar01a, Gil93, Hal89, Hal93a, Hen93, Imb97, Klo92, Kuk02, Leo98, McC02, McK94, McK98b, Mil89b, Mil91, Nei92, Per10, Per91, Riz91, Sab91, Sto09, Wee03]. pretest [GK12]. prevalence [vHBB22]. previously [Sem18]. Price [ABH^{+}17, Bar95, BKL17, BHM16, BH15, BPW16, CMR15, CDKS17, Dav95, HH15, Lor91, Mad97, Mad98, McK98b, Nei92, Phi89, SSS16, Smi95, Tur98, Wee02, WKN17b, AK10, AG96, BBGS24, BP13, BSS08, CPV03, CEK09, Cla06, DL95, Djo24, DH10, GVZ19, GS07, GKZ18, GHK20, HST10, HKS14, HM03, Hor92, KOW21, KLLO18, LS93c, MM88a, McC02, MR90, Nym91, Pak18, PHK07, RW05, Sal91, SF09, Sar99, She89, SZ06, Szr94, Xu13]. price-earnings [BBGS24]. price-inflation [Nym91]. Price-Setting [BHM16]. Prices [ABT15, AGHKS16, BBE15, GL17, Hur17, MFM17, Mel88, MS17, Mil15, RM17, AK19, ABP21, BW09, BL14a, Bec01, BDKK12, BT19, BPPB10, Bra11, CHP23, CFG22, CC07, CH11, Cut86, DTV96, DH10, DK13, EHS09, Fei87, GP22b, HPY19, JL08, KZ20, KZ22, KSN23, KG96, MST12, Pak18, Par10, Par11b, PSS04, SP18, SS23b, Vog92, Wil06b]. Pricing [BS16, MS17, AG89, AFHI95, BG94, CL09b, DDL23b, DKB10, FO18, GMRO19, HLV02, LMM05, ORW02, Pro14, RS08b, SZ06, TW17, Vea86]. primal [KT11]. primarily [Hub14]. primary [LD22]. prime [HSWW07]. prime-age [HSWW07]. Principal [FMN05, HJC10, Wes20]. prior [Cam01, CJZ22, HMSS14, LS09b, Wri13]. prioribus [Phi91a]. Priors [CT15, CCM21, EPR11, KM91, MH14, Vil09]. Private [BP18, Mil19, CPV03, Dep18, GS07, LO08, MS90, MQ09, Nyh03, Sar99]. private-sector [LO08]. privately [FLS22]. privileged [BBC20]. Prize [Ano12b, Ano12c, Pes01c, Pes07a, Pes09b, Pes09c, Pes10b, Ano10b, Ano11, Ano13b]. Probabilistic [GNNS12, DLR06, DGM11, KSN23]. probabilities [DS91, JL24]. Probability [Nüe91, CH10b, Hub14, KV14, Orm14, PT13, McC02]. Probit [EHSJ17, ABK23, AG13, BH13, HL19, PF00, PS00]. probits [BEK22, Cre07]. problem [AGM90, Brü13, JPPS14, Sto91, Woo05, Hen93]. Problems [HMS16, HS90, vOW11]. procedure [BK86, iKLS19]. procedures [SM95]. Process [HMT15, Cam01, DS91, FS01, Hen06, PP06, SC90, Vea92]. processes [ABC16, GL14b, Hil07, MR01, Pro96, Rac97, SN06]. processing [DP13]. procurement [JL08, Xu13]. Procyclical [BT19]. produce [BG94, MRW09]. Product [PMM17, DM13, DFM21, Kal12, dHW22, AGMA22]. Production [GKV15, GB89, MW23, Oku93, ABH^{+}07, Bai91, Ber89, DKP22, DM13, DM18, GL89, KK03, KL01, KT11, MS95, Mor94, ÖO00, Orn06, Ped07, RT88, Ras95, Ros93, SHK11, Van07, ZR98, iKLS19]. productive [CMBS15]. Productivity [BEK16, DK12, MT21, Bas09, Coe96, DS04, FS09, Gre92, HF00, LLP23, MZK20, RZH03, SS98, Wan06]. profession [PS00]. Professional [ARST16, BSW15, FM16, Mar98, DHKB22, Esc95, NS21, Woh96]. Professor [Hal88, Tob97]. profile [MRS10, Tre98]. Program [Ano91g, GVB17, Ano90i, CK14, Czi04, EO14, HCW12, MT13b, Pol02]. Programme [Ano07b, Ano94e, Pes03b]. programmes [Rid86]. programming [And92, Bai09, BBG04, CN97, CNJ97, CNZ99, Edd96, Eis19, KPS04, Mrk01]. programs [BFS09, CFJvdB18, Kon01, MvdKH20]. progresa [LS14b, SV22]. progress [Bai91]. Progressive [Din91]. Projection [DS04, LFZC23]. projections [CPBS20, dHW22]. projects [Bai03]. promote [Edw21]. promotion [PS00]. promotional [HKS14]. propensities [CMS24]. propensity [HHLL20, SSX22]. proper [DM20]. properties [BL91, BD99, CI91, GPA89, MMOL08, RS92, Söd94, TWW89]. property [AC87, MSH19]. proportional [ACG04, McC94, Szy19]. prosperity [Edw21, MP08]. prostate [Bas14]. protecting [DS91]. Protestantism [Edw21]. provision [GMH95, Hug88]. proxies [KLNT24]. prudence [CM09, DRT21]. Pseudo [BBE15]. Pseudo-Panel [BBE15]. Public [KKS16, MQ09, Cha02, Dep18, GB89, KV10, TL03]. public/private [Dep18]. publication [Tri93]. Publishers [Gre93]. Publishing [Ano90d, Cle94, Pep05, Wee02]. Pump [BKL17]. purchase [FP09, VB02]. Purchasing [KFHO15, Pat90, PC15, BF05, EK87, PP06, Shi06, Sol05]. pursuit [DS04, Ger88, WPS18]. puzzle [CK13, GM09, JA17, LN07, MP01]. puzzling [CK13, Ker19]. Python [CS09].

q [Gal88]. quadratic [BM06, BNS02, EH94]. quadrature [DS12]. Quah [Her19]. qualitative [Mad18, MSL89, MSW13, SM95]. Quality [GPW17, BS07, CTU13, DK12, Gra96, Lou19, MM05a, Scr22, Tho19, BLV16]. Quantifying [BHLS23, KZ02, Wol07, BH13, Li12]. Quantile [BG17, HL17, LM17, BDT18, BV12, Buc98, CM24a, CL18, HLP20, HKS14, LJ10, MM00, MM05b, MV06, MSH19, PH24]. Quantiles [MS17, BDT18, Giu11, HV18, Kor06]. quantitative [EK87, Ger88, SM95, Mad97]. quantity [ACS99, BSB86, Bro87]. quarterly [MM03]. Quasi [HM03, KL01, HV13]. Quasi-fixed [KL01]. Quasi-rational [HM03]. questions [KV14]. quickly [ILL06]. quiet [RTK14]. quite [RTK14]. quote [BNP11].

R [Ano90b, Ano90c, Ano90d, Ano90f, Ano90h, Arb93, Kus91, McK97, Mil89a, Nüe91, Oxl96, Riz91, Smi95, Son89a, Szr94, Vis88, CNZ99, DP13, Fin10, Har08, HHJC11, HHJC14, Mil15, NT22, RH02, Rac12]. R&D [BCC15, Cin97, CEF07, Hus08, Mil19, vOvGvGB19]. race [AJ19, Li12, PS00]. Raiders [CL23]. raise [KL12, VV98]. ran [Cre07]. Random [BBE15, PI09, CFK95, Cho99a, CO89, Cou24, DDL23b, Hoq91, Joc23, MP96, MR06, Mur12, MBR09, Tsi02, TK14, Ver20]. random-coefficients [MBR09]. random-effects [DDL23b, MP96]. randomization [ALR19]. randomized [BBvdH09]. randomized-response [BBvdH09]. randomness [Fei87]. range [DGM03]. Rank [KP94, BLM10, CKM11, FPV12, GRLM08, Lew03, Una06]. Rank-ordered [KP94, FPV12]. Ranking [AHM21, FPV12]. rankings [Bal99]. Rare [CDG14]. rate [Ala87, AL06, BS04, BCO04, BKKS20, BK19, BMV00, Cho99b, CD05, CLM96, CP97, DN89, FNR08, GHHP93, GMH14, GKPP07, HMW97, Hal87, HD91, HSF05, HP04b, JA17, KN93, Lan06, LVG19, Li99, LMZ23, Lui00, MSP13, Pra10, RS08a, RvdB13, RV23, Sal99, SN06, SW86, SPS05, Tse98, VV98, WZ14, WY23, WX20, vN96]. ratemaking [DV92]. Rates [WBL16, ANT94, ABKK13, AM11, Bek99, BL14b, Ber14, CCM21, Cho99a, CC07, De 94, DLM06, Doh13, EFR11, EZ23, FLS22, FO95, FMMR18, Fuk06, GL12, GR08, HP04a, HMSU23, HM17, JKvdW14, KHS14, Kil99, KO14, KL95, Lan00, LT91, Miz92, MP07, PW96, SLKN04, SXY20, TH04, VV13, WFBM01, vDKvdWW14]. Ratings [DK16, MS93]. ratio [Han92, Han96, MHM99, Sli23, Smi87]. rational [BW09, DL95, FT90, GV87, HM03, KR90, Lew03, She89]. Rationality [RS16, CT86b, Iva92, Mad18]. rationing [Gal96, HRV94, Mee90]. ratios [BBGS24, GGS08, RW05]. RATS [Lar90, Mcc97, Tuf98]. RBC [FL94]. Re [LHL17, BM06, Bek99, Coo09, LS14b, Sch90, VV13]. re-employment [VV13]. re-evaluation [Bek99]. Re-Examination [LHL17, Coo09, Sch90]. re-interpretation [BM06]. reaction [Cho99b, DR13]. Real [AFHI95, CG13, GVZ19, HLST21, Hon22, JA17, LMZ16, LFZC23, Lui00, NS17, AM23, ASS21, Bek99, BL14b, Cho99a, CO96, CLM96, DHKB22, Djo20, Djo24, DK13, GP00, KHS14, KN93, LMST22, MM01, MSP13, PS10, PR04, QW23, SS23b, SLKN04, Söd94, SXY20, Wri13]. Real-time [AFHI95, CG13, GVZ19, HLST21, LFZC23, AM23, DHKB22, PR04, QW23, SXY20]. Realising [SS10]. Realized [BDT18, BNS18, CKM17, CPA15, HHS12, HLV14, ABFN10, BNS02, CV11, FB02, HKO12, JMY19, LM09, LM18, Lof19, Med02]. really [FLM13, HLW14, MM88a]. Reanalyzing [KS16]. Reappraisal [BCC15, CLM89]. reasoning [GH11]. Reassessing [CCV23, CR24, CM16]. reassessment [ALS24, CR22, EK87]. rebus [Phi91a]. recall [DL22]. received [LT06]. receiver [BDNS02]. Recession [FG16, BT23]. Recessions [CM16, CDG14, Gal06, KMP05]. recidivism [BC07a]. recipients [HSWW07]. reconciliation [GJSvN23]. Reconciling [Rop11, AP11]. reconsideration [KHS14]. reconsidered [BB22, DTV96]. recorded [Hon13]. recording [VV93]. recovery [TZ14]. recruitment [Rid86]. Recurrent [NTK22, KL95]. Recursive [PI09]. redistribution [BT23]. reduce [HMRC05]. Reduced [BK21a, CKM11, EKL10]. Reduced-form [BK21a, EKL10]. reducing [Hil06]. Reduction [Ano90a]. Redux [LHJC17]. reemployment [Bil00]. Reevaluating [DRT21, Scr22]. reexamination [DKT12, MP08]. Referees [Ano07a, Ano09d, Ano10a, Ano13a, Par11a, Ano89a, Ano92a, Ano92b, Ano93a, Ano94a, Ano95a, Ano05a, Ano08a, Ano09a, Ano09b, Ano12a]. refinement [CS21a]. refinements [BDT18]. Refining [DW15, Zho20]. reform [FHKN22, Hua09, Win04, ZDE12, dHW22]. reforms [DFM21, KO14]. Regime [Anu15, BRiS11, DWW17, vN96, CHP23, FS01, GMT98, HPS97, Nyh03, RR95, SV05, Son14]. regime-switching [GMT98, Nyh03]. Regimes [FRF15, ABCF19, AM11, DJ95, HLST21, MMOL08]. region [GMS11]. Regional [BHM16, KMMP20, SNVCXDE17]. regions [GRLR20, PZ06]. register [JRA21]. Regnery [Pep05]. Regression [CCE23, CJ16, CLP22, Don15, HL17, LHL17, MRY15, SSS16, AM20, BV12, Buc98, CTU13, CJ97, CL18, CGS18, CMM09, DS04, Gew86, Gio89, Gur97, HGK18, HLP20, Hen10, HN89, HS21, JV22, KUW19, KT01, Kor06, LJ10, LMPTR23, LS09b, MM00, MM05b, MV06, NO21, Pac19, Son89b, TP97, WDBR93, YT06, BCMM16, Ano90g, Son89a]. Regressions [BG17, KKM16, ATV19, BGR10, BO12, CN04, FLS01, HMW97, Hen10, JLR14, Kil99, Kra12, LS99, PW20, PH24, WZP19, WW13]. regressor [EWB09]. Regressors [PY15, BFK19, BD05b, CP18, CPY18, HL19, HIT20, HR12, KS19, Ng95b]. regularity [Des93, FS09]. regularization [HKO12, Juo22]. regularized [HGK18]. Regulation [MR16, CFK95]. regulatory [RR95, SS98]. rehabilitation [FHL04]. relation [Blo02, Gal96, Pat90, She89]. relations [BPPB10]. Relationship [BEK15, CHW17, EV16, KKS16, BH20, Blo92, Cor12, CEF07, Dem89, GT09, Hil07, LK09, PRS05, Raj95, RRB92]. relationships [DMP21, GL89, ORD01, PSS01, Tay89, Vog92, McK94]. Relative [Brz15, CM16, GS07, GHK20, HF00, KOO21, Yue94]. relevance [LR04, LMZ23]. relevant [KS17, Wei96]. reliability [Lil00, McC99b, PR04, QW23, She89]. reliable [MNW23a]. religion [DKT12]. Remained [ACCM17]. remedies [DP11]. Remixed [GK17]. renegotiation [GLS07]. rent [HP04a]. rents [DD19]. repeated [HSWW07, SV22]. replacement [Fer23b, MS95]. Replacing [HH15]. Replicating [DT20, Fu19, Gou22, Gra17, NT22, Non16, RH21, DG03, Edw21, Fer23b, Tro24]. Replication [AF17, Bal10, BG17, Cha05, EZB17, Far12, GP11, JRA21, JZ16, Kul15, LB19, Mil15, Pak18, Sav16, SHM17, Scr22, SP18, Spi16, Ste18, Sum17, WG23, iKLS19, BKK20, CCN22, CM24b, DV13, DJK18, Nak10, NO21, Pes03a, Ros22]. Replications [Ano09e]. Reply [BMW90a, DM06b, DW09b, McC00, Pap11]. reported [BSBC^{+}04]. Reporting [Tho04, LK09]. reproducible [KZ09, MR09]. resampled [Leu22]. Resampling [DS91]. Research [Ano95o, BLM10, GMH95, HP05, KZ09, MR09, Pag94]. Reserve [Den06, Ilb12, RS16, CJS07, EKL10, JL08, MST12, KZ20, QW23]. reserves [MS07, She13]. Residential [WKN17b, AG89, AFHI95, Hug88, MSH19]. residual [DKB10]. residual-based [DKB10]. residuals [Che08, DIW00]. resistance [BEK12]. resource [BA98, McK97]. respecting [YL08]. respondents [Hen06]. response [ANT94, AK19, BBvdH09, BD05a, CH11, FFHL20, Ger96, KV09, KT08, LS14a, Lee04, MT00, MSL89, PW96, PR06, RR95, SW18a, SM95, Sue95, Whi88, WM01, Yam19]. Responses [Jar10, BF09, EZ23, HV13, TZ14]. REStat [CM24b]. restricted [Ho06, Nüe91]. restriction [Fuj11, MT13b]. restrictions [DS06a, DZ18, GV87, GK12, Kra12, LN14, Rom99, Wea92]. Results [CF15, Gra17, Kie88, Non16, Cha05, CD96, DG03, Edw21, Fu19, HHJ86, LB19, Mis88, NT22, Pak18, RH21, TH04]. retail [HKS14]. retention [CPB19]. Retesting [Wel03]. Rethinking [MP01]. Retirement [MAL04, ACG04, BG06, GS04, Hey04, HSZ04, JMM07, KPS04, VV14, Wol18]. retrospective [SE07]. Return [LM17, SV14, BRRT21, BP06, Isa04, JA17, MSS06, RTÅ98]. Returns [CLU17, CJ16, KS16, Uys15, ABFN10, Bai91, Cha05, CK02, CF94, CO96, CR22, DHR11, DG11, DM11, DJK18, FW03, FVP02, GA11, Giu11, GRLM08, Gon05, GH11, GT06, HHS12, HD91, KLT20, KLNT24, KT04, LO08, LGR19, Lof19, MKV98, Mil19, SV05, Thi20, Wat99, WC04, Wri99, ZR98, ZHJJ20]. revealed [Cha02, HM13]. Revealing [Ilb12]. Reverse [WW13]. reversion [BG94, SLKN04]. Review [Abe94, And93, Ano90a, Ano90b, Ano90c, Ano90e, Ano90d, Ano90f, Ano90g, Ano90h, Ano08b, Arb93, Bai93, Bar95, Bla95, Bol92, Bro94, Car88, Car92, Cif88, Cle94, Col92, Dav95, DT94, De 92, De 93, Fie96, Fin10, Gan88b, Gan92, Gar01a, Gil93, Gio89, Gre93, Hal89, Hal93a, Hal93b, Hay04, Imb97, Klo92, Klo93, Koo92c, Kuk02, Kus91, Leo98, Mal88, Mar98, McC02, McK94, McK98b, Mel92, Mil89a, Mil89b, Mil91, Mil92, Mof93, Nei92, Ng95a, Nüe91, Oxl96, Per10, Per91, Phi89, Pis92, Riz91, Sab91, Smi95, Son89a, Son89b, Sor89, Sto09, Szr94, Vin00, Vin03, Vis88, Wee01, Wee03, And92, Aus94, Bai89, Bew88, Bri89, Byr87, DS90, Die89, Din91, Edd96, EB88, Esc95, Fer94, Hal88, Har93, Har08, HL92, HH95, Kid92]. review [Kil93, Kol01, Lar90, Mad97, Mad98, Mar02, McC99a, McC06, McK93, Mca97, Mcc97, Mel87, Paa94, Pet91, Pol02, Rac06, Sch05, Sen99, Sil97, Ter88, Tri88, Tur98, Vea91, Wat89, Whi89, Woh96]. Reviews [Hen93, McK97, Wee02, EB88, Lar90, Tri87, Tri88, Woh96]. revision [Lea89]. revisions [AGMA22, CG13, FZ12, KY10]. Revisited [MM95b, MT16, RV23, Bla96, EHP12, KÜ24, Mil19, MH14, SS22, Wil06b, vOvGvGB19]. Revisiting [BHLTM24, BT23, FHKN22, Her19, KOO21, Lea97]. revolution [FS23, SW20]. Reweighting [EGH15]. rich [KLS19]. Richard [Ano12c, Pes01c, Pes09c, Pes10b]. right [Hua09]. Rijckevorsel [Ano90a]. rise [BMW24]. Risk [Brz15, BHH15, FO18, GL12, HPS18, LSZ17, MM09, RM17, SKL17, ALM22, ATV19, BP06, BJM^{+}08, BK86, BC19, BMV00, CI91, CCM22a, CR13, CFG08, EJ20, EP22, GP22b, GL03, GM13, Gra02, GS20, HH90, HLW14, Kom07, KL05, LP08, NSA98, OSS18, PU88, PH24, Sch05, WZ14, dRS10, vOZ19]. risk-averse [BJM^{+}08]. Risk-neutral [FO18]. risk-return [BP06]. Risks [DL17, BC07a, Doh13, MP96]. Risky [San15, DST14]. road [MP08]. roads [Li12]. Robertson [Nüe91]. Robust [BVV16, CF15, CHW17, CJS23, DHKB22, DFM21, Gia07, KKM16, MM05a, MRGB16, PMM96, Uys15, BDKK12, CS21a, CGS18, CKZ10, KZ24, LOW17, Leu22, MNW23a, RS08b, ZDE12, iKLS19, TZ14, Son89b]. Robustbase [Fin10]. Robustness [GY21, Tem98, AC87, CLSM10, CJZ22, FZ12]. Rodney [DW09b]. Rohwedder [JRA21]. Role [AGHKS16, LMZ16, Ahn23, BP13, BHLS23, CD20, CNT22, DRT21, DKT12, Hor92, JS04, KM14a, Lof19, Pro14, RS86a, TWW89]. Roles [EGH15]. rolling [Hil07]. Ron [McC99a]. root [BBP07, Blo92, CP97, HPS99, Kap07, Koo92b, Mac96, MN14, Pes07b, Sam91, Shi01, SLKN04, Sto91, Wri99]. roots [CGP99, Lan00, Phi91a, SV91]. Rotterdam [Ano92j, BS08]. Rough [MP08]. round [JL08]. Rounding [Don15, KV14]. Rousseeuw [Son89b]. routes [Phi91a, SV91]. Routledge [Mel92, Mil92]. Roy [Mun22]. RStudio [Rac12]. Rule [KFHO15, BT19, KOS05]. rules [DPJF06, OvDvdW17]. run [And91, APVY16, BPW16, BPPB10, CN04, Doh13, EIT01, Fac07, FW03, GP22a, Her19, Hil07, HK22, KL01, NC94, PF00, Pat90, QR86, QR87, Raj95, Ros93, Bar95]. Running [Don15, DK23]. Rural [MSN86, ALR19, BA98, BB22, GvSZ05, JR02]. Rust [Fer23b].

S [And93, Ano90b, Ano90g, Ano03, DG03, Gan92, Gil93, Hen93, Imb97, Mil91, CN97, Hal93b, Kon04]. S&P500 [LMM05]. S-PLUS [Kon04, CN97, Hal93b]. Safer [JZ16]. safety [GB89, Li12]. Saharan [CLS21]. Sales [BL14b, BM06, CH11, GL89, HKS14, Hon13]. Salvador [CPUBUW24]. Sample [CJ24, GPW17, HH15, HM15, LM17, Aka12, BIKY07, BD99, BC24, CGS18, Dev07, DKB10, FL07, FLS12, FN06, Koh89, MV06, MR01, Mar01, Pac19, PR06, SW13, SW18a, TL03, Vel92]. sampler [DR93]. samples [Blo92, CI20, DP11, DPJF06, Zhe04]. sampling [HS14, ST21]. sanctions [ALV13, LP23, VV13]. Sargan [Phi89, HP01]. SAS [Sil96]. Satiation [LHL17]. satisfaction [JS11]. satisfying [FS09]. saturation [Ano86d]. saving [Fuk06]. savings [And91, CLH04, MQ09]. Scale [Yat00, Bai91, BMM17, CF94, FH12, HKF19, KG96, TWW89, WW18, WC04, ZR98]. scales [Pen05, SSW06, Wil06a]. scanner [HKS14]. Scholars [Ano07b, Pes03b]. school [CPB19, DD21, DS06b, FHKN22, GR08, LS14b, Li06, MT13b, MM89, NR19, SK07, TL03]. Schooling [CLU17, CJ16, Uys15, BH20, Cha05, CR22, DHR11, DP11, DJK18, EF13, HV13, KLT20, KLNT24, KT04, Tro24]. schools [BBC20, DD21, TL03]. Schularick [Sum17]. Science [Ano95o, Abe94]. Scientific [JZ16, DV13]. Scilab [Mrk01]. Scope [SL15]. Scopus [Ros22]. score [BC19, CKL13, HHLL20, Sil01]. scores [BK08, SSX22, Wil19]. scoring [OvDvdW17]. sealed [LP08, Sar99]. sealed-bid [LP08, Sar99]. Search [FS17, GP06, Hon22, HSWW07, JKLGS13, KvdK22, Kie88, KRV95, MGSW13, Pra10, Rom99]. search-matching [Pra10]. season [COQD19]. Seasonal [JK90, Bus06, Cub99, GAR01b, MMOL08, Pro96, Hen93]. Seasonality [Osb88, HPS18, Scr22, Hen93]. seasonals [MB02]. Seat [JZ16]. seatbelt [ALS24]. Seats [JZ16, Pol02]. secession [RV22]. second [Ano90h, Ter02, AAQB17]. Secondary [CLU17, LD22]. secret [JL08]. Section [BiS15, RW16, AH99, BP07, Pes03a, Pes07b, Wes20]. Sectional [Ano16, BKP16, BHP16, BCC15, DH16, EM17, Ver16, Byr92, GRLM08, Hei17, HK21, Kap07, Oku93]. Sectionally [MSS16, FN23]. Sector [LSZ17, GS07, GS20, GHK20, LO08, MS90, MZK17, TA92b]. Sectoral [BHM16, DS06a, EPW07, JL24]. sectors [EUW23, FvDF05]. secular [Han93]. Security [BBG04, HSZ04, DR11, GS04]. seekers [MvdKH20]. segment [VB02]. segment-level [VB02]. segmentation [MV07]. segregation [HMRC05]. selected [DPJF06]. Selecting [FGLL19]. Selection [BM86, GvdK12, HMT15, HM15, Lee04, Spi16, AKS19, AG89, AK07, BKK20, BC24, CJ24, CD96, DPJF06, ERS03, FLS12, Fos17a, FN06, GP97, GKPP07, Hus08, JKLGS13, Kor13, KMS13, LLL12, LPT04, Mac17, Mar01, NO21, Pri02, Rac97, SW13, SW18a, Sli23, Vea92, Vel92, YHLL21]. Selectivity [HL16, BH20]. Self [AG89, GNS17, MFM17, Sem18, BSBC^{+}04, Gil88, RS86b, SW18a]. self-employed [Gil88]. Self-employment [Sem18, RS86b]. Self-Exciting [MFM17]. Self-Perturbed [GNS17]. self-reported [BSBC^{+}04]. Self-selection [AG89, SW18a]. sellers [RLS00]. Semi [BC07a, CK14, Gur97, Lee95b, MGSW13, PS07, Sok16, Wil06a, ABKK13, Ger96, KL12, MV96, Nag22]. semi-linear [KL12]. Semi-nonparametric [BC07a, MGSW13, Nag22]. Semi-Parametric [Sok16, CK14, Gur97, Lee95b, Wil06a, Ger96, MV96]. Semi-structural [PS07]. Semiparametric [AG96, BEK15, BDP98, BSS08, BHMW98, BHH15, CEG13, HL17, KPT05, LZ08, New90, Pen05, PS10, Rob88, RPSF05, YHLL21, BA98, Cam01, CLH04, DM98, GvSZ05, HLV02, HN89, HLMvS98, HZ19, Hus08, JMY19, KV09, LOR13, LS99, MSH19, MSS06, Mar01, Moo89, Ols98, Rom99, Sch98, SSW06, SZ06, Yat00]. semper [Phi91a]. SEMs [JW05]. sender [BDNS02]. sender-receiver [BDNS02]. Senegal [CGS19]. Sense [Ano09e, NO21]. sensitive [DGM11, VV14]. Sensitivity [EPS17, Mal88, SL15, SSS16, Ash09, IMN08, Lee04, Ano90g]. Separability [Lee95a, And91, BL91]. separation [BMW90b, BMW90a, GP06, Got90]. Sequential [Bil00, Ble16, Hei08, HS14, BH18, CP97, DPR06]. Sequentially [CP18]. serial [AH99, DKB10]. Serially [Sti00]. Series [Ano01d, Ano01e, BKKR15, FPR15, McK98b, Nüe91, Wat07, AH99, BW11, Bai93, BF94, BB19, CJ97, CI91, CT92, CO96, Cub99, DD00, DS91, Die89, DGM03, GRLR20, HT21, KvD06, KSN23, KN93, LL96, LGH92, LMSM18, LK04, MS18a, MM03, Oku93, Pat06, Phi95, Phi01, RdM18, RTÅ98, Sab91, Sam91, She93, Smi93, Smi05, Tey05, Tow92, Tuf98, Wei96, Car92, Leo98]. Serlenga [Bal10]. service [FJS21]. services [And91, EUW23, HSWW07]. set [Ano90b, KLT20, RV23]. set-identified [RV23]. SETAR [CS99]. sets [BS07, FMN05, HW10]. Setting [BHM16]. Several [CPQP15]. severity [HGK18, KPST11]. shadow [NPT18]. shale [BNR21, FS23, SW20]. shape [BE92, JS11, Rom99, SdV18]. shares [Kal12, Ped07, TW21]. sharing [ATV19, CFG08, EP22]. Sharp [HLM17]. sharpen [QT23]. sharpness [MW11]. Shazam [Byr87, Vea91, Whi88, McC99b, Sil96]. Shields [Sav16]. shift [RR95]. shifting [SW20, vDKvdWW14]. Shifts [Anu15, BRiS11, iSG23, EPW07, FVP02, LS93a, SV05]. Shin [Bal10]. shipping [OSS18]. shock [BZ98, CT86b, Gia07, PT13, YH22]. shock-absorber [CT86b]. Shocks [BHM16, FG16, GM17, AK19, BF94, BM19, BM98, CHL14, Can05, CLSM10, Cle22, CFJvdB18, CDG14, Djo24, DE22, FG18, GHK20, HR20, Her19, Jar10, JKR22, KMG22, Koo92a, LVG19, LMZ23, LN14, MU09, MSP21, NPT18, NSA98, Pro96, Rav97, RS21, SW05, WM01]. shopping [HL07]. Short [APVY16, CPQ10, CML04, HPS23, AM11, FN06, GL03, PF00, QR86, QR87, TZ22]. short-dated [TZ22]. Short-run [APVY16, PF00, QR87]. Short-term [CPQ10, CML04, AM11]. Should [HKF19, LT18, Fer23a]. shrinkage [HHO21, HP21, HK23, SXY20]. Sibling [NR19]. siblings [BBC20]. sick [BKS18, Tho04]. sickness [FHL04, ZK14]. side [Wol07].

Sided [Sok16, AF07]. Siegel [HGRSS15, Kri15]. sieve [CHP24]. Sign [FM97, Ala87, Fuj11, LN14]. Sign- [FM97]. signal [LK21, Tho19]. signaling [LKL12]. Signals [CPQP15, GHL04]. significance [BCDJ99]. Silver [kHcL16]. SIMPC [Bla95]. Simple [KT15, Vel92, Woo05, BS95, EKL10, IS02, Mro12, PS03a, Pes07b, Rot92]. simplification [BKK20]. simplified [Smi87]. Sims [Sim91]. simulated [BK08, MS93, Smi93]. Simulation [Ano92j, Cha11, FP11, LS93b, LK21, XL16, Ano91b, FT90, Haj94, Hod99, Lar90, SC90, Wee03]. Simulation-based [FP11, LS93b, LK21]. simulations [AKS19]. simultaneity [Brü13]. Simultaneous [OPM19, And03, IP12, KL93, Lee95b, RPSF05, VD98]. simultaneous-equations [KL93, VD98]. Since [ACCM17, CCV23]. Singapore [De 93, FLS22, LT91]. single [FMMR18, YHLL21]. single-index [YHLL21]. singular [CFM14]. SIPP [BD05a]. SIR [PY22]. sisters [PS03b]. site [Ano93c]. six [KMS13]. size [ANT94, Ano03, BZ98, BBDF14, GLUW03, GK12, HJC10, JLR14, KLS23, Oku93, YT06]. Sizes [MW17, MM00]. Skew [FH12, SGK12]. skewed [ABS22, ALL24, KT08]. Skewness [RM17, Dov13]. skill [CKT19, DKP22]. skills [Gon05, KLT20]. slave [DV13]. slope [AS14, Ala87, Moe12]. Slowdown [Gra17, Pee05]. slowdowns [JL24]. slowly [NS21]. Sluggish [Bro87]. Small [BD99, Dev07, MM01, PR06, Ros22, ANT94, AGM90, BF94, BL06, CI20, EZ23, HW10, JJVW01, JP10, MVW92, Nym91]. Small-sample [Dev07, PR06]. Small-system [MM01]. smaller [BS04]. smelters [GN00]. smoking [Abr06]. Smooth [HKS14, HST16, JKvdW14, LHJC17, AM11, BO12, CLH04, Des08, MSS06, MS93, NH01, Sol05, TA92a]. Smoothed [Kor06]. smoothing [CI20, Ros93, TM94]. Snooks [Per91]. Social [BBG04, Col92, CM17b, GS04, HL16, HL21, HSZ04, Ped07, CGR04, CCLZ18, DR11, DM03, KH95, LfL09, RS09, SK07]. Society [Dav95, Gre93, Hal89, Sto86, Hal89]. Socio [CT02]. Socio-economic [CT02]. socioeconomic [GB89, JS11]. Software [Edd96, Fer94, Har08, Kol01, Lar90, Pet91, Rac06, Swa01, Tri87, Tri88, BD03, BFC02, Bri89, BBP03, CL11, Fin10, GAA^{+}89, Koo99, Lil00, McC99b, Sen99, Tri91, YL08]. Softwear [EB88]. soil [CPUBUW24]. Solow [DIW00, LPS97, MP04, Tem98]. Solution [KP17, Eis19, GM09, HS90, LN07]. solutions [McC99a, Woo05]. Some [Cho88, Gil88, Gra02, HHJ86, Mis88, RS92, CT86a, Cin97, GP00, Koo94, PP06, RS86a, FNR08]. son [Min03]. Sonnberger [Gio89]. Sons [Ano90f, Ano90g, Cif88, De 93, Nüe91, Son89a, Son89b]. SORITEC [Mar02]. source [BSS08, Kas98, LR08]. Sources [BM16, CMR15, CH10a, DIW00, KMP08, MP08, Son89a, Gan88b]. South [Lee16, Wee02, Pro95, MV06]. south-east [Pro95]. South-Western [Wee02]. Sovereign [DK16, DZ18]. Soviet [Bai91]. Space [BHKL17, WCC17, ACF22, Bra11, FGLS18, Hei08, HMRC05, SP18, Car92]. Spacing [MM17]. Spanish [CM06, DGM03]. Spanos [McC02]. Sparse [DLRS21, FPR15, BR23, KS17]. sparsity [BK21a, DD23, HHO21]. Spatial [BBE15, BEK16, CF13, EHSJ17, FH13, KP16, ABP21, BBP07, BEK12, CB17, CT02, EK07, FLS12, HM13, MSH19, PL08, Sci23, BCMM16, fLY16]. Spatio [LRV17, Mil15, BL14a, BHP16]. Spatio-Temporal [LRV17, BHP16]. Special [Ano86d, Ano08b, Ano16, CGR04, DM03, Ano92d, Ano93b, CFP09, DIM08, HP01]. specialization [DD21]. specialty [BLM10]. specific [AD09, Bal10, CMM09, Djo20, SS07]. Specification [CCM15, DS12, HN89, Jag91, KT15, BS08, Ber89, Dav95, DM98, FS19, KT11, PL08, Yue94]. Specifications [CR15, ILOS96, SC90, TWW89]. specifying [MM88b]. Spectral [FO95]. spectrum [CEG22]. Speculation [JP15, GMH14]. speculative [CNT22, KM14a]. speed [CF13]. spells [CFJvdB18, NS93, TA92b]. Spending [CD16, BP18, KW21, MSP21]. spillover [ALR19, NR19]. Spillovers [BEK16, BCC15, DZ18, SNVCXDE17, BH13, Cin97, HH02, JKR22, KW14, LMZ23, Mil19, PL08, SV05]. spiral [MR90]. Spline [MRY15, HHJC14]. split [MO04]. Sponsorship [Pes01a]. sporadic [AH99]. sporting [Tho04]. spot [Bec01, HMW97, Vog92]. Spotting [War17]. spousal [Ols98]. Spread [CM16, Gal06, HH02]. spreads [DF97, DMP00]. Springer [Bai93, Car92, McC99a, Mof93, Vis88, De 92]. Springer-Verlag [Bai93, McC99a, Mof93, De 92]. Spurious [EV16, CD05, TW17]. Squares [FPR15, KPP90, Rez08]. SST [Tri88, Tri91]. St. [RS86a]. Stability [Ber01, EPS17, DG92, LS10b, LTW99, Raj95, Tay89]. stabilize [KZ20]. Stable [ACCM17, BM98, SdV18, LB95]. staff [CPBS20, EKL10]. Stage [BHP16]. staircase [BJJ11]. STAMP [Tey05]. stand [BS08]. Standard [ACCM17, AG17, kHcL16, FLM13, MS93]. Standardized [GNS17, Che08]. standardized-residuals-based [Che08]. standards [MV06]. Stanley [Vis88]. STAR [CM02, FvDF05]. startups [CD20]. Stata [Fer94, Kol01, Pet91]. State [BHKL17, CT14, DK16, MS17, WCC17, ACF22, Bie09, CLLY06, DL95, GA19, GL14b, Hei08, HSWW07, HPY19, Lou19, MSP21, RV22, Vil09, Car92]. state-level [GA19, MSP21]. State-Space [WCC17, Car92]. stated [BSS11, Hen06, VV14, Kuk02]. statement [Pes86]. States [BHKL17, BL14a, BP18, Eva89, JS87, KLS23, Pro95, SLR97, VNZ97]. statgraphics [DT94]. static [LLD13]. stationarity [Coo09, DW91, Gre92, KvD93, Sol05]. stationary [BD05b, iSG23, Cub99, Eic09, GP06, Phi01, Rac97, Sam91, Shi01]. statistic [Smi87]. Statistical [Can94, DPU97, GPA89, MZ97, SZ07, Tri88, Tri91, FL94, Kus91, MR01, McC02, Nüe91, Ter88, Tey98, BK21b, De 92]. Statistically [ALL24, JO20]. Statistics [Col92, McC99a, Nüe91, Fie96, Klo93, Leu22]. Statistique [Klo93]. status [AB00, Bie09, BG18a, DST14, Gil88, KP08]. statutory [ZK14]. Stay [HKS16, HKF19, SW21]. stayer [FK03]. stayers [Ver20]. Steady [Lou19, Vil09]. Steady-state [Lou19, Vil09]. Steel [DW09b]. step [Hus08, RPSF05]. Stepping [GW15]. Stepping-Stone [GW15]. sterling [GHHP93, GKWZ16]. sterling-dollar [GHHP93]. stick [VV13]. Stickiness [DK16]. sticky [DH10]. Still [JZ16]. sting [CPQ10]. Stochastic [DFF12, GUW16, Gor97, JKLGS13, KT11, Leo98, LJ00, MKT16, MSS16, MVW92, Sen99, Tsi02, And03, BP10, CHP24, CPUBUW24, CE18, CEHK20, FT90, HM12, HP04a, HS90, HW22, HP21, Hug88, KS91, KU02, pLK20, LPS97, LK21, MS98, MY07, NH01, PY22, Phi91b, Poi91, Sea90, She93, SC90, Tsi06, Wat99, WVG14]. Stock [CL15, HL96, HKW11, HVW11, AW96, ABFN10, AT06, BW09, BRiS11, CT92, CK02, CC07, Cho88, CT19, Dav87a, Dav87b, FM97, GT06, KU02, KLLO18, MS88, Par10, Par11b, PSS04, RW05, SV05, Sen11, SVW08, Wat99, YHLW06, ZHJJ20]. stock-flow [BRiS11, Dav87a, Dav87b]. Stocks [Mad97, EFR11]. Stone [Ano12c, GW15, Pes01c, Pes09c, Pes10b, RH21]. Storage [GL17]. Strachan [DW09b]. strange [RS92]. Strategic [HMSS14, KW89, KZ20]. strategies [Gal20]. strategy [WZP19]. stratified [BHLTM24, For95]. street [DMM05]. strength [BKP21]. stress [CM24a]. stretching [GPA89]. strike [Dej93, Jag91, KW89]. strikes [GM09]. Strong [EM17, AK07, BHP16]. Stronger [AK19, DW91]. Structural [BFK19, BiS15, BMM17, BKKR15, BH03, Die89, FGS19, Gal06, KRV95, KP17, LS93a, Luc15, Per10, PY15, RS08a, Sci23, SZ18, Tey05, Van17, ABC16, AF19, Ano23a, ALL24, BP03, BR23, Bur92, CH96, CFLP02, Czi04, DGG13, De 92, DLR06, FGLL19, FM14, GP06, GKP13, Her19, Hey04, Ken94, KPST11, KM14b, LPU22, LN14, MG08, MS18a, Mee90, MR90, PS07, Sab91, SD03, Smi12, Son14, ZLKH05, ZK05, De 93, Mal88]. Structural-Form [Van17]. Structure [AB16, PP16, AH99, Ber89, CMV18, CCM21, DLM06, DG11, Doh13, GN00, GL12, JKvdW14, Kau10, Lan00, LT91, MQ04, Mis88, Moe12, Mul97, Oku93, PL08, Ras95, Tow92, TA92b]. structures [Riz91]. Student [GPW17, BC23, BBDF14, BFS98, Gew93, HKMR03, SdV18]. student- [Gew93]. students [ERS03]. Studies [Ano90d, AKS19, AC87, HP01, HMRC05, SS23a]. Study [EZB17, Spi16, And91, And04, Blo02, BBvdH09, CS99, Fac07, Hil07, HV13, KV10, Lea97, Lee95b, LfL09, RS08b, SLR97, ZK05]. studying [VB02]. stumpage [HP04a]. style [Che23]. Stylized [DW15, RTÅ98, BL14b, HJ93]. Sub [CLS21]. Sub-Saharan [CLS21]. Subdivision [WKN17b]. subject [ORD01, Wea92]. Subjective [LHL17, BM11, Bla96, EMvS98, GH11, HSZ04, KV14, LfL09, LK09, Zaf11]. submarkets [AG13, BCMM16]. Subprime [Anu15]. Subsampling [CC07]. subsequent [Bel01, PS03b]. subset [FLHS07]. subsidies [CEF07, Fio12, Hus08]. Subspace [HK23]. Substitution [GP11, NSA98, Bai91, DU98, GL12, KP87, Nis05]. Successful [JZ16]. suffer [LPT04]. sugarcane [Ras95]. suggested [HP05]. suggestion [Kau10]. Sullivan [DJK18]. superior [MRW09]. superiority [CJS23]. Supplemental [Ano90b]. Supply [BNR21, DJ16, GM17, XL16, ACS99, AW96, ABS99, BK08, BM86, Cha11, CS21b, DS06a, Dev07, EMvS98, FS23, FJS21, HR20, HM03, Kum12, LS21, Lee95b, Lee95a, LN14, MS88, MT21, Mof93, Ste18, Wol07]. supply-side [Wol07]. support [HPH07]. Surprises [WBL16, Moe12]. Survey [ARST16, AG17, CJR04, Iva92, RS16, AGM90, AGR17, BLR06, CS14, Cra87, JRA21, KLT20, MSW13, PV89, RRB92, Rob88, SW08, SW21, WL86, BMW24]. survey-administrative [KLT20]. surveys [CS21b, ILL06, Sor89]. Survival [GAA^{+}89, HSZ04]. sustainability [BRiS11, Dav05, Mar00]. sustainable [SPS13]. SVAR [Lak19]. SVARs [JW05, OPM19, RV23]. SVIX [Lof19]. Swait [Kuk02]. Sweave [Rac12, MR09]. Sweden [FHKN22, FHL04, KPS04]. Swedish [KH95, RvdB13, ST99]. Switching [BC19, BCRV16, DWW17, HGRSS15, Hur17, BH18, CHP23, Des08, FS01, FM97, GMT98, HHP92, HPS99, Han92, Han96, LM18, Nyh03, OPS22, PRS05, SS05, Son14, SPS05, vN96]. Switzerland [JA17]. sword [GY21]. Symansky [Ano90b]. symmetric [BM98, EHV94, GL89, PT13, Ras95, Sar99]. symmetry [HLV02]. synchronization [SS05]. Synthetic [GVB17, BK17, Che20, Dev07]. System [MKT16, And92, BS08, BL06, Cha92, FFS01, HR96, HH95, KV10, KN90, Lew03, LR08, Mac99, MM01, Mrk01, RH21, ZK14]. systematic [BK86]. Systemic [vOZ19, BC19]. Systems [FHP16, Sil96, VGCBHJC18, BL91, BRiS11, BR99, CH95, Des93, GP06, GH11, KT11, Ng95b, PS10, Pro14, RPSF05, Sea90, Ste18, Una06, VNZ97].

T [Ano90c, Hal89, McK94, Mil89b, Nüe91, Wes20, Gew93, SGK12, SdV18, WPN19]. tables [DFF12]. Tail [DL17, LSZ17, CCM22a, HW22, HPY19, PSW19, PH24, SdV18]. tailed [ALL24, FNR08, KLS23, LJ00]. tails [PSW19, Thi20, TW17]. tale [KOO21]. tales [PSW19]. TANF [HSWW07]. target [De 94, DPJF06, HP04b, KSD98, Li99]. targeting [AW96, JJVW01, MS88]. targets [PPQ23, RM00]. taste [NSA98]. Tax [BM19, CDKS17, DR13, ABKK13, DK11, EZ23]. taxable [AJL24]. taxation [ABKK13]. taxes [ACS99, Kum12, SS91]. Taylor [And93, BEK16, EP04, KFHO15, Sum17]. teach [RH02]. Teacher [GPW17]. Teaching [MS06, BD02, GMH95]. technical [Bai91, FH12, TK14, MSS16]. technically [FLHS07]. Techniques [Ano92j, AP12, Ano91b, Ano92d, Ano93b, BF05, Par87, And93]. technological [Cin97, Coe96, EK07, GN00]. technologies [ABH^{+}07, BPR00]. Technology [EM17, MKT16, BH13, CLSM10, IT94, ILOS96, MP04, Nak86, Ras95, SS98]. teen [SK07]. telecommunications [Ber89]. tell [BD05a, CS21b, FZ12]. Telling [PSW19]. Temporal [LRV17, Mil15, PP06, BL14a, EMW11, BHP16]. Temporary [GW15, MO04, CH11, IMN08]. Tenreyro [Scr22]. tensions [MKP22]. Teräsvirta. [McK94]. Term [CL15, GW15, LT91, Moe12, AM11, CPQ10, CML04, CMV18, CCM21, DLM06, Doh13, FHL04, GL12, HMSS14, JKvdW14, KMG22, Lan00, Mis88, Nak93]. terms [MH14, PU88, Pap11]. terrorism [MM09]. Terry [Mil92]. Test [BS16, CF15, HM15, KP16, Kul15, PY15, BEH04, BFC02, BDKK12, Djo20, DH10, HPS99, Han92, Han96, Hen10, Hoq91, JMPW18, JM98, KP87, Lee95a, MR88, NT22, Pes07b, Ros93, Rot92, Shi01, Sil01, WH11, Wil19, vN96, Gio89]. testable [ABFN10]. Testing [ABK23, ATW08, BF05, BK19, Bia97, BM12, BCDJ99, iSG23, Cho99a, CCM18, CT86b, DM98, EMvS98, FW03, FRSRAF05, GUW16, GAR01b, HHP92, Han92, Han96, HLV02, HS21, JO20, Joc20, Joc23, Mad18, McC94, MM88b, MS93, NC94, Ng95b, NH01, Smi87, SPS05, Tur09, VFL99, WKN17a, Wri99, Bil00, CM24a, CP18, CJS23, DS12, HM13, HPH07, HN89, LS14b, LSV23, MY07, NO21, PSS01, Sal99, Vog92]. Tests [Bus06, CM14, CM17a, DH16, ER89, GMRO19, HN00, MR16, PC15, RS16, And03, BRRT21, BBP07, Bia97, Blo92, BD05b, BK21b, BH21, CT86a, Che08, CC07, CH10b, Des93, DM20, DKB10, FP11, FS19, FO95, Gal88, GV87, GHL04, GK12, HPR08, Hil07, HK22, HW22, Jag91, Kap07, Klo92, Koo92b, Mac96, MHM99, MN14, PV89, PMM96, QW23, RS08b, SPS13, SN06, SS91, SLKN04, Tem98, Tim95, Vel92, Wes08]. TESTU01 [McC06]. text [KTR^{+}22]. text. [McC99a]. textbooks [Wee01]. Texts [Arb93]. Textual [NS17, BCM24]. their [BHLS23, DK13, EP22, IMN08, MM88a, PS03b]. Theoretical [Kri15, Med02]. theories [BP10]. Theory [Ano90c, BKP21, Dem89, FPS94, FJS21, FL15, Gan88b, Hal93a, Hor92, Kom07, OPM19, Sha86, WPS18, YZ15, Ano90h, AP11, BP07, BHMW98, BFS98, CM02, CKKK18, Day92, DK01, EK07, FM97, JL08, Kie88, Koe88, LT06, McC02, Mor94, PY22, Rom99, Sch90, SV96, Mil91, Phi89]. there [BP06, Bie09, CP97, CR22]. third [Car88]. Thisted [Kus91]. Thomas [Mil89a, Vis88]. thought [AK19, DW91, HK22, Sem18]. threat [Dav05]. three [RS08b, RPSF05]. three-factor [RS08b]. three-step [RPSF05]. Threshold [LHL17, RZ93, Gal96, Gal06, HH20, HKF19, LL96, MKV98, Rot92]. thresholds [AT06]. Tickets’ [AF17]. tiered [Cha11, PY96, XL16]. tightness [Sto94]. tile [Sea90]. timber [HP04a, LP08, Pro95]. Time [BE92, CGK14, Car92, CGPX21, CKP16, CK02, CD16, CR15, Des03, EHSJ17, FPR15, LMZ16, Leo98, LMSM18, MS15, MS17, MS09, MSP13, Oku93, PEM16, AG89, AH99, AM23, ABFN10, Arb93, AFHI95, BW11, Bai93, BF94, Bal10, BB19, BP13, BEH04, Bek99, BK19, BDKK12, BC23, BDG20, Bra11, CKT19, CI91, CT92, Cha92, CE17, CE18, Che98, CS21b, CG13, CO96, CKZ10, Cub99, DD00, DHKB22, DS91, Die89, DJK18, ER89, Fer17, FHHP23, FK03, GMRO19, GVZ19, GRLR20, HH02, HT21, HLST21, HPS23, Hei17, HMSU23, HP21, KY10, KMV19, KvD06, KSN23, KSD98, KS19, KN93, LL96, LMZ23, LFZC23, MS18a, NO21, ORD01, Pat06, Phi95, Phi01, PR04, PT13, QW23, RdM18, Sab91, Sam91, SS07, SP18, She93, Smi93, Smi05]. time [SW21, SXY20, Tey05, Tim95, Tow92, Tuf98, TP97, Wei96, Wri13, ANT94, McK98b]. time-invariant [KS19]. time-of-use [AG89, ANT94]. Time-series [Oku93, CI91, CT92, DS91, KN93, Sam91, She93, Smi93, Tow92]. time-specific [Bal10, SS07]. Time-Varying [CR15, CGPX21, Des03, MS09, MSP13, Bek99, BDKK12, BDG20, CE17, CE18, CKZ10, DHKB22, ER89, FHHP23, GMRO19, HH02, HMSU23, HP21, LMZ23, ORD01, SW21, Tim95]. timely [KMMP20]. times [CCN22, GQZ23]. Timing [DLR06, AH99, BBG04, Fio12, FP09, PPSV21, VB02]. Tinbergen [Ano92j]. tip [Orm14]. tobacco [JL03]. Tobin [Gal88, Lea97]. Tobit [Cha11, LZ08, Moo89, XL16]. Tokyo [De 92, Vis88]. Tom [Wee01]. toolkit [Kid92, Rac00]. tools [CHL14, Rac00, Tri88, Tri91]. top [TW21]. Toronto [De 93]. total [Coe96]. TOU [ANT94]. tour [FIM05]. tourism [Hel06]. TR [Rot92]. Tracking [DD21]. Trade [DMVR22, EHP12, BP06, Bal10, BMR23, BO94, BG18b, DV13, DE22, EP04, EPW07, EP22, GNNS12, IRT08, KL93, MKP22, SS07, VD98]. trade-off [BP06]. trades [Nyh03]. Trading [YZ15, BRRT21, EHP12, Fis94, GL03, KM14a, LGR19, PT13, BLV16]. Traffic [Li12, ALS24]. Train [Wee03]. training [BFS09, CFJvdB18, jLjL05, LO08, RvdB13, Rid86, RSU22, RS09]. TRAMO [Pol02]. transaction [AP11, Fis94]. transactions [HET08]. transfer [HV13]. transforms [CP18]. Transient [KK96, TK14]. Transition [HST16, LRS17, SHM17, BO12, Des08, FK03, PS09, RvdB13, RR18, Sol05, TA92a]. Transitions [ABCP02, EHSJ17, Blo02, CJ04]. Transitory [RS21, Ano03, GRW06, GLUW03, KA07, LVG19, Rav97, Sen11]. translog [DKP22, ILOS96]. Transmission [Dah17, FS17, AGM07, Can05, EFR11, GS20, JL24, Jus01, Pos19, SS98]. transport [MV07]. traps [JR02]. Treasury [AS06]. Treated [HLM17]. Treatment [HLM17, AK07, Bas14, BG18a, CPUBUW24, FFHL20, Gew93, HS21, HL10, KT01, KPST11, KT08, Lee04, LS14b, LOW17, Lee24, Mee90, MT13b, PW20, PI09, RvdB13, RSU22, Sak20, Sli23, Ver20]. treatment-response [KT08]. Treatments [Uys15, Bas14, CCE23, HHLL20]. Tree [War17]. Trend [CKP16, Non16, Shi01, DW91, EUW23, KS91, LK04, MB02, NS21, Poi91, Sam91, Sol05, vHBB22]. trend-cycle [LK04]. trend-stationarity [DW91, Sol05]. Trend-stationary [Shi01]. Trending [GL17]. Trends [GUW16, Leo98, CH05, DJK18, Fac07, Gor97, MM05a, MVW92, NH01, Phi91b]. triangular [Pro14]. Trimming [HH15]. trivariate [Hil07]. true [CD05, FP89, Tan10, TK14]. truncated [GC91]. trust [Fer23a]. TSMod [FIM05]. TSP [McC99b, Sil96, Sil97]. TSW [Pol02]. Turkish [EGI05]. turning [Gor97, Kau10]. turnover [Kie88]. TV [HL10]. TVAR [FRF15]. twin [FGV18]. Twins [GPW17, Isa04]. Two [CK20, FM19, Sok16, XL16, AAB09, Ano90b, AF07, Cha11, CGS18, DMT06, GHK20, Hus08, JW05, JV22, MM89, Pac19, PY96, SS23a, BHP16]. two-part [DMT06]. two-sample [CGS18, Pac19]. two-sector [GHK20]. Two-Sided [Sok16, AF07]. Two-Stage [BHP16]. two-step [Hus08]. Two-tiered [XL16, Cha11, PY96]. two-way [JV22, MM89, SS23a]. type [JLR14, SZ07, Che08]. types [Kal12, Vog92].

U.K. [Cut86, CT86b, RS86b, WL86]. U.S. [AV01, CF15, GA19, MC90, Sch90]. U.S.A. [TW87]. UCLA [Ano90i, Ano91g]. UK [Abe94, Mil89b, BKK20, BLV16, CMV18, CLM96, DMW94, Dav87a, Dav87b, DW88, GAR01b, GMH95, HHP92, Hen01, JW05, JKLGS13, MSL89, Mee90, NT22, ÖO00, Rav97, TE01]. Unbalanced [BBE15, Sea90]. unbiased [Fai96, SPS05]. Uncertain [AGMA22, CM10, JMV10, Kra12]. Uncertainty [ARST16, ABCF19, Ano08b, BSW15, CCN22, MB16, Sav16, Sch16, AM95, AF19, Ano23a, BM19, CC23, CCM20, CHP23, CH10a, CG23, CMM09, CF13, DMTvdK11, DR11, DK13, DV10, ELW10, EHP12, FLS01, FG18, Gan92, Gia07, GP00, GHOS04, HMSS14, JP10, KZ02, KK22, LL06, LS10a, MS18b, RRB92, Sam91, ZHJJ20]. unconditional [BG17]. Uncovering [RM00, WZ14, Tan10]. under-age [Kuo12]. undergraduate [McK97, MS06, McK97]. underlying [GH96]. Understanding [AJ19, EUW23, HMW97, HGK18, ST21, SP04, Bro94]. Undesirable [MKT16]. unemployed [RS09]. Unemployment [Bel01, BZ98, FG16, HI07, Pol13, QR86, QR87, Sto94, Van17, Ahn23, Ala87, AV01, ALV13, BE92, CTU13, CM01, CD05, CT02, CFJvdB18, Des08, Eva89, FP02, GP06, KRV95, Li06, MM01, NS93, Sal99, Ste07, Szy19, TA92b, VVR00, Web95]. unified [Che08]. uniform [LOW17]. Union [AB00, BO94, MZK17, WZ14]. unionism [VV98]. Unions [MS90, VV98]. Unit [CGP99, BBP07, Blo92, CP97, DP12, DPR06, EHV94, HPS99, Kap07, Koo92b, Lan00, Mac96, MN14, Pes07b, Phi91a, Sam91, SV91, Shi01, SLKN04, Sto91, Wri99]. United [BL14a, BP18, DSW01, EK87, Eva89, JL24, JS87, KLS23, KMMP20, MB02, Pro95, SLR97, VNZ97]. units [Fac07, Han93, JS87]. univariate [BM98, De 94, Sol05]. universities [ERS03, GMH95]. University [Ano92j, Arb93, Bro94, Col92, Dav95, Gan92, Gar01a, Gil93, Hal89, Hal93a, Hen93, Imb97, Klo92, Kuk02, Leo98, McC02, McK94, McK98b, Mil89b, Nei92, Riz91, Sab91, Smi95, Sto09, Wee03, GMH95]. Unknown [ABH^{+}17, AB16, LHL17, SKG21]. unmeasured [McC94]. unobservable [DL95]. Unobserved [BD15, BHH15, DJ16, LK21, Mac17, Ahn23, Bal10, Bas09, Bel01, Bus06, FPV12, KG96, LZ08, RvdB13, Sak20, SS07, Woo05]. Unraveling [BEK15]. Unravelling [DM07]. unreliability [QW23]. unrestricted [PW12]. unstable [Cor12, GR10]. update [ASS21, EL91, Tri91]. Updated [Ano95d, Ano95c]. Upper [CLU17]. Urban [LRV17, CGS19]. Urga [Ano03]. US$60 [Mad97]. US$74.95 [Mad98]. US$95 [Mad97]. US/UK [CLM96]. USA [Abe94, Gre93, Mil15, Alb03, AV97, Buc98, DPU97, DO14, EPS17, HMS01, HPY19, KO14, MSP21, Sok16]. usage [Pri02]. use [AG89, BJJ11, CS21b, KH95, LT18, Pap11, vOW11, ANT94]. Used [CM17a, EHS09, Zaf11]. useful [CLSS22, DF97]. users [Ano94i, Ano94j, Ano95i, Ano95c, Ano95d]. Using [Adk11, BLM10, BCRV16, FP16, FM16, GH11, HGRSS15, LHL17, LGH92, MFM17, Mil15, MT16, Non16, PL08, PY15, QT23, RH02, Van17, Ver16, Abr06, ATW08, ALR19, AGR17, Ano91b, Ano92j, AHT94, ABKK13, BD02, Bai03, Bal06, BNS02, BÇCV14, Bas14, BBS97, BEK12, BSS11, BFS09, BBDF14, CJ97, CS21a, CT14, CLM18, CP18, Cho88, CG09, CK20, CI20, CKZ10, DJ00, De 94, Edd96, For95, FSPWWE12, Gal06, GL89, HHJ86, HT21, HPR08, HV18, HR12, Isa04, KTR^{+}22, KPST11, Kor13, KL95, LL06, pLK20, Lak19, LP23, LU05, Leb92, Leu22, LMST22, LFZC23, LM18, MM05b, MRV95, MP08, MT11, MBR09, NT22, OSS18, OvDvdW17, ORW02, PY22, PW12, Sak20, ST21, Sch90, SS06, Smi87, Smi93, SGK12, SV22, SPS05, SXY20, TM94]. using [TA92a, TZ22, Tur09, Wei96, Wil19, WL86, YT06, AKS19]. Utility [BPC15, BK08, Cha00, Mil91, Rac06]. Utility-Consistent [BPC15, BK08]. utilization [Gur97, Sha86].

V [Ano90d]. v.1 [Vin03]. v4.2 [Mcc97]. vacancies [Fuj11]. vacation [MV96, VK87]. Validating [ZK05]. validatory [Rac97]. Valuation [RW05, SL15, BF09, BS07]. valuations [Cho88]. Value [GL03, AC87, EMvS98, Fév08, KP22, MM09, Sar99, She89, SZ91, Tim95]. Value-at-risk [GL03]. valued [GRLR20]. Values [CM17b, Kul15, CPV03, GSS20, MSH19, Sar99, Tur09]. valuing [DG11]. Var [Kor13, MR90, Wri13, ALL24, BK21a, BCRV16, DdMPS07, DWW17, DLRS21, FP11, Hil07, JJVW01, KK97, KW14, PW12, Sci23, WCC17]. VAR-Exploiting [BK21a]. VAR-models [KK97]. variability [Dem89]. Variable [CT15, Don15, FL15, Spi16, BKK20, CL18, CLP22, DR93, Dej93, DK23, JKLGS13, JV22, KZ24, Koh89, Kor13, NO21, Smi87, YHLL21]. Variables [CG16, FHP16, AIK99, Ash09, AM11, BKA90, Bas14, CFM14, CGS18, CSS12, CR88, Czi04, DL95, DG03, EH99, FP09, IP12, KS17, Kra12, LS93b, Lee95b, Nak10, PF00, PW96, SPS05, Sti00, Wee01, De 93]. variance [BNS02, FW03, LM18, LS93c, MP10, PL08, She93]. variances [Sam91, YH22]. Variation [BM16, CKP16, CD16, MS17, PEM16, BNS02, BK19, BBDF14, CGK14, Che98, EMW11, KY10, TL03]. variations [DGM11]. various [Vog92]. VARMA [APVY16, KT15]. VARMAs [CE17]. vars [LN14, ACCM17, Koo13, BR23, CCM15, CE18, CEHK20, CM10, FGS19, Gal06, JMV10, KMV19, Ker19]. vary [NS93]. Varying [CR15, LHJC17, BP13, Bek99, BDKK12, BDG20, CGPX21, CE17, CE18, CKZ10, DHKB22, Des03, ER89, FHHP23, GMRO19, HH02, Hen06, HMSU23, HP21, KMV19, LOR13, LMZ23, MS09, MSP13, ORD01, PT13, SW21, Tim95]. varying-coefficient [LOR13]. Vector [CM17a, CG23, CFH16, FM16, LL16, ABC16, AF19, Ano23a, AAQB17, BGR10, CCN22, CM24a, Che23, CEG13, FHHP23, GL14b, HHO21, HK23, JKLGS13, LP22, LB19, MW20, Pee05, Smi93, Vil09, Yam19, YH22]. vehicle [Li12]. vehicles [DDL23b]. velocity [Raj95]. verification [DS06b]. Verlag [Bai93, Car92, Gio89, McC99a, Mof93, Vis88, De 92]. Version [EL91, Pie96, Tri91, Aus94, HL92, Paa94, Pol02, Sep98, Tri88, Tuf98, Woh96]. Versus [ABT15, GW15, BW09, BLM10, BS08, Bel01, CJ20, DIW00, Djo20, ELT22, FVRR05, JA17, KMS13, LJ00, PKH12, RvdB13]. vessels [FLHS07]. via [Ash09, APVY16, CCM18, DD23, Edw21, FRSRAF05, LMPTR23, MWWW24, MZ97, SSX22, YH22]. Victorian [CLM89]. video [CNS09]. video-game [CNS09]. vii [Car88]. viii [Ano90b, Gre93, Hal89, Mil89a]. violence [AAQB17]. visiting [Van11]. visits [Win04]. Vislie [Bol92]. ViSta [BD02]. Visual [BD02, BFM15, DGM11]. vivos [HO09]. Vol [Ano03, DG03]. volatilities [ABFN10, CT92, CCM21]. Volatility [BH15, CMR15, Cho88, CR15, CL15, CK20, ES16, EV16, MFM17, AG02, ABCF19, AT06, BRRT21, BCO04, BP13, BHP13, BDT18, CCM22b, CE18, CEHK20, CV11, CSS12, CT19, DHKB22, DN89, FS01, FM97, FB02, FM02, HH02, HL96, HL05, HHS12, HLV14, Hos12, HP21, KU02, Leb92, LK21, LJ00, LM09, LS93c, Lof19, MS98, MRW09, Med02, NSS12, RZ93, RS08a, SS10, Smi05, TZ22, Wat99, Wri99]. volatility-switching [FM97]. Volume [Ano90b, BRRT21, BCO04, Hal89]. Volumes [Hal89, Ano90b, Klo93]. voluntary [GvdK12]. voter [KP94]. voting [HM13]. vs [GRW06, HSF05, KM91]. vulnerability [CR24].

W [Ano90b, Ano90c, Ano90h, Cif88, Cle94, Gio89, McK94, McK97, Mil89b, Pis92]. Wage [Car00, Sal91, Ber01, BKN95, Buc98, CR22, Dep18, EGP11, GT09, Jal18, LO08, MM05b, Mac17, MR90, Mor08, Mor23, Nag22, Nym91, ST21, Sch98, Sli23, SV14, Ste07, TL03, VV98]. wage- [Nym91]. wage-price [MR90]. wages [CK14, CO96, EGP11, Gil88, Hos12, KA07, KT08, MM01, Mel88, VV98]. wait [DMM05]. waiting [TP97]. wake [SVW08]. Wal [HL07]. Wal-Mart [HL07]. Wald [GV87]. Walk [DMM05, Cho99a, CO89]. Walters [Ano03]. Wansbeek [Wee01, JRA21]. wants [HLW14]. Was [FLM13, Fer23b]. Wascher [Rop11]. Washington [Ano90b, HSWW07]. wave [Wol18]. waves [Tow92]. way [JV22, MM89, SS23a]. Weak [CGS18, EM17, GP11, MB16, BD99, Far12, FL07, BHP16]. Weak-instrument [CGS18]. weakly [CPY18]. Wealth [Hil06, Kra17, Blo02, GP22a, Pud93]. wed [Dup21]. weekly [CCM22a, LMST22, Ols98, WG23]. Weeks [Str09]. weight [KL12]. Weighted [BO12, HPR08]. weighting [Hub14, HHLL20]. Welfare [ELW10, And04, BS95, Cha02, CS91, HPH07, Ho06, HV18, IS02, MY07, MR06, VV13]. Welfare-maximizing [ELW10]. Well [LHL17, BG94, HMS01, SS05]. Well-Being [LHL17, HMS01]. well-fitted [BG94]. Welsch [Son89a]. were [BS04]. west [Jar10, BHMW98, Nak86, Tre98]. Western [Wee02, Coe96]. Westview [Mil91]. Wheatsheaf [Oxl96]. whether [LGH92]. White [Dav95]. Who [BBC20, EO14, HLW14, And03, MQ09, McC99a]. wholesale [Wol07]. Whose [VV98]. wide [CM24b, Ros22, Smi12]. wide-body [Smi12]. Wild [MW17, FM19]. Wildly [MW17]. Wiley [Ano90a, Ano90c, Ano90f, Ano90g, Ano90h, Cif88, De 93, Fie96, Mad97, Mal88, McK97, Nüe91, Son89a, Son89b, Szr94, Tur98]. Will [FZ12, BS08]. William [Szr94]. willingness [BSS11]. Willman [Ste18]. Windmeijer [Far12]. window [Hil07]. Windows [Mar02, Mar98, Pol02, Rac00, Vin00, Woh96]. wine [CMM09]. within [ALR19, AJ19, GS04, KOO21, MB02, Sar99]. within-cluster [ALR19]. without [BSB86, LfL09, MT13b]. wives [Ols98]. Woessmann [Edw21].

Woittiez [Mof93]. Women [XL16, Cha11, jLjL05, ST21, Sem18]. won’t [BD05a]. Wooldridge [Wee02]. words [MKP22]. Work [CM17b, GW15, Dup21, FMN05, Ger88, LK09]. worked [Han93, Ols98]. worker [CGK14, Fuj11, RZH03]. workers [KPS04, SV14]. workhorse [Zho20]. Workplace [Mor23]. Workstation [Whi89]. World [Hal89, SKL17, AGM07, NS90, Ros22]. worldwide [GN00]. worse [HK22]. Wright [Nüe91]. wrong [Pep05].

X [Ano90h, Imb97, Riz91, Arb93, Cle94, Dav95, McK94]. xi [Car92, Gar01a, Phi89]. xii [Ano90a, Ano90e, Ano90d, Ano90f, Klo92, Leo98, Mil91, Pis92, Riz91]. xiv [Ano90g, Col92, De 93, Hal93a, McK97, Mil92, Son89b]. xix [Bol92, De 92, Vis88]. Xplore [NS90, Tey98]. Xplore’-ing [NS90]. xv [Bar95, McK98b, Per91, Son89a]. xvi [Sab91, Tur98]. xvii [Ano90b]. xviii [McC99a]. xx [Kus91]. xxi [Gan88b]. xxiii [Cif88]. xxiv [Mil89b]. xxix [Ano90h]. xxviii [McC02]. xxxv [Ano90c, Sor89]. XYP19.95 [Gil93].

year [KA07]. years [Sta20]. yellow [TP97]. Yen [CI91, Tse98]. yen-dollar [Tse98]. Yen/Dollar [CI91]. Yes [KW14]. Yield [CM16, HGRSS15, Kri15, AGR17, DIL14, DMP00, MS09, WX20]. yields [FHHP23]. Yilmaz [KW14]. Yogo [GP11]. York [Ano90a, Ano90c, Ano90g, Ano90h, Bai93, Car88, Cif88, Col92, De 92, De 93, Fie96, Gan92, Hal89, Hal93a, Kus91, Mad97, Mad98, Mal88, McC99a, McK97, Mel92, Smi95, Son89a, Son89b, Szr94, Tur98, Vis88]. young [ABCP02, AB00, VV98]. youth [HV13, Li06]. you’ve [Pep05].

Zellner [Fie96]. Zero [HI16, KS16, Kee86, CHP24, CMV18, KW21, WY23]. zero-inefficiency [CHP24]. zeros [Mul97]. Zipf [KLS23]. Zone [War17, CH05, De 94, KSD98, Li99]. Zurcher [Fer23b]. Zurich [Sor89]. Zvi [Pak00].

References

Alan:2009:EEE

[AAB09] Sule Alan, Orazio Attanasio, and Martin Browning. Estimating Euler equations with noisy data: two exact GMM estimators. Journal of Applied Econometrics, 24(2):309–324, March 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Asali:2017:CVS

[AAQB17] Muhammad Asali, Aamer S. Abu-Qarn, and Michael Beenstock. The cycle of violence in the Second Intifada: Causality in nonlinear vector autoregressive models. Journal of Applied Econometrics, 32(6):1197–1205, September–October 2017. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anderson:2007:NAL

[AAV07] Heather M. Anderson, George Athanasopoulos, and Farshid Vahid. Nonlinear autoregressive leading indicator models of output in G-7 countries. Journal of Applied Econometrics, 22 (1):63–87, January–February 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

An:2023:IIM

[AAZS23] Zidong An, Salem Abo-Zaid, and Xuguang Simon Sheng. Inattention and the impact of monetary policy. Journal of Applied Econometrics, 38(4):623–643, June–July 2023. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Arulampalam:2000:USY

[AB00] Wiji Arulampalam and Alison L. Booth. Union status of young men in Britain: a decade of change. Journal of Applied Econometrics, 15(3):289–310, May–June 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ando:2016:PDM

[AB16] Tomohiro Ando and Jushan Bai. Panel Data Models with Grouped Factor Structure Under Unknown Group Membership. Journal of Applied Econometrics, 31(1):163–191, January–February 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ahelegbey:2016:BGM

[ABC16] Daniel Felix Ahelegbey, Monica Billio, and Roberto Casarin. Bayesian graphical models for structural vector autoregressive processes. Journal of Applied Econometrics, 31(2):357–386, March 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Amengual:2021:EHG

[ABC21] Dante Amengual, Jesús Bueren, and Julio A. Crego. Endogenous health groups and heterogeneous dynamics of the elderly. Journal of Applied Econometrics, 36(7):878–897, November– December 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Angelini:2019:UAV

[ABCF19] Giovanni Angelini, Emanuele Bacchiocchi, Giovanni Caggiano, and Luca Fanelli. Uncertainty across volatility regimes. Journal of Applied Econometrics, 34(3):437–455, April–May 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Aassve:2002:THM

[ABCP02] Arnstein Aassve, Simon Burgess, Andrew Chesher, and Carol Propper. Transitions from home to marriage of young Americans. Journal of Applied Econometrics, 17(1):1–23, January– February 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Abeysinghe:1994:BRE

[Abe94] Tilak Abeysinghe. Book review: Econometrics: Alchemy or science? essays in econometric methodology, David F. Hendry. Oxford, UK and Cambridge, USA: Blackwell, 1993, pp. 518, ISBN 1-55786-264-8, £50-00 (hardbound). Journal of Applied Econometrics, 9(3):343–344, July–September 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Andersen:2010:CTM

[ABFN10] Torben G. Andersen, Tim Bollerslev, Per Frederiksen, and Morten Ørregaard Nielsen. Continuous-time models, realized volatilities, and testable distributional implications for daily stock returns. Journal of Applied Econometrics, 25(2):233– 261, March 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Allon:2007:NEC

[ABH^{+}07] Gad Allon, Michael Beenstock, Steven Hackman, Ury Passy, and Alexander Shapiro. Nonparametric estimation of concave production technologies by entropic methods. Journal of Applied Econometrics, 22(4):795–816, June–July 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

An:2017:IEO

[ABH^{+}17] Yonghong An, Michael R. Baye, Yingyao Hu, John Morgan, and Matt Shum. Identification and estimation of online price competition with an unknown number of firms. Journal of Applied Econometrics, 32(1):80–102, January–February 2017.

CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Acerenza:2023:TIA

[ABK23] Santiago Acerenza, Otávio Bartalotti, and Désiré Kédagni. Testing identifying assumptions in bivariate probit models. Journal of Applied Econometrics, 38(3):407–422, April–May 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Arin:2013:NGE

[ABKK13] K. Peren Arin, Michael Berlemann, Faik Koray, and Torben Kuhlenkasper. Nonlinear growth effects of taxation: a semi– parametric approach using average marginal tax rates. Journal of Applied Econometrics, 28(5):883–899, August 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Aquaro:2021:EIS

[ABP21] Michele Aquaro, Natalia Bailey, and M. Hashem Pesaran. Estimation and inference for spatial models with heterogeneous coefficients: An application to US house prices. Journal of Applied Econometrics, 36(1):18–44, January–February 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Abrevaya:2006:EES

[Abr06] Jason Abrevaya. Estimating the effect of smoking on birth outcomes using a matched panel data approach. Journal of Applied Econometrics, 21(4):489–519, May–June 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Andrews:1993:BEM

[ABS93] R. W. Andrews, J. O. Berger, and M. H. Smith. Bayesian estimation of manufacturing effects in a fuel economy model. Journal of Applied Econometrics, 8(S1):S5–S18, December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Aronsson:1999:IIB

[ABS99] Thomas Aronsson, Sören Blomquist, and Hans Sacklén. Identifying interdependent behaviour in an empirical model of labour supply. Journal of Applied Econometrics, 14(6):607– 626, November–December 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Amengual:2022:NS

[ABS22] Dante Amengual, Xinyue Bei, and Enrique Sentana. Normal but skewed? Journal of Applied Econometrics, 37(7):1295– 1313, November–December 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Aastveit:2015:WDO

[ABT15] Knut Are Aastveit, Hilde C. Bjørnland, and Leif Anders Thorsrud. What drives oil prices? Emerging versus developed economies. Journal of Applied Econometrics, 30(7):1013–1028, November–December 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Atkinson:1987:BAA

[AC87] Scott E. Atkinson and Thomas D. Crocker. A Bayesian approach to assessing the robustness of hedonic property value studies. Journal of Applied Econometrics, 2(1):27–45, January 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Aastveit:2017:SVR

[ACCM17] Knut Are Aastveit, Andrea Carriero, Todd E. Clark, and Massimiliano Marcellino. Have standard VARS remained stable since the crisis? Journal of Applied Econometrics, 32(5):931– 951, August 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Angelini:2022:BID

[ACF22] Giovanni Angelini, Giuseppe Cavaliere, and Luca Fanelli. Bootstrap inference and diagnostics in state space models: With applications to dynamic macro models. Journal of Applied Econometrics, 37(1):3–22, January–February 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

An:2004:MMP

[ACG04] Mark Y. An, Bent Jesper Christensen, and Nabanita Datta Gupta. Multivariate mixed proportional hazard modelling of the joint retirement of married couples. Journal of Applied Econometrics, 19(6):687–704, 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Aaberge:1999:LSI

[ACS99] Rolf Aaberge, Ugo Colombino, and Steinar Strøm. Labour supply in Italy: an empirical analysis of joint household decisions, with taxes and quantity constraints. Journal of Applied Econometrics, 14(4):403–422, July–August 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ackerberg:2006:CCA

[AD06] Daniel A. Ackerberg and Paul J. Devereux. Comment on ‘The case against JIVE’. Journal of Applied Econometrics, 21(6): 835–838, September–October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [DM06a].

Atkinson:2009:FEF

[AD09] Scott E. Atkinson and Jeffrey H. Dorfman. Feasible estimation of firm-specific allocative inefficiency through Bayesian numerical methods. Journal of Applied Econometrics, 24(4):675–697, June–July 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Adebayo:2004:BGM

[Ade04] Samson B. Adebayo. Bayesian geoadditive modelling of breastfeeding initiation in Nigeria. Journal of Applied Econometrics, 19(2):267–281, March–April 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Adkins:2011:UGM

[Adk11] Lee C. Adkins. Using gretl for Monte Carlo experiments. Journal of Applied Econometrics, 26(5):880–885, August 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Argentesi:2007:EMP

[AF07] Elena Argentesi and Lapo Filistrucchi. Estimating market power in a two-sided market: The case of newspapers. Journal of Applied Econometrics, 22(7):1247–1266, December 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Albergaria:2017:NRF

[AF17] Matheus Albergaria and Luiz Paulo Fávero. Narrow replication of Fisman and Miguel’s (2007a) ‘Corruption, Norms, and Legal Enforcement: Evidence from Diplomatic Parking Tickets’. Journal of Applied Econometrics, 32(4):919–922, June– July 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Angelini:2019:EUI

[AF19] Giovanni Angelini and Luca Fanelli. Exogenous uncertainty and the identification of structural vector autoregressions with external instruments. Journal of Applied Econometrics, 34 (6):951–971, September–October 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See correction [Ano23a].

Aubin:1995:RTP

[AFHI95] Christophe Aubin, Denis Fougère, Emmanuel Husson, and Marc Ivaldi. Real-time pricing of electricity for residential customers: Econometric analysis of an experiment. Journal of Applied Econometrics, 10(S1):S171–S191, December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Armantier:2008:ANE

[AFR08] Olivier Armantier, Jean-Pierre Florens, and Jean-François Richard. Approximation of Nash equilibria in Bayesian games. Journal of Applied Econometrics, 23(7):965–981, November– December 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Aastveit:2017:DFM

[AFR17] Knut Are Aastveit, Claudia Foroni, and Francesco Ravazzolo. Density forecasts with Midas models. Journal of Applied Econometrics, 32(4):783–801, June–July 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Aigner:1989:SSR

[AG89] Dennis J. Aigner and Khalifa Ghali. Self-selection in the residential electricity time-of-use pricing experiments. Journal of Applied Econometrics, 4(S1):S131–S144, December 1989.

CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anglin:1996:SEH

[AG96] Paul M. Anglin and Ramazan Gençay. Semiparametric estimation of a hedonic price function. Journal of Applied Econometrics, 11(6):633–648, November 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Andreou:2002:DMB

[AG02] Elena Andreou and Eric Ghysels. Detecting multiple breaks in financial market volatility dynamics. Journal of Applied Econometrics, 17(5):579–600, September–October 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Amisano:2013:EPS

[AG13] Gianni Amisano and Maria Letizia Giorgetti. Entry into pharmaceutical submarkets: a Bayesian panel probit analysis. Journal of Applied Econometrics, 28(4):667–701, June–July 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Altavilla:2017:ENS

[AG17] Carlo Altavilla and Domenico Giannone. The effectiveness of non-standard monetary policy measures: Evidence from survey data. Journal of Applied Econometrics, 32(5):952–964, August 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anundsen:2016:BCR

[AGHKS16] André K. Anundsen, Karsten Gerdrup, Frank Hansen, and Kasper Kragh-Sørensen. Bubbles and crises: The role of house prices and credit. Journal of Applied Econometrics, 31(7):1291– 1311, November–December 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Alessie:1990:POS

[AGM90] Rob Alessie, Raymond Gradus, and Bertrand Melenberg. The problem of not observing small expenditures in a consumer expenditure survey. Journal of Applied Econometrics, 5(2): 151–166, April–June 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Artis:2007:TMC

[AGM07] Michael Artis, Ana Beatriz Galvão, and Massimiliano Marcellino. The transmission mechanism in a changing world. Journal of Applied Econometrics, 22(1):39–61, January–February 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anesti:2022:UKN

[AGMA22] Nikoleta Anesti, Ana Beatriz Galvão, and Silvia Miranda-Agrippino. Uncertain kingdom: Nowcasting Gross Domestic Product and its revisions. Journal of Applied Econometrics, 37 (1):42–62, January–February 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Altavilla:2017:AYC

[AGR17] Carlo Altavilla, Raffaella Giacomini, and Giuseppe Ragusa. Anchoring the yield curve using survey expectations. Journal of Applied Econometrics, 32(6):1055–1068, September–October 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Amacher:1999:EST

[AH99] Gregory S. Amacher and Daniel Hellerstein. The error structure of time series cross-section hedonic models with sporadic event timing and serial correlation. Journal of Applied Econometrics, 14(3):233–252, May–June 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anenberg:2019:MMC

[AHKM19] Elliot Anenberg, Aurel Hizmo, Edward Kung, and Raven Molloy. Measuring mortgage credit availability: A frontier estimation approach. Journal of Applied Econometrics, 34(6):865– 882, September–October 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Aaberge:2021:RID

[AHM21] Rolf Aaberge, Tarjei Havnes, and Magne Mogstad. Ranking intersecting distribution functions. Journal of Applied Econometrics, 36(6):639–662, September–October 2021. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Abrigo:2022:EHI

[AHM22] Michael R. M. Abrigo, Timothy J. Halliday, and Teresa Molina. Expanding health insurance for the elderly of the Philippines. Journal of Applied Econometrics, 37(3):500–520, April–May 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ahn:2023:ROU

[Ahn23] Hie Joo Ahn. The role of observed and unobserved heterogeneity in the duration of unemployment. Journal of Applied Econometrics, 38(1):3–23, January–February 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Arguea:1994:ECP

[AHT94] N. M. Arguea, C. Hsiao, and G. A. Taylor. Estimating consumer preferences using market data — an application to us automobile demand. Journal of Applied Econometrics, 9(1):1– 18, January–March 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Angrist:1999:JIV

[AIK99] J. D. Angrist, G. W. Imbens, and A. B. Krueger. Jackknife instrumental variables estimation. Journal of Applied Econometrics, 14(1):57–67, January–February 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See replication report [Nak10].

Avdic:2017:AGM

[AJ17] Daniel Avdic and Per Johansson. Absenteeism, gender and the morbidity–mortality paradox. Journal of Applied Econometrics, 32(2):440–462, March 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Aucejo:2019:CGU

[AJ19] Esteban M. Aucejo and Jonathan James. Catching up to girls: Understanding the gender imbalance in educational attainment within race. Journal of Applied Econometrics, 34(4):502–525, June–July 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Aronsson:2024:MLB

[AJL24] Thomas Aronsson, Katharina Jenderny, and Gauthier Lanot. A maximum likelihood bunching estimator of the elasticity of taxable income. Journal of Applied Econometrics, 39(1):200– 216, January–February 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Augurzky:2007:APM

[AK07] Boris Augurzky and Jochen Kluve. Assessing the performance of matching algorithms when selection into treatment is strong. Journal of Applied Econometrics, 22(3):533–557, April–May 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Alquist:2010:WDW

[AK10] Ron Alquist and Lutz Kilian. What do we learn from the price of crude oil futures? Journal of Applied Econometrics, 25(4): 539–573, June–July 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Alessi:2019:RAP

[AK19] Lucia Alessi and Mark Kerssenfischer. The response of asset prices to monetary policy shocks: Stronger than thought. Journal of Applied Econometrics, 34(5):661–672, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Akay:2012:FSC

[Aka12] Alpaslan Akay. Finite-sample comparison of alternative methods for estimating dynamic panel data models. Journal of Applied Econometrics, 27(7):1189–1204, November–December 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Advani:2019:MHS

[AKS19] Arun Advani, Toru Kitagawa, and Tymon Słoczyński. Mostly harmless simulations? Using Monte Carlo studies for estimator selection. Journal of Applied Econometrics, 34(6):893–910, September–October 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Alexander:2006:NMG

[AL06] Carol Alexander and Emese Lazar. Normal mixture GARCH(1,1): applications to exchange rate modelling. Journal of Applied Econometrics, 21(3):307–336, April 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Alaouze:1987:EES

[Ala87] Chris M. Alaouze. Empirical evidence on the sign of the slope of the hazard rate from unemployment from a fixed effects model. Journal of Applied Econometrics, 2(2):159–168, April 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Albuquerque:2003:PLL

[Alb03] Pedro H. Albuquerque. A practical log-linear aggregation method with examples: heterogeneous income growth in the USA. Journal of Applied Econometrics, 18(6):665–678, November–December 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anttonen:2024:SIS

[ALL24] Jetro Anttonen, Markku Lanne, and Jani Luoto. Statistically identified structural VAR model with potentially skewed and fat-tailed errors. Journal of Applied Econometrics, 39(3):422– 437, April–May 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Angelopoulos:2022:CLI

[ALM22] Konstantinos Angelopoulos, Spyridon Lazarakis, and James Malley. Cyclical labour income risk in Great Britain. Journal of Applied Econometrics, 37(1):116–130, January–February 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Alik-Lagrange:2019:EWC

[ALR19] Arthur Alik-Lagrange and Martin Ravallion. Estimating within-cluster spillover effects using a cluster randomization with application to knowledge diffusion in rural India. Journal of Applied Econometrics, 34(1):110–128, January–February 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anderson:2024:MSL

[ALS24] D. Mark Anderson, Yang Liang, and Joseph J. Sabia. Mandatory seatbelt laws and traffic fatalities: a reassessment. Journal of Applied Econometrics, 39(3):513–521, April–May 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Arni:2013:HEU

[ALV13] Patrick Arni, Rafael Lalive, and Jan C. Van Ours. How effective are unemployment benefit sanctions? Looking beyond unemployment exit. Journal of Applied Econometrics, 28(7):1153– 1178, November–December 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Adang:1995:NCI

[AM95] Pim Adang and Bertrand Melenberg. Nonnegativity constraints and intratemporal uncertainty in a multi-good lifecycle model. Journal of Applied Econometrics, 10(1):1–15, January–March 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Audrino:2011:MFS

[AM11] Francesco Audrino and Marcelo C. Medeiros. Modeling and forecasting short-term interest rates: The benefits of smooth regimes, macroeconomic variables, and bagging. Journal of Applied Econometrics, 26(6):999–1022, September–October 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Abrevaya:2020:ICR

[AM20] Jason Abrevaya and Chris Muris. Interval censored regression with fixed effects. Journal of Applied Econometrics, 35(2):198– 216, March 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Amburgey:2023:RTP

[AM23] Aaron J. Amburgey and Michael W. McCracken. On the realtime predictive content of financial condition indices for growth. Journal of Applied Econometrics, 38(2):137–163, March 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anderson:1991:EAA

[And91] G. J. Anderson. Expenditure allocation across nondurables, services, durables and savings: An empirical study of separability in the long run. Journal of Applied Econometrics, 6(2): 153–168, April–June 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anderson:1992:GPS

[And92] Richard G. Anderson. The GAUSS programming system: A review. Journal of Applied Econometrics, 7(2):215–219, April– June 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Andrews:1993:BRA

[And93] Martyn Andrews. Book review: Applied econometric techniques, K. Cuthbertson, S. G. Hall, and M. P. Taylor. Philip Allan, Hemel Hempstead, 1992, ISBN 0-86003-084-9, £45 hardback. Journal of Applied Econometrics, 8(4):423–424, October– December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anderson:2003:PAW

[And03] Gordon Anderson. Poverty in America 1970–1990: who did gain ground? An application of stochastic dominance criteria employing simultaneous inequality tests in a partial panel. Journal of Applied Econometrics, 18(6):621–640, November– December 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anderson:2004:MIA

[And04] Gordon Anderson. Making inferences about the polarization, welfare and poverty of nations: a study of 101 countries 1970–1995. Journal of Applied Econometrics, 19(5):537–550, September–October 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Andreasen:2013:NLD

[And13] Martin M. Andreasen. Non-linear DSGE models and the central difference Kalman filter. Journal of Applied Econometrics, 28(6):929–955, September–October 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1986:CEa

[Ano86a] Anonymous. Calendar of events. Journal of Applied Econometrics, 1(2):209–210, April 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1986:CEb

[Ano86b] Anonymous. Calendar of events. Journal of Applied Econometrics, 1(3):295–296, June 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1986:CEc

[Ano86c] Anonymous. Calender of events. Journal of Applied Econometrics, 1(4):366–367, October 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1986:CPJ

[Ano86d] Anonymous. Call for paper: Journal of forecasting: Special issue on predicting saturation and logistic growth. Journal of Applied Econometrics, 1(3):296, June 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1986:CP

[Ano86e] Anonymous. Call for papers. Journal of Applied Econometrics, 1(4):368, October 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1986:Ma

[Ano86f] Anonymous. Masthead. Journal of Applied Econometrics, 1 (1):fmi, January 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1986:Mb

[Ano86g] Anonymous. Masthead. Journal of Applied Econometrics, 1(2): fmi, April 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1986:Mc

[Ano86h] Anonymous. Masthead. Journal of Applied Econometrics, 1(3): fmi, June 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1986:Md

[Ano86i] Anonymous. Masthead. Journal of Applied Econometrics, 1 (4):fmi, October 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1987:Aa

[Ano87a] Anonymous. Announcement. Journal of Applied Econometrics, 2(2):i, April 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1987:Ab

[Ano87b] Anonymous. Announcement. Journal of Applied Econometrics, 2(3):i, July 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1987:Ac

[Ano87c] Anonymous. Announcement. Journal of Applied Econometrics, 2(4):i, October 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1987:CEa

[Ano87d] Anonymous. Calender of events. Journal of Applied Econometrics, 2(1):83–84, January 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1987:CEb

[Ano87e] Anonymous. Calender of events. Journal of Applied Econometrics, 2(2):169–170, April 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1987:CEc

[Ano87f] Anonymous. Calender of events. Journal of Applied Econometrics, 2(3):257–258, July 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1987:CEd

[Ano87g] Anonymous. Calender of events. Journal of Applied Econometrics, 2(4):351–352, October 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1987:Ma

[Ano87h] Anonymous. Masthead. Journal of Applied Econometrics, 2 (1):fmi, January 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1987:Mb

[Ano87i] Anonymous. Masthead. Journal of Applied Econometrics, 2(2): fmi, April 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1987:Mc

[Ano87j] Anonymous. Masthead. Journal of Applied Econometrics, 2 (3):fmi, July 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1987:Md

[Ano87k] Anonymous. Masthead. Journal of Applied Econometrics, 2 (4):fmi, October 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1988:Ma

[Ano88a] Anonymous. Masthead. Journal of Applied Econometrics, 3 (1):fmi, January 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1988:Mb

[Ano88b] Anonymous. Masthead. Journal of Applied Econometrics, 3(2): fmi, April 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1988:Mc

[Ano88c] Anonymous. Masthead. Journal of Applied Econometrics, 3(3): fmi, July–September 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1988:Md

[Ano88d] Anonymous. Masthead. Journal of Applied Econometrics, 3 (4):fmi, October–December 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1989:AR

[Ano89a] Anonymous. Acknowledgement of referees. Journal of Applied Econometrics, 4(4):315, October–December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1989:Ma

[Ano89b] Anonymous. Masthead. Journal of Applied Econometrics, 4(1): fmi, January–March 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1989:Mb

[Ano89c] Anonymous. Masthead. Journal of Applied Econometrics, 4 (2):fmi, April–June 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1989:Mc

[Ano89d] Anonymous. Masthead. Journal of Applied Econometrics, 4(3): fmi, July–September 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1989:Md

[Ano89e] Anonymous. Masthead. Journal of Applied Econometrics, 4 (4):fmi, October–December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1989:Me

[Ano89f] Anonymous. Masthead. Journal of Applied Econometrics, 4 (S1):fmi, December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1990:BRC

[Ano90a] Anonymous. Book review: Component and Correspondence Analysis: Dimension Reduction by Functional Approximation, edited by J. L. Van Rijckevorsel and J. De Leeuw. Wiley, New York, 1988. ISBN 0-471-9847-4, cloth. £19.95. pp. xii + 146. Journal of Applied Econometrics, 5(4):409–410, September– December 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1990:BRE

[Ano90b] Anonymous. Book review: Empirical macroeconomics for interdependent economies, (two volumes). Edited by R. C.

Bryant, D. W. Henderson, G. Holtham. Hooper and S. Symansky. (The Brookings Institution, Washington, DC, 1988. pp. xvii — 342 and viii + 347, ISBN 0-8157-1 140-9 and 0-8157-1 139-5 (Supplemental Volume). £30.50 the set). Journal of Applied Econometrics, 5(1):96–98, January–March 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1990:BRI

[Ano90c] Anonymous. Book review: Introduction to the Theory and Practice of Econometrics, 2nd ed. G. G. Judge, R. C. Hill, W. E. Griffiths, H. Lutkepohl, and T. Lee. John Wiley, New York, 1988, ISBN 0-471-62414-4, cloth. £85.40. pp. xxxv + 1024. Journal of Applied Econometrics, 5(3):306–308, July– September 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1990:BRMb

[Ano90d] Anonymous. Book review: Macroeconometric modelling of the Indian economy. Studies on inflation and growth, K. Krishnamurty and V. Pandit. Forword by L. R. Klein. (Hindustan Publishing Corporation, Dehli, 1985, pp. xii + 159, ISBN, cloth, £). Journal of Applied Econometrics, 5(2):199–202, April–June 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1990:BRMa

[Ano90e] Anonymous. Book review: Modelling the Labour Market, edited by M. Beenstock. (Chapman and Hall, London, 1988, pp. xii + 206, ISBN 0-412-28830-3, £32.50, hardcover). Journal of Applied Econometrics, 5(2):197–199, April–June 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1990:BRP

[Ano90f] Anonymous. Book review: Praclical methods of optimization, 2nd ed, R. Fletcher. John Wiley and ISBN 0-471-86187, £24.50. pp. xii + 436. Sons, Chichester, 1987. Journal of Applied Econometrics, 5(4):407–409, September–December 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1990:BRS

[Ano90g] Anonymous. Book review: Sensitivity analysis in linear regression, S. Chatterjee and A. S. Hadi. John Wiley and Sons, New York, 1988. ISBN 0-471-82216-7, $25.95 hardbound, pp. xiv + 315. Journal of Applied Econometrics, 5(1):93–96, January– March 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1990:BRT

[Ano90h] Anonymous. Book review: The theory and practice of econometrics, O. G. Judge, W. E. Griffith, R. C. Hill, C. H. Lee, and H. Lütkepohl, second edition, Wiley, New York, 1985. ISBN 0-471-89530-X, £45.65 Hardcover. Pp. xxix + 1019. Journal of Applied Econometrics, 5(3):305–306, July–September 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1990:CPU

[Ano90i] Anonymous. Call for papers. UCLA program on applied econometrics and the journal of applied econometrics. Journal of Applied Econometrics, 5(4):409–410, September–December 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1990:Ma

[Ano90j] Anonymous. Masthead. Journal of Applied Econometrics, 5(1): fmi, January–March 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1990:Mb

[Ano90k] Anonymous. Masthead. Journal of Applied Econometrics, 5 (2):fmi, April–June 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1990:Mc

[Ano90l] Anonymous. Masthead. Journal of Applied Econometrics, 5(3): fmi, July–September 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1990:Md

[Ano90m] Anonymous. Masthead. Journal of Applied Econometrics, 5 (4):fmi, September–December 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1991:EP

[Ano91a] Anonymous. Editorial preface. Journal of Applied Econometrics, 6(4):i, October–December 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1991:ICE

[Ano91b] Anonymous. International conference on: Econometric inference using simulation techniques. Journal of Applied Econometrics, 6(4):478, October–December 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1991:Ma

[Ano91c] Anonymous. Masthead. Journal of Applied Econometrics, 6(1): fmi, January–March 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1991:Mb

[Ano91d] Anonymous. Masthead. Journal of Applied Econometrics, 6 (2):fmi, April–June 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1991:Mc

[Ano91e] Anonymous. Masthead. Journal of Applied Econometrics, 6(3): fmi, July–September 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1991:Md

[Ano91f] Anonymous. Masthead. Journal of Applied Econometrics, 6 (4):fmi, October–December 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1991:UPA

[Ano91g] Anonymous. UCLA Program in Applied Econometrics Conference on Nonlinear Dynamics and Econometrics. April 5–6, 1991. Journal of Applied Econometrics, 6(4):475–477, October– December 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1992:ARa

[Ano92a] Anonymous. Acknowledgement to referees. Journal of Applied Econometrics, 7(1):3, January–March 1992. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1992:ARb

[Ano92b] Anonymous. Acknowledgement to referees. Journal of Applied Econometrics, 7(4):337, October–December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1992:E

[Ano92c] Anonymous. Erratum. Journal of Applied Econometrics, 7(2): 211–213, April–June 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1992:SIC

[Ano92d] Anonymous. For a special issue on calibration techniques and econometrics. Journal of Applied Econometrics, 7(S1):S201, December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1992:Ma

[Ano92e] Anonymous. Masthead. Journal of Applied Econometrics, 7(1): fmi, January–March 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1992:Mb

[Ano92f] Anonymous. Masthead. Journal of Applied Econometrics, 7 (2):fmi, April–June 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1992:Mc

[Ano92g] Anonymous. Masthead. Journal of Applied Econometrics, 7(3): fmi, July–September 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1992:Md

[Ano92h] Anonymous. Masthead. Journal of Applied Econometrics, 7 (4):fmi, October–December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1992:Me

[Ano92i] Anonymous. Masthead. Journal of Applied Econometrics, 7 (S1):fmi, December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1992:TIJ

[Ano92j] Anonymous. Tinbergen Institute and Journal of Applied Econometrics Conference on: Econometric Inference using Simulation Techniques. Erasmus University Rotterdam, The Netherlands 5–6 June, 1992. Journal of Applied Econometrics, 7(S1):S197–S199, December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1993:AR

[Ano93a] Anonymous. Acknowledgement to referees. Journal of Applied Econometrics, 8(4):335, October–December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1993:ASI

[Ano93b] Anonymous. Announcement. A special issue on calibration techniques techniques and econometric. Journal of Applied Econometrics, 8(S1):S175, December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1993:AJA

[Ano93c] Anonymous. Announcement. Journal of Applied Econometrics FTP archive site. Journal of Applied Econometrics, 8(4):339– 340, October–December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1993:CP

[Ano93d] Anonymous. Call for papers. Journal of Applied Econometrics, 8(S1):S177, December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1993:FA

[Ano93e] Anonymous. First announcement. Journal of Applied Econometrics, 8(3):333, July–September 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1993:Ma

[Ano93f] Anonymous. Masthead. Journal of Applied Econometrics, 8(1): fmi, January–March 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1993:Mb

[Ano93g] Anonymous. Masthead. Journal of Applied Econometrics, 8 (2):fmi, April–June 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1993:Mc

[Ano93h] Anonymous. Masthead. Journal of Applied Econometrics, 8(3): fmi, July–September 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1993:Md

[Ano93i] Anonymous. Masthead. Journal of Applied Econometrics, 8 (4):fmi, October–December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1993:Me

[Ano93j] Anonymous. Masthead. Journal of Applied Econometrics, 8 (S1):fmi, December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:AR

[Ano94a] Anonymous. Acknowledgement to referees. Journal of Applied Econometrics, 9(4):349, October–December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:ACP

[Ano94b] Anonymous. Announcement and call for papers. Journal of Applied Econometrics, 9(2):228, April–June 1994. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:A

[Ano94c] Anonymous. Announcements. Journal of Applied Econometrics, 9(4):483, October–December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:CPa

[Ano94d] Anonymous. Call for papers. Journal of Applied Econometrics, 9(1):105, January–March 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:CPb

[Ano94e] Anonymous. Conference programme. Journal of Applied Econometrics, 9(S1):S163, December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:FPa

[Ano94f] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 9(2):227, April–June 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:FPb

[Ano94g] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 9(3):345, July–September 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:FPc

[Ano94h] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 9(4):348, October–December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:GUa

[Ano94i] Anonymous. Guidelines for users. Journal of Applied Econometrics, 9(1):106–107, January–March 1994. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:GUb

[Ano94j] Anonymous. Guidelines for users. Journal of Applied Econometrics, 9(2):229–230, April–June 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:Ma

[Ano94k] Anonymous. Masthead. Journal of Applied Econometrics, 9(1): fmi, January–March 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:Mb

[Ano94l] Anonymous. Masthead. Journal of Applied Econometrics, 9 (2):fmi, April–June 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:Mc

[Ano94m] Anonymous. Masthead. Journal of Applied Econometrics, 9(3): fmi, July–September 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:Md

[Ano94n] Anonymous. Masthead. Journal of Applied Econometrics, 9 (4):fmi, October–December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1994:Me

[Ano94o] Anonymous. Masthead. Journal of Applied Econometrics, 9 (S1):fmi, December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:AR

[Ano95a] Anonymous. Acknowledge to referees. Journal of Applied Econometrics, 10(4):481–482, October–December 1995. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:ACP

[Ano95b] Anonymous. Announcement and call for papers. Journal of Applied Econometrics, 10(1):96, January–March 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:JAEa

[Ano95c] Anonymous. Journal of Applied Econometrics Data Archive: updated information for users. Journal of Applied Econometrics, 10(3):342, July–September 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:JAEb

[Ano95d] Anonymous. Journal of Applied Econometrics Data Archive: Updated information for users. Journal of Applied Econometrics, 10(4):484, October–December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:FPa

[Ano95e] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 10(1):96, January–March 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:FPb

[Ano95f] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 10(2):219, April–June 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:FPc

[Ano95g] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 10(3):343, July–September 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:FPd

[Ano95h] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 10(4):483, October–December 1995. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:GUJ

[Ano95i] Anonymous. Guidelines for users: Journal of Applied Econometrics Data Archive. Journal of Applied Econometrics, 10(2): 217–218, April–June 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:Ma

[Ano95j] Anonymous. Masthead. Journal of Applied Econometrics, 10 (1):fmi, January–March 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:Mb

[Ano95k] Anonymous. Masthead. Journal of Applied Econometrics, 10 (2):fmi, April–June 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:Mc

[Ano95l] Anonymous. Masthead. Journal of Applied Econometrics, 10 (3):fmi, July–September 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:Md

[Ano95m] Anonymous. Masthead. Journal of Applied Econometrics, 10 (4):fmi, October–December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:Me

[Ano95n] Anonymous. Masthead. Journal of Applied Econometrics, 10 (S1):fmi, December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1995:NBE

[Ano95o] Anonymous. National Bureau of Economic Research National Science Foundation Journal of Applied Econometrics: Conference on forecasting. Journal of Applied Econometrics, 10(S1): S193, December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:A

[Ano96a] Anonymous. Announcement. Journal of Applied Econometrics, 11(6):599, November 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:ACP

[Ano96b] Anonymous. Announcements and calls for papers. Journal of Applied Econometrics, 11(1):117, January 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:JAEa

[Ano96c] Anonymous. Journal of Applied Econometrics Data Archive. Journal of Applied Econometrics, 11(2):218, March 1996. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:JAEb

[Ano96d] Anonymous. Journal of Applied Econometrics Data Archive. Journal of Applied Econometrics, 11(3):342, May 1996. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:JAEc

[Ano96e] Anonymous. Journal of Applied Econometrics Data Archive. Journal of Applied Econometrics, 11(4):452, July 1996. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:JAEd

[Ano96f] Anonymous. Journal of Applied Econometrics Data Archive. Journal of Applied Econometrics, 11(5):597, September 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:FPa

[Ano96g] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 11(2):217, March 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:FPb

[Ano96h] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 11(3):341, May 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:FPc

[Ano96i] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 11(4):451, July 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:FPd

[Ano96j] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 11(5):595–596, September 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:Ma

[Ano96k] Anonymous. Masthead. Journal of Applied Econometrics, 11 (1):fmi, January 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:Mb

[Ano96l] Anonymous. Masthead. Journal of Applied Econometrics, 11(2):fmi, March 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:Mc

[Ano96m] Anonymous. Masthead. Journal of Applied Econometrics, 11 (3):fmi, May 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:Md

[Ano96n] Anonymous. Masthead. Journal of Applied Econometrics, 11 (4):fmi, July 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:Me

[Ano96o] Anonymous. Masthead. Journal of Applied Econometrics, 11 (5):fmi, September 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:1996:Mf

[Ano96p] Anonymous. Masthead. Journal of Applied Econometrics, 11 (6):fmi, November 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2001:FPa

[Ano01a] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 16(2):196, March–April 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2001:FPb

[Ano01b] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 16(4):562, July–August 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2001:FPc

[Ano01c] Anonymous. Forthcoming papers. Journal of Applied Econometrics, 16(5):655, September–October 2001. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2001:JAEa

[Ano01d] Anonymous. Journal of applied econometrics annual lecture series. Journal of Applied Econometrics, 16(1):94, January– February 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2001:JAEb

[Ano01e] Anonymous. Journal of applied econometrics annual lecture series. Journal of Applied Econometrics, 16(2):195, March– April 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2003:EJA

[Ano03] Anonymous. Erratum: Journal of Applied Econometrics Vol 18 No 1 2003, pages 47–59. Are differences in firm size transitory or permanent? P. A. Geroski, S. Lazarova, G. Urga and C. F. Walters. Journal of Applied Econometrics, 18(2):249, March– April 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [GLUW03].

Anonymous:2005:AR

[Ano05a] Anonymous. Acknowledgement to referees. Journal of Applied Econometrics, 20(7):947–949, December 2005. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2005:A

[Ano05b] Anonymous. Announcement. Journal of Applied Econometrics, 20(7):943–944, December 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2006:JAE

[Ano06] Anonymous. Journal of applied econometrics distinguished authors. Journal of Applied Econometrics, 21(2):273–274, March 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2007:AR

[Ano07a] Anonymous. Acknowledgement to referees. Journal of Applied Econometrics, 22(7):1397–1398, December 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2007:JAE

[Ano07b] Anonymous. Journal of applied econometrics scholars programme. Journal of Applied Econometrics, 22(3):702, April– May 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2008:AR

[Ano08a] Anonymous. Acknowledgement to referees. Journal of Applied Econometrics, 23(5):697–698, August 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2008:BRJ

[Ano08b] Anonymous. Book review: Journal of Applied Econometrics: Call for papers for special issue: “Forecast Uncertainty in Macroeconomics and Finance”. Journal of Applied Econometrics, 23(1):135, January–February 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2009:ARa

[Ano09a] Anonymous. Acknowledgement to referees. Journal of Applied Econometrics, 24(4):705–707, June–July 2009. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2009:ARb

[Ano09b] Anonymous. Acknowledgement to referees. Journal of Applied Econometrics, 24(7):1215–1217, November–December 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2009:E

[Ano09c] Anonymous. Erratum. Journal of Applied Econometrics, 24 (6):1055, September–October 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2009:EAR

[Ano09d] Anonymous. Erratum: Acknowledgement to referees. Journal of Applied Econometrics, 24(5):867–869, August 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2009:RNS

[Ano09e] Anonymous. Replications in the ‘narrow sense’. Journal of Applied Econometrics, 24(6):1054, September–October 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2010:AR

[Ano10a] Anonymous. Acknowledgement to referees. Journal of Applied Econometrics, 25(6):1069–1072, September–October 2010. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2010:JAE

[Ano10b] Anonymous. Journal of applied econometrics dissertation prize. Journal of Applied Econometrics, 25(3):507, April–May 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2011:JAE

[Ano11] Anonymous. Journal of applied econometrics dissertation prize. Journal of Applied Econometrics, 26(3):547, April–May 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2012:AR

[Ano12a] Anonymous. Acknowledgement to referees. Journal of Applied Econometrics, 27(4):700–703, June–July 2012. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2012:JAE

[Ano12b] Anonymous. Journal of applied econometrics dissertation prize. Journal of Applied Econometrics, 27(3):523, April–May 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2012:RSP

[Ano12c] Anonymous. The Richard stone prize in applied econometrics. Journal of Applied Econometrics, 27(7):1211–1212, November– December 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2013:AR

[Ano13a] Anonymous. Acknowledgement to referees. Journal of Applied Econometrics, 28(7):1231–1234, November–December 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2013:JAE

[Ano13b] Anonymous. Journal of applied econometrics dissertation prize. Journal of Applied Econometrics, 28(3):526, April–May 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2016:CSD

[Ano16] Anonymous. Cross-sectional Dependence in Panel Data Models: A Special Issue. Journal of Applied Econometrics, 31(1): 1–3, January–February 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2018:IIa

[Ano18a] Anonymous. Issue information. Journal of Applied Econometrics, 33(1):i–iv, January–February 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2018:IIb

[Ano18b] Anonymous. Issue information. Journal of Applied Econometrics, 33(2):i–iv, March 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2018:IIc

[Ano18c] Anonymous. Issue information. Journal of Applied Econometrics, 33(3):i–iv, April–May 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2018:IId

[Ano18d] Anonymous. Issue information. Journal of Applied Econometrics, 33(4):i–iv, June–July 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2018:IIe

[Ano18e] Anonymous. Issue information. Journal of Applied Econometrics, 33(5):i–iv, August 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2018:IIf

[Ano18f] Anonymous. Issue information. Journal of Applied Econometrics, 33(6):i–iv, September–October 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2018:IIg

[Ano18g] Anonymous. Issue information. Journal of Applied Econometrics, 33(7):i–iv, November–December 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2019:IIa

[Ano19a] Anonymous. Issue information. Journal of Applied Econometrics, 34(1):i–iv, January–February 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2019:IIb

[Ano19b] Anonymous. Issue information. Journal of Applied Econometrics, 34(2):i–iv, March 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2019:IIc

[Ano19c] Anonymous. Issue information. Journal of Applied Econometrics, 34(3):i–iv, April–May 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2019:IId

[Ano19d] Anonymous. Issue information. Journal of Applied Econometrics, 34(4):i–iv, June–July 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2019:IIe

[Ano19e] Anonymous. Issue information. Journal of Applied Econometrics, 34(5):i–iv, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2019:II

[Ano19f] Anonymous. Issue information. Journal of Applied Econometrics, 34(7):1–4, November–December 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2020:IIa

[Ano20a] Anonymous. Issue information. Journal of Applied Econometrics, 35(1):i–iv, January–February 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2020:IIb

[Ano20b] Anonymous. Issue information. Journal of Applied Econometrics, 35(2):i–iv, March 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2020:IIc

[Ano20c] Anonymous. Issue information. Journal of Applied Econometrics, 35(3):i–iv, April–May 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2020:IId

[Ano20d] Anonymous. Issue information. Journal of Applied Econometrics, 35(4):i–iv, June–July 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2020:IIe

[Ano20e] Anonymous. Issue information. Journal of Applied Econometrics, 35(5):i–iv, August 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2020:IIf

[Ano20f] Anonymous. Issue information. Journal of Applied Econometrics, 35(6):i–iv, September–October 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2020:IIg

[Ano20g] Anonymous. Issue information. Journal of Applied Econometrics, 35(7):i–iv, November–December 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2021:IIa

[Ano21a] Anonymous. Issue information. Journal of Applied Econometrics, 36(1):1–4, January–February 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2021:IIb

[Ano21b] Anonymous. Issue information. Journal of Applied Econometrics, 36(2):163, March 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2021:IIc

[Ano21c] Anonymous. Issue information. Journal of Applied Econometrics, 36(3):271, April–May 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2021:IId

[Ano21d] Anonymous. Issue information. Journal of Applied Econometrics, 36(4):367, June–July 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2021:IIe

[Ano21e] Anonymous. Issue information. Journal of Applied Econometrics, 36(5):493, August 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2021:IIf

[Ano21f] Anonymous. Issue information. Journal of Applied Econometrics, 36(6):637, September–October 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2021:IIg

[Ano21g] Anonymous. Issue information. Journal of Applied Econometrics, 36(7):859, November–December 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2022:C

[Ano22a] Anonymous. Correction. Journal of Applied Econometrics, 37 (4):829–830, June–July 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2022:IIa

[Ano22b] Anonymous. Issue information. Journal of Applied Econometrics, 37(1):i–ii, January–February 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2022:IIb

[Ano22c] Anonymous. Issue information. Journal of Applied Econometrics, 37(2):225, March 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2022:IIc

[Ano22d] Anonymous. Issue information. Journal of Applied Econometrics, 37(3):459, April–May 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2022:IId

[Ano22e] Anonymous. Issue information. Journal of Applied Econometrics, 37(4):665, June–July 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2022:IIe

[Ano22f] Anonymous. Issue information. Journal of Applied Econometrics, 37(5):841, August 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2022:IIf

[Ano22g] Anonymous. Issue information. Journal of Applied Econometrics, 37(6):1091, September–October 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2022:IIg

[Ano22h] Anonymous. Issue information. Journal of Applied Econometrics, 37(7):1275, November–December 2022. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2023:CEU

[Ano23a] Anonymous. Correction to “Exogenous uncertainty and the identification of structural vector autoregressions with external instruments”. Journal of Applied Econometrics, 38(5):795–797, August 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [AF19].

Anonymous:2023:IIa

[Ano23b] Anonymous. Issue information. Journal of Applied Econometrics, 38(2):135, March 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2023:IIb

[Ano23c] Anonymous. Issue information. Journal of Applied Econometrics, 38(3):269, April–May 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2023:IIc

[Ano23d] Anonymous. Issue information. Journal of Applied Econometrics, 38(4):451, June–July 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2023:IId

[Ano23e] Anonymous. Issue information. Journal of Applied Econometrics, 38(5):669, August 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2023:IIe

[Ano23f] Anonymous. Issue information. Journal of Applied Econometrics, 38(6):799, September–October 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2023:IIf

[Ano23g] Anonymous. Issue information. Journal of Applied Econometrics, 38(7):987, November–December 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2024:IIa

[Ano24a] Anonymous. Issue information. Journal of Applied Econometrics, 39(2):235, March 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anonymous:2024:IIb

[Ano24b] Anonymous. Issue information. Journal of Applied Econometrics, 39(3):363, April–May 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Aigner:1994:RSM

[ANT94] D. J. Aigner, J. Newman, and A. Tishler. The response of small and medium-size business customers to Time-of-Use (TOU) electricity rates in Israel. Journal of Applied Econometrics, 9(3):283–304, July–September 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anundsen:2015:ERS

[Anu15] André K. Anundsen. Econometric regime shifts and the US subprime bubble. Journal of Applied Econometrics, 30(1):145– 169, January–February 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Attanasio:2011:ICC

[AP11] Orazio P. Attanasio and Monica Paiella. Intertemporal consumption choices, transaction costs and limited participation in financial markets: reconciling data and theory. Journal of Applied Econometrics, 26(2):322–343, March 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Amini:2012:CMA

[AP12] Shahram M. Amini and Christopher F. Parmeter. Comparison of model averaging techniques: assessing growth determinants. Journal of Applied Econometrics, 27(5):870–876, August 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anzuini:2015:MEP

[APP15] Alessio Anzuini, Patrizio Pagano, and Massimiliano Pisani. Macroeconomic effects of precautionary demand for oil. Journal of Applied Econometrics, 30(6):968–986, September–October 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Athanasopoulos:2016:DLR

[APVY16] George Athanasopoulos, Donald S. Poskitt, Farshid Vahid, and Wenying Yao. Determination of long-run and short-run dynamics in EC–VARMA models via canonical correlations. Journal of Applied Econometrics, 31(6):1100–1119, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Arbia:1993:BRC

[Arb93] Giuseppe Arbia. Book review: Continuous time econometric modelling, A. R. Bergstrom, Advanced Texts In Econometrics, Oxford University Press, Oxford, 1990, ISBN 0-19-828367-9, Paperback $17.50; ISBN 0-19-829340-7, hardback £40.00; pp. x + 326. Journal of Applied Econometrics, 8(2):227–229, April– June 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Abel:2016:MBU

[ARST16] Joshua Abel, Robert Rich, Joseph Song, and Joseph Tracy. The measurement and behavior of uncertainty: Evidence from the ECB survey of professional forecasters. Journal of Applied Econometrics, 31(3):533–550, April–May 2016. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Armantier:2006:ECF

[AS06] Olivier Armantier and Erwann SbaÏ. Estimation and comparison of [French] Treasury auction formats when bidders are asymmetric. Journal of Applied Econometrics, 21(6):745–779, September–October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Abrevaya:2014:ECP

[AS14] Jason Abrevaya and Shu Shen. Estimation of censored panel– data models with slope heterogeneity. Journal of Applied Econometrics, 29(4):523–548, June–July 2014. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ashley:2009:ACI

[Ash09] Richard Ashley. Assessing the credibility of instrumental variables inference with imperfect instruments via sensitivity analysis. Journal of Applied Econometrics, 24(2):325–337, March 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anatolyev:2021:HDF

[ASS21] Stanislav Anatolyev, Sergei Seleznev, and Veronika Selezneva. How does the financial market update beliefs about the real economy? Evidence from the oil market. Journal of Applied Econometrics, 36(7):938–961, November–December 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Audrino:2006:EPM

[AT06] Francesco Audrino and Fabio Trojani. Estimating and predicting multivariate volatility thresholds in global stock markets. Journal of Applied Econometrics, 21(3):345–369, April 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Asdrubali:2019:HRA

[ATV19] Pierfederico Asdrubali, Simone Tedeschi, and Luigi Ventura. Heterogeneity in risk aversion and risk sharing regressions. Journal of Applied Econometrics, 34(5):827–835, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Alfo:2008:TCH

[ATW08] Marco Alfò, Giovanni Trovato, and Robert J. Waldmann. Testing for country heterogeneity in growth models using a finite mixture approach. Journal of Applied Econometrics, 23 (4):487–514, June 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Austin:1994:SVR

[Aus94] G. Austin. SC version 1.107: A review. Journal of Applied Econometrics, 9(2):219–223, April–June 1994. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Anderson:1997:CBI

[AV97] Heather M. Anderson and Farshid Vahid. On the correspondence between individual and aggregate food consumption functions: evidence from the USA and The Netherlands. Journal of Applied Econometrics, 12(5):477–507, September– October 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Altissimo:2001:NLD

[AV01] Filippo Altissimo and Giovanni L. Violante. The non-linear dynamics of output and unemployment in the U.S. Journal of Applied Econometrics, 16(4):461–486, July–August 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Amano:1996:MST

[AW96] Robert A. Amano and Tony S. Wirjanto. Money stock targeting and money supply: A closer examination of the data. Journal of Applied Econometrics, 11(1):93–104, January 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bhalotra:1998:IRA

[BA98] Sonia Bhalotra and Cliff Attfield. Intrahousehold resource allocation in rural Pakistan: a semiparametric analysis. Journal of Applied Econometrics, 13(5):463–480, September–October 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baillie:1989:FMR

[Bai89] Richard T. Baillie. Forecast master: A review. Journal of Applied Econometrics, 4(3):305–307, July–September 1989.

CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bairam:1991:EST

[Bai91] Erkin Bairam. Elasticity of substitution, technical progress and returns to scale in branches of Soviet industry: A new CES production function approach. Journal of Applied Econometrics, 6(1):91–96, January–March 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baillie:1993:BRI

[Bai93] Richard T. Baillie. Book review: Introduction to multiple time series analysis, H. LÜTKEPOHL. Springer-Verlag. Berlin And New York, 1991. ISBN 0-387-53194-7, cloth £31, pp. 1 + 545. Journal of Applied Econometrics, 8(3):325–326, July– September 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baiocchi:2003:MEP

[Bai03] Giovanni Baiocchi. Managing econometric projects using Perl. Journal of Applied Econometrics, 18(3):371–378, May–June 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baiocchi:2009:POO

[Bai09] Giovanni Baiocchi. PDL: an object-oriented programming environment for econometrics. Journal of Applied Econometrics, 24(5):849–856, August 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baltagi:1999:AER

[Bal99] Badi H. Baltagi. Applied econometrics rankings: 1989–1995. Journal of Applied Econometrics, 14(4):423–441, July–August 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baltagi:2006:EEM

[Bal06] Badi H. Baltagi. Estimating an economic model of crime using panel data from North Carolina. Journal of Applied Econometrics, 21(4):543–547, May–June 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baltagi:2010:NRS

[Bal10] Badi H. Baltagi. Narrow replication of Serlenga and Shin (2007) Gravity models of intra-EU trade: application of the CCEP– HT estimation in heterogeneous panels with unobserved common time-specific factors. Journal of Applied Econometrics, 25 (3):505–506, April–May 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [SS07].

Barker:1995:BRM

[Bar95] Terry Barker. Book review: Macroeconomic modelling of the long run, Colin Hargreaves (ed.) Edward Elgar, Aldershot, 1992, ISBN 1-85278-657-4, pp. xv + 394. Price £49.95 (hb). Journal of Applied Econometrics, 10(2):211–212, April–June 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Basistha:2009:HPC

[Bas09] Arabinda Basistha. Hours per capita and productivity: evidence from correlated unobserved components models. Journal of Applied Econometrics, 24(1):187–206, January–February 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Basu:2014:EPC

[Bas14] Anirban Basu. Estimating person-centered treatment (pet) effects using instrumental variables: an application to evaluating prostate cancer treatments. Journal of Applied Econometrics, 29(4):671–691, June–July 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Barigozzi:2019:NNE

[BB19] Matteo Barigozzi and Christian Brownlees. NETS: Network estimation for time series. Journal of Applied Econometrics, 34 (3):347–364, April–May 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Buliskeria:2022:DRB

[BB22] Nino Buliskeria and Jaromir Baxa. Do rural banks matter that much? Burgess and Pande (2005) reconsidered. Journal of Applied Econometrics, 37(6):1266–1274, September–October 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bertoni:2020:WBP

[BBC20] Marco Bertoni, Giorgio Brunello, and Lorenzo Cappellari. Who benefits from privileged peers? Evidence from siblings in schools. Journal of Applied Econometrics, 35(7):893–916, November–December 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Boucher:2014:DPA

[BBDF14] Vincent Boucher, Yann Bramoullé, Habiba Djebbari, and Bernard Fortin. Do peers affect student achievement? Evidence from Canada using group size variation. Journal of Applied Econometrics, 29(1):91–109, January–February 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baltagi:2015:HHP

[BBE15] Badi H. Baltagi, Georges Bresson, and Jean-Michel Etienne. Hedonic housing prices in Paris: An unbalanced spatial lag pseudo-panel model with nested random effects. Journal of Applied Econometrics, 30(3):509–528, April–May 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Burkhauser:2004:DPM

[BBG04] Richard V. Burkhauser, J. S. Butler, and Gulcin Gumus. Dynamic programming model estimates of social security disability insurance application timing. Journal of Applied Econometrics, 19(6):671–685, 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Babii:2024:PDN

[BBGS24] Andrii Babii, Ryan T. Ball, Eric Ghysels, and Jonas Striaukas. Panel data nowcasting: The case of price-earnings ratios. Journal of Applied Econometrics, 39(2):292–307, March 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Brooks:2003:MGM

[BBP03] Chris Brooks, Simon P. Burke, and Gita Persand. Multivariate GARCH models: software choice and estimation issues. Journal of Applied Econometrics, 18(6):725–734, November– December 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baltagi:2007:PUR

[BBP07] Badi H. Baltagi, Georges Bresson, and Alain Pirotte. Panel unit root tests and spatial dependence. Journal of Applied Econometrics, 22(2):339–360, March 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bearse:1997:EEM

[BBS97] Peter M. Bearse, Hamparsum Bozdogan, and Alan M. Schlottmann. Empirical econometric modelling of food consumption using a new informational complexity approach. Journal of Applied Econometrics, 12(5):563–592, September– October 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Boswijk:2019:CD

[BBS19] H. Peter Boswijk, Maurice J. G. Bun, and Maarten Pieter Schinkel. Cartel dating. Journal of Applied Econometrics, 34 (1):26–42, January–February 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bockenholt:2009:DRR

[BBvdH09] Ulf Böckenholt, Sema Barlas, and Peter G. M. van der Heijden. Do randomized-response designs eliminate response biases? An empirical study of non-compliance behavior. Journal of Applied Econometrics, 24(3):377–392, April–May 2009. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bierens:2007:SNC

[BC07a] Herman J. Bierens and Jose R. Carvalho. Semi-nonparametric competing risks analysis of recidivism. Journal of Applied Econometrics, 22(5):971–993, August 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Browning:2007:HHD

[BC07b] Martin Browning and M. Dolores Collado. Habits and heterogeneity in demands: a panel data analysis. Journal of Applied Econometrics, 22(3):625–640, April–May 2007. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bernardi:2019:SGA

[BC19] Mauro Bernardi and Leopoldo Catania. Switching generalized autoregressive score copula models with application to systemic risk. Journal of Applied Econometrics, 34(1):43–65, January– February 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bietenbeck:2023:NEI

[BC23] Jan Bietenbeck and Matthew Collins. New evidence on the importance of instruction time for student achievement on international assessments. Journal of Applied Econometrics, 38 (3):423–431, April–May 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Brewer:2024:ASS

[BC24] Dylan Brewer and Alyssa Carlson. Addressing sample selection bias for machine learning methods. Journal of Applied Econometrics, 39(3):383–400, April–May 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bottasso:2015:RIK

[BCC15] Anna Bottasso, Carolina Castagnetti, and Maurizio Conti. R&D, innovation and knowledge spillovers: A reappraisal of Bottazzi and Peri (2007) in the presence of cross-sectional dependence. Journal of Applied Econometrics, 30(2):350–352, March 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bai:2022:MFM

[BCCM22] Yu Bai, Andrea Carriero, Todd E. Clark, and Massimiliano Marcellino. Macroeconomic forecasting in a multi-country context. Journal of Applied Econometrics, 37(6):1230–1255, September–October 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Basturk:2014:PPE

[BÇCV14] Nalan Bastürk, Cem Çakmakli, S. Pinar Ceyhan, and Herman K. Van Dijk. Posterior-predictive evidence on US inflation using extended new Keynesian Phillips curve models with nonfiltered data. Journal of Applied Econometrics, 29(7):1164– 1182, November–December 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Burkhauser:1999:TSI

[BCDJ99] Richard V. Burkhauser, Amy Crews Cutts, Mary C. Daly, and Stephen P. Jenkins. Testing the significance of income distribution changes over the 1980s business cycle: a cross-national comparison. Journal of Applied Econometrics, 14(3):253–272, May–June 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Barbaglia:2024:FGE

[BCM24] Luca Barbaglia, Sergio Consoli, and Sebastiano Manzan. Forecasting GDP in Europe with textual data. Journal of Applied Econometrics, 39(2):338–355, March 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bhattacharjee:2016:ESR

[BCMM16] Arnab Bhattacharjee, Eduardo Castro, Taps Maiti, and João Marques. Endogenous Spatial Regression and Delineation of Submarkets: A New Framework with Application to Housing Markets. Journal of Applied Econometrics, 31(1):32–57, January–February 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baum:2004:NEE

[BCO04] Christopher F. Baum, Mustafa Caglayan, and Neslihan Ozkan. Nonlinear effects of exchange rate volatility on the volume of bilateral exports. Journal of Applied Econometrics, 19(1):1–23, January–February 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Banerjee:2001:AIM

[BCR01] Anindya Banerjee, Lynne Cockerell, and Bill Russell. An I(2) analysis of inflation and the markup. Journal of Applied Econometrics, 16(3):221–240, May–June 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Billio:2016:IBE

[BCRV16] Monica Billio, Roberto Casarin, Francesco Ravazzolo, and Herman K. Van Dijk. Interconnections between Eurozone and US booms and busts using a Bayesian panel Markov-switching VAR model. Journal of Applied Econometrics, 31(7):1352– 1370, November–December 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baillie:1996:AIF

[BCT96] Richard T. Baillie, Ching-Fan Chung, and Margie A. Tieslau. Analysing inflation by the fractionally integrated ARFIMA– GARCH model. Journal of Applied Econometrics, 11(1):23– 40, January 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bernard:1995:CIO

[BD95] Andrew B. Bernard and Steven N. Durlauf. Convergence in international output. Journal of Applied Econometrics, 10(2): 97–108, April–June 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Blomquist:1999:SSP

[BD99] Sören Blomquist and Matz Dahlberg. Small sample properties of LIML and jackknife IV estimators: experiments with weak instruments. Journal of Applied Econometrics, 14(1):69–88, January–February 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baiocchi:2002:VET

[BD02] Giovanni Baiocchi and Walter Distaso. Visual econometrics: teaching and practising econometrics using ViSta. Journal of Applied Econometrics, 17(4):405–414, July–August 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baiocchi:2003:GES

[BD03] Giovanni Baiocchi and Walter Distaso. GRETL: Econometric software for the GNU generation. Journal of Applied Econometrics, 18(1):105–110, January–February 2003. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bollinger:2005:DTW

[BD05a] Christopher R. Bollinger and Martin H. David. I didn’t tell, and I won’t tell: dynamic response error in the SIPP. Journal of Applied Econometrics, 20(4):563–569, May–June 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Boswijk:2005:DAC

[BD05b] H. Peter Boswijk and Jurgen A. Doornik. Distribution approximations for cointegration tests with stationary exogenous regressors. Journal of Applied Econometrics, 20(6):797–810, September–October 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Blomquist:2006:CAJ

[BD06] Sören Blomquist and Matz Dahlberg. The case against JIVE: a comment. Journal of Applied Econometrics, 21(6):839–841, September–October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [DM06a].

Battistin:2015:IEE

[BD15] Erich Battistin and Michele De Nadai. Identification and estimation of Engel curves with endogenous and unobserved expenditures. Journal of Applied Econometrics, 30(3):487–508, April–May 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Boldea:2020:CPE

[BDG20] Otilia Boldea, Bettina Drepper, and Zhuojiong Gan. Change point estimation in panel data with time-varying individual effects. Journal of Applied Econometrics, 35(6):712–727, September–October 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bae:2007:MDF

[BdJ07] Youngsoo Bae and Robert M. de Jong. Money demand function estimation by nonlinear cointegration. Journal of Applied Econometrics, 22(4):767–793, June–July 2007. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bernard:2012:IRT

[BDKK12] Jean-Thomas Bernard, Jean-Marie Dufour, Lynda Khalaf, and Maral Kichian. An identification-robust test for time-varying parameters in the dynamics of energy prices. Journal of Applied Econometrics, 27(4):603–624, June–July 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Blume:2002:LCS

[BDNS02] Andreas Blume, Douglas V. DeJong, George R. Neumann, and N. E. Savin. Learning and communication in senderreceiver games: an econometric investigation. Journal of Applied Econometrics, 17(3):225–247, May–June 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Blundell:1998:SEC

[BDP98] Richard Blundell, Alan Duncan, and Krishna Pendakur. Semiparametric estimation and consumer demand. Journal of Applied Econometrics, 13(5):435–461, September–October 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bee:2018:REQ

[BDT18] Marco Bee, Debbie J. Dupuis, and Luca Trapin. Realized extreme quantile: A joint model for conditional quantiles and measures of volatility with EVT refinements. Journal of Applied Econometrics, 33(3):398–415, April–May 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bergstrom:1992:TAD

[BE92] R. Bergström and P.-A. Edin. Time aggregation and the distributional shape of unemployment duration. Journal of Applied Econometrics, 7(1):5–30, January–March 1992. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Beck:2001:ACH

[Bec01] Stacie Beck. Autoregressive conditional heteroscedasticity in commodity spot prices. Journal of Applied Econometrics, 16 (2):115–132, March–April 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Becker:2004:GTT

[BEH04] Ralf Becker, Walter Enders, and Stan Hurn. A general test for time dependence in parameters. Journal of Applied Econometrics, 19(7):899–906, November–December 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bekdache:1999:TVB

[Bek99] Basma Bekdache. The time-varying behaviour of real interest rates: a re-evaluation of the recent evidence. Journal of Applied Econometrics, 14(2):171–190, March–April 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Behrens:2012:DGU

[BEK12] Kristian Behrens, Cem Ertur, and Wilfried Koch. ‘dual’ gravity: using spatial econometrics to control for multilateral resistance. Journal of Applied Econometrics, 27(5):773–794, August 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Berlemann:2015:URB

[BEK15] Michael Berlemann, Sören Enkelmann, and Torben Kuhlenkasper. Unraveling the relationship between Presidential approval and the economy: A multidimensional semiparametric approach. Journal of Applied Econometrics, 30(3):468–486, April–May 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baltagi:2016:FLP

[BEK16] Badi H. Baltagi, Peter H. Egger, and Michaela Kesina. Firmlevel productivity spillovers in China’s chemical industry: A spatial Hausman–Taylor approach. Journal of Applied Econometrics, 31(1):214–248, January–February 2016. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baltagi:2022:BEM

[BEK22] Badi H. Baltagi, Peter H. Egger, and Michaela Kesina. Bayesian estimation of multivariate panel probits with higherorder network interdependence and an application to firms’ global market participation in Guangdong. Journal of Applied Econometrics, 37(7):1356–1378, November–December 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Belzil:2001:UIS

[Bel01] Christian Belzil. Unemployment insurance and subsequent job duration: job matching versus unobserved heterogeneity. Journal of Applied Econometrics, 16(5):619–636, September– October 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bauwens:2007:EIO

[BEL07] Luc Bauwens, Alvaro Escribano, and Michel Lubrano. The econometrics of industrial organization. Journal of Applied Econometrics, 22(7):1153–1156, December 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bernstein:1989:EES

[Ber89] Jeffrey I. Bernstein. An examination of the equilibrium specification and structure of production for Canadian telecommunications. Journal of Applied Econometrics, 4(3):265–282, July– September 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bergstrom:2001:SWA

[Ber01] A. R. Bergstrom. Stability and wage acceleration in macroeconomic models of cyclical growth. Journal of Applied Econometrics, 16(3):327–340, May–June 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Berge:2014:FDE

[Ber14] Travis J. Berge. Forecasting disconnected exchange rates. Journal of Applied Econometrics, 29(5):713–735, August 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bewley:1988:AR

[Bew88] Ronald Bewley. Autobox: A review. Journal of Applied Econometrics, 3(3):240–244, July–September 1988. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Beyer:1998:MMD

[Bey98] Andreas Beyer. Modelling money demand in Germany. Journal of Applied Econometrics, 13(1):57–76, January–February 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Balke:1994:LSS

[BF94] Nathan S. Balke and Thomas B. Fomby. Large shocks, small shocks, and economic fluctuations: Outliers in macroeconomic time series. Journal of Applied Econometrics, 9(2):181–200, April–June 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bacchiocchi:2005:TPP

[BF05] Emanuele Bacchiocchi and Luca Fanelli. Testing the purchasing power parity through I(2) cointegration techniques. Journal of Applied Econometrics, 20(6):749–770, September–October 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Balcombe:2009:DCC

[BF09] Kelvin Balcombe and Iain Fraser. Dichotomous-choice contingent valuation with ‘dont know’ responses and misreporting. Journal of Applied Econometrics, 24(7):1137–1152, November– December 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Belaire-Franch:2002:HCB

[BFC02] Jorge Belaire-Franch and Dulce Contreras. How to compute the BDS test: a software comparison. Journal of Applied Econometrics, 17(6):691–699, December 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baltagi:2019:SCH

[BFK19] Badi H. Baltagi, Qu Feng, and Chihwa Kao. Structural changes in heterogeneous panels with endogenous regressors. Journal of Applied Econometrics, 34(6):883–892, September–October 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Balcombe:2015:VAA

[BFM15] Kelvin Balcombe, Iain Fraser, and Eugene McSorley. Visual attention and attribute attendance in multi-attribute choice experiments. Journal of Applied Econometrics, 30(3):447–467, April–May 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Butler:1998:DMC

[BFS98] J. S. Butler, T. Aldrich Finegan, and John J. Siegfried. Does more calculus improve student learning in intermediate microand macroeconomic theory? Journal of Applied Econometrics, 13(2):185–202, March–April 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bergemann:2009:EDE

[BFS09] Annette Bergemann, Bernd Fitzenberger, and Stefan Speckesser. Evaluating the dynamic employment effects of training programs in East Germany using conditional difference-in-differences. Journal of Applied Econometrics, 24(5):797–823, August 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bonomo:1994:CWF

[BG94] M. Bonomo and R. Garcia. Can a well-fitted equilibrium assetpricing model produce mean reversion? Journal of Applied Econometrics, 9(1):19–29, January–March 1994. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Blau:2006:HIR

[BG06] David M. Blau and Donna B. Gilleskie. Health insurance and retirement of married couples. Journal of Applied Econometrics, 21(7):935–953, November 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baltagi:2017:RUQ

[BG17] Badi H. Baltagi and Pallab Kumar Ghosh. Replication of unconditional quantile regressions by Firpo, Fortin and Lemieux (2009). Journal of Applied Econometrics, 32(1):218–223, January–February 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Botosaru:2018:DDW

[BG18a] Irene Botosaru and Federico H. Gutierrez. Difference-in-differences when the treatment status is observed in only one period. Journal of Applied Econometrics, 33(1):73–90, January–February 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bove:2018:GDT

[BG18b] Vincenzo Bove and Gunes Gokmen. Genetic distance, trade, and the diffusion of development. Journal of Applied Econometrics, 33(4):617–623, June–July 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bluhm:2021:FCE

[BGLS21] Richard Bluhm, Martin Gassebner, Sarah Langlotz, and Paul Schaudt. Fueling conflict? (De)escalation and bilateral aid. Journal of Applied Econometrics, 36(2):244–261, March 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Banbura:2010:LBV

[BGR10] Marta Bańbura, Domenico Giannone, and Lucrezia Reichlin. Large Bayesian vector auto regressions. Journal of Applied Econometrics, 25(1):71–92, January–February 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baltagi:1998:ECP

[BGV98] Badi H. Baltagi, James M. Griffin, and Sharada R. Vadali. Excess capacity: a permanent characteristic of US airlines? Journal of Applied Econometrics, 13(6):645–657, November– December 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Belzil:2003:SEI

[BH03] Christian Belzil and Jörgen Hansen. Structural estimates of the intergenerational education correlation. Journal of Applied Econometrics, 18(6):679–696, November–December 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Burda:2013:PPF

[BH13] Martin Burda and Matthew Harding. Panel probit with flexible correlated effects: quantifying technology spillovers in the presence of latent heterogeneity. Journal of Applied Econometrics, 28(6):956–981, September–October 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bocart:2015:VPI

[BH15] Fabian Y. R. P. Bocart and Christian M. Hafner. Volatility of price indices for heterogeneous goods with applications to the fine art market. Journal of Applied Econometrics, 30(2):291– 312, March 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bognanni:2018:SMC

[BH18] Mark Bognanni and Edward Herbst. A sequential Monte Carlo approach to inference in multiple-equation Markov-switching models. Journal of Applied Econometrics, 33(1):126–140, January–February 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Belzil:2020:EUF

[BH20] Christian Belzil and Jörgen Hansen. The evolution of the US family income–schooling relationship and educational selectivity. Journal of Applied Econometrics, 35(7):841–859, November–December 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bugni:2021:PTE

[BH21] Federico A. Bugni and Joel L. Horowitz. Permutation tests for equality of distributions of functional data. Journal of Applied Econometrics, 36(7):861–877, November–December 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Burda:2015:BSC

[BHH15] Martin Burda, Matthew Harding, and Jerry Hausman. A Bayesian semiparametric competing risk model with unobserved heterogeneity. Journal of Applied Econometrics, 30(3): 353–376, April–May 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Barra:2017:JBA

[BHKL17] István Barra, Lennart Hoogerheide, Siem Jan Koopman, and André Lucas. Joint Bayesian analysis of parameters and states in nonlinear non-Gaussian state space models. Journal of Applied Econometrics, 32(5):1003–1026, August 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Borup:2023:QIN

[BHLS23] Daniel Borup, Jorge Wolfgang Hansen, Benjamin Dybro Liengaard, and Erik Christian Montes Schütte. Quantifying investor narratives and their role during COVID-19. Journal of Applied Econometrics, 38(4):512–532, June–July 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bai:2024:RAM

[BHLTM24] Yuehao Bai, Meng Hsuan Hsieh, Jizhou Liu, and Max Tabord-Meehan. Revisiting the analysis of matched-pair and stratified experiments in the presence of attrition. Journal of Applied Econometrics, 39(2):256–268, March 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Beck:2016:ISR

[BHM16] Guenter W. Beck, Kirstin Hubrich, and Massimiliano Marcellino. On the importance of sectoral and regional shocks for price-setting. Journal of Applied Econometrics, 31(7):1234– 1253, November–December 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Burda:1998:SAG

[BHMW98] Michael C. Burda, Wolfgang Härdle, Marlene Müller, and Axel Werwatz. Semiparametric analysis of German East-West migration intentions: facts and theory. Journal of Applied Econometrics, 13(5):525–541, September–October 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bauwens:2013:MVM

[BHP13] Luc Bauwens, Christian M. Hafner, and Diane Pierret. Multivariate volatility modeling of electricity futures. Journal of Applied Econometrics, 28(5):743–761, August 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bailey:2016:TSA

[BHP16] Natalia Bailey, Sean Holly, and M. Hashem Pesaran. A Two-Stage Approach to Spatio-Temporal Analysis with Strong and Weak Cross-Sectional Dependence. Journal of Applied Econometrics, 31(1):249–280, January–February 2016. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bianchi:1997:TCE

[Bia97] Marco Bianchi. Testing for convergence: evidence from nonparametric multimodality tests. Journal of Applied Econometrics, 12(4):393–409, July 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Biewen:2009:MSD

[Bie09] Martin Biewen. Measuring state dependence in individual poverty histories when there is feedback to employment status and household composition. Journal of Applied Econometrics, 24(7):1095–1116, November–December 2009. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bernard:2007:FSI

[BIKY07] Jean-Thomas Bernard, Nadhem Idoudi, Lynda Khalaf, and Clément Yélou. Finite sample inference methods for dynamic energy demand models. Journal of Applied Econometrics, 22 (7):1211–1226, December 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bilias:2000:STD

[Bil00] Yannis Bilias. Sequential testing of duration data: the case of the Pennsylvania ‘reemployment bonus’ experiment. Journal of Applied Econometrics, 15(6):575–594, November–December 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Banerjee:2015:CPD

[BiS15] Anindya Banerjee and Josep Lluís Carrion i Silvestre. Cointegration in panel data with structural breaks and cross-section dependence. Journal of Applied Econometrics, 30(1):1–23, January–February 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bonhomme:2009:PAC

[BJ09] Stéphane Bonhomme and Grégory Jolivet. The pervasive absence of compensating differentials. Journal of Applied Econometrics, 24(5):763–795, August 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bretteville-Jensen:2011:CDS

[BJJ11] Anne Line Bretteville-Jensen and Liana Jacobi. Climbing the drug staircase: a Bayesian analysis of the initiation of hard drug use. Journal of Applied Econometrics, 26(7):1157–1186, November–December 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

BenMansour:2008:RAA

[BJM^{+}08] Selima Ben Mansour, Elyès Jouini, Jean-Michel Marin, Clotilde Napp, and Christian Robert. Are risk-averse agents more optimistic? A Bayesian estimation approach. Journal of Applied Econometrics, 23(6):843–860, September–October 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bera:1986:APP

[BK86] Anil K. Bera and Srinivasan Kannan. An adjustment procedure for predicting systematic risk. Journal of Applied Econometrics, 1(4):317–332, October 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bloemen:2008:EUC

[BK08] Hans G. Bloemen and Arie Kapteyn. The estimation of utilityconsistent labor supply models by means of simulated scores. Journal of Applied Econometrics, 23(4):395–422, June 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Becker:2017:EEC

[BK17] Martin Becker and Stefan Klößner. Estimating the economic costs of organized crime by synthetic control methods. Journal of Applied Econometrics, 32(7):1367–1369, November– December 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Berger:2019:TTV

[BK19] Tino Berger and Bernd Kempa. Testing for time variation in the natural rate of interest. Journal of Applied Econometrics, 34(5):836–842, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Beyeler:2021:RFF

[BK21a] Simon Beyeler and Sylvia Kaufmann. Reduced-form factor augmented VAR-Exploiting sparsity to include meaningful factors. Journal of Applied Econometrics, 36(7):989–1012, November– December 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Breitung:2021:HFC

[BK21b] Jörg Breitung and Malte Knüppel. How far can we forecast? Statistical tests of the predictive content. Journal of Applied Econometrics, 36(4):369–392, June–July 2021. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baltagi:1990:EEP

[BKA90] Badi H. Baltagi and Sophon Khanti-Akom. On efficient estimation with panel data: An empirical comparison of instrumental variables estimators. Journal of Applied Econometrics, 5(4):401–406, September–December 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See replication report [DG03].

Blazejowski:2020:MSV

[BKK20] Marcin Błazejowski, Paweł Kufel, and Jacek Kwiatkowski. Model simplification and variable selection: a replication of the UK inflation model by Hendry (2001). Journal of Applied Econometrics, 35(5):645–652, August 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bauwens:2015:CSB

[BKKR15] Luc Bauwens, Gary Koop, Dimitris Korobilis, and Jeroen V. K. Rombouts. The contribution of structural break models to forecasting macroeconomic series. Journal of Applied Econometrics, 30(4):596–620, June–July 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Beckmann:2020:ERP

[BKKS20] Joscha Beckmann, Gary Koop, Dimitris Korobilis, and Rainer Alexander Schüssler. Exchange rate predictability and dynamic Bayesian learning. Journal of Applied Econometrics, 35(4):410–421, June–July 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baumeister:2017:ICB

[BKL17] Christiane Baumeister, Lutz Kilian, and Thomas K. Lee. Inside the crystal ball: New approaches to predicting the gasoline price at the pump. Journal of Applied Econometrics, 32(2): 275–295, March 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bowlus:1995:EEW

[BKN95] Audra J. Bowlus, Nicholas M. Kiefer, and George R. Neumann. Estimation of equilibrium wage distributions with heterogeneity. Journal of Applied Econometrics, 10(S1):S119–S131, December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bailey:2016:ECS

[BKP16] Natalia Bailey, George Kapetanios, and M. Hashem Pesaran. Exponent of cross-sectional dependence: Estimation and inference. Journal of Applied Econometrics, 31(6):929–960, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bailey:2021:MFS

[BKP21] Natalia Bailey, George Kapetanios, and M. Hashem Pesaran. Measurement of factor strength: Theory and practice. Journal of Applied Econometrics, 36(5):587–613, August 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bockerman:2018:KMY

[BKS18] Petri Böckerman, Ohto Kanninen, and Ilpo Suoniemi. A kink that makes you sick: The effect of sick pay on absence. Journal of Applied Econometrics, 33(4):568–579, June–July 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baccouche:1991:DSP

[BL91] Rafiq Baccouche and François Laisney. Describing the separability properties of empirical demand systems. Journal of Applied Econometrics, 6(2):181–206, April–June 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bruggemann:2006:SMS

[BL06] Ralf Brüggemann and Helmut Lütkepohl. A small monetary system for the euro area based on German data. Journal of Applied Econometrics, 21(6):683–702, September–October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baltagi:2014:FES

[BL14a] Badi H. Baltagi and Jing Li. Further evidence on the spatio– temporal model of house prices in the United States. Journal of Applied Econometrics, 29(3):515–522, April–May 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See replication report [Mil15].

Benati:2014:SIR

[BL14b] Luca Benati and Thomas A. Lubik. Sales, inventories and real interest rates: a century of stylized facts. Journal of Applied Econometrics, 29(7):1210–1222, November–December 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Boneva:2017:DCM

[BL17] Lena Boneva and Oliver Linton. A discrete-choice model for large heterogeneous panels with interactive fixed effects with an application to the determinants of corporate bond issuance. Journal of Applied Econometrics, 32(7):1226–1243, November– December 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Black:1995:RS

[Bla95] R. J. Black. Review of SIMPC 3.81. Journal of Applied Econometrics, 10(4):469–476, October–December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Blattenberger:1996:MDR

[Bla96] Gail Blattenberger. Money demand revisited: An operational subjective approach. Journal of Applied Econometrics, 11(2): 153–168, March 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Blevins:2016:SMC

[Ble16] Jason R. Blevins. Sequential Monte Carlo methods for estimating dynamic microeconomic models. Journal of Applied Econometrics, 31(5):773–804, August 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bao:2010:GIV

[BLM10] Yong Bao, Melody Lo, and Franklin G. Mixon, Jr. Generalinterest versus specialty journals: Using intellectual influence of econometrics research to rank economics journals and articles. Journal of Applied Econometrics, 25(2):345–353, March 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Blough:1992:RBP

[Blo92] Stephen R. Blough. The relationship between power and level for generic unit root tests in finite samples. Journal of Applied Econometrics, 7(3):295–308, July–September 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bloemen:2002:RBW

[Blo02] Hans G. Bloemen. The relation between wealth and labour market transitions: an empirical study for the Netherlands.

Journal of Applied Econometrics, 17(3):249–268, May–June 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bauwens:2006:MGM

[BLR06] Luc Bauwens, Sébastien Laurent, and Jeroen V. K. Rombouts. Multivariate GARCH models: a survey. Journal of Applied Econometrics, 21(1):79–109, January–February 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bond:2010:CAG

[BLS10] Steve Bond, Asli Leblebicioğlu, and Fabio Schiantarelli. Capital accumulation and growth: a new look at the empirical evidence. Journal of Applied Econometrics, 25(7):1073–1099, November– December 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Boneva:2016:EFT

[BLV16] Lena Boneva, Oliver Linton, and Michael Vogt. The Effect of Fragmentation in Trading on Market Quality in the UK Equity Market. Journal of Applied Econometrics, 31(1):192–213, January–February 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Blundell:1986:SCM

[BM86] Richard Blundell and Costas Meghir. Selection criteria for a microeconometric model of labour supply. Journal of Applied Econometrics, 1(1):55–80, January 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baillie:1991:BGE

[BM91] Richard T. Baillie and Robert J. Myers. Bivariate GARCH estimation of the optimal commodity futures hedge. Journal of Applied Econometrics, 6(2):109–124, April–June 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bidarkota:1998:OUI

[BM98] Prasad V. Bidarkota and J. Huston McCulloch. Optimal univariate inflation forecasting with symmetric stable shocks. Journal of Applied Econometrics, 13(6):659–670, November– December 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Banerjee:2006:RIL

[BM06] Anindya Banerjee and Paul Mizen. A re-interpretation of the linear quadratic model when inventories and sales are polynomially cointegrated. Journal of Applied Econometrics, 21(8): 1249–1264, December 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bellemare:2011:IMA

[BM11] Charles Bellemare and Charles F. Manski. Introduction: ‘measurement and analysis of subjective expectations’. Journal of Applied Econometrics, 26(3):351, April–May 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bontemps:2012:TDA

[BM12] Christian Bontemps and Nour Meddahi. Testing distributional assumptions: A GMM approach. Journal of Applied Econometrics, 27(6):978–1012, September–October 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Banbura:2014:MLE

[BM14] Marta Bańbura and Michele Modugno. Maximum likelihood estimation of factor models on datasets with arbitrary pattern of missing data. Journal of Applied Econometrics, 29(1):133– 160, January–February 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Barigozzi:2016:IIS

[BM16] Matteo Barigozzi and Alessio Moneta. Identifying the independent sources of consumption variation. Journal of Applied Econometrics, 31(2):420–449, March 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bertolotti:2019:TSH

[BM19] Fabio Bertolotti and Massimiliano Marcellino. Tax shocks with high and low uncertainty. Journal of Applied Econometrics, 34(6):972–993, September–October 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Banerjee:2017:SFC

[BMM17] Anindya Banerjee, Massimiliano Marcellino, and Igor Masten. Structural FECM: Cointegration in large-scale structural FAVAR models. Journal of Applied Econometrics, 32(6):1069– 1086, September–October 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Boer:2023:MIU

[BMR23] Lukas Boer, Lukas Menkhoff, and Malte Rieth. The multifaceted impact of US trade policy on financial markets. Journal of Applied Econometrics, 38(3):388–406, April–May 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Brown:1993:I

[BMV93] Bryan W. Brown, Alain Monfort, and Herman K. Van Dijk. Introduction. Journal of Applied Econometrics, 8(S1):S1–S3, December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bos:2000:DER

[BMV00] Charles S. Bos, Ronald J. Mahieu, and Herman K. Van Dijk. Daily exchange rate behaviour and hedging of currency risk. Journal of Applied Econometrics, 15(6):671–696, November– December 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Black:1990:RCG

[BMW90a] Matthew Black, Robert Moffitt, and John Warner. Reply to comment by Glenn Gotz on “The dynamics of job separation: The case of federal employees”. Journal of Applied Econometrics, 5(3):269–272, July–September 1990. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [BMW90b, Got90].

Black:1990:DJS

[BMW90b] Matthew Black, Robert Moffitt, and John T. Warner. The dynamics of job separation: The case of federal employees. Journal of Applied Econometrics, 5(3):245–262, July–September 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bernhardt:2024:HDD

[BMW24] Robert Bernhardt, David Munro, and Erin L. Wolcott. How does the dramatic rise of nonresponse in the Current Population Survey impact labor market indicators? Journal of Applied Econometrics, 39(3):498–512, April–May 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bai:2009:BDI

[BN09] Jushan Bai and Serena Ng. Boosting diffusion indices. Journal of Applied Econometrics, 24(4):607–629, June–July 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bien:2011:IMI

[BNP11] Katarzyna Bien, Ingmar Nolte, and Winfried Pohlmeier. An inflated multivariate integer count hurdle model: an application to bid and ask quote dynamics. Journal of Applied Econometrics, 26(4):669–707, June–July 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bjornland:2021:SFS

[BNR21] Hilde C. Bjørnland, Frode Martin Nordvik, and Maximilian Rohrer. Supply flexibility in the shale patch: Evidence from North Dakota. Journal of Applied Econometrics, 36(3):273– 292, April–May 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Barndorff-Nielsen:2002:EQV

[BNS02] Ole E. Barndorff-Nielsen and Neil Shephard. Estimating quadratic variation using realized variance. Journal of Applied Econometrics, 17(5):457–477, September–October 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Brownlees:2018:RN

[BNS18] Christian Brownlees, Eulàlia Nualart, and Yucheng Sun. Realized networks. Journal of Applied Econometrics, 33(7):986– 1006, November–December 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Borland:1994:DAT

[BO94] J. Borland and S. Ouliaris. The determinants of Australian trade union membership. Journal of Applied Econometrics, 9(4):453–468, October–December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Becker:2012:WST

[BO12] Ralf Becker and Denise R. Osborn. Weighted smooth transition regressions. Journal of Applied Econometrics, 27(5):795–811, August 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bollino:1992:BRR

[Bol92] Carlo Andrea Bollino. Book review: Recent modelling approaches in applied energy economics, O. Bjerkholt, O. Olsen and J. Vislie (eds) Chapman and Hall, London, 1990, pp. xix + 268, ISBN 0-412-35340-7,£40 hardbound. Journal of Applied Econometrics, 7(4):436–439, October–December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bai:2003:CAM

[BP03] Jushan Bai and Pierre Perron. Computation and analysis of multiple structural change models. Journal of Applied Econometrics, 18(1):1–22, January–February 2003. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bali:2006:TRR

[BP06] Turan G. Bali and Lin Peng. Is there a risk-return trade-off? Evidence from high-frequency data. Journal of Applied Econometrics, 21(8):1169–1198, December 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baltagi:2007:HCS

[BP07] Badi H. Baltagi and M. Hashem Pesaran. Heterogeneity and cross section dependence in panel data models: theory and applications introduction. Journal of Applied Econometrics, 22 (2):229–232, March 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Blavatskyy:2010:MSC

[BP10] Pavlo R. Blavatskyy and Ganna Pogrebna. Models of stochastic choice and decision theories: why both are important for analyzing decisions. Journal of Applied Econometrics, 25(6): 963–986, September–October 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Baumeister:2013:RTV

[BP13] Christiane Baumeister and Gert Peersman. The role of time– varying price elasticities in accounting for volatility changes in the crude oil market. Journal of Applied Econometrics, 28 (7):1087–1109, November–December 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bernardini:2018:PDO

[BP18] Marco Bernardini and Gert Peersman. Private debt overhang and the government spending multiplier: Evidence for the United States. Journal of Applied Econometrics, 33(4): 485–508, June–July 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bhat:2015:NUC

[BPC15] Chandra R. Bhat, Rajesh Paleti, and Marisol Castro. A new utility-consistent econometric approach to multivariate count data modeling. Journal of Applied Econometrics, 30(5):806– 825, August 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bosco:2010:LRR

[BPPB10] Bruno Bosco, Lucia Parisio, Matteo Pelagatti, and Fabio Baldi. Long-run relations in European electricity prices. Journal of Applied Econometrics, 25(5):805–832, August 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bollino:2000:LHT

[BPR00] C. Andrea Bollino, Federico Perali, and Nicola Rossi. Linear household technologies. Journal of Applied Econometrics, 15 (3):275–287, May–June 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bollerslev:2016:DHP

[BPW16] Tim Bollerslev, Andrew J. Patton, and Wenjing Wang. Daily house price indices: Construction, modeling, and longer-run predictions. Journal of Applied Econometrics, 31(6):1005–1025, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Blundell:1999:ELD

[BR99] Richard Blundell and Jean Marc Robin. Estimation in large and disaggregated demand systems: an estimator for conditionally linear systems. Journal of Applied Econometrics, 14 (3):209–232, May–June 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Budnik:2023:ISV

[BR23] Katarzyna Budnik and Gerhard Rünstler. Identifying structural VARs from sparse narrative instruments: Dynamic effects of US macroprudential policies. Journal of Applied Econometrics, 38(2):186–201, March 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Brady:2011:MDH

[Bra11] Ryan R. Brady. Measuring the diffusion of housing prices across space and over time. Journal of Applied Econometrics, 26(2): 213–231, March 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Brillet:1989:EMM

[Bri89] Jean-Louis Brillet. Econometric modelling on microcomputers: A review of major software packages. Journal of Applied Econometrics, 4(1):73–92, January–March 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Berenguer-Rico:2011:RSS

[BRiS11] Vanessa Berenguer-Rico and Josep Lluís Carrion i Silvestre. Regime shifts in stock-flow I(2)–I(1) systems: the case of US fiscal sustainability. Journal of Applied Econometrics, 26(2): 298–321, March 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Browne:1987:SQA

[Bro87] F. X. Browne. Sluggish quantity adjustment in a non-clearing market — a disequilibrium econometric application to the loan market. Journal of Applied Econometrics, 2(4):335–349, October 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Browning:1994:BRU

[Bro94] Martin Browning. Book review: Understanding consumption, Angus Deaton, Oxford University Press, Oxford, 1992, ISBN 0-19-828759-3, £17.95, hard cover, pp. 240. Journal of Applied Econometrics, 9(1):101–103, January–March 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Balboa:2021:MFI

[BRRT21] Marina Balboa, Paulo M. M. Rodrigues, Antonio Rubia, and A. M. Robert Taylor. Multivariate fractional integration tests allowing for conditional heteroskedasticity with an application to return volatility and trading volume. Journal of Applied Econometrics, 36(5):544–565, August 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bruckner:2013:SPA

[Brü13] Markus Brückner. On the simultaneity problem in the aid and growth debate. Journal of Applied Econometrics, 28(1):126– 150, January–February 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bruze:2018:IMN

[Bru18] Gustaf Bruze. Intergenerational mobility: New evidence from consumption data. Journal of Applied Econometrics, 33(4): 580–593, June–July 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Brzezinski:2015:RRA

[Brz15] Michał Brzezinski. Relative risk aversion and power-law distribution of macroeconomic disasters. Journal of Applied Econometrics, 30(1):170–175, January–February 2015. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Breslaw:1995:SEM

[BS95] Jon A. Breslaw and J. Barry Smith. A simple and efficient method for estimating the magnitude and precision of welfare changes. Journal of Applied Econometrics, 10(3):313–327, July–September 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Basistha:2004:WWC

[BS04] Arabinda Basistha and Richard Startz. Why were changes in the federal funds rate smaller in the 1990s? Journal of Applied Econometrics, 19(3):339–354, May–June 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Banzhaf:2007:MAM

[BS07] H. Spencer Banzhaf and V. Kerry Smith. Meta-analysis in model implementation: choice sets and the valuation of air quality improvements. Journal of Applied Econometrics, 22 (6):1013–1031, September–October 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Barnett:2008:RMV

[BS08] William A. Barnett and Ousmane Seck. Rotterdam model versus almost ideal demand system: will the best specification please stand up? Journal of Applied Econometrics, 23(6):795– 824, September–October 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Beare:2016:ETP

[BS16] Brendan K. Beare and Lawrence D. W. Schmidt. An empirical test of pricing kernel monotonicity. Journal of Applied Econometrics, 31(2):338–356, March 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bronsard:1986:CAD

[BSB86] Camille Bronsard and Lise Salvas-Bronsard. Commodity and asset demands with and without quantity constraints in the labour market. Journal of Applied Econometrics, 1(2):185– 208, April 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Benitez-Silva:2004:HLB

[BSBC^{+}04] Hugo Benítez-Silva, Moshe Buchinsky, Hiu Man Chan, Sofia Cheidvasser, and John Rust. How large is the bias in selfreported disability? Journal of Applied Econometrics, 19(6): 649–670, 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bontemps:2008:SHP

[BSS08] Christophe Bontemps, Michel Simioni, and Yves Surry. Semiparametric hedonic price models: assessing the effects of agricultural nonpoint source pollution. Journal of Applied Econometrics, 23(6):825–842, September–October 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bellemare:2011:MWP

[BSS11] Charles Bellemare, Alexander Sebald, and Martin Strobel. Measuring the willingness to pay to avoid guilt: estimation using equilibrium and stated belief models. Journal of Applied Econometrics, 26(3):437–453, April–May 2011. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Boero:2015:MCP

[BSW15] Gianna Boero, Jeremy Smith, and Kenneth F. Wallis. The measurement and characteristics of professional forecasters’ uncertainty. Journal of Applied Econometrics, 30(7):1029–1046, November–December 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bi:2014:EFL

[BT14] Huixin Bi and Nora Traum. Estimating fiscal limits: the case of Greece. Journal of Applied Econometrics, 29(7):1053–1072, November–December 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bjornland:2019:CPF

[BT19] Hilde C. Bjørnland and Leif Anders Thorsrud. Commodity prices and fiscal policy design: Procyclical despite a rule. Journal of Applied Econometrics, 34(2):161–180, March 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bietenbeck:2023:REG

[BT23] Jan Bietenbeck and Petra Thiemann. Revisiting the effect of growing up in a recession on attitudes towards redistribution. Journal of Applied Econometrics, 38(5):786–794, August 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Buchinsky:1998:DCF

[Buc98] Moshe Buchinsky. The dynamics of changes in the female wage distribution in the USA: a quantile regression approach. Journal of Applied Econometrics, 13(1):1–30, January–February 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Burgess:1992:NDS

[Bur92] S. M. Burgess. Nonlinear dynamics in a structural model of employment. Journal of Applied Econometrics, 7(S1):S101– S118, December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Busetti:2006:TSI

[Bus06] Fabio Busetti. Tests of seasonal integration and cointegration in multivariate unobserved component models. Journal of Applied Econometrics, 21(4):419–438, May–June 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Benoit:2012:BQR

[BV12] Dries F. Benoit and Dirk Van den Poel. Binary quantile regression: a Bayesian approach based on the asymmetric Laplace distribution. Journal of Applied Econometrics, 27(7):1174– 1188, November–December 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Benoit:2016:ORB

[BVV16] Dries F. Benoit, Stefan Van Aelst, and Dirk Van den Poel. Outlier-robust Bayesian multinomial choice modeling. Journal of Applied Econometrics, 31(7):1445–1466, November– December 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Balke:2009:MFV

[BW09] Nathan S. Balke and Mark E. Wohar. Market fundamentals versus rational bubbles in stock prices: a Bayesian perspective. Journal of Applied Econometrics, 24(1):35–75, January– February 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bai:2011:CMC

[BW11] Jushan Bai and Peng Wang. Conditional Markov chain and its application in economic time series analysis. Journal of Applied Econometrics, 26(5):715–734, August 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Byron:1987:SR

[Byr87] R. P. Byron. Shazam: A review. Journal of Applied Econometrics, 2(1):79–82, January 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Byron:1992:PAC

[Byr92] R. P. Byron. Polynomial approximations in cross-sectional models. Journal of Applied Econometrics, 7(3):309–322, July– September 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Bianchi:1998:UPD

[BZ98] Marco Bianchi and Gylfi Zoega. Unemployment persistence: does the size of the shock matter? Journal of Applied Econometrics, 13(3):283–304, May–June 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Campolieti:2001:BSE

[Cam01] Michele Campolieti. Bayesian semiparametric estimation of discrete duration models: an application of the Dirichlet process prior. Journal of Applied Econometrics, 16(1):1–22, January–February 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Campbell:2011:CLM

[Cam11] Jeffrey R. Campbell. Competition in large markets. Journal of Applied Econometrics, 26(7):1113–1136, November–December 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Canova:1994:SIC

[Can94] Fabio Canova. Statistical inference in calibrated models. Journal of Applied Econometrics, 9(S1):S123–S144, December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Canova:2005:TUS

[Can05] Fabio Canova. The transmission of US shocks to Latin America. Journal of Applied Econometrics, 20(2):229–251, 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carlucci:1988:BRE

[Car88] Francesco Carlucci. Book review: Econometric Methods, third edition, By J. Johnston. (McGraw-Hill, New York, 1984, pp. vii + 568, ISBN 0-07-032685, $28.95). Journal of Applied Econometrics, 3(3):249–254, July–September 1988. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carlucci:1992:BRS

[Car92] Francesco Carlucci. Book review: State-Space Modeling of Time Series, Masanao Aoki. Springer Verlag, Berlin, 1987. ISBN 3-540-17257-2, paperback, DM 49.80, pp. xi + 314. Journal of Applied Econometrics, 7(1):109–110, January–March 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cardoso:2000:WDA

[Car00] A. R. Cardoso. Wage differentials across firms: an application of multilevel modelling. Journal of Applied Econometrics, 15 (4):343–354, July–August 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Catania:2017:DSA

[CB17] Leopoldo Catania and Anna Gloria Billé. Dynamic spatial autoregressive models with autoregressive and heteroskedastic disturbances. Journal of Applied Econometrics, 32(6):1178– 1196, September–October 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cuba-Borda:2019:LEM

[CBGIZ19] Pablo Cuba-Borda, Luca Guerrieri, Matteo Iacoviello, and Molin Zhong. Likelihood evaluation of models with occasionally binding constraints. Journal of Applied Econometrics, 34 (7):1073–1085, November–December 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Choi:2007:SHT

[CC07] In Choi and Timothy K. Chue. Subsampling hypothesis tests for nonstationary panels with applications to exchange rates and stock prices. Journal of Applied Econometrics, 22(2):233– 264, March 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Caggiano:2023:GFU

[CC23] Giovanni Caggiano and Efrem Castelnuovo. Global financial uncertainty. Journal of Applied Econometrics, 38(3):432–449, April–May 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Caetano:2023:RDD

[CCE23] Carolina Caetano, Gregorio Caetano, and Juan Carlos Escanciano. Regression discontinuity design with multivalued treatments. Journal of Applied Econometrics, 38(6):840–856, September–October 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cohen-Cole:2018:MCI

[CCLZ18] Ethan Cohen-Cole, Xiaodong Liu, and Yves Zenou. Multivariate choices and identification of social interactions. Journal of Applied Econometrics, 33(2):165–178, March 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carriero:2015:BVS

[CCM15] Andrea Carriero, Todd E. Clark, and Massimiliano Marcellino. Bayesian VARs: Specification choices and forecast accuracy. Journal of Applied Econometrics, 30(1):46–73, January– February 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Coroneo:2018:TOM

[CCM18] Laura Coroneo, Valentina Corradi, and Paulo Santos Monteiro. Testing for optimal monetary policy via moment inequalities. Journal of Applied Econometrics, 33(6):780–796, September– October 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carriero:2020:AIC

[CCM20] Andrea Carriero, Todd E. Clark, and Massimiliano Marcellino. Assessing international commonality in macroeconomic uncertainty and its effects. Journal of Applied Econometrics, 35(3): 273–293, April–May 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carriero:2021:NAP

[CCM21] Andrea Carriero, Todd E. Clark, and Massimiliano Marcellino. No-arbitrage priors, drifting volatilities, and the term structure of interest rates. Journal of Applied Econometrics, 36(5):495– 516, August 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carriero:2022:NTR

[CCM22a] Andrea Carriero, Todd E. Clark, and Massimiliano Marcellino. Nowcasting tail risk to economic activity at a weekly frequency. Journal of Applied Econometrics, 37(5):843–866, August 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carriero:2022:GCI

[CCM22b] Andrea Carriero, Francesco Corsello, and Massimiliano Marcellino. The global component of inflation volatility. Journal of Applied Econometrics, 37(4):700–721, June–July 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Caggiano:2022:UMP

[CCN22] Giovanni Caggiano, Efrem Castelnuovo, and Gabriela Nodari. Uncertainty and monetary policy in good and bad times: a replication of the vector autoregressive investigation by Bloom (2009). Journal of Applied Econometrics, 37(1):210– 217, January–February 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chernis:2023:RDB

[CCV23] Tony Chernis, Patrick J. Coe, and Shaun P. Vahey. Reassessing the dependence between economic growth and financial conditions since 1973. Journal of Applied Econometrics, 38(2):260– 267, March 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Christoffersen:1996:FRF

[CD96] Peter F. Christoffersen and Francis X. Diebold. Further results on forecasting and model selection under asymmetric loss. Journal of Applied Econometrics, 11(5):561–571, September 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Crepon:1997:EIF

[CD97] Bruno Crépon and Emmanuel Duguet. Estimating the innovation function from patent numbers: GMM on count panel data. Journal of Applied Econometrics, 12(3):243–263, May 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cockx:2005:DDE

[CD05] Bart Cockx and Muriel Dejemeppe. Duration dependence in the exit rate out of unemployment in Belgium. Is it true or spurious? Journal of Applied Econometrics, 20(1):1–23, January– February 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cimadomo:2016:CTV

[CD16] Jacopo Cimadomo and Antonello D’Agostino. Combining time variation and mixed frequencies: an analysis of government spending multipliers in Italy. Journal of Applied Econometrics, 31(7):1276–1290, November–December 2016. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chang:2018:GFI

[CD18] Minsu Chang and Francis J. DiTraglia. A generalized focused information criterion for GMM. Journal of Applied Econometrics, 33(3):378–397, April–May 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carlino:2020:RSL

[CD20] Gerald Carlino and Thorsten Drautzburg. The role of startups for local labor markets. Journal of Applied Econometrics, 35(6):751–775, September–October 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Curdia:2014:RSG

[CDG14] Vasco Cúrdia, Marco Del Negro, and Daniel L. Greenwald. Rare shocks, great recessions. Journal of Applied Econometrics, 29(7):1031–1052, November–December 2014. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Coglianese:2017:ATA

[CDKS17] John Coglianese, Lucas W. Davis, Lutz Kilian, and James H. Stock. Anticipation, tax avoidance, and the price elasticity of gasoline demand. Journal of Applied Econometrics, 32(1):1–15, January–February 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chan:2017:EEB

[CE17] Joshua C. C. Chan and Eric Eisenstat. Efficient estimation of Bayesian VARMAs with time-varying coefficients. Journal of Applied Econometrics, 32(7):1277–1297, November–December 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chan:2018:BMC

[CE18] Joshua C. C. Chan and Eric Eisenstat. Bayesian model comparison for time-varying parameter VARs with stochastic volatility. Journal of Applied Econometrics, 33(4):509–532, June–July 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Czarnitzki:2007:RBR

[CEF07] Dirk Czarnitzki, Bernd Ebersberger, and Andreas Fier. The relationship between R&D collaboration, subsidies and R&D performance: Empirical evidence from Finland and Germany. Journal of Applied Econometrics, 22(7):1347–1366, December 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cipollini:2013:SVM

[CEG13] Fabrizio Cipollini, Robert F. Engle, and Giampiero M. Gallo. Semiparametric vector mem. Journal of Applied Econometrics, 28(7):1067–1086, November–December 2013. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Choi:2022:GBS

[CEG22] Jinho Choi, Juan Carlos Escanciano, and Junjie Guo. Generalized band spectrum estimation with an application to the New Keynesian Phillips curve. Journal of Applied Econometrics, 37(5):1055–1078, August 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chan:2020:CLM

[CEHK20] Joshua C. C. Chan, Eric Eisenstat, Chenghan Hou, and Gary Koop. Composite likelihood methods for large Bayesian VARs with stochastic volatility. Journal of Applied Econometrics, 35(6):692–711, September–October 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ching:2009:PCM

[CEK09] Andrew Ching, Tülin Erdem, and Michael Keane. The price consideration model of brand choice. Journal of Applied Econometrics, 24(3):393–420, April–May 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chirinko:1994:EFM

[CF94] R. S. Chirinko and S. M. Fazzari. Economic fluctuations, market power, and returns to scale: Evidence from firm-level data. Journal of Applied Econometrics, 9(1):47–69, January–March 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cuaresma:2013:SFM

[CF13] Jesús Crespo Cuaresma and Martin Feldkircher. Spatial filtering, model uncertainty and the speed of income convergence in Europe. Journal of Applied Econometrics, 28(4):720–741, June–July 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Castelnuovo:2015:MPI

[CF15] Efrem Castelnuovo and Luca Fanelli. Monetary policy indeterminacy and identification failures in the U.S.: Results from a robust test. Journal of Applied Econometrics, 30(6):924–947, September–October 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cavaliere:2008:IDR

[CFG08] Giuseppe Cavaliere, Luca Fanelli, and Attilio Gardini. International dynamic risk sharing. Journal of Applied Econometrics, 23(1):1–16, January–February 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chiaie:2022:CFC

[CFG22] Simona Delle Chiaie, Laurent Ferrara, and Domenico Giannone. Common factors of commodity prices. Journal of Applied Econometrics, 37(3):461–476, April–May 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cuaresma:2016:FGV

[CFH16] Jesús Crespo Cuaresma, Martin Feldkircher, and Florian Huber. Forecasting with global vector autoregressive models: a Bayesian approach. Journal of Applied Econometrics, 31 (7):1371–1391, November–December 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Crepon:2018:ISE

[CFJvdB18] Bruno Crépon, Marc Ferracci, Gregory Jolivet, and Gerard J. van den Berg. Information shocks and the empirical evaluation of training programs during unemployment spells. Journal of Applied Econometrics, 33(4):594–616, June–July 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Caudill:1995:CNR

[CFK95] Steven B. Caudill, Jon M. Ford, and David L. Kaserman. Certificate-of-need regulation and the diffusion of innovations: A random coefficient model. Journal of Applied Econometrics, 10(1):73–78, January–March 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Collard:2002:SMU

[CFLP02] Fabrice Collard, Patrick Fève, François Langot, and Corinne Perraudin. A structural model of US aggregate job flows. Journal of Applied Econometrics, 17(3):197–223, May–June 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Canova:2014:CVE

[CFM14] Fabio Canova, Filippo Ferroni, and Christian Matthes. Choosing the variables to estimate singular DSGE models. Journal of Applied Econometrics, 29(7):1099–1117, November–December 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chintagunta:2009:ISI

[CFP09] Pradeep Chintagunta, Philip Hans Franses, and Richard Paap. Introduction to the special issue on new econometric models in marketing. Journal of Applied Econometrics, 24(3):375–376, April–May 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clements:2009:FUO

[CG09] Michael P. Clements and Ana Beatriz Galvão. Forecasting US output growth using leading indicators: an appraisal using MI-DAS models. Journal of Applied Econometrics, 24(7):1187– 1206, November–December 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clements:2013:RTF

[CG13] Michael P. Clements and Ana Beatriz Galvão. Real-time forecasting of inflation and output growth with autoregressive models in the presence of data revisions. Journal of Applied Econometrics, 28(3):458–477, April–May 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chaudhuri:2016:GMM

[CG16] Saraswata Chaudhuri and David K. Guilkey. GMM with multiple missing variables. Journal of Applied Econometrics, 31(4): 678–706, June–July 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clements:2023:DFB

[CG23] Michael P. Clements and Ana Beatriz Galvão. Density forecasting with Bayesian Vector Autoregressive models under macroeconomic data uncertainty. Journal of Applied Econometrics, 38 (2):164–185, March 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Campolieti:2014:TVD

[CGK14] Michele Campolieti, Deborah Gefang, and Gary Koop. Time variation in the dynamics of worker flows: evidence from North America and Europe. Journal of Applied Econometrics, 29(2): 265–290, March 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cati:1999:URP

[CGP99] Regina Celia Cati, Marcio G. P. Garcia, and Pierre Perron. Unit roots in the presence of abrupt governmental interventions with an application to Brazilian data. Journal of Applied Econometrics, 14(1):27–56, January–February 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Casas:2021:TVI

[CGPX21] Isabel Casas, Jiti Gao, Bin Peng, and Shangyu Xie. Timevarying income elasticities of healthcare expenditure for the OECD and Eurozone. Journal of Applied Econometrics, 36(3): 328–345, April–May 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Christensen:2004:SIE

[CGR04] Bent Jesper Christensen, Nabanita Datta Gupta, and John Rust. Special issue on the econometrics of social insurance. Journal of Applied Econometrics, 19(6):647–648, 2004. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Choi:2018:WIR

[CGS18] Jaerim Choi, Jiaying Gu, and Shu Shen. Weak-instrument robust inference for two-sample instrumental variables regression. Journal of Applied Econometrics, 33(1):109–125, January– February 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cronin:2019:MED

[CGS19] Christopher J. Cronin, David K. Guilkey, and Ilene S. Speizer. Measurement error in discrete health facility choice models: An example from urban Senegal. Journal of Applied Econometrics, 34(7):1102–1120, November–December 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clements:1995:FCS

[CH95] Michael P. Clements and David F. Hendry. Forecasting in cointegrated systems. Journal of Applied Econometrics, 10(2):127– 146, April–June 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clements:1996:ICS

[CH96] Michael P. Clements and David F. Hendry. Intercept corrections and structural change. Journal of Applied Econometrics, 11(5):475–494, September 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carvalho:2005:CTC

[CH05] Vasco M. Carvalho and Andrew C. Harvey. Convergence in the trends and cycles of Euro-zone income. Journal of Applied Econometrics, 20(2):275–289, 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chung:2007:MFE

[CH07] Jaehun Chung and Yongmiao Hong. Model-free evaluation of directional predictability in foreign exchange markets. Journal of Applied Econometrics, 22(5):855–889, August 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ciccarelli:2010:FUS

[CH10a] Matteo Ciccarelli and Kirstin Hubrich. Forecast uncertainty: sources, measurement and evaluation. Journal of Applied Econometrics, 25(4):509–513, June–July 2010. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clements:2010:FET

[CH10b] Michael P. Clements and David I. Harvey. Forecast encompassing tests and probability forecasts. Journal of Applied Econometrics, 25(6):1028–1062, September–October 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Copeland:2011:RPS

[CH11] Adam Copeland and George Hall. The response of prices, sales, and output to temporary changes in demand. Journal of Applied Econometrics, 26(2):232–269, March 2011. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chambers:1992:ECT

[Cha92] Marcus J. Chambers. Estimation of a continuous-time dynamic demand system. Journal of Applied Econometrics, 7(1):53– 64, January–March 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chamberlain:2000:EAM

[Cha00] Gary Chamberlain. Econometric applications of maxmin expected utility. Journal of Applied Econometrics, 15(6):625–644, November–December 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chattopadhyay:2002:DAH

[Cha02] Sudip Chattopadhyay. Divergence in alternative Hicksian welfare measures: the case of revealed preference for public amenities. Journal of Applied Econometrics, 17(6):641–666, December 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chan:2005:RRL

[Cha05] Joshua C. C. Chan. Replication of the results in ‘learning about heterogeneity in returns to schooling’. Journal of Applied Econometrics, 20(3):439–443, March–April 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chang:2011:SET

[Cha11] Sheng-Kai Chang. Simulation estimation of two-tiered dynamic panel tobit models with an application to the labor supply of married women. Journal of Applied Econometrics, 26(5):854– 871, August 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chakravarty:2015:GBB

[Cha15] Abhishek Chakravarty. Gender-biased breastfeeding in Egypt: Examining the fertility preference hypotheses of Jayachandran and Kuziemko (2011). Journal of Applied Econometrics, 30(5): 848–855, August 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chesher:1998:IDH

[Che98] Andrew Chesher. Individual demands from household aggregates: time and age variation in the composition of diet.

Journal of Applied Econometrics, 13(5):505–524, September– October 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chen:2008:UAS

[Che08] Yi-Ting Chen. A unified approach to standardized-residuals-based correlation tests for GARCH-type models. Journal of Applied Econometrics, 23(1):111–133, January–February 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chen:2020:DSC

[Che20] Yi-Ting Chen. A distributional synthetic control method for policy evaluation. Journal of Applied Econometrics, 35(5):505– 525, August 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chen:2023:DAK

[Che23] Shiu-Sheng Chen. A direct approach to Kilian–Lewis style counterfactual analysis in vector autoregression models. Journal of Applied Econometrics, 38(7):1068–1076, November– December 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Caldara:2014:PTP

[CHL14] Dario Caldara, Richard Harrison, and Anna Lipińska. Practical tools for policy analysis in DSGE models with missing shocks. Journal of Applied Econometrics, 29(7):1145–1163, November– December 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chou:1988:VPS

[Cho88] Ray Yeutien Chou. Volatility persistence and stock valuations: Some empirical evidence using GARCH. Journal of Applied Econometrics, 3(4):279–294, October–December 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Choi:1999:TRW

[Cho99a] In Choi. Testing the random walk hypothesis for real exchange rates. Journal of Applied Econometrics, 14(3):293–308, May– June 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Choi:1999:EDR

[Cho99b] Woon Gyu Choi. Estimating the discount rate policy reaction function of the monetary authority. Journal of Applied Econometrics, 14(4):379–401, July–August 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cho:2011:EMM

[Cho11] Sung-Jin Cho. An empirical model of mainframe computer investment. Journal of Applied Econometrics, 26(1):122–150, January–February 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chang:2023:OPU

[CHP23] Yoosoon Chang, Ana María Herrera, and Elena Pesavento. Oil prices uncertainty, endogenous regime switching, and inflation anchoring. Journal of Applied Econometrics, 38(6):820–839, September–October 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cai:2024:PSE

[CHP24] Jun Cai, William C. Horrace, and Christopher F. Parmeter. Penalized sieve estimation of zero-inefficiency stochastic frontiers. Journal of Applied Econometrics, 39(1):41–65, January– February 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chu:2017:RRB

[CHW17] Chi-Yang Chu, Daniel J. Henderson, and Le Wang. The robust relationship between US food aid and civil conflict. Journal of Applied Econometrics, 32(5):1027–1032, August 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Canova:1991:TSP

[CI91] Fabio Canova and Takatoshi Ito. The time-series properties of the risk premium in the yen/dollar exchange market. Journal of Applied Econometrics, 6(2):125–142, April–June 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Coroneo:2020:CPA

[CI20] Laura Coroneo and Fabrizio Iacone. Comparing predictive accuracy in small samples using fixed-smoothing asymptotics.

Journal of Applied Econometrics, 35(4):391–409, June–July 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cifarelli:1988:BRM

[Cif88] Giulio Cifarelli. Book review: Measurement error models, W. A. Fuller, John Wiley and Sons, New York, 1987. ISBN 0-471-86187-1. $44.95, pp. xxiii + 440. Journal of Applied Econometrics, 3(4):315–317, October–December 1988. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cincera:1997:PRT

[Cin97] Michele Cincera. Patents, R&D, and technological spillovers at the firm level: some evidence from econometric count models for panel data. Journal of Applied Econometrics, 12(3):265– 280, May 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cameron:1997:CDR

[CJ97] A. Colin Cameron and Per Johansson. Count data regression using series expansions: with applications. Journal of Applied Econometrics, 12(3):203–223, May 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cappellari:2004:MLI

[CJ04] Lorenzo Cappellari and Stephen P. Jenkins. Modelling low income transitions. Journal of Applied Econometrics, 19(5): 593–610, September–October 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chib:2016:BFR

[CJ16] Siddhartha Chib and Liana Jacobi. Bayesian fuzzy regression discontinuity analysis and returns to compulsory schooling. Journal of Applied Econometrics, 31(6):1026–1047, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Choi:2020:DVN

[CJ20] In Choi and Hanbat Jeong. Differencing versus nondifferencing in factor-based forecasting. Journal of Applied Econometrics, 35(6):728–750, September–October 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carlson:2024:SSL

[CJ24] Alyssa Carlson and Riju Joshi. Sample selection in linear panel data models with heterogeneous coefficients. Journal of Applied Econometrics, 39(2):237–255, March 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Contoyannis:2004:DHB

[CJR04] Paul Contoyannis, Andrew M. Jones, and Nigel Rice. The dynamics of health in the British household panel survey. Journal of Applied Econometrics, 19(4):473–503, July–August 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clements:2007:EFF

[CJS07] Michael P. Clements, Fred Joutz, and Herman O. Stekler. An evaluation of the forecasts of the federal reserve: a pooled approach. Journal of Applied Econometrics, 22(1):121–136, January–February 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Corradi:2023:RFS

[CJS23] Valentina Corradi, Sainan Jin, and Norman R. Swanson. Robust forecast superiority testing with an application to assessing pools of expert forecasters. Journal of Applied Econometrics, 38(4):596–622, June–July 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chan:2022:APR

[CJZ22] Joshua C. C. Chan, Liana Jacobi, and Dan Zhu. An automated prior robustness analysis in Bayesian model comparison. Journal of Applied Econometrics, 37(3):583–602, April–May 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chen:2002:TIE

[CK02] Yi-Ting Chen and Chung-Ming Kuan. Time irreversibility and EGARCH effects in US stock index returns. Journal of Applied Econometrics, 17(5):565–578, September–October 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chortareas:2013:HPP

[CK13] Georgios Chortareas and George Kapetanios. How puzzling is the PPP puzzle? An alternative half-life measure of convergence to PPP. Journal of Applied Econometrics, 28(3):435–457, April–May 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chen:2014:SPE

[CK14] Stacey H. Chen and Shakeeb Khan. Semi-parametric estimation of program impacts on dispersion of potential wages. Journal of Applied Econometrics, 29(6):901–919, September– October 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Conrad:2020:I

[CK20] Christian Conrad and Onno Kleen. Two are better than one: Volatility forecasting using multiplicative component GARCH– MIDAS mode. Journal of Applied Econometrics, 35(1):19–45, January–February 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Choi:2018:MFM

[CKKK18] In Choi, Dukpa Kim, Yun Jung Kim, and Noh-Sun Kwark. A multilevel factor model: Identification, asymptotic theory and applications. Journal of Applied Econometrics, 33(3):355–377, April–May 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Creal:2013:GAS

[CKL13] Drew Creal, Siem Jan Koopman, and André Lucas. Generalized autoregressive score models with applications. Journal of Applied Econometrics, 28(5):777–795, August 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carriero:2011:FLD

[CKM11] Andrea Carriero, George Kapetanios, and Massimiliano Marcellino. Forecasting large datasets with Bayesian reduced rank multivariate models. Journal of Applied Econometrics, 26(5): 735–761, August 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Callot:2017:MFL

[CKM17] Laurent A. F. Callot, Anders B. Kock, and Marcelo C. Medeiros. Modeling and forecasting large realized covariance matrices and portfolio choice. Journal of Applied Econometrics, 32(1):140–158, January–February 2017. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chan:2016:BMT

[CKP16] Joshua C. C. Chan, Gary Koop, and Simon M. Potter. A bounded model of time variation in trend inflation, Nairu and the Phillips Curve. Journal of Applied Econometrics, 31(3): 551–565, April–May 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Caetano:2019:TCE

[CKT19] Gregorio Caetano, Josh Kinsler, and Hao Teng. Towards causal estimates of children’s time allocation on skill development. Journal of Applied Econometrics, 34(4):588–605, June–July 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Creal:2010:ERU

[CKZ10] Drew Creal, Siem Jan Koopman, and Eric Zivot. Extracting a robust US business cycle using a time-varying multivariate model-based bandpass filter. Journal of Applied Econometrics, 25(4):695–719, June–July 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Caggiano:2009:IOC

[CL09a] Giovanni Caggiano and Leone Leonida. International output convergence: evidence from an autocorrelation function approach. Journal of Applied Econometrics, 24(1):139–162, January–February 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chen:2009:LAE

[CL09b] Xiaohong Chen and Sydney C. Ludvigson. Land of addicts? An empirical investigation of habit-based asset pricing models. Journal of Applied Econometrics, 24(7):1057–1093, November– December 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chang:2011:MLM

[CL11] Jae Bong Chang and Jayson L. Lusk. Mixed logit models: accuracy and software choice. Journal of Applied Econometrics, 26 (1):167–172, January–February 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Conrad:2015:ALT

[CL15] Christian Conrad and Karin Loch. Anticipating long-term stock market volatility. Journal of Applied Econometrics, 30 (7):1090–1114, November–December 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chen:2018:ECG

[CL18] Le-Yu Chen and Sokbae Lee. Exact computation of GMM estimators for instrumental variable quantile regression models. Journal of Applied Econometrics, 33(4):553–567, June–July 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chang:2023:RLH

[CL23] Andrew C. Chang and Trace J. Levinson. Raiders of the lost high-frequency forecasts: New data and evidence on the efficiency of the Fed’s forecasting. Journal of Applied Econometrics, 38(1):88–104, January–February 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clark:2006:DEP

[Cla06] Todd E. Clark. Disaggregate evidence on the persistence of consumer price inflation. Journal of Applied Econometrics, 21 (5):563–587, July–August 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clements:1994:BRN

[Cle94] Michael P. Clements. Book review: New directions in econometric practice, W. W. Charemza and D. F. Deadman, Edward Elgar Publishing. Aldershot, 1992. ISBN 1-85278-461-x, hardcover, £45.00, pp. 370. Journal of Applied Econometrics, 9 (1):99–101, January–March 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clements:2022:IFP

[Cle22] Michael P. Clements. Individual forecaster perceptions of the persistence of shocks to GDP. Journal of Applied Econometrics, 37(3):640–656, April–May 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chou:2004:HIS

[CLH04] Shin-Yi Chou, Jin-Tan Liu, and Cliff J. Huang. Health insurance and savings over the life cycle — a semiparametric smooth coefficient estimation. Journal of Applied Econometrics, 19(3): 295–322, May–June 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Co:2006:ISD

[CLLY06] Catherine Y. Co, John S. Landon-Lane, and Myeong-Su Yun. Inter-state dynamics of invention activities, 1930–2000. Journal of Applied Econometrics, 21(8):1111–1134, December 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Crafts:1989:CLV

[CLM89] N. F. R. Crafts, S. J. Leybourne, and T. C. Mills. The climacteric in Late Victorian Britain and France: A reappraisal of the evidence. Journal of Applied Econometrics, 4(2):103–117, April–June 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Creedy:1996:NLM

[CLM96] John Creedy, Jenny Lye, and Vance L. Martin. A non-linear model of the real US/UK exchange rate. Journal of Applied Econometrics, 11(6):669–686, November 1996. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chang:2018:CCC

[CLM18] Andrew C. Chang, Phillip Li, and Shawn M. Martin. Comparing cross-country estimates of Lorenz curves using a Dirichlet distribution across estimators and datasets. Journal of Applied Econometrics, 33(3):473–478, April–May 2018. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Crossley:2022:RID

[CLP22] Thomas F. Crossley, Peter Levell, and Stavros Poupakis. Regression with an imputed dependent variable. Journal of Applied Econometrics, 37(7):1277–1294, November–December 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chicoine:2021:IHA

[CLS21] Luke Chicoine, Emily Lyons, and Alexia Sahue. The impact of HIV/AIDS on human capital investment in Sub-Saharan Africa: New evidence. Journal of Applied Econometrics, 36(6): 842–852, September–October 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Canova:2010:ETS

[CLSM10] Fabio Canova, David Lopez-Salido, and Claudio Michelacci. The effects of technology shocks on hours and output: a robustness analysis. Journal of Applied Econometrics, 25(5):755– 773, August 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Coulombe:2022:HML

[CLSS22] Philippe Goulet Coulombe, Maxime Leroux, Dalibor Stevanovic, and Stéphane Surprenant. How is machine learning useful for macroeconomic forecasting? Journal of Applied Econometrics, 37(5):920–964, August 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carneiro:2017:AMR

[CLU17] Pedro Carneiro, Michael Lokshin, and Nithin Umapathi. Average and marginal returns to upper secondary schooling in Indonesia. Journal of Applied Econometrics, 32(1):16–36, January–February 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cameron:2001:EUH

[CM01] Gavin Cameron and John Muellbauer. Earnings, unemployment, and housing in Britain. Journal of Applied Econometrics, 16(3):203–220, May–June 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chan:2002:MLE

[CM02] Felix Chan and Michael McAleer. Maximum likelihood estimation of STAR and STAR–GARCH models: theory and Monte Carlo evidence. Journal of Applied Econometrics, 17(5):509– 534, September–October 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carro:2006:DMC

[CM06] Jesús M. Carro and Pedro Mira. A dynamic model of contraceptive choice of Spanish couples. Journal of Applied Econometrics, 21(7):955–980, November 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Christodoulakis:2009:APE

[CM09] G. A. Christodoulakis and E. C. Mamatzakis. Assessing the prudence of economic forecasts in the EU. Journal of Applied Econometrics, 24(4):583–606, June–July 2009. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clark:2010:AFV

[CM10] Todd E. Clark and Michael W. McCracken. Averaging forecasts from VARs with uncertain instabilities. Journal of Applied Econometrics, 25(1):5–29, January–February 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clark:2014:TEF

[CM14] Todd E. Clark and Michael W. Mccracken. Tests of equal forecast accuracy for overlapping models. Journal of Applied Econometrics, 29(3):415–430, April–May 2014. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Croushore:2016:RRP

[CM16] Dean Croushore and Katherine Marsten. Reassessing the relative power of the yield spread in forecasting recessions. Journal of Applied Econometrics, 31(6):1183–1191, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clark:2017:TPA

[CM17a] Todd E. Clark and Michael W. McCracken. Tests of predictive ability for vector autoregressions used for conditional forecasting. Journal of Applied Econometrics, 32(3):533–553, April– May 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Colvin:2017:WES

[CM17b] Christopher L. Colvin and Matthew McCracken. Work ethic, social ethic, no ethic: Measuring the economic values of modern Christians. Journal of Applied Econometrics, 32(5):1043–1053, August 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chavleishvili:2024:FST

[CM24a] Sulkhan Chavleishvili and Simone Manganelli. Forecasting and stress testing with quantile vector autoregression. Journal of Applied Econometrics, 39(1):66–85, January–February 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Crudu:2024:NWR

[CM24b] Federico Crudu and Advait Moharir. Narrow and wide replication of Chalfin and McCrary (REStat, 2018). Journal of Applied Econometrics, 39(1):217–224, January–February 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Calderon:2015:ICP

[CMBS15] César Calderón, Enrique Moral-Benito, and Luis Servén. Is infrastructure capital productive? A dynamic heterogeneous approach. Journal of Applied Econometrics, 30(2):177–198, March 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Camba-Mendez:2004:STM

[CML04] Gonzalo Camba-Mendez and Ana Lamo. Short-term monitoring of fiscal policy discipline. Journal of Applied Econometrics, 19(2):247–265, March–April 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Costanigro:2009:ECS

[CMM09] Marco Costanigro, Ron C. Mittelhammer, and Jill J. Mc-Cluskey. Estimating class-specific parametric models under class uncertainty: local polynomial regression clustering in an hedonic analysis of wine markets. Journal of Applied Econometrics, 24(7):1117–1135, November–December 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cavalcanti:2015:CPV

[CMR15] Tiago V. De V. Cavalcanti, Kamiar Mohaddes, and Mehdi Raissi. Commodity price volatility and the sources of growth. Journal of Applied Econometrics, 30(6):857–873, September– October 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chatterjee:2021:EHC

[CMS21] Arpita Chatterjee, James Morley, and Aarti Singh. Estimating household consumption insurance. Journal of Applied Econometrics, 36(5):628–635, August 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cho:2024:DMP

[CMS24] Yunho Cho, James Morley, and Aarti Singh. Did marginal propensities to consume change with the housing boom and bust? Journal of Applied Econometrics, 39(1):174–199, January–February 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carriero:2018:UTS

[CMV18] Andrea Carriero, Sarah Mouabbi, and Elisabetta Vangelista. UK term structure decompositions at the zero lower bound. Journal of Applied Econometrics, 33(5):643–661, August 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cribari-Neto:1997:EPE

[CN97] Francisco Cribari-Neto. Econometric programming environments: GAUSS, Ox and S-Plus. Journal of Applied Econometrics, 12(1):77–89, January 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Coe:2004:LRM

[CN04] Patrick J. Coe and James M. Nason. Long-run monetary neutrality and long-horizon regressions. Journal of Applied Econometrics, 19(3):355–373, May–June 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cribari-Neto:1997:MEP

[CNJ97] Francisco Cribari-Neto and Mark J. Jensen. MATLAB as an econometric programming environment. Journal of Applied Econometrics, 12(6):735–744, November–December 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chintagunta:2009:MMM

[CNS09] Pradeep K. Chintagunta, Harikesh S. Nair, and R. Sukumar. Measuring marketing-mix effects in the 32/64 bit video-game console market. Journal of Applied Econometrics, 24(3):421– 445, April–May 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cross:2022:RPS

[CNT22] Jamie L. Cross, Bao H. Nguyen, and Trung Duc Tran. The role of precautionary and speculative demand in the global market for crude oil. Journal of Applied Econometrics, 37(5):882–895, August 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cribari-Neto:1999:RYA

[CNZ99] Francisco Cribari-Neto and Spyros G. Zarkos. R: yet another econometric programming environment. Journal of Applied Econometrics, 14(3):319–329, May–June 1999. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Corbae:1989:RWT

[CO89] Dean Corbae and Sam Ouliaris. A random walk through the Gibson paradox. Journal of Applied Econometrics, 4(3):295– 303, July–September 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cooley:1996:TSA

[CO96] Thomas F. Cooley and Masao Ogaki. A time series analysis of real wages, consumption and asset returns. Journal of Applied Econometrics, 11(2):119–134, March 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Coelli:1996:MTF

[Coe96] T. J. Coelli. Measurement of total factor productivity growth and biases in technological change in Western Australian agriculture. Journal of Applied Econometrics, 11(1):77–91, January 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Coleman:1992:BRC

[Col92] William Coleman. Book review: Conceptual anomalies in Economics and Statistics. Lessons from the Social Experiment, L. G. Neuberg, first edition, Cambridge University Press, New York, 1989, ISBN 0521-30444X, Hardcover, $54.40. Pp. xiv + 365. Journal of Applied Econometrics, 7(3):333–334, July– September 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cook:2009:RES

[Coo09] Steven Cook. A re-examination of the stationarity of inflation. Journal of Applied Econometrics, 24(6):1047–1053, September–October 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clarke:2019:DSB

[COQD19] Damian Clarke, Sonia Oreffice, and Climent Quintana-Domeque. The demand for season of birth. Journal of Applied Econometrics, 34(5):707–723, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Correa:2012:ICU

[Cor12] Juan A. Correa. Innovation and competition: An unstable relationship. Journal of Applied Econometrics, 27(1):160–166, January–February 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Coulombe:2024:MRF

[Cou24] Philippe Goulet Coulombe. The macroeconomy as a random forest. Journal of Applied Econometrics, 39(3):401–421, April– May 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cotterman:1992:CAA

[CP92] R. Cotterman and F. Peracchi. Classification and aggregation: An application to industrial classification in CPS data. Journal of Applied Econometrics, 7(1):31–51, January–March 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Culver:1997:TUR

[CP97] Sarah E. Culver and David H. Papell. Is there a unit root in the inflation rate? Evidence from sequential break and panel data models. Journal of Applied Econometrics, 12(4):435–444, July 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cho:2018:STP

[CP18] Jin Seo Cho and Peter C. B. Phillips. Sequentially testing polynomial model hypotheses using power transforms of regressors. Journal of Applied Econometrics, 33(1):141–159, January–February 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Corsi:2015:MAK

[CPA15] Fulvio Corsi, Stefano Peluso, and Francesco Audrino. Missing in asynchronicity: A Kalman-em approach for multivariate realized covariance estimation. Journal of Applied Econometrics, 30(3):377–397, April–May 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cockx:2019:MEG

[CPB19] Bart Cockx, Matteo Picchio, and Stijn Baert. Modeling the effects of grade retention in high school. Journal of Applied Econometrics, 34(3):403–424, April–May 2019. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Champagne:2020:IBC

[CPBS20] Julien Champagne, Guillaume Poulin-Bellisle, and Rodrigo Sekkel. Introducing the Bank of Canada staff economic projections database. Journal of Applied Econometrics, 35(1): 114–129, January–February 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Camacho:2002:WLI

[CPQ02] Maximo Camacho and Gabriel Perez-Quiros. This is what the leading indicators lead. Journal of Applied Econometrics, 17 (1):61–80, January–February 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Camacho:2010:IES

[CPQ10] Maximo Camacho and Gabriel Perez-Quiros. Introducing the euro-sting: Short-term indicator of euro area growth. Journal of Applied Econometrics, 25(4):663–694, June–July 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Camacho:2015:ENS

[CPQP15] Maximo Camacho, Gabriel Perez-Quiros, and Pilar Poncela. Extracting nonlinear signals from several economic indicators. Journal of Applied Econometrics, 30(7):1073–1089, November– December 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Centorrino:2024:BET

[CPUBUW24] Samuele Centorrino, María Pérez-Urdiales, Boris Bravo-Ureta, and Alan Wall. Binary endogenous treatment in stochastic frontier models with an application to soil conservation in El Salvador. Journal of Applied Econometrics, 39(3):365–382, April–May 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Campo:2003:AFP

[CPV03] Sandra Campo, Isabelle Perrigne, and Quang Vuong. Asymmetry in first-price auctions with affiliated private values. Journal of Applied Econometrics, 18(2):179–207, March–April 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chudik:2018:HPJ

[CPY18] Alexander Chudik, M. Hashem Pesaran, and Jui-Chung Yang. Half-panel jackknife fixed-effects estimation of linear panels with weakly exogenous regressors. Journal of Applied Econometrics, 33(6):816–836, September–October 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cornwell:1988:EEP

[CR88] Christopher Cornwell and Peter Rupert. Efficient estimation with panel data: An empirical comparison of instrumental variables estimators. Journal of Applied Econometrics, 3(2):149– 155, April 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Castagnetti:2013:ECB

[CR13] Carolina Castagnetti and Eduardo Rossi. Euro corporate bond risk factors. Journal of Applied Econometrics, 28(3):372–391, April–May 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clark:2015:MFP

[CR15] Todd E. Clark and Francesco Ravazzolo. Macroeconomic forecasting performance under alternative specifications of timevarying volatility. Journal of Applied Econometrics, 30(4): 551–575, June–July 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cygan-Rehm:2022:TNW

[CR22] Kamila Cygan-Rehm. Are there no wage returns to compulsory schooling in Germany? A reassessment. Journal of Applied Econometrics, 37(1):218–223, January–February 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cho:2024:RGV

[CR24] Dooyeon Cho and Seunghwa Rho. Reassessing growth vulnerability. Journal of Applied Econometrics, 39(1):225–234, January–February 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Crafts:1987:CS

[Cra87] N. F. R. Crafts. Cliometrics, 1971–1986: A survey. Journal of Applied Econometrics, 2(3):171–192, July 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Creel:2007:RFM

[Cre07] Michael Creel. I ran four million probits last night: HPC clustering with ParallelKnoppix. Journal of Applied Econometrics, 22(1):215–223, January–February 2007. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Conrad:1991:DDN

[CS91] K. Conrad and M. Schröder. Demand for durable and nondurable goods, environmental policy and consumer welfare. Journal of Applied Econometrics, 6(3):271–286, July– September 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clements:1999:MCS

[CS99] Michael P. Clements and Jeremy Smith. A Monte Carlo study of the forecasting performance of empirical SETAR models. Journal of Applied Econometrics, 14(2):123–141, March–April 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Choirat:2009:EP

[CS09] Christine Choirat and Raffello Seri. Econometrics with Python. Journal of Applied Econometrics, 24(4):698–704, June–July 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chang:2014:DGE

[CS14] Dongfeng Chang and Apostolos Serletis. The demand for gasoline: evidence from household survey data. Journal of Applied Econometrics, 29(2):291–313, March 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Canova:2015:EMP

[CS15]  Fabio Canova and Alain Schlaepfer.  Has the Euro–

Mediterranean partnership affected Mediterranean business cycles? Journal of Applied Econometrics, 30(2):241–262, March 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Canen:2021:CAU

[CS21a] Nathan Canen and Kyungchul Song. Counterfactual analysis under partial identification using locally robust refinement. Journal of Applied Econometrics, 36(4):416–436, June–July 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chou:2021:WTU

[CS21b] Cheng Chou and Ruoyao Shi. What time use surveys can (and cannot) tell us about labor supply. Journal of Applied Econometrics, 36(7):917–937, November–December 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Christiansen:2012:CLF

[CSS12] Charlotte Christiansen, Maik Schmeling, and Andreas Schrimpf. A comprehensive look at financial volatility prediction by economic variables. Journal of Applied Econometrics, 27(6):956– 977, September–October 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Canova:2014:IRA

[CSvD14] Fabio Canova, Frank Schorfheide, and Herman van Dijk. Introduction to recent advances in methods and applications for DSGE models. Journal of Applied Econometrics, 29(7):1029– 1030, November–December 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cameron:1986:EMB

[CT86a] A. Colin Cameron and Pravin K. Trivedi. Econometric models based on count data. Comparisons and applications of some estimators and tests. Journal of Applied Econometrics, 1(1):29– 53, January 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cuthbertson:1986:MAD

[CT86b] Keith Cuthbertson and Mark P. Taylor. Monetary anticipation and the demand for money in the U.K.: Testing rationality in the shock-absorber hypothesis. Journal of Applied Econometrics, 1(4):355–365, October 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cao:1992:NTS

[CT92] C. Q. Cao and R. S. Tsay. Nonlinear time-series analysis of stock volatilities. Journal of Applied Econometrics, 7(S1): S165–S185, December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Conley:2002:SED

[CT02] Timothy G. Conley and Giorgio Topa. Socio-economic distance and spatial patterns in unemployment. Journal of Applied Econometrics, 17(4):303–327, July–August 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Clements:2003:EIF

[CT03a] Michael P. Clements and Nick Taylor. Evaluating interval forecasts of high-frequency financial data. Journal of Applied Econometrics, 18(4):445–456, July–August 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Conley:2003:ILI

[CT03b] Timothy G. Conley and Giorgio Topa. Identification of local interaction models with imperfect location data. Journal of Applied Econometrics, 18(5):605–618, September–October 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carro:2014:SDH

[CT14] Jesus M. Carro and Alejandra Traferri. State dependence and heterogeneity in health using a bias–corrected fixed–effects estimator. Journal of Applied Econometrics, 29(2):181–207, March 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chan:2015:PPC

[CT15] Joshua C. C. Chan and Justin L. Tobias. Priors and posterior computation in linear endogenous variable models with imperfect instruments. Journal of Applied Econometrics, 30 (4):650–674, June–July 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chua:2019:IFS

[CT19] Chew Lian Chua and Sarantis Tsiaplias. Information flows and stock market volatility. Journal of Applied Econometrics, 34(1): 129–148, January–February 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Caliendo:2013:BDU

[CTU13] Marco Caliendo, Konstantinos Tatsiramos, and Arne Uhlendorff. Benefit duration, unemployment duration and job match quality: a regression–discontinuity approach. Journal of Applied Econometrics, 28(4):604–627, June–July 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cuaresma:2011:HDA

[Cua11] Jesus Crespo Cuaresma. How different is Africa? A comment on Masanjala and Papageorgiou. Journal of Applied Econometrics, 26(6):1041–1047, September–October 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [MP08] and comment [Pap11].

Cubadda:1999:CCS

[Cub99] Gianluca Cubadda. Common cycles in seasonal non-stationary time series. Journal of Applied Econometrics, 14(3):273–291, May–June 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cuthbertson:1986:BUK

[Cut86] Keith Cuthbertson. The behaviour of U.K. export prices of manufactured goods 1970–1983. Journal of Applied Econometrics, 1(3):255–275, June 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Coenen:2001:DME

[CV01] G. Coenen and J.-L. Vega. The demand for M3 in the euro area. Journal of Applied Econometrics, 16(6):727–748, November– December 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chiriac:2011:MFM

[CV11] Roxana Chiriac and Valeri Voev. Modelling and forecasting multivariate realized volatility. Journal of Applied Econometrics, 26(6):922–947, September–October 2011. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cziraky:2004:LPS

[Czi04] Dario Cziráky. LISREL 8.54: A program for structural equation modelling with latent variables. Journal of Applied Econometrics, 19(1):135–141, January–February 2004. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dahlhaus:2017:CMP

[Dah17] Tatjana Dahlhaus. Conventional monetary policy transmission during financial crises: An empirical analysis. Journal of Applied Econometrics, 32(2):401–421, March 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Davis:1987:SFCa

[Dav87a] E. P. Davis. A stock-flow consistent macro-econometric model of the UK economy — Part I. Journal of Applied Econometrics, 2(2):111–132, April 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Davis:1987:SFCb

[Dav87b] E. P. Davis. A stock-flow consistent macro-econometric model of the UK economy — Part II. Journal of Applied Econometrics, 2(4):259–307, October 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Davidson:1995:BRE

[Dav95] James Davidson. Book review: Estimation, inference and specification analysis, by H. White. Econometric Society Monographs No. 22, Cambridge University Press, 1994, pp. x + 380. ISBN 0-521-25280-6; Price £30 (hardback). Journal of Applied Econometrics, 10(4):477–479, October–December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Davig:2005:PED

[Dav05] Troy Davig. Periodically expanding discounted debt: a threat to fiscal policy sustainability? Journal of Applied Econometrics, 20(7):829–840, December 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Day:1992:CED

[Day92] R. H. Day. Complex economic dynamics: Obvious in history, generic in theory, elusive in data. Journal of Applied Econometrics, 7(S1):S9–S23, December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Demery:2000:IIT

[DD00] David Demery and Nigel W. Duck. Incomplete information and the time series behaviour of consumption. Journal of Applied Econometrics, 15(4):355–366, July–August 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dias:2019:MPH

[DD19] Daniel A. Dias and João B. Duarte. Monetary policy, housing rents, and inflation dynamics. Journal of Applied Econometrics, 34(5):673–687, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DeGroote:2021:TSH

[DD21] Olivier De Groote and Koen Declercq. Tracking and specialization of high schools: Heterogeneous effects of school choice. Journal of Applied Econometrics, 36(7):898–916, November– December 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Despois:2023:IIF

[DD23] Thomas Despois and Catherine Doz. Identifying and interpreting the factors in factor models via sparsity: Different approaches. Journal of Applied Econometrics, 38(4):533–555, June–July 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Desjardins:2023:FC

[DDL23a] Denise Desjardins, Georges Dionne, and Yang Lu. Featured cover. Journal of Applied Econometrics, 38(2):??, March 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Desjardins:2023:HRE

[DDL23b] Denise Desjardins, Georges Dionne, and Yang Lu. Hierarchical random-effects model for the insurance pricing of vehicles belonging to a fleet. Journal of Applied Econometrics, 38(2): 242–259, March 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Demirer:2018:EGB

[DDLY18] Mert Demirer, Francis X. Diebold, Laura Liu, and Kamil Yilmaz. Estimating global bank network connectedness. Journal of Applied Econometrics, 33(1):1–15, January–February 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dees:2007:EIL

[DdMPS07] Stephane Dees, Filippo di Mauro, M. Hashem Pesaran, and L. Vanessa Smith. Exploring the international linkages of the euro area: a global VAR analysis. Journal of Applied Econometrics, 22(1):1–38, January–February 2007. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DeArcangelis:1992:BRS

[De 92] Giuseppe De Arcangelis. Book review: Statistical analysis and forecasting of economic structural change. P. Hack (ed.) Springer-Verlag, Berlin, Heidelberg, New York, London, Paris, Tokyo, Hong Kong, 1989, pp. xix + 488. ISBN 0-387-51454-6; hardbound. Journal of Applied Econometrics, 7(4):435–436, October–December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DeArcangelis:1993:BRS

[De 93] Giuseppe De Arcangelis. Book review: Structural Equations with Latent Variables, Kenneth A. Bollen. John Wiley and Sons, Inc., New York, Chichester, Brisbane, Toronto, Singapore, 1989, ISBN 0-471-01171-1, £39.90 hardbound, pp. xiv + 514. Journal of Applied Econometrics, 8(1):111–113, January– March 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DeJong:1994:UAE

[De 94] F. De Jong. A univariate analysis of EMS exchange rates using a target zone model. Journal of Applied Econometrics, 9(1):31– 45, January–March 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Douch:2022:BTE

[DE22] Mustapha Douch and Terence Huw Edwards. The bilateral trade effects of announcement shocks: Brexit as a natural field experiment. Journal of Applied Econometrics, 37(2):305–329, March 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dejong:1993:BIL

[Dej93] D. N. Dejong. Bayesian inference in limited dependent variable models. An application to measuring strike duration. Journal of Applied Econometrics, 8(2):115–128, April–June 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Demetriades:1989:RBL

[Dem89] Panikos O. Demetriades. The relationship between the level and variability of inflation: Theory and evidence. Journal of Applied Econometrics, 4(3):239–250, July–September 1989.

CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dennis:2006:PPU

[Den06] Richard Dennis. The policy preferences of the US Federal reserve. Journal of Applied Econometrics, 21(1):55–77, January– February 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Depalo:2018:IIP

[Dep18] Domenico Depalo. Identification issues in the public/private wage gap, with an application to Italy. Journal of Applied Econometrics, 33(3):435–456, April–May 2018. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Deschamps:1993:JTR

[Des93] P. J. Deschamps. Joint tests for regularity and autocorrelation in allocation systems. Journal of Applied Econometrics, 8(2): 195–211, April–June 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Deschamps:2003:TVI

[Des03] Philippe J. Deschamps. Time-varying intercepts and equilibrium analysis: an extension of the dynamic almost ideal demand model. Journal of Applied Econometrics, 18(2):209–236, March–April 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Deschamps:2008:CST

[Des08] Philippe J. Deschamps. Comparing smooth transition and Markov switching autoregressive models of US unemployment. Journal of Applied Econometrics, 23(4):435–462, June 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Devereux:2007:SSB

[Dev07] Paul J. Devereux. Small-sample bias in synthetic cohort models of labor supply. Journal of Applied Econometrics, 22(4): 839–848, June–July 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Davis:1997:FSU

[DF97] E. Philip Davis and Gabriel Fagan. Are financial spreads useful indicators of future inflation and output growth in EU countries? Journal of Applied Econometrics, 12(6):701–714, November–December 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dardanoni:2012:SMI

[DFF12] Valentino Dardanoni, Mario Fiorini, and Antonio Forcina. Stochastic monotonicity in intergenerational mobility tables. Journal of Applied Econometrics, 27(1):85–107, January– February 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Duval:2021:RPE

[DFM21] Romain Duval, Davide Furceri, and Jakob Miethe. Robust political economy correlates of major product and labor market reforms in advanced economies: Evidence from BAMLE for logit models. Journal of Applied Econometrics, 36(1):98–124, January–February 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dechert:1992:LEN

[DG92] W. D. Dechert and R. Gencay. Lyapunov exponents as a nonparametric diagnostic for stability analysis. Journal of Applied Econometrics, 7(S1):S41–S60, December 1992. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Drukker:2003:RRE

[DG03] David M. Drukker and Weihua Guan. “Replicating the results in ‘On efficient estimation with panel data: an empirical comparison of instrumental variables estimators’ ”: by Baltagi, B. H. and Khanti-Akom, S. (1990), Journal of Applied Econometrics, Vol 5, pages 401–406. Journal of Applied Econometrics, 18(1):119, January–February 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [BKA90].

DiezDeLosRios:2011:AVN

[DG11] Antonio Diez De Los Rios and René Garcia. Assessing and valuing the nonlinear structure of hedge fund returns. Journal of Applied Econometrics, 26(2):193–212, March 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DAgostino:2013:MFS

[DGG13] Antonello D’Agostino, Luca Gambetti, and Domenico Giannone. Macroeconomic forecasting and structural change. Journal of Applied Econometrics, 28(1):82–101, January–February 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dolado:2003:LRD

[DGM03] Juan J. Dolado, Jesús Gonzalo, and Laura Mayoral. Longrange dependence in Spanish political opinion poll series. Journal of Applied Econometrics, 18(2):137–155, March–April 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Delavande:2011:EPE

[DGM11] Adeline Delavande, Xavier Giné, and David McKenzie. Eliciting probabilistic expectations with visual aids in developing countries: how sensitive are answers to variations in elicitation design? Journal of Applied Econometrics, 26(3):479–497, April–May 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Domowitz:1990:IEC

[DH90] Ian Domowitz and Craig S. Hakkio. Interpreting an error correction model: Partial adjustment, forward-looking behaviour, and dynamic international money demand. Journal of Applied Econometrics, 5(1):29–46, January–March 1990. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Douglas:2010:WGP

[DH10] Christopher Douglas and Ana María Herrera. Why are gasoline prices sticky? A test of alternative models of price adjustment. Journal of Applied Econometrics, 25(6):903–928, September– October 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Demetrescu:2016:DTN

[DH16] Matei Demetrescu and Ulrich Homm. Directed tests of no crosssectional correlation in large-N panel data models. Journal of Applied Econometrics, 31(1):4–31, January–February 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Demetrescu:2022:RIU

[DHKB22] Matei Demetrescu, Christoph Hanck, and Robinson Kruse-Becher. Robust inference under time-varying volatility: a realtime evaluation of professional forecasters. Journal of Applied Econometrics, 37(5):1010–1030, August 2022. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dagsvik:2011:ERS

[DHR11] John K. Dagsvik, Torbjørn Hægeland, and Arvid Raknerud. Estimating the returns to schooling: a likelihood approach based on normal mixtures. Journal of Applied Econometrics, 26(4):613–640, June–July 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

deHaan:2022:IPL

[dHW22] Jakob de Haan and Rasmus Wiese. The impact of product and labour market reform on growth: Evidence for OECD countries based on local projections. Journal of Applied Econometrics, 37(4):746–770, June–July 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Diebold:1989:STS

[Die89] Francis X. Diebold. Structural time series analysis and modelling package: A review. Journal of Applied Econometrics, 4 (2):195–204, April–June 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dewachter:2014:IYC

[DIL14] Hans Dewachter, Leonardo Iania, and Marco Lyrio. Information in the yield curve: a macro–finance approach. Journal of Applied Econometrics, 29(1):42–64, January–February 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dubois:2008:ISI

[DIM08] Pierre Dubois, Marc Ivaldi, and Thierry Magnac. Introduction to the special issue on the econometrics of auctions. Journal of Applied Econometrics, 23(7):867–869, November–December 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dineen:1991:PEM

[Din91] Chris Dineen. Progressive econometric modelling (PERM): A review. Journal of Applied Econometrics, 6(2):213–217, April– June 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DeJong:2000:KIV

[DIW00] David N. DeJong, Beth F. Ingram, and Charles H. Whiteman. Keynesian impulses versus Solow residuals: identifying sources of business cycle fluctuations. Journal of Applied Econometrics, 15(3):311–329, May–June 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Durlauf:1995:MRC

[DJ95] Steven N. Durlauf and Paul A. Johnson. Multiple regimes and cross-country growth behaviour. Journal of Applied Econometrics, 10(4):365–384, October–December 1995. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dahlberg:2000:EDB

[DJ00] Matz Dahlberg and Eva Johansson. An examination of the dynamic behaviour of local governments using GMM bootstrapping methods. Journal of Applied Econometrics, 15(4):401– 416, July–August 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dagsvik:2016:LSC

[DJ16] John K. Dagsvik and Zhiyang Jia. Labor supply as a choice among latent jobs: Unobserved heterogeneity and identification. Journal of Applied Econometrics, 31(3):487–506, April– May 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dynarski:2018:HIF

[DJK18] Susan Dynarski, Brian Jacob, and Daniel Kreisman. How important are fixed effects and time trends in estimating returns to schooling? Evidence from a replication of Jacobson, Lalonde, and Sullivan, 2005. Journal of Applied Econometrics, 33(7):1098–1108, November–December 2018. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Djogbenou:2020:CRA

[Djo20] Antoine A. Djogbenou. Comovements in the real activity of developed and emerging economies: A test of global versus specific international factors. Journal of Applied Econometrics, 35 (3):344–370, April–May 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Djogbenou:2024:IOP

[Djo24] Antoine A. Djogbenou. Identifying oil price shocks with global, developed, and emerging latent real economy activity factors. Journal of Applied Econometrics, 39(1):128–149, January–February 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Diebold:2001:MPT

[DK01] Francis X. Diebold and Lutz Kilian. Measuring predictability: theory and macroeconomic applications. Journal of Applied Econometrics, 16(6):657–669, November–December 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Davis:2011:EEG

[DK11] Lucas W. Davis and Lutz Kilian. Estimating the effect of a gasoline tax on carbon emissions. Journal of Applied Econometrics, 26(7):1187–1214, November–December 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Das:2012:PED

[DK12] Abhiman Das and Subal C. Kumbhakar. Productivity and efficiency dynamics in Indian banking: An input distance function approach incorporating quality of inputs and outputs. Journal of Applied Econometrics, 27(2):205–234, March 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Drakos:2013:IDM

[DK13] Konstantinos Drakos and Panagiotis Th. Konstantinou. Investment decisions in manufacturing: assessing the effects of real oil prices and their uncertainty. Journal of Applied Econometrics, 28(1):151–165, January–February 2013. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dimitrakopoulos:2016:SDS

[DK16] Stefanos Dimitrakopoulos and Michalis Kolossiatis. State dependence and stickiness of sovereign credit ratings: Evidence from a panel of countries. Journal of Applied Econometrics, 31 (6):1065–1082, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dong:2023:WCW

[DK23] Yingying Dong and Michal Kolesár. When can we ignore measurement error in the running variable? Journal of Applied Econometrics, 38(5):735–750, August 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dufour:2010:MRB

[DKB10] Jean-Marie Dufour, Lynda Khalaf, and Marie-Claude Beaulieu. Multivariate residual-based finite-sample tests for serial dependence and ARCH effects with applications to asset pricing models. Journal of Applied Econometrics, 25(2):263–285, March 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DelBono:2022:IDL

[DKP22] Emilia Del Bono, Josh Kinsler, and Ronni Pavan. Identification of dynamic latent factor models of skill formation with translog production. Journal of Applied Econometrics, 37(6):1256–1265, September–October 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Durlauf:2012:GDR

[DKT12] Steven N. Durlauf, Andros Kourtellos, and Chih Ming Tan. Is God in the details? A reexamination of the role of religion in economic growth. Journal of Applied Econometrics, 27 (7):1059–1075, November–December 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Deaton:1995:ENR

[DL95] Angus Deaton and Guy Laroque. Estimating a nonlinear rational expectations commodity price model with unobservable state variables. Journal of Applied Econometrics, 10(S1):S9– S40, December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DeNicolo:2017:FTR

[DL17] Gianni De Nicolò and Marcella Lucchetta. Forecasting tail risks. Journal of Applied Econometrics, 32(1):159–170, January–February 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Deng:2022:ELF

[DL22] Zichen Deng and Maarten Lindeboom. Early-life famine exposure, hunger recall, and later-life health. Journal of Applied Econometrics, 37(4):771–787, June–July 2022. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dewachter:2006:JMT

[DLM06] Hans Dewachter, Marco Lyrio, and Konstantijn Maes. A joint model for the term structure of interest rates and the macroeconomy. Journal of Applied Econometrics, 21(4):439– 462, May–June 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DInnocenzo:2024:HDN

[DLOZ24] Enzo D’Innocenzo, André Lucas, Anne Opschoor, and Xingmin Zhang. Heterogeneity and dynamics in network models. Journal of Applied Econometrics, 39(1):150–173, January–February 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DenVanBerg:1994:ALP

[DLR94] G. J. Den Van Berg, M. Lindeboom, and G. Ridder. Attrition in longitudinal panel data and the empirical analysis of dynamic labour market behaviour. Journal of Applied Econometrics, 9(4):421–435, October–December 1994. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DeJong:2006:TSC

[DLR06] David N. DeJong, Roman Liesenfeld, and Jean-François Richard. Timing structural change: a conditional probabilistic approach. Journal of Applied Econometrics, 21(2):175–190, March 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dufays:2021:SCP

[DLRS21] Arnaud Dufays, Zhuo Li, Jeroen V. K. Rombouts, and Yong Song. Sparse change-point VAR models. Journal of Applied Econometrics, 36(6):703–727, September–October 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Delgado:1998:TNN

[DM98] Miguel A. Delgado and Juan Mora. Testing non-nested semiparametric models: an application to Engel curves specification. Journal of Applied Econometrics, 13(2):145–162, March– April 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Durlauf:2003:SIE

[DM03] Steven N. Durlauf and Robert A. Moffitt. Special issue on empirical analysis of social interactions. Journal of Applied Econometrics, 18(5):499, September–October 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Davidson:2006:CAJ

[DM06a] Russell Davidson and James G. MacKinnon. The case against JIVE. Journal of Applied Econometrics, 21(6):827–833, September–October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [AD06, BD06] and reply [DM06b].

Davidson:2006:RAD

[DM06b] Russell Davidson and James G. MacKinnon. Reply to Ackerberg and Devereux and Blomquist and Dahlberg on ‘The case against JIVE’. Journal of Applied Econometrics, 21(6):843– 844, September–October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dungey:2007:UFM

[DM07] Mardi Dungey and Vance L. Martin. Unravelling financial market linkages during crises. Journal of Applied Econometrics, 22 (1):89–119, January–February 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dominitz:2011:MIE

[DM11] Jeff Dominitz and Charles F. Manski. Measuring and interpreting expectations of equity returns. Journal of Applied Econometrics, 26(3):352–370, April–May 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dobbelaere:2013:PDE

[DM13] Sabien Dobbelaere and Jacques Mairesse. Panel data estimates of the production function and product and labor market imperfections. Journal of Applied Econometrics, 28(1):1–46, January–February 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Donni:2018:CDH

[DM18] Olivier Donni and Eleonora Matteazzi. Collective decisions, household production, and labor force participation. Journal of Applied Econometrics, 33(7):1064–1080, November–December 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dovern:2020:OIT

[DM20] Jonas Dovern and Hans Manner. Order-invariant tests for proper calibration of multivariate density forecasts. Journal of Applied Econometrics, 35(4):440–456, June–July 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Das:2005:WWE

[DMM05] Sanghamitra Das, Charles F. Manski, and Mark D. Manuszak. Walk or wait? An empirical analysis of street crossing decisions. Journal of Applied Econometrics, 20(4):529–548, May– June 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dungey:2000:MLF

[DMP00] Mardi Dungey, Vance L. Martin, and Adrian R. Pagan. A multivariate latent factor decomposition of international bond yield spreads. Journal of Applied Econometrics, 15(6):697–715, November–December 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Deij:2021:WIG

[DMP21] Sabine Deij, Jakob B. Madsen, and Laura Puzzello. When are instruments generated from geographic characteristics in bilateral relationships invalid? Journal of Applied Econometrics, 36 (4):437–452, June–July 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Deb:2006:BAT

[DMT06] Partha Deb, Murat K. Munkin, and Pravin K. Trivedi. Bayesian analysis of the two-part model with endogeneity: application to health care expenditure. Journal of Applied Econometrics, 21(7):1081–1099, November 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DeBruin:2011:MCU

[DMTvdK11] Wändi Bruine De Bruin, Charles F. Manski, Giorgio Topa, and Wilbert van der Klaauw. Measuring consumer uncertainty about future inflation. Journal of Applied Econometrics, 26(3): 454–478, April–May 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Duernecker:2022:TOG

[DMVR22] Georg Duernecker, Moritz Meyer, and Fernando Vega-Redondo. Trade openness and growth: a network-based approach. Journal of Applied Econometrics, 37(6):1182–1203, September–October 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Davidson:1994:MUG

[DMW94] J. Davidson, G. Madonia, and P. Westaway. Modelling the UK gilt-edged market. Journal of Applied Econometrics, 9(3):231– 253, July–September 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Diebold:1989:DER

[DN89] Francis X. Diebold and Marc Nerlove. The dynamics of exchange rate volatility: A multivariate latent factor ARCH model. Journal of Applied Econometrics, 4(1):1–21, January– March 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Doser:2023:IEN

[DNRS23] Alexander Doser, Ricardo Nunes, Nikhil Rao, and Viacheslav Sheremirov. Inflation expectations and nonlinearities in the Phillips curve. Journal of Applied Econometrics, 38(4):453– 471, June–July 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dungey:2014:MLO

[DO14] Mardi Dungey and Denise R. Osborn. Modelling large open economies with international linkages: the USA and euro area. Journal of Applied Econometrics, 29(3):377–393, April–May 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Doh:2013:LRR

[Doh13] Taeyoung Doh. Long-run risks in the term structure of interest rates: estimation. Journal of Applied Econometrics, 28(3):478– 497, April–May 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dong:2015:RDA

[Don15] Yingying Dong. Regression discontinuity applications with rounding errors in the running variable. Journal of Applied Econometrics, 30(3):422–446, April–May 2015. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dovonon:2013:CHF

[Dov13] Prosper Dovonon. Conditionally heteroskedastic factor models with skewness and leverage effects. Journal of Applied Econometrics, 28(7):1110–1137, November–December 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DeHaan:2011:EIS

[DP11] Monique De Haan and Erik Plug. Estimating intergenerational schooling mobility on censored samples: consequences and remedies. Journal of Applied Econometrics, 26(1):151–166, January–February 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DeLuca:2012:EEC

[DP12] Giuseppe De Luca and Franco Peracchi. Estimating Engel curves under unit and item nonresponse. Journal of Applied Econometrics, 27(7):1076–1099, November–December 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Delgado:2013:EEE

[DP13] Michael S. Delgado and Christopher F. Parmeter. Embarrassingly easy embarrassingly parallel processing in R. Journal of Applied Econometrics, 28(7):1224–1230, November–December 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Donkers:2006:DTS

[DPJF06] Bas Donkers, Richard Paap, Jedid-Jah Jonker, and Philip Hans Franses. Deriving target selection rules from endogenously selected samples. Journal of Applied Econometrics, 21(5):549– 562, July–August 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Donald:2006:EMM

[DPR06] Stephen G. Donald, Harry J. Paarsch, and Jacques Robert. An empirical model of the multi-unit, sequential, clock auction. Journal of Applied Econometrics, 21(8):1221–1247, December 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DeCrombrugghe:1997:SDF

[DPU97] Denis De Crombrugghe, Franz C. Palm, and Jean-Pierre Urbain. Statistical demand functions for food in the USA and the Netherlands. Journal of Applied Econometrics, 12(5):615–645, September–October 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Danielsson:1993:AGI

[DR93] J. Danielsson and J.-F. Richard. Accelerated Gaussian importance sampler with application to dynamic latent variable models. Journal of Applied Econometrics, 8(S1):S153–S173, December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Delavande:2011:IUA

[DR11] Adeline Delavande and Susann Rohwedder. Individuals’ uncertainty about future social security benefits and portfolio choice.

Journal of Applied Econometrics, 26(3):498–519, April–May 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DiPorto:2013:TLR

[DR13] Edoardo Di Porto and Federico Revelli. Tax–limited reaction functions. Journal of Applied Econometrics, 28(5):823–839, August 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dungey:2018:IC

[DR18] Mardi Dungey and Eric Renault. Identifying contagion. Journal of Applied Econometrics, 33(2):227–250, March 2018. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

deRoos:2010:DMU

[dRS10] Nicolas de Roos and Yianis Sarafidis. Decision making under risk in deal or no deal. Journal of Applied Econometrics, 25(6): 987–1027, September–October 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Demetrescu:2021:RPE

[DRT21] Matei Demetrescu, Christoph Roling, and Anna Titova. Reevaluating the prudence of economic forecasts in the EU: The role of instrument persistence. Journal of Applied Econometrics, 36(1):151–161, January–February 2021. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Delgado:1990:KR

[DS90] Miguel A. Delgado and Thanasis Stengos. N-kernel: A review. Journal of Applied Econometrics, 5(3):299–304, July– September 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Denton:1991:RTS

[DS91] F. T. Denton and B. G. Spencer. Resampling a time-series process: A method of estimating the probabilities associated with alternative plans for protecting pensions against inflation. Journal of Applied Econometrics, 6(3):303–314, July– September 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Das:2004:PPR

[DS04] Sanghamitra Das and Ramprasad Sengupta. Projection pursuit regression and disaggregate productivity effects: the case of the Indian blast furnaces. Journal of Applied Econometrics, 19(3): 397–418, May–June 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dagsvik:2006:SLS

[DS06a] John K. Dagsvik and Steinar StrØm. Sectoral labour supply, choice restrictions and functional form. Journal of Applied Econometrics, 21(6):803–826, September–October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dominitz:2006:IEB

[DS06b] Jeff Dominitz and Robert P. Sherman. Identification and estimation of bounds on school performance measures: a nonparametric analysis of a mixture model with verification. Journal of Applied Econometrics, 21(8):1295–1326, December 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dhaene:2012:STM

[DS12] Geert Dhaene and J. M. C. Santos Silva. Specification and testing of models estimated by quadrature. Journal of Applied Econometrics, 27(2):322–332, March 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DePaula:2014:HBA

[DST14] Áureo De Paula, Gil Shapira, and Petra E. Todd. How beliefs about HIV status affect risky behaviors: evidence from Malawi. Journal of Applied Econometrics, 29(6):944–964, September– October 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dutta:2001:IDI

[DSW01] Jayasri Dutta, J. A. Sefton, and M. R. Weale. Income distribution and income dynamics in the United Kingdom. Journal of Applied Econometrics, 16(5):599–617, September–October 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Davies:1994:RS

[DT94] N. Davies and A. R. Tremayne. Review of statgraphics. Journal of Applied Econometrics, 9(3):335–341, July–September 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Deb:1997:DMC

[DT97] Partha Deb and Pravin K. Trivedi. Demand for medical care by the elderly: a finite mixture approach. Journal of Applied Econometrics, 12(3):313–336, May 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dunn:2020:RLP

[DT20] Richard A. Dunn and Nathan W. Tefft. Replicating the Levitt and Porter estimates of drunk driving. Journal of Applied Econometrics, 35(6):786–796, September–October 2020. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Deb:1996:ECM

[DTV96] Partha Deb, Pravin K. Trivedi, and Panayotis Varangis. The excess co-movement of commodity prices reconsidered. Journal of Applied Econometrics, 11(3):275–291, May 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DeLaCroix:1998:ISI

[DU98] David De La Croix and Jean-Pierre Urbain. Intertemporal substitution in import demand and habit formation. Journal of Applied Econometrics, 13(6):589–612, November–December 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dupuy:2021:MCW

[Dup21] Arnaud Dupuy. Migration in China: To work or to wed? Journal of Applied Econometrics, 36(4):393–415, June–July 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dionne:1992:AIR

[DV92] G. Dionne and C. Vanasse. Automobile insurance ratemaking in the presence of asymmetrical information. Journal of Applied Econometrics, 7(2):149–165, April–June 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Durlauf:2010:IMU

[DV10] Steven N. Durlauf and Shaun P. Vahey. Introduction: ‘model uncertainty and macroeconomics’. Journal of Applied Econometrics, 25(1):1–3, January–February 2010. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Deconinck:2013:NSR

[DV13] Koen Deconinck and Marijke Verpoorten. Narrow and scientific replication of ‘The slave trade and the origins of mistrust in Africa’. Journal of Applied Econometrics, 28(1):166–169, January–February 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Desai:1988:KME

[DW88] Meghnad Desai and Guglielmo Weber. A Keynesian macroeconometric model of the UK: 1955–1984. Journal of Applied Econometrics, 3(1):1–33, January 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dejong:1991:CTS

[DW91] David N. Dejong and Charles H. Whiteman. The case for trend-stationarity is stronger than we thought. Journal of Applied Econometrics, 6(4):413–421, October–December 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Diebold:1996:IEF

[DW96] Francis X. Diebold and Mark W. Watson. Introduction: Econometric forecasting. Journal of Applied Econometrics, 11(5): 453–454, September 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Doppelhofer:2009:JGDa

[DW09a] Gernot Doppelhofer and Melvyn Weeks. Jointness of growth determinants. Journal of Applied Econometrics, 24(2):209– 244, March 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See comments [Str09, LS09a] and reply [DW09b].

Doppelhofer:2009:JGDb

[DW09b] Gernot Doppelhofer and Melvyn Weeks. ‘Jointness of growth determinants’: Reply to comments by Rodney Strachan, Eduardo Ley and Mark F. J. Steel. Journal of Applied Econometrics, 24(2):252–256, March 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [DW09a, Str09, LS09a].

DeGraeve:2015:RSF

[DW15] Ferre De Graeve and Karl Walentin. Refining stylized facts from factor models of inflation. Journal of Applied Econometrics, 30(7):1192–1209, November–December 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Diks:2016:NGC

[DW16] Cees Diks and Marcin Wolski. Nonlinear Granger causality: Guidelines for multivariate analysis. Journal of Applied Econometrics, 31(7):1333–1351, November–December 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Droumaguet:2017:GCR

[DWW17] Matthieu Droumaguet, Anders Warne, and Tomasz Woźniak. Granger causality and regime inference in Markov switching VAR models with Bayesian methods. Journal of Applied Econometrics, 32(4):802–818, June–July 2017. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

DeSantis:2018:SAS

[DZ18] Roberto A. De Santis and Srecko Zimic. Spillovers among sovereign debt markets: Identification through absolute magnitude restrictions. Journal of Applied Econometrics, 33(5): 727–747, August 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ericsson:1988:SRR

[EB88] Neil R. Ericsson and K. N. Berk. Softwear reviews: A review of data-fit: An interactive econometric modelling package for IBM-compatible PCs: 1. Introduction and methodological overview. Journal of Applied Econometrics, 3(4):319–332, October–December 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eddelbuttel:1996:SRO

[Edd96] Dirk Eddelbüttel. Software review: Object-oriented econometrics: Matrix programming in C++ using GCC and newmat. Journal of Applied Econometrics, 11(2):199–209, March 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eddelbuttel:2000:EO

[Edd00] Dirk Eddelbüttel. Econometrics with Octave. Journal of Applied Econometrics, 15(5):531–542, September–October 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Edwards:2021:DPP

[Edw21] Jeremy Edwards. Did Protestantism promote prosperity via higher human capital? Replicating the Becker–Woessmann (2009) results. Journal of Applied Econometrics, 36(6):853– 858, September–October 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ermisch:2013:EPE

[EF13] John Ermisch and Marco Francesconi. The effect of parental employment on child schooling. Journal of Applied Econometrics, 28(5):796–822, August 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ehrmann:2011:SBM

[EFR11] Michael Ehrmann, Marcel Fratzscher, and Roberto Rigobon. Stocks, bonds, money markets and exchange rates: measuring international financial transmission. Journal of Applied Econometrics, 26(6):948–974, September–October 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Elder:2015:IRI

[EGH15] Todd E. Elder, John H. Goddeeris, and Steven J. Haider. Isolating the roles of individual covariates in reweighting estimation. Journal of Applied Econometrics, 30(7):1169–1191, November– December 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

El-Gamal:2005:IHT

[EGI05] Mahmoud A. El-Gamal and Hulusi Inanoglu. Inefficiency and heterogeneity in Turkish banking: 1990–2000. Journal of Applied Econometrics, 20(5):641–664, July–August 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eckstein:2011:JWM

[EGP11] Zvi Eckstein, Suqin Ge, and Barbara Petrongolo. Job and wage mobility with minimum wages and imperfect compliance. Journal of Applied Econometrics, 26(4):580–612, June–July 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Engsted:1994:LQA

[EH94] Tom Engsted and Niels Haldrup. The linear quadratic adjustment cost model and the demand for labour. Journal of Applied Econometrics, 9(S1):S145–S159, December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Engsted:1999:ELM

[EH99] Tom Engsted and Niels Haldrup. Estimating the LQAC model with I(2) variables. Journal of Applied Econometrics, 14(2): 155–170, March–April 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eicher:2012:TCD

[EHP12] Theo S. Eicher, Christian Henn, and Chris Papageorgiou. Trade creation and diversion revisited: Accounting for model uncertainty and natural trading partner effects. Journal of Applied Econometrics, 27(2):296–321, March 2012. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Engers:2009:AMD

[EHS09] Maxim Engers, Monica Hartmann, and Steven Stern. Annual miles drive used car prices. Journal of Applied Econometrics, 24(1):1–33, January–February 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Elhorst:2017:TDM

[EHSJ17] J. Paul Elhorst, Pim Heijnen, Anna Samarina, and Jan P. A. M. Jacobs. Transitions at different moments in time: A spatial probit approach. Journal of Applied Econometrics, 32(2):422– 439, March 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eggink:1994:SAL

[EHV94] E. Eggink, J. P. Hop, and B. M. S. Van Praag. A symmetric approach to the labour market with the household as the unit of observation. Journal of Applied Econometrics, 9(2):133–161, April–June 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eickmeier:2009:CHE

[Eic09] Sandra Eickmeier. Comovements and heterogeneity in the euro area analyzed in a non-stationary dynamic factor model. Journal of Applied Econometrics, 24(6):933–959, September– October 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eisenhauer:2019:ASF

[Eis19] Philipp Eisenhauer. The approximate solution of finite-horizon discrete-choice dynamic programming models. Journal of Applied Econometrics, 34(1):149–154, January–February 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ericsson:2001:OIL

[EIT01] Neil R. Ericsson, John S. Irons, and Ralph W. Tryon. Output and inflation in the long run. Journal of Applied Econometrics, 16(3):241–253, May–June 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ejrnaes:2020:FPL

[EJ20] Mette Ejrnæs and Thomas H. Jørgensen. Family planning in a life-cycle model with income risk. Journal of Applied Econometrics, 35(5):567–586, August 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Edison:1987:QRP

[EK87] Hali J. Edison and Jan Tore Klovland. A quantitative reassessment of the purchasing power parity hypothesis: Evidence from Norway and the United Kingdom. Journal of Applied Econometrics, 2(4):309–333, October 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ertur:2007:GTI

[EK07] Cem Ertur and Wilfried Koch. Growth, technological interdependence and spatial externalities: theory and evidence. Journal of Applied Econometrics, 22(6):1033–1062, September– October 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Edge:2010:CFP

[EKL10] Rochelle M. Edge, Michael T. Kiley, and Jean-Philippe Laforte. A comparison of forecast performance between federal reserve staff forecasts, simple reduced-form models, and a DSGE model. Journal of Applied Econometrics, 25(4):720–754, June– July 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ericsson:1991:UPG

[EL91] Neil R. Ericsson and Heidi Lyss. An update to PC-give: Version 6.01. Journal of Applied Econometrics, 6(3):321–325, July– September 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ellingsen:2022:NMV

[ELT22] Jon Ellingsen, Vegard H. Larsen, and Leif Anders Thorsrud. News media versus FRED-MD for macroeconomic forecasting. Journal of Applied Econometrics, 37(1):63–81, January– February 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Edge:2010:WMM

[ELW10] Rochelle M. Edge, Thomas Laubach, and John C. Williams. Welfare-maximizing monetary policy under parameter uncertainty. Journal of Applied Econometrics, 25(1):129–143, January–February 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ertur:2017:WSC

[EM17] Cem Ertur and Antonio Musolesi. Weak and strong crosssectional dependence: A panel data analysis of international technology diffusion. Journal of Applied Econometrics, 32(3): 477–503, April–May 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Esfahani:2014:EGM

[EMP14] Hadi Salehi Esfahani, Kamiar Mohaddes, and M. Hashem Pesaran. An empirical growth model for major oil exporters. Journal of Applied Econometrics, 29(1):1–21, January–February 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Euwals:1998:TPV

[EMvS98] Rob Euwals, Bertrand Melenberg, and Arthur van Soest. Testing the predictive value of subjective labour supply data. Journal of Applied Econometrics, 13(5):567–585, September– October 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Engelberg:2011:ATV

[EMW11] Joseph Engelberg, Charles F. Manski, and Jared Williams. Assessing the temporal variation of macroeconomic forecasts by a panel of changing composition. Journal of Applied Econometrics, 26(7):1059–1078, November–December 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Engle:2002:NFA

[Eng02] Robert Engle. New frontiers for ARCH models. Journal of Applied Econometrics, 17(5):425–446, September–October 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eren:2014:WBJ

[EO14] Ozkan Eren and Serkan Ozbeklik. Who benefits from job corps? A distributional analysis of an active labor market program. Journal of Applied Econometrics, 29(4):586–611, June–July 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Egger:2004:DTF

[EP04] Peter Egger and Michael Pfaffermayr. Distance, trade and FDI: a Hausman–Taylor SUR approach. Journal of Applied Econometrics, 19(2):227–246, March–April 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Everaert:2014:PAC

[EP14] Gerdie Everaert and Lorenzo Pozzi. The predictability of aggregate consumption growth in OECD countries: a panel data analysis. Journal of Applied Econometrics, 29(3):431–453, April–May 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Everaert:2022:EMI

[EP22] Gerdie Everaert and Lorenzo Pozzi. Encompassing measures of international consumption risk sharing and their link with trade and financial globalization. Journal of Applied Econometrics, 37(2):433–449, March 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eicher:2011:DPP

[EPR11] Theo S. Eicher, Chris Papageorgiou, and Adrian E. Raftery. Default priors and predictive performance in Bayesian model averaging, with application to growth determinants. Journal of Applied Econometrics, 26(1):30–55, January–February 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Everaert:2017:SES

[EPS17] Gerdie Everaert, Lorenzo Pozzi, and Ruben Schoonackers. On the stability of the excess sensitivity of aggregate consumption growth in the USA. Journal of Applied Econometrics, 32(4): 819–840, June–July 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Egger:2007:SAE

[EPW07] Peter Egger, Michael Pfaffermayr, and Andrea Weber. Sectoral adjustment of employment to shifts in outsourcing and trade: evidence from a dynamic fixed effects multinomial logit model. Journal of Applied Econometrics, 22(3):559–580, April– May 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Engel:1989:TIC

[ER89] Charles Engel and Anthony P. Rodrigues. Tests of international CAPM with time-varying covariances. Journal of Applied Econometrics, 4(2):119–138, April–June 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Epple:2003:PEF

[ERS03] Dennis Epple, Richard Romano, and Holger Sieg. Peer effects, financial aid and selection of students into colleges and universities: an empirical analysis. Journal of Applied Econometrics, 18(5):501–525, September–October 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eisenstat:2016:MIV

[ES16] Eric Eisenstat and Rodney W. Strachan. Modelling inflation volatility. Journal of Applied Econometrics, 31(5):805–820, August 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Escribano:1995:PPR

[Esc95] Alvaro Escribano. PCgive professional 8: A review. Journal of Applied Econometrics, 10(1):79–86, January–March 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eo:2023:UTI

[EUW23] Yunjong Eo, Luis Uzeda, and Benjamin Wong. Understanding trend inflation through the lens of the goods and services sectors. Journal of Applied Econometrics, 38(5):751–766, August 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Everaert:2016:DBC

[EV16] Gerdie Everaert and Hauke Vierke. Demographics and business cycle volatility: A spurious relationship? Journal of Applied Econometrics, 31(7):1467–1477, November–December 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Evans:1989:OUD

[Eva89] George W. Evans. Output and unemployment dynamics in the United States: 1950–1985. Journal of Applied Econometrics, 4 (3):213–237, July–September 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ebbes:2009:FIA

[EWB09] Peter Ebbes, Michel Wedel, and Ulf Böckenholt. Frugal IV alternatives to identify the parameter for an endogenous regressor. Journal of Applied Econometrics, 24(3):446–468, April– May 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Eskandari:2023:HRC

[EZ23] Ruhollah Eskandari and Morteza Zamanian. Heterogeneous responses to corporate marginal tax rates: Evidence from small and large firms. Journal of Applied Econometrics, 38(7):1018– 1047, November–December 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Elshiewy:2017:DBC

[EZB17] Ossama Elshiewy, German Zenetti, and Yasemin Boztug. Differences between classical and Bayesian estimates for mixed logit models: A replication study. Journal of Applied Econometrics, 32(2):470–476, March 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fachin:2007:LRT

[Fac07] Stefano Fachin. Long-run trends in internal migrations in Italy: a study in panel cointegration with dependent units. Journal of Applied Econometrics, 22(2):401–428, March 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fair:1996:CMU

[Fai96] Ray C. Fair. Computing median unbiased estimates in macroeconometric models. Journal of Applied Econometrics, 11(4): 431–435, July 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Farbmacher:2012:GMW

[Far12] Helmut Farbmacher. GMM with many weak moment conditions: Replication and application of Newey and Windmeijer (2009). Journal of Applied Econometrics, 27(2):343–346, March 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Forsberg:2002:BGB

[FB02] Lars Forsberg and Tim Bollerslev. Bridging the gap between the distribution of realized (ECU) volatility and ARCH modelling (of the euro): the GARCH–NIG model. Journal of Applied Econometrics, 17(5):535–548, September–October 2002.

CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fiorentini:1996:ADC

[FCP96] Gabriele Fiorentini, Giorgio Calzolari, and Lorenzo Panattoni. Analytic derivatives and the computation of GARCH estimates. Journal of Applied Econometrics, 11(4):399–417, July 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Feinstone:1987:MME

[Fei87] Lauren J. Feinstone. Minute by minute: Efficiency, normality, and randomness in intra-daily asset prices. Journal of Applied Econometrics, 2(3):193–214, July 1987. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ferrall:1994:SRR

[Fer94] C. Ferrall. Software review: a review of Stata 3.1. Journal of Applied Econometrics, 9(4):469–477, October–December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fernholz:2017:NML

[Fer17] Ricardo T. Fernholz. Nonparametric methods and localtime-based estimation for dynamic power law distributions. Journal of Applied Econometrics, 32(7):1244–1260, November– December 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ferman:2023:IDD

[Fer23a] Bruno Ferman. Inference in difference-in-differences: How much should we trust in independent clusters? Journal of Applied Econometrics, 38(3):358–369, April–May 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ferrall:2023:WHZ

[Fer23b] Christopher Ferrall. Was Harold Zurcher myopic after all? Replicating Rust’s engine replacement estimates. Journal of Applied Econometrics, 38(7):1093–1100, November–December 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fevrier:2008:NIE

[Fév08] Philippe Février. Nonparametric identification and estimation of a class of common value auction models. Journal of Applied Econometrics, 23(7):949–964, November–December 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ferreira:2019:ACG

[FFG19] Francisco H. G. Ferreira, Sergio Firpo, and Antonio F. Galvao. Actual and counterfactual growth incidence and delta Lorenz curves: Estimation and inference. Journal of Applied Econometrics, 34(3):385–402, April–May 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fricke:2020:ENR

[FFHL20] Hans Fricke, Markus Frölich, Martin Huber, and Michael Lechner. Endogeneity and non-response bias in treatment evaluation — nonparametric identification of causal effects by instruments. Journal of Applied Econometrics, 35(5):481–504, August 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fisher:2001:ECF

[FFS01] Douglas Fisher, Adrian R. Fleissig, and Apostolos Serletis. An empirical comparison of flexible demand system functional forms. Journal of Applied Econometrics, 16(1):59–80, January– February 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Furlanetto:2016:MSU

[FG16] Francesco Furlanetto and Nicolas Groshenny. Mismatch shocks and unemployment during the great recession. Journal of Applied Econometrics, 31(7):1197–1214, November–December 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ferrara:2018:WME

[FG18] Laurent Ferrara and Pierre Guérin. What are the macroeconomic effects of high-frequency uncertainty shocks? Journal of Applied Econometrics, 33(5):662–679, August 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ferroni:2019:SSI

[FGLL19] Filippo Ferroni, Stefano Grassi, and Miguel A. León-Ledesma. Selecting structural innovations in DSGE models. Journal of Applied Econometrics, 34(2):205–220, March 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Forni:2018:DFM

[FGLS18] Mario Forni, Alessandro Giovannelli, Marco Lippi, and Stefano Soccorsi. Dynamic factor model with infinite-dimensional factor space: Forecasting. Journal of Applied Econometrics, 33 (5):625–642, August 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Forni:2019:SVN

[FGS19] Mario Forni, Luca Gambetti, and Luca Sala. Structural VARs and noninvertible macroeconomic models. Journal of Applied Econometrics, 34(2):221–246, March 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Farbmacher:2018:ICT

[FGV18] Helmut Farbmacher, Raphael Guber, and Johan Vikström. Increasing the credibility of the twin birth instrument. Journal of Applied Econometrics, 33(3):457–472, April–May 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Feng:2012:ATE

[FH12] Qu Feng and William C. Horrace. Alternative technical efficiency measures: Skew, bias and scale. Journal of Applied Econometrics, 27(2):253–268, March 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fell:2013:SCC

[FH13] Harrison Fell and Alan C. Haynie. Spatial competition with changing market institutions. Journal of Applied Econometrics, 28(4):702–719, June–July 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fischer:2023:GBT

[FHHP23] Manfred M. Fischer, Niko Hauzenberger, Florian Huber, and Michael Pfarrhofer. General Bayesian time-varying parameter vector autoregressions for modeling government bond yields. Journal of Applied Econometrics, 38(1):69–87, January– February 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fischer:2022:RSC

[FHKN22] Martin Fischer, Gawain Heckley, Martin Karlsson, and Therese Nilsson. Revisiting Sweden’s comprehensive school reform: Effects on education and earnings. Journal of Applied Econometrics, 37(4):811–819, June–July 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Frolich:2004:MER

[FHL04] Markus Frölich, Almas Heshmati, and Michael Lechner. A microeconometric evaluation of rehabilitation of long-term sickness in Sweden. Journal of Applied Econometrics, 19(3):375– 396, May–June 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fisher:2016:EMM

[FHP16] Lance A. Fisher, Hyeon-Seung Huh, and Adrian R. Pagan. Econometric methods for modelling systems with a mixture of i (1) and i (0) variables. Journal of Applied Econometrics, 31 (5):892–911, August 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fiebig:1996:BRB

[Fie96] Denzil G. Fiebig. Book review: Book review: Bayesian analysis in statistics and econometrics: Essays in honor of Arnold Zellner. Edited by Donald A. Berry, Kathryn M. Challoner and John Geweke. John Wiley, New York, 1996, ISBN 0-471-11856-7, hardbound, £90, pp.xxii + 577. Journal of Applied Econometrics, 11(6):695–697, November 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fomby:2013:GAN

[FIL13] Thomas Fomby, Yuki Ikeda, and Norman V. Loayza. The growth aftermath of natural disasters. Journal of Applied Econometrics, 28(3):412–434, April–May 2013. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fuertes:2005:GTT

[FIM05] Ana-Maria Fuertes, Marwan Izzeldin, and Anthony Murphy. A guided tour of TSMod 4.03. Journal of Applied Econometrics, 20(5):691–698, July–August 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Finger:2010:RRS

[Fin10] Robert Finger. Review of ‘robustbase’ software for R. Journal of Applied Econometrics, 25(7):1205–1210, November– December 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fiorini:2012:FEE

[Fio12] Mario Fiorini. Fostering educational enrolment through subsidies: the issue of timing. Journal of Applied Econometrics, 27 (5):741–772, August 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fisher:1994:ATT

[Fis94] Stephen J. Fisher. Asset trading, transaction costs and the equity premium. Journal of Applied Econometrics, 9(S1):S71– S94, December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

French:2004:DDH

[FJ04] Eric French and John Bailey Jones. On the distribution and dynamics of health care costs. Journal of Applied Econometrics, 19(6):705–721, 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fortin:2021:LSS

[FJS21] Bernard Fortin, Nicolas Jacquemet, and Bruce Shearer. Labour supply, service intensity, and contracts: Theory and evidence on physicians. Journal of Applied Econometrics, 36(6):686–702, September–October 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fougere:2003:BIM

[FK03] Denis Fougère and Thierry Kamionka. Bayesian inference for the mover-stayer model in continuous time with an application to labour market transition data. Journal of Applied Econometrics, 18(6):697–723, November–December 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Feve:1994:RMT

[FL94] Patrick Féve and François Langot. The RBC models through statistical inference: An application with French data. Journal of Applied Econometrics, 9(S1):S11–S35, December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Flores-Lagunes:2007:FSE

[FL07] Alfonso Flores-Lagunes. Finite sample evidence of IV estimators under weak instruments. Journal of Applied Econometrics, 22(3):677–694, April–May 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Frolich:2015:CMN

[FL15] Markus Frölich and Michael Lechner. Combining matching and nonparametric instrumental variable estimation: Theory and an application to the evaluation of active labour market policies. Journal of Applied Econometrics, 30(5):718–738, August 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fleissig:1997:DLF

[Fle97] Adrian R. Fleissig. The dynamic Laurent flexible form and the demand for money. Journal of Applied Econometrics, 12(6): 687–699, November–December 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Flores-Lagunes:2007:ITE

[FLHS07] Alfonso Flores-Lagunes, William C. Horrace, and Kurt E. Schnier. Identifying technically efficient fishing vessels: a nonempty, minimal subset approach. Journal of Applied Econometrics, 22(4):729–745, June–July 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fagan:2013:WGS

[FLM13] Gabriel Fagan, James R. Lothian, and Paul D. Mcnelis. Was the gold standard really destabilizing? Journal of Applied Econometrics, 28(2):231–249, March 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fernandez:2001:MUC

[FLS01] Carmen Fernández, Eduardo Ley, and Mark F. J. Steel. Model uncertainty in cross-country growth regressions. Journal of Applied Econometrics, 16(5):563–576, September–October 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Flores-Lagunes:2012:ESS

[FLS12] Alfonso Flores-Lagunes and Kurt Erik Schnier. Estimation of sample selection models with spatial dependence. Journal of Applied Econometrics, 27(2):173–204, March 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fesselmeyer:2022:DDR

[FLS22] Eric Fesselmeyer, Haoming Liu, and Alberto Salvo. Declining discount rates in Singapore’s market for privately developed apartments. Journal of Applied Econometrics, 37(2):330–350, March 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:2016:ISD

[fLY16] Lung fei Lee and Jihai Yu. Identification of Spatial Durbin Panel Models. Journal of Applied Econometrics, 31(1):133–162, January–February 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fornari:1997:SVS

[FM97] Fabio Fornari and Antonio Mele. Sign- and volatility-switching ARCH models: theory and applications to international stock markets. Journal of Applied Econometrics, 12(1):49–65, January 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Franses:2002:FVI

[FM02] Philip Hans Franses and Michael McAleer. Financial volatility: an introduction. Journal of Applied Econometrics, 17(5):419– 424, September–October 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Foroni:2014:MFS

[FM14] Claudia Foroni and Massimiliano Marcellino. Mixed-frequency structural models: identification, estimation, and policy analysis. Journal of Applied Econometrics, 29(7):1118–1144, November–December 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Frey:2016:FBV

[FM16] Christoph Frey and Frieder Mokinski. Forecasting with Bayesian vector autoregressions estimated using professional forecasts. Journal of Applied Econometrics, 31(6):1083–1099, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Finlay:2019:TAW

[FM19] Keith Finlay and Leandro M. Magnusson. Two applications of wild bootstrap methods to improve inference in cluster-IV models. Journal of Applied Econometrics, 34(6):911–933, September–October 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fries:2018:NNR

[FMMR18] Sébastien Fries, Jean-Stéphane Mésonnier, Sarah Mouabbi, and Jean-Paul Renne. National natural rates of interest and the single monetary policy in the euro area. Journal of Applied Econometrics, 33(6):763–779, September–October 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Favero:2005:PCW

[FMN05] Carlo A. Favero, Massimiliano Marcellino, and Francesca Neglia. Principal components at work: the empirical analysis of monetary policy with large data sets. Journal of Applied Econometrics, 20(5):603–620, July–August 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Foroni:2019:MFM

[FMS19] Claudia Foroni, Massimiliano Marcellino, and Dalibor Stevanovic. Mixed-frequency models with moving-average components. Journal of Applied Econometrics, 34(5):688–706, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Francesconi:2006:IMS

[FN06] Marco Francesconi and Cheti Nicoletti. Intergenerational mobility and sample selection in short panels. Journal of Applied Econometrics, 21(8):1265–1293, December 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fosten:2023:NCS

[FN23] Jack Fosten and Shaoni Nandi. Nowcasting from crosssectionally dependent panels. Journal of Applied Econometrics, 38(6):898–919, September–October 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fagiolo:2008:OGR

[FNR08] Giorgio Fagiolo, Mauro Napoletano, and Andrea Roventini. Are output growth-rate distributions fat-tailed? Some evidence from OECD countries. Journal of Applied Econometrics, 23 (5):639–669, August 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fong:1995:STM

[FO95] Wai Mun Fong and Sam Ouliaris. Spectral tests of the martingale hypothesis for exchange rates. Journal of Applied Econometrics, 10(3):255–271, July–September 1995. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Feunou:2018:RNM

[FO18] Bruno Feunou and Cédric Okou. Risk-neutral moment-based estimation of affine option pricing models*. Journal of Applied Econometrics, 33(7):1007–1025, November–December 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fortin:1995:HBD

[For95] Nicole M. Fortin. Heterogeneity biases, distributional effects, and aggregate consumption: An empirical analysis using stratified microdata. Journal of Applied Econometrics, 10(3):287– 311, July–September 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fosten:2017:MSE

[Fos17a] Jack Fosten. Model selection with estimated factors and idiosyncratic components. Journal of Applied Econometrics, 32 (6):1087–1106, September–October 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Francis:2017:ECF

[FOS17b] Neville Francis, Michael T. Owyang, and Ozge Savascin. An endogenously clustered factor approach to international business cycles. Journal of Applied Econometrics, 32(7):1261–1276, November–December 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fry:1989:CTC

[FP89] Vanessa Fry and Panos Pashardes. Constructing the true cost of living index from the Engel curves of the PIGLOG model. Journal of Applied Econometrics, 4(1):41–56, January–March 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Franses:2002:CLE

[FP02] Philip Hans Franses and Richard Paap. Censored latent effects autoregression, with an application to US unemployment. Journal of Applied Econometrics, 17(4):347–366, July–August 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fok:2009:MCL

[FP09] Dennis Fok and Richard Paap. Modeling category-level purchase timing with brand-level marketing variables. Journal of Applied Econometrics, 24(3):469–489, April–May 2009. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fukac:2010:LIE

[FP10] Martin Fukac and Adrian Pagan. Limited information estimation and evaluation of DSGE models. Journal of Applied Econometrics, 25(1):55–70, January–February 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fanelli:2011:SBT

[FP11] Luca Fanelli and Giulio Palomba. Simulation-based tests of forward-looking models under VAR learning dynamics. Journal of Applied Econometrics, 26(5):762–782, August 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Firpo:2016:IED

[FP16] Sergio Firpo and Cristine Pinto. Identification and estimation of distributional impacts of interventions using changes in inequality measures. Journal of Applied Econometrics, 31(3): 457–486, April–May 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fuentes:2015:SPL

[FPR15] Julieta Fuentes, Pilar Poncela, and Julio Rodríguez. Sparse partial least squares in time series for macroeconomic forecasting. Journal of Applied Econometrics, 30(4):576–595, June– July 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Favero:1994:DMI

[FPS94] Carlo A. Favero, M. Hashem Pesaran, and Sunil Sharma. A duration model of irreversible oil investment: Theory and empirical evidence. Journal of Applied Econometrics, 9(S1):S95– S112, December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fok:2012:ROL

[FPV12] Dennis Fok, Richard Paap, and Bram Van Dijk. a rank-ordered logit model with unobserved heterogeneity in ranking capabilities. Journal of Applied Econometrics, 27(5):831–846, August 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ferraresi:2015:FPC

[FRF15] Tommaso Ferraresi, Andrea Roventini, and Giorgio Fagiolo. Fiscal policies and credit regimes: A TVAR approach. Journal of Applied Econometrics, 30(7):1047–1072, November– December 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fernandez-Rodriguez:2005:TCD

[FRSRAF05] Fernando Fernández-Rodríguez, Simón Sosvilla-Rivero, and Julián Andrada-Félix. Testing chaotic dynamics via Lyapunov exponents. Journal of Applied Econometrics, 20(7):911–930, December 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fong:2001:MCV

[FS01] Wai Mun Fong and Kim Hock See. Modelling the conditional volatility of commodity index futures as a regime switching process. Journal of Applied Econometrics, 16(2):133–163, March– April 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Feng:2009:EPU

[FS09] Guohua Feng and Apostolos Serletis. Efficiency and productivity of the US banking industry, 1998–2005: evidence from the Fourier cost function satisfying global regularity conditions. Journal of Applied Econometrics, 24(1):105–138, January– February 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Feve:2017:STM

[FS17] Patrick Fève and Jean-Guillaume Sahuc. In search of the transmission mechanism of fiscal policy in the Euro area. Journal of Applied Econometrics, 32(3):704–718, April–May 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fiorentini:2019:DST

[FS19] Gabriele Fiorentini and Enrique Sentana. Dynamic specification tests for dynamic factor models. Journal of Applied Econometrics, 34(3):325–346, April–May 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Foroni:2023:SOR

[FS23] Claudia Foroni and Livio Stracca. The shale oil revolution and the global oil supply curve. Journal of Applied Econometrics, 38(3):370–387, April–May 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fruhwirth-Schnatter:2006:HDC

[FSK06] Sylvia Frühwirth-Schnatter and Sylvia Kaufmann. How do changes in monetary policy affect bank lending? An analysis of Austrian bank data. Journal of Applied Econometrics, 21(3):275–305, April 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fruhwirth-Schnatter:2012:LME

[FSPWWE12] Sylvia Frühwirth-Schnatter, Christoph Pamminger, Andrea Weber, and Rudolf Winter-Ebmer. Labor market entry and earnings dynamics: Bayesian inference using mixtures-of-experts Markov chain clustering. Journal of Applied Econometrics, 27(7):1116–1137, November–December 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fair:1990:FIE

[FT90] Ray C. Fair and John B. Taylor. Full information estimation and stochastic simulation of models with rational expectations. Journal of Applied Econometrics, 5(4):381–392, September– December 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fu:2019:BCR

[Fu19] Bowen Fu. Bubbles and crises: Replicating the Anundsen et al. (2016) results. Journal of Applied Econometrics, 34(5):822– 826, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fujita:2011:DWF

[Fuj11] Shigeru Fujita. Dynamics of worker flows and vacancies: evidence from the sign restriction approach. Journal of Applied Econometrics, 26(1):89–121, January–February 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fukuda:2006:APC

[Fuk06] Kosei Fukuda. Age-period-cohort decomposition of aggregate data: an application to US and Japanese household saving rates. Journal of Applied Econometrics, 21(7):981–998, November 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fok:2005:MLP

[FvDF05] Dennis Fok, Dick van Dijk, and Philip Hans Franses. A multilevel panel STAR model for US manufacturing sectors. Journal of Applied Econometrics, 20(6):811–827, September–October 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Franses:2002:MFL

[FVP02] Philip Hans Franses, Marco Van Der Leij, and Richard Paap. Modelling and forecasting level shifts in absolute returns. Journal of Applied Econometrics, 17(5):601–616, September– October 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Fernandez-Villaverde:2005:EDE

[FVRR05] Jesús Fernández-Villaverde and Juan F. Rubio-Ramírez. Estimating dynamic equilibrium economies: linear versus nonlinear likelihood. Journal of Applied Econometrics, 20(7):891–910, December 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Falk:2003:TLR

[FW03] Barry Falk and Chun-Hsuan Wang. Testing long-run PPP with infinite-variance returns. Journal of Applied Econometrics, 18 (4):471–484, July–August 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Feldkircher:2012:IDR

[FZ12] Martin Feldkircher and Stefan Zeugner. The impact of data revisions on the robustness of growth determinants — a note on ‘determinants of economic growth: Will data tell?’. Journal of Applied Econometrics, 27(4):686–694, June–July 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Geweke:2011:HMN

[GA11] John Geweke and Gianni Amisano. Hierarchical Markov normal mixture models with applications to financial asset returns. Journal of Applied Econometrics, 26(1):1–29, January– February 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gonzalez-Astudillo:2019:EUO

[GA19] Manuel González-Astudillo. Estimating the U.S. output gap with state-level data. Journal of Applied Econometrics, 34(5): 795–810, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Goldstein:1989:SAS

[GAA^{+}89] Richard Goldstein, Jennifer Anderson, Arlene Ash, Ben Craig, David Harrington, and Marcello Pagano. Survival analysis software on MS/PC–DOS computers. Journal of Applied Econometrics, 4(4):393–414, October–December 1989. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Galeotti:1988:TMQ

[Gal88] Marzio Galeotti. Tobin’s marginal ‘ q ’ and tests of the firm’s dynamic equilibrium. Journal of Applied Econometrics, 3(4): 267–277, October–December 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Galbraith:1996:CRT

[Gal96] John W. Galbraith. Credit rationing and threshold effects in the relation between money and output. Journal of Applied Econometrics, 11(4):419–429, July 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Galvao:2006:SBT

[Gal06] Ana Beatriz C. Galvão. Structural break threshold VARs for predicting US recessions using the spread. Journal of Applied Econometrics, 21(4):463–487, May–June 2006. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gallant:2020:CBM

[Gal20] A. Ronald Gallant. Complementary Bayesian method of moments strategies. Journal of Applied Econometrics, 35(4):422– 439, June–July 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gandolfo:1988:A

[Gan88a] Giancarlo Gandolfo. Announcement. Journal of Applied Econometrics, 3(1):81, January 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gandolfo:1988:BRS

[Gan88b] Giancarlo Gandolfo. Book review: Sources of international comparative advantage: Theory and evidence. By E. E. Leamer. (Cambridge, Massachusetts: MIT Press, 1984. Pp. xxi + 353. $60.75 cloth; $16.95 paper). Journal of Applied Econometrics, 3(1):83–85, January 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gandolfo:1992:BRO

[Gan92] Giancarlo Gandolfo. Book review: Optimal control, expectations and uncertainty, S. Holly and A. Hughes Hallett. Cambridge University Press, Cambridge and New York, 1989. ISBN 0-521-2644-8, cloth, £25. Pp. i-ix + 244. Journal of Applied Econometrics, 7(1):110–114, January–March 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Garratt:2001:BRA

[Gar01a] Anthony Garratt. Book review: Applied macroeconometrics, Carlo A. Favero, Oxford University Press, Oxford, 2001, ISBN 0-19-877583-0 (hardback), pp. xi + 282, £40.00. Journal of Applied Econometrics, 16(5):647–652, September–October 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gil-Alana:2001:TSF

[GAR01b] L. A. Gil-Alaña and P. M. Robinson. Testing of seasonal fractional integration in UK and Japanese consumption and income. Journal of Applied Econometrics, 16(2):95–114, March– April 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gyimah-Brempong:1989:PPS

[GB89] Kwabena Gyimah-Brempong. Production of public safety: Are socioeconomic characteristics of local communities important factors? Journal of Applied Econometrics, 4(1):57–71, January–March 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Grogger:1991:MTC

[GC91] J. T. Grogger and R. T. Carson. Models for truncated counts. Journal of Applied Econometrics, 6(3):225–238, July– September 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Geroski:1988:PMP

[Ger88] P. A. Geroski. In pursuit of monopoly power: Recent quantitative work in industrial: Economics. Journal of Applied Econometrics, 3(2):107–123, April 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gerfin:1996:PSP

[Ger96] Michael Gerfin. Parametric and semi-parametric estimation of the binary response model of labour market participation. Journal of Applied Econometrics, 11(3):321–339, May 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Geweke:1986:EII

[Gew86] John Geweke. Exact inference in the inequality constrained normal linear regression model. Journal of Applied Econometrics, 1(2):127–141, April 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Geweke:1993:BTI

[Gew93] J. Geweke. Bayesian treatment of the independent student- t linear model. Journal of Applied Econometrics, 8(S1):S19–S40, December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gallizo:2008:MPA

[GGS08] Jose Luis Gallizo, Pilar Gargallo, and Manuel Salvador. Multivariate partial adjustment of financial ratios: a Bayesian hierarchical approach. Journal of Applied Econometrics, 23(1): 43–64, January–February 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Garratt:1996:MUE

[GH96] Anthony Garratt and Stephen G. Hall. Measuring underlying economic activity. Journal of Applied Econometrics, 11(2):135– 151, March 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gouret:2011:WKM

[GH11] Fabian Gouret and Guillaume Hollard. When Kahneman meets Manski: Using dual systems of reasoning to interpret subjective expectations of equity returns. Journal of Applied Econometrics, 26(3):371–392, April–May 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Goodhart:1993:NEH

[GHHP93] C. A. E. Goodhart, S. G. Hall, S. G. B. Henry, and B. Pesaran. News effects in a high-frequency model of the sterlingdollar exchange rate. Journal of Applied Econometrics, 8(1): 1–13, January–March 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Guerrieri:2020:I

[GHK20] Luca Guerrieri, Dale Henderson, and Jinill Kim. Interpreting shocks to the relative price of investment with a twosector model. Journal of Applied Econometrics, 35(1):82–98, January–February 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gregory:2004:MSA

[GHL04] Allan W. Gregory, Alfred A. Haug, and Nicoletta Lomuto. Mixed signals among tests for cointegration. Journal of Applied Econometrics, 19(1):89–98, January–February 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Grier:2004:AEU

[GHOS04] Kevin B. Grier, Ólan T. Henry, Nilss Olekalns, and Kalvinder Shields. The asymmetric effects of uncertainty on inflation and output growth. Journal of Applied Econometrics, 19(5): 551–565, September–October 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See replication report [Sav16].

Grammig:2005:DCM

[GHS05] Joachim Grammig, Reinhard Hujer, and Michael Scheidler. Discrete choice modelling in airline network management. Journal of Applied Econometrics, 20(4):467–486, May–June 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Giannoni:2007:ROM

[Gia07] Marc P. Giannoni. Robust optimal monetary policy in a forward-looking model with parameter and shock uncertainty. Journal of Applied Econometrics, 22(1):179–213, January– February 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gill:1988:CES

[Gil88] Andrew M. Gill. Choice of employment status and the wages of employees and the self-employed: Some further evidence. Journal of Applied Econometrics, 3(3):229–234, July–September 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gill:1993:BRC

[Gil93] Leonard Gill. Book review: A course in econometrics, by A. S. Goldberger, Harvard University Press, 1991, ISBN 0-674-17544-1, XYP19.95 hardbound. pp. 405. Journal of Applied Econometrics, 8(2):229–230, April–June 1993. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gilbert:1995:MMF

[Gil95] Christopher L. Gilbert. Modelling market fundamentals: A model of the aluminium market. Journal of Applied Econometrics, 10(4):385–410, October–December 1995. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Giovannetti:1989:BRL

[Gio89] Giorgia Giovannetti. Book review: The linear regression model under test, by W. Krämer and H. Sonnberger. Physica Verlag, Heidelberg, 1986, ISBN 3-7908-0356-1 cloth, D.M. 98.00, pp. i–ix + 189. Journal of Applied Econometrics, 4(2):209–211, April–June 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Giustinelli:2011:NPB

[Giu11] Pamela Giustinelli. Non-parametric bounds on quantiles under monotonicity assumptions: with an application to the Italian education returns. Journal of Applied Econometrics, 26(5):783– 824, August 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Goto:2023:ERN

[GJSvN23] Eiji Goto, Jan P. A. M. Jacobs, Tara M. Sinclair, and Simon van Norden. Employment reconciliation and nowcasting. Journal of Applied Econometrics, 38(7):1007–1017, November–December 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Guggenberger:2012:SDT

[GK12] Patrik Guggenberger and Gitanjali Kumar. On the size distortion of tests after an overidentifying restrictions pretest. Journal of Applied Econometrics, 27(7):1138–1160, November– December 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gu:2017:EBR

[GK17] Jiaying Gu and Roger Koenker. Empirical Bayesball remixed: Empirical Bayes methods for longitudinal data. Journal of Applied Econometrics, 32(3):575–599, April–May 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Groen:2013:MMM

[GKP13] Jan J. J. Groen, George Kapetanios, and Simon Price. Multivariate methods for monitoring structural change. Journal of Applied Econometrics, 28(2):250–274, March 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Groom:2007:DDF

[GKPP07] Ben Groom, Phoebe Koundouri, Ekaterini Panopoulou, and Theologos Pantelidis. Discounting the distant future: How much does model selection affect the certainty equivalent rate? Journal of Applied Econometrics, 22(3):641–656, April–May 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gaynor:2015:AHP

[GKV15] Martin S. Gaynor, Samuel A. Kleiner, and William B. Vogt. Analysis of hospital production: An output index approach. Journal of Applied Econometrics, 30(3):398–421, April–May 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Giraitis:2016:EDP

[GKWZ16] Liudas Giraitis, George Kapetanios, Anne Wetherilt, and Filip Zikes. Estimating the Dynamics and Persistence of Financial Networks, with an Application to the Sterling Money Market. Journal of Applied Econometrics, 31(1):58–84, January– February 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Glushenkova:2018:BPC

[GKZ18] Marina Glushenkova, Andros Kourtellos, and Marios Zachariadis. Barriers to price convergence. Journal of Applied Econometrics, 33(7):1081–1097, November–December 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Granger:1989:IPS

[GL89] C. W. J. Granger and T. H. Lee. Investigation of production, sales and inventory relationships using multicointegration and non-symmetric error correction models. Journal of Applied Econometrics, 4(S1):S145–S159, December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Giot:2003:VRL

[GL03] Pierre Giot and Sébastien Laurent. Value-at-risk for long and short trading positions. Journal of Applied Econometrics, 18 (6):641–663, November–December 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Garcia:2012:RAI

[GL12] René Garcia and Richard Luger. Risk aversion, intertemporal substitution, and the term structure of interest rates. Journal of Applied Econometrics, 27(6):1013–1036, September–October 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gaglianone:2014:COD

[GL14a] Wagner Piazza Gaglianone and Luiz Renato Lima. Constructing optimal density forecasts from point forecast combinations. Journal of Applied Econometrics, 29(5):736–757, August 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gospodinov:2014:MMM

[GL14b] Nikolay Gospodinov and Damba Lkhagvasuren. A momentmatching method for approximating vector autoregressive processes by finite-state Markov chains. Journal of Applied Econometrics, 29(5):843–859, August 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gouel:2017:ECS

[GL17] Christophe Gouel and Nicolas Legrand. Estimating the competitive storage model with trending commodity prices. Journal of Applied Econometrics, 32(4):744–763, June–July 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Guasch:2007:CIL

[GLS07] J. Luis Guasch, Jean-Jacques Laffont, and Stéphane Straub. Concessions of infrastructure in Latin America: Governmentled renegotiation. Journal of Applied Econometrics, 22(7): 1267–1294, December 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Geroski:2003:DFS

[GLUW03] P. A. Geroski, S. Lazarova, G. Urga, and C. F. Walters. Are differences in firm size transitory or permanent? Journal of Applied Econometrics, 18(1):47–59, January–February 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See erratum [Ano03].

Green:1991:TEC

[GM91] Steven L. Green and Knut Anton Mork. Toward efficiency in the crude-oil market. Journal of Applied Econometrics, 6(1): 45–66, January–March 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gadea:2009:ASP

[GM09] M. Dolores Gadea and Laura Mayoral. Aggregation is not the solution: the PPP puzzle strikes back. Journal of Applied Econometrics, 24(6):875–894, September–October 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Goldman:2013:MER

[GM13] Dana Goldman and Nicole Maestas. Medical expenditure risk and household portfolio choice. Journal of Applied Econometrics, 28(4):527–550, June–July 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gambetti:2017:LSS

[GM17] Luca Gambetti and Alberto Musso. Loan supply shocks and the business cycle. Journal of Applied Econometrics, 32(4): 764–782, June–July 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gillitzer:2019:DGI

[GM19] Christian Gillitzer and Martin McCarthy. Does global inflation help forecast inflation in industrialized countries? Journal of Applied Econometrics, 34(5):850–857, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Glass:1995:EPU

[GMH95] J. C. Glass, D. G. McKillop, and N. Hyndman. Efficiency in the provision of university teaching and research: An empirical analysis of UK universities. Journal of Applied Econometrics, 10(1):61–72, January–March 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gramacy:2014:ERF

[GMH14] Robert Gramacy, Samuel W. Malone, and Enrique Ter Horst. Exchange rate fundamentals, forecasting, and speculation: Bayesian models in black markets. Journal of Applied Econometrics, 29(1):22–41, January–February 2014. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gourieroux:1993:II

[GMR93] C. Gourieroux, A. Monfort, and E. Renault. Indirect inference. Journal of Applied Econometrics, 8(S1):S85–S118, December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Galvao:2019:TAP

[GMRO19] Antonio F. Galvao, Gabriel Montes-Rojas, and Jose Olmo. Tests of asset pricing with time-varying factor loads. Journal of Applied Econometrics, 34(5):762–778, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Geil:1997:EIH

[GMRZ97] Peter Geil, Andreas Million, Ralph Rotte, and Klaus F. Zimmermann. Economic incentives and hospitalization in Germany. Journal of Applied Econometrics, 12(3):295–311, May 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gobillon:2011:ELF

[GMS11] Laurent Gobillon, Thierry Magnac, and Harris Selod. The effect of location on finding a job in the Paris region. Journal of Applied Econometrics, 26(7):1079–1112, November–December 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ghysels:1998:BIP

[GMT98] Eric Ghysels, Robert E. McCulloch, and Ruey S. Tsay. Bayesian inference for periodic regime-switching models. Journal of Applied Econometrics, 13(2):129–143, March–April 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gagne:2000:CTS

[GN00] Robert Gagné and Carmine Nappi. The cost and technological structure of aluminium smelters worldwide. Journal of Applied Econometrics, 15(4):417–432, July–August 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Greenwood-Nimmo:2012:PFO

[GNNS12] Matthew Greenwood-Nimmo, Viet Hoang Nguyen, and Yongcheol Shin. Probabilistic forecasting of output growth, inflation and the balance of trade in a GVAR framework. Journal of Applied Econometrics, 27(4):554–573, June–July 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Grassi:2017:FSS

[GNS17] Stefano Grassi, Nima Nonejad, and Paolo Santucci De Magistris. Forecasting with the standardized self-perturbed Kalman filter. Journal of Applied Econometrics, 32(2):318–341, March 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Galvao:2022:FLF

[GO22] Ana Beatriz Galvão and Michael Owyang. Forecasting lowfrequency macroeconomic events with high-frequency data.

Journal of Applied Econometrics, 37(7):1314–1333, November– December 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gonzalez:2005:NBR

[Gon05] Libertad González. Nonparametric bounds on the returns to language skills. Journal of Applied Econometrics, 20(6):771– 795, September–October 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gordon:1997:STD

[Gor97] Stephen Gordon. Stochastic trends, deterministic trends, and business cycle turning points. Journal of Applied Econometrics, 12(4):411–434, July 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gotz:1990:CDJ

[Got90] Glenn A. Gotz. Comment on ‘The dynamics of job separation: The case of federal employees’. Journal of Applied Econometrics, 5(3):263–268, July–September 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [BMW90b] and reply [BMW90a].

Gouvea:2022:LDI

[Gou22] Raphael Gouvêa. Large devaluations and inflation inequality: Replicating Cravino and Levchenko (2017) with evidence from Brazil. Journal of Applied Econometrics, 37(3):657–664, April– May 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Greenberg:1997:PAM

[GP97] Edward Greenberg and Robert P. Parks. A predictive approach to model selection and multicollinearity. Journal of Applied Econometrics, 12(1):67–75, January 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Grier:2000:ERN

[GP00] Kevin B. Grier and Mark J. Perry. The effects of real and nominal uncertainty on inflation and output growth: some GARCH-M evidence. Journal of Applied Econometrics, 15(1):45–58, January–February 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Garcia-Perez:2006:JSN

[GP06] J. Ignacio García-Pérez. Job separation in a non-stationary search model: a structural estimation to evaluate alternative unemployment insurance systems. Journal of Applied Econometrics, 21(2):245–272, March 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gomes:2011:NRY

[GP11] Fábio Augusto Reis Gomes and Lourenço S. Paz. Narrow replication of yogo (2004) estimating the elasticity of intertemporal substitution when instruments are weak. Journal of Applied Econometrics, 26(7):1215–1216, November–December 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gardberg:2022:ACW

[GP22a] Malin Gardberg and Lorenzo Pozzi. Aggregate consumption and wealth in the long run: The impact of financial liberalization. Journal of Applied Econometrics, 37(1):161–186, January–February 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Garratt:2022:CPI

[GP22b] Anthony Garratt and Ivan Petrella. Commodity prices and inflation risk. Journal of Applied Econometrics, 37(2):392–414, March 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Greenberg:1989:SPD

[GPA89] E. Greenberg, W. A. Pollard, and W. T. Alpert. Statistical properties of data stretching. Journal of Applied Econometrics, 4(4):383–391, October–December 1989. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gerritsen:2017:TQS

[GPW17] Sander Gerritsen, Erik Plug, and Dinand Webbink. Teacher quality and student achievement: Evidence from a sample of Dutch twins. Journal of Applied Econometrics, 32(3):643–660, April–May 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Guerron-Quintana:2010:WYM

[GQ10] Pablo A. Guerron-Quintana. What you match does matter: the effects of data on DSGE estimation. Journal of Applied Econometrics, 25(5):774–804, August 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Guerroon-Quintana:2023:MFT

[GQZ23] Pablo Guerróon-Quintana and Molin Zhong. Macroeconomic forecasting in times of crises. Journal of Applied Econometrics, 38(3):295–320, April–May 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gorgens:2008:BAS

[GR08] Tue Gørgens and Chris Ryan. A bounds analysis of school completion rates in Australia. Journal of Applied Econometrics, 23 (3):287–304, April 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Giacomini:2010:FCU

[GR10] Raffaella Giacomini and Barbara Rossi. Forecast comparisons in unstable environments. Journal of Applied Econometrics, 25 (4):595–620, June–July 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Granger:1996:CWI

[Gra96] Clive W. J. Granger. Can we improve the perceived quality of economic forecasts? Journal of Applied Econometrics, 11(5): 455–473, September 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Granger:2002:SCR

[Gra02] Clive W. J. Granger. Some comments on risk. Journal of Applied Econometrics, 17(5):447–456, September–October 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Grant:2017:EMS

[Gra17] Angelia L. Grant. The early millennium slowdown: Replicating the Peersman (2005) results. Journal of Applied Econometrics, 32(1):224–232, January–February 2017. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Greasley:1992:SBE

[Gre92] D. Greasley. The stationarity of British economic and productivity growth 1856–1913. Journal of Applied Econometrics, 7 (2):203–209, April–June 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Greene:1993:BRC

[Gre93] William Greene. Book review: Computational economics and econometrics, (papers presented at the annual meeting of the Society For Economic Dynamics and Control, Minneapolis, Minnesota, USA, June, 1990), Edited By H. Amman, D. Belsley, and L. Pau. Kluwer Academic Publishers, Boston, USA, 1992, ISBN 0-7923-1287-2, hardback, pp. viii + 171. Journal of Applied Econometrics, 8(3):328–330, July–September 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gonzalez-Rivera:2008:JCS

[GRLM08] Gloria González-Rivera, Tae-Hwy Lee, and Santosh Mishra. Jumps in cross-sectional rank and expected returns: a mixture model. Journal of Applied Econometrics, 23(5):585–606, August 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gonzalez-Rivera:2020:PRI

[GRLR20] Gloria Gonzalez-Rivera, Yun Luo, and Esther Ruiz. Prediction regions for interval-valued time series. Journal of Applied Econometrics, 35(4):373–390, June–July 2020. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Geweke:2000:IID

[GRV00] John Geweke, John Rust, and Herman K. Van Dijk. Introduction: inference and decision making. Journal of Applied Econometrics, 15(6):545–546, November–December 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Garratt:2006:PVT

[GRW06] Anthony Garratt, Donald Robertson, and Stephen Wright. Permanent vs transitory components and economic fundamentals. Journal of Applied Econometrics, 21(4):521–542, May–June 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gustman:2004:SSP

[GS04] Alan L. Gustman and Thomas L. Steinmeier. Social security, pensions and retirement behaviour within the family. Journal of Applied Econometrics, 19(6):723–737, 2004. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Getachew:2007:PER

[GS07] Lullit Getachew and Robin C. Sickles. The policy environment and relative price efficiency of Egyptian private sector manufacturing: 1987/88–1995/96. Journal of Applied Econometrics, 22(4):703–728, June–July 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gross:2020:I

[GS20] Christian Gross and Pierre L. Siklos. Analyzing credit risk transmission to the nonfinancial sector in Europe: A network approach. Journal of Applied Econometrics, 35(1):61–81, January–February 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gong:2020:PAO

[GSS20] Yifan Gong, Todd Stinebrickner, and Ralph Stinebrickner. Perceived and actual option values of college enrollment. Journal of Applied Econometrics, 35(7):940–959, November–December 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Guidolin:2006:EMN

[GT06] Massimo Guidolin and Allan Timmermann. An econometric model of nonlinear dynamics in the joint distribution of stock and bond returns. Journal of Applied Econometrics, 21(1): 1–22, January–February 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gladden:2009:RBW

[GT09] Tricia Gladden and Christopher Taber. The relationship between wage growth and wage levels. Journal of Applied Econometrics, 24(6):914–932, September–October 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gurmu:1997:SPE

[Gur97] Shiferaw Gurmu. Semi-parametric estimation of Hurdle regression models with an application to medicaid3 utilization. Journal of Applied Econometrics, 12(3):225–242, May 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gengenbach:2016:ECT

[GUW16] Christian Gengenbach, Jean-Pierre Urbain, and Joakim Westerlund. Error correction testing in panels with common stochastic trends. Journal of Applied Econometrics, 31(6):982–1004, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gregory:1987:FWT

[GV87] Allan W. Gregory and Michael R. Veall. Formulating Wald tests of the restrictions implied by the rational expectations hypothesis. Journal of Applied Econometrics, 2(1):61–68, January 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gardeazabal:2017:ECB

[GVB17] Javier Gardeazabal and Ainhoa Vega-Bayo. An empirical comparison between the synthetic control method and HSIAO et al. ’s panel data approach to program evaluation. Journal of Applied Econometrics, 32(5):983–1002, August 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gautier:2012:SFE

[GvdK12] Pieter A. Gautier and Bas van der Klaauw. Selection in a field experiment with voluntary participation. Journal of Applied Econometrics, 27(1):63–84, January–February 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gustafsson:2023:BOH

[GVS23] Oskar Gustafsson, Mattias Villani, and Pär Stockhammar. Bayesian optimization of hyperparameters from noisy marginal likelihood estimates. Journal of Applied Econometrics, 38(4): 577–595, June–July 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Gong:2005:EGC

[GvSZ05] Xiaodong Gong, Arthur van Soest, and Ping Zhang. The effects of the gender of children on expenditure patterns in rural China: a semiparametric analysis. Journal of Applied Econometrics, 20(4):509–527, May–June 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Garratt:2019:RTF

[GVZ19] Anthony Garratt, Shaun P. Vahey, and Yunyi Zhang. Realtime forecast combinations for the oil price. Journal of Applied Econometrics, 34(3):456–462, April–May 2019. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Givord:2015:WDS

[GW15] Pauline Givord and Lionel Wilner. When does the steppingstone work? Fixed-term contracts versus temporary agency work in changing economic conditions. Journal of Applied Econometrics, 30(5):787–805, August 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Grant:2021:DES

[GY21] Everett Grant and Julieta Yung. The double-edged sword of global integration: Robustness, fragility, and contagion in the international firm network. Journal of Applied Econometrics, 36(6):760–783, September–October 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hajivassiliou:1994:SEA

[Haj94] V. A. Hajivassiliou. A simulation estimation analysis of the external debt crises of developing countries. Journal of Applied Econometrics, 9(2):109–131, April–June 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hall:1987:FLM

[Hal87] S. G. Hall. A forward looking model of the exchange rate. Journal of Applied Econometrics, 2(1):47–60, January 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hallett:1988:PCE

[Hal88] Andrew Hughes Hallett. Professor Chow’s Econometrics: A review. Journal of Applied Econometrics, 3(2):157–164, April 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hallett:1989:BRA

[Hal89] Andrew Hughes Hallett. Book review: Advances in econometrics, Fifth World Congress of the Econometric Society, Volumes I and II. Edited by T. F. Bewley, Cambridge University Press, Cambridge and New York, 1988 (Econometric Society Monographs 13 and 14), ISBN 0-521-34430-1 and 0-521-34552, £27.50 per volume (hardcover), pp. viii + 313 and viii + 259. Journal of Applied Econometrics, 4(2):207–209, April– June 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hallett:1993:BRC

[Hal93a] Andrew Hughes Hallett. Book review: Control Theory and Dynamic Games in Economic Policy Analysis, M. L. Petit, Cambridge University Press, Cambridge and New York, 1990, £30.00, ISBN 0521 385 237, pp. xiv + 338. Journal of Applied Econometrics, 8(3):326–327, July–September 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hallman:1993:RP

[Hal93b] J. Hallman. Review of S-Plus. Journal of Applied Econometrics, 8(2):213–219, April–June 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hamilton:2021:MGE

[Ham21] James D. Hamilton. Measuring global economic activity. Journal of Applied Econometrics, 36(3):293–303, April–May 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hansen:1992:LRT

[Han92] B. E. Hansen. The likelihood ratio test under nonstandard conditions: Testing the Markov switching model of GNP. Journal of Applied Econometrics, 7(S1):S61–S82, December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hanse:1993:CSB

[Han93] G. D. Hanse. The cyclical and secular behaviour of the labour input: Comparing efficiency units and hours worked. Journal of Applied Econometrics, 8(1):71–80, January–March 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hansen:1996:ELR

[Han96] Bruce E. Hansen. Erratum: The likelihood ratio test under nonstandard conditions: Testing the Markov switching model of GNP. Journal of Applied Econometrics, 11(2):195–198, March 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Harrison:1993:GR

[Har93] A. Harrison. Gaim 1.1: A review. Journal of Applied Econometrics, 8(1):101–107, January–March 1993. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Harrison:2008:SRR

[Har08] Teresa D. Harrison. Software review: Review of np software for R. Journal of Applied Econometrics, 23(6):861–865, September–October 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hayo:2004:RP

[Hay04] Bernd Hayo. Review of PcGive 10. Journal of Applied Econometrics, 19(4):525–531, July–August 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hsiao:2012:PDA

[HCW12] Cheng Hsiao, H. Steve Ching, and Shui Ki Wan. A panel data approach for program evaluation: measuring the benefits of political and economic integration of Hong Kong with Mainland China. Journal of Applied Econometrics, 27(5):705–740, August 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hols:1991:LDE

[HD91] Martien C. A. B. Hols and Casper G. De Vries. The limiting distribution of extremal exchange rate returns. Journal of Applied Econometrics, 6(3):287–302, July–September 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Heiss:2008:SNI

[Hei08] Florian Heiss. Sequential numerical integration in nonlinear state space models for microeconometric panel data. Journal of Applied Econometrics, 23(3):373–389, April 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Heinemann:2017:EEF

[Hei17] Alexander Heinemann. Efficient estimation of factor models with time and cross-sectional dependence. Journal of Applied Econometrics, 32(6):1107–1122, September–October 2017. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hellstrom:2006:BCD

[Hel06] Jörgen Hellström. A bivariate count data model for household tourism demand. Journal of Applied Econometrics, 21(2):213– 226, March 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Henry:1993:BRS

[Hen93] S. G. B. Henry. Book reviews: Seasonal adjustment as a practical problem, F. A. G. Den Butter and M. M. G. Fase. North-Holland, 1991, ISBN 90-267-1264-3, US $94.50/DFL 165, pp.

226. Modelling Seasonality, Edited by S. Hylleberg. Oxford University Press, 1992, ISBN 0-19-877317, $45, pp. 476. Journal of Applied Econometrics, 8(4):421–423, October–December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hendry:2001:MUI

[Hen01] David F. Hendry. Modelling UK inflation, 1875–1991. Journal of Applied Econometrics, 16(3):255–275, May–June 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hensher:2006:HDR

[Hen06] David A. Hensher. How do respondents process stated choice experiments? Attribute consideration under varying information load. Journal of Applied Econometrics, 21(6):861–878, September–October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Henderson:2010:TMR

[Hen10] Daniel J. Henderson. A test for multimodality of regression derivatives with application to nonparametric growth regressions. Journal of Applied Econometrics, 25(3):458–480, April– May 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Herwartz:2019:LRN

[Her19] Helmut Herwartz. Long-run neutrality of demand shocks: Revisiting Blanchard and Quah (1989) with independent structural shocks. Journal of Applied Econometrics, 34(5):811–819, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Holloway:2008:BED

[HET08] Garth Holloway, Simeon Ehui, and Amare Teklu. Bayes estimates of distance-to-market: transactions costs, cooperatives and milk-market development in the Ethiopian highlands. Journal of Applied Econometrics, 23(5):683–696, August 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Heyma:2004:SDA

[Hey04] Arjan Heyma. A structural dynamic analysis of retirement behaviour in the Netherlands. Journal of Applied Econometrics, 19(6):739–759, 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hobijn:2000:APR

[HF00] Bart Hobijn and Philip Hans Franses. Asymptotically perfect and relative convergence of productivity. Journal of Applied Econometrics, 15(1):59–81, January–February 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hambuckers:2018:UED

[HGK18] Julien Hambuckers, Andreas Groll, and Thomas Kneib. Understanding the economic determinants of the severity of operational losses: A regularized generalized Pareto regression approach. Journal of Applied Econometrics, 33(6):898–935, September–October 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hevia:2015:EFY

[HGRSS15] Constantino Hevia, Martin Gonzalez-Rozada, Martin Sola, and Fabio Spagnolo. Estimating and forecasting the yield curve using a Markov switching dynamic Nelson and Siegel model. Journal of Applied Econometrics, 30(6):987–1009, September– October 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Han:1990:FPE

[HH90] Aaron Han and Jerry A. Hausman. Flexible parametric estimation of duration and competing risk models. Journal of Applied Econometrics, 5(1):1–28, January–March 1990. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hutton:1995:MCA

[HH95] John Hutton and James Hutton. The Maple computer algebra system: A review. Journal of Applied Econometrics, 10(3):329– 337, July–September 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Haigh:2002:CSH

[HH02] Michael S. Haigh and Matthew T. Holt. Crack spread hedging: accounting for time-varying volatility spillovers in the energy futures markets. Journal of Applied Econometrics, 17(3): 269–289, May–June 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hickman:2015:RST

[HH15] Brent R. Hickman and Timothy P. Hubbard. Replacing sample trimming with boundary correction in nonparametric estimation of first-price auctions. Journal of Applied Econometrics, 30 (5):739–762, August 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hautsch:2020:MDI

[HH20] Nikolaus Hautsch and Rodrigo Herrera. Multivariate dynamic intensity peaks-over-threshold models. Journal of Applied Econometrics, 35(2):248–272, March 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hall:1986:FEM

[HHJ86] S. G. Hall, S. G. B. Henry, and C. B. Johns. Forecasting with an econometric model: Some recent results using the national institute model. Journal of Applied Econometrics, 1(2):163– 183, April 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ho:2011:NPP

[HHJC11] Anson T. Y. Ho, Kim P. Huynh, and David T. Jacho-Chávez. npRmpi: A package for parallel distributed kernel estimation in R. Journal of Applied Econometrics, 26(2):344–349, March 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ho:2014:CPN

[HHJC14] Anson T. Y. Ho, Kim P. Huynh, and David T. Jacho-Chávez. CRS: a package for nonparametric spline estimation in R. Journal of Applied Econometrics, 29(2):348–352, March 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ho:2016:FEC

[HHJC16] Anson T. Y. Ho, Kim P. Huynh, and David T. Jacho-Chávez. Flexible estimation of copulas: An application to the US housing crisis. Journal of Applied Econometrics, 31(3):603–610, April–May 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Huber:2020:DIE

[HHLL20] Martin Huber, Yu-Chin Hsu, Ying-Ying Lee, and Layal Lettry. Direct and indirect effects of continuous treatments based on generalized propensity score weighting. Journal of Applied Econometrics, 35(7):814–840, November–December 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hardle:2015:LAM

[HHM15] Wolfgang K. Härdle, Nikolaus Hautsch, and Andrija Mihoci. Local adaptive multiplicative error models for high-frequency forecasts. Journal of Applied Econometrics, 30(4):529–550, June–July 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hauzenberger:2021:CSS

[HHO21] Niko Hauzenberger, Florian Huber, and Luca Onorante. Combining shrinkage and sparsity in conjugate vector autoregressive models. Journal of Applied Econometrics, 36(3):304–327, April–May 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hall:1992:TDS

[HHP92] S. G. Hall, S. G. B. Henry, and M. Pemberton. Testing a discrete switching disequilibrium model of the UK labour market. Journal of Applied Econometrics, 7(1):83–91, January–March 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hansen:2012:RGJ

[HHS12] Peter Reinhard Hansen, Zhuo Huang, and Howard Howan Shek. Realized GARCH: a joint model for returns and realized measures of volatility. Journal of Applied Econometrics, 27(6):877–906, September–October 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hajivassiliou:2007:ULC

[HI07] Vassilis A. Hajivassiliou and Yannis M. Ioannides. Unemployment and liquidity constraints. Journal of Applied Econometrics, 22(3):479–510, April–May 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hirose:2016:ZLB

[HI16] Yasuo Hirose and Atsushi Inoue. The zero lower bound and parameter bias in an estimated DSGE model. Journal of Applied Econometrics, 31(4):630–651, June–July 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hill:2006:WDR

[Hil06] Daniel H. Hill. Wealth dynamics: reducing noise in panel data. Journal of Applied Econometrics, 21(6):845–860, September– October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hill:2007:ETL

[Hil07] Jonathan B. Hill. Efficient tests of long-run causation in trivariate VAR processes with a rolling window study of the moneyincome relationship. Journal of Applied Econometrics, 22(4): 747–765, June–July 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hecq:2020:MCN

[HIT20] Alain Hecq, Joao Victor Issler, and Sean Telg. Mixed causalnoncausal autoregressions with exogenous regressors. Journal of Applied Econometrics, 35(3):328–343, April–May 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Harvey:1993:DSF

[HJ93] A. C. Harvey and A. Jaeger. Detrending, stylized facts and the business cycle. Journal of Applied Econometrics, 8(3):231– 247, July–September 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Huynh:2010:FSD

[HJC10] Kim P. Huynh and David T. Jacho-Chávez. Firm size distributions through the lens of functional principal components analysis. Journal of Applied Econometrics, 25(7):1211–1214, November–December 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Honore:2020:IEM

[HJdP20] Bo Honoré, Thomas Jørgensen, and Áureo de Paula. The informativeness of estimation moments. Journal of Applied Econometrics, 35(7):797–813, November–December 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hoke:2021:CCE

[HK21] Sinem Hacioğlu Hoke and George Kapetanios. Common correlated effect cross-sectional dependence corrections for nonlinear conditional mean panel models. Journal of Applied Econometrics, 36(1):125–150, January–February 2021. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hjalmarsson:2022:LRP

[HK22] Erik Hjalmarsson and Tamas Kiss. Long-run predictability tests are even worse than you thought. Journal of Applied Econometrics, 37(7):1334–1355, November–December 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Huber:2023:SSC

[HK23] Florian Huber and Gary Koop. Subspace shrinkage in conjugate Bayesian vector autoregressions. Journal of Applied Econometrics, 38(4):556–576, June–July 2023. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Huber:2019:SSS

[HKF19] Florian Huber, Gregor Kastner, and Martin Feldkircher. Should I stay or should I go? A latent threshold approach to large-scale mixture innovation models. Journal of Applied Econometrics, 34(5):621–640, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hautsch:2015:DHF

[HKM15] Nikolaus Hautsch, Lada M. Kyj, and Peter Malec. Do highfrequency data improve high-dimensional portfolio allocations? Journal of Applied Econometrics, 30(2):263–290, March 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hanushek:2003:DPA

[HKMR03] Eric A. Hanushek, John F. Kain, Jacob M. Markman, and Steven G. Rivkin. Does peer ability affect student achievement? Journal of Applied Econometrics, 18(5):527–544, September– October 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hautsch:2012:BRA

[HKO12] Nikolaus Hautsch, Lada M. Kyj, and Roel C. A. Oomen. A blocking and regularization approach to high-dimensional realized covariance estimation. Journal of Applied Econometrics, 27(4):625–645, June–July 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Haupt:2014:SQB

[HKS14] Harry Haupt, Kathrin Kagerer, and Winfried J. Steiner. Smooth quantile–based modeling of brand sales, price and promotional effects from retail scanner panels. Journal of Applied Econometrics, 29(6):1007–1028, September–October 2014. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Herwartz:2016:MHA

[HKS16] Helmut Herwartz, Nadja Klein, and Christoph Strumann. Modelling hospital admission and length of stay by means of generalised count data models. Journal of Applied Econometrics, 31(6):1159–1182, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hudomiet:2011:SMC

[HKW11] Péter Hudomiet, Gábor Kézdi, and Robert J. Willis. Stock market crash and expectations of American households. Journal of Applied Econometrics, 26(3):393–415, April–May 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Higgins:1992:MVR

[HL92] Matthew L. Higgins and Nabil Ltaifa. Microtsp version 7.0: A review. Journal of Applied Econometrics, 7(4):423–429, October–December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hamilton:1996:SMV

[HL96] James D. Hamilton and Gang Lin. Stock market volatility and the business cycle. Journal of Applied Econometrics, 11(5): 573–593, September 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hansen:2005:FCV

[HL05] Peter R. Hansen and Asger Lunde. A forecast comparison of volatility models: does anything beat a GARCH(1,1)? Journal of Applied Econometrics, 20(7):873–889, December 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hausman:2007:CBI

[HL07] Jerry Hausman and Ephraim Leibtag. Consumer benefits from increased competition in shopping outlets: Measuring the effect of Wal-Mart. Journal of Applied Econometrics, 22(7):1157– 1177, December 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Huang:2010:DTE

[HL10] Fali Huang and Myoung-Jae Lee. Dynamic treatment effect analysis of TV effects on child cognitive development. Journal of Applied Econometrics, 25(3):392–419, April–May 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hsieh:2016:SIM

[HL16] Chih-Sheng Hsieh and Lung Fei Lee. A social interactions model with endogenous friendship formation and selectivity. Journal of Applied Econometrics, 31(2):301–319, March 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Harding:2017:PQR

[HL17] Matthew Harding and Carlos Lamarche. Penalized quantile regression with semiparametric correlated effects: An application with heterogeneous preferences. Journal of Applied Econometrics, 32(2):342–358, March 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Han:2019:EGB

[HL19] Sukjin Han and Sungwon Lee. Estimation in a generalization of bivariate probit models with dummy endogenous regressors. Journal of Applied Econometrics, 34(6):994–1015, September– October 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hsieh:2021:SIS

[HL21] Chih-Sheng Hsieh and Xu Lin. Social interactions and social preferences in social networks. Journal of Applied Econometrics, 36(2):165–189, March 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Huber:2017:SIB

[HLM17] Martin Huber, Lukas Laffers, and Giovanni Mellace. Sharp IV bounds on average treatment effects on the treated and other populations under endogeneity and noncompliance. Journal of Applied Econometrics, 32(1):56–79, January–February 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Horowitz:1998:IAS

[HLMvS98] Joel Horowitz, Myoung-Jae Lee, Bertrand Melenberg, and Arthur van Soest. Introduction: application of semiparametric methods for micro-data. Journal of Applied Econometrics, 13(5):431–433, September–October 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Harding:2020:CCE

[HLP20] Matthew Harding, Carlos Lamarche, and M. Hashem Pesaran. Common correlated effects estimation of heterogeneous dynamic panel quantile regression models. Journal of Applied Econometrics, 35(3):294–314, April–May 2020. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Harvey:2021:RTD

[HLST21] David I. Harvey, Stephen J. Leybourne, Robert Sollis, and A. M. Robert Taylor. Real-time detection of regimes of predictability in the US equity premium. Journal of Applied Econometrics, 36(1):45–70, January–February 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hodgson:2002:TCA

[HLV02] Douglas J. Hodgson, Oliver Linton, and Keith Vorkink. Testing the capital asset pricing model efficiently under elliptical symmetry: a semiparametric approach. Journal of Applied Econometrics, 17(6):617–639, December 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hansen:2014:RBG

[HLV14] Peter Reinhard Hansen, Asger Lunde, and Valeri Voev. Realized beta GARCH: a multivariate GARCH model with realized measures of volatility. Journal of Applied Econometrics, 29(5): 774–799, August 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hartley:2014:WRW

[HLW14] Roger Hartley, Gauthier Lanot, and Ian Walker. Who really wants to be a millionaire? Estimates of risk aversion from gameshow data. Journal of Applied Econometrics, 29(6):861– 879, September–October 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Holt:2003:QRE

[HM03] Matthew T. Holt and Andrew M. McKenzie. Quasi-rational and ex ante price expectations in commodity supply models: an empirical analysis of the US broiler market. Journal of Applied Econometrics, 18(4):407–426, July–August 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Henderson:2008:GL

[HM08] Daniel J. Henderson and Daniel L. Millimet. Is gravity linear? Journal of Applied Econometrics, 23(2):137–172, March 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hafner:2012:DSC

[HM12] Christian M. Hafner and Hans Manner. Dynamic stochastic copula models: estimation, inference and applications. Journal of Applied Econometrics, 27(2):269–295, March 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Henry:2013:ERP

[HM13] Marc Henry and Ismael Mourifié. Euclidean revealed preferences: testing the spatial voting model. Journal of Applied Econometrics, 28(4):650–666, June–July 2013. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Huber:2015:TCI

[HM15] Martin Huber and Blaise Melly. A test of the conditional independence assumption in sample selection models. Journal of Applied Econometrics, 30(7):1144–1168, November–December 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hummel:2017:LFP

[HM17] Patrick Hummel and R. Preston McAfee. Loss functions for predicted click-through rates in auctions for online advertising. Journal of Applied Econometrics, 32(7):1314–1328, November– December 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hunt-McCool:1994:EDM

[HMKN94] Janet Hunt-McCool, B. F. Kiker, and Ying Chu Ng. Estimates of the demand for medical care under different functional forms. Journal of Applied Econometrics, 9(2):201–218, April–June 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Herranz:2005:ARO

[HMRC05] Neus Herranz, Ricardo Mora, and Javier Ruiz-Castillo. An algorithm to reduce the occupational space in gender segregation studies. Journal of Applied Econometrics, 20(1):25–37, January–February 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hirschberg:2001:CAW

[HMS01] Joseph G. Hirschberg, Esfandiar Maasoumi, and Daniel J. Slottje. Clusters of attributes and well-being in the USA. Journal of Applied Econometrics, 16(3):445–460, May–June 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hansen:2016:EBD

[HMS16] Stephen Hansen, Michael McMahon, and Sorawoot Srisuma. Estimating Bayesian decision problems with heterogeneous expertise. Journal of Applied Econometrics, 31(4):762–771, June– July 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hoevenaars:2014:SAA

[HMSS14] Roy P. P. M. Hoevenaars, Roderick D. J. Molenaar, Peter C. Schotman, and Tom B. M. Steenkamp. Strategic asset allocation for long–term investors: parameter uncertainty and prior information. Journal of Applied Econometrics, 29(3):353–376, April–May 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hillebrand:2023:ERM

[HMSU23] Eric Hillebrand, Jakob Guldbæk Mikkelsen, Lars Spreng, and Giovanni Urga. Exchange rates and macroeconomic fundamentals: Evidence of instabilities from time-varying factor loadings. Journal of Applied Econometrics, 38(6):857–877, September– October 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hu:2015:EIS

[HMT15] Xuequn Hu, Murat K. Munkin, and Pravin K. Trivedi. Estimating incentive and selection effects in the Medigap insurance market: An application with Dirichlet process mixture model. Journal of Applied Econometrics, 30(7):1115–1143, November– December 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hai:1997:USF

[HMW97] Weike Hai, Nelson C. Mark, and Yangru Wu. Understanding spot and forward exchange rate regressions. Journal of Applied Econometrics, 12(6):715–734, November–December 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Horowitz:1989:STC

[HN89] Joel L. Horowitz and George R. Neumann. Specification testing in censored regression models: Parametric and semiparametric methods. Journal of Applied Econometrics, 4(S1):S61–S86, December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Harvey:2000:TMF

[HN00] David Harvey and Paul Newbold. Tests for multiple forecast encompassing. Journal of Applied Econometrics, 15(5):471– 482, September–October 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ho:2006:WER

[Ho06] Katherine Ho. The welfare effects of restricted hospital choice in the US medical care market. Journal of Applied Econometrics, 21(7):1039–1079, November 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hochguertel:2009:CIV

[HO09] Stefan Hochguertel and Henry Ohlsson. Compensatory inter vivos gifts. Journal of Applied Econometrics, 24(6):993–1023, September–October 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hochguertel:2003:PMP

[Hoc03] Stefan Hochguertel. Precautionary motives and portfolio decisions. Journal of Applied Econometrics, 18(1):61–77, January– February 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hodgson:1999:AEC

[Hod99] Douglas J. Hodgson. Adaptive estimation of cointegrated models: simulation evidence and an application to the forward exchange market. Journal of Applied Econometrics, 14(6):627– 650, November–December 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hong:2013:MEN

[Hon13] Seung-Hyun Hong. Measuring the effect of Napster on recorded music sales: difference-in-differences estimates under compositional changes. Journal of Applied Econometrics, 28(2):297– 324, March 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hong:2022:REA

[Hon22] Seung-Hyun Hong. Real estate agents’ influence on housing search. Journal of Applied Econometrics, 37(3):563–582, April– May 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hoque:1991:ATR

[Hoq91] Asraul Hoque. An application and test for a random coefficient model in Bangladesh agriculture. Journal of Applied Econometrics, 6(1):77–90, January–March 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Horowitz:1992:RLP

[Hor92] Joel L. Horowitz. The role of the list price in housing markets: Theory and an econometric model. Journal of Applied Econometrics, 7(2):115–129, April–June 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hospido:2012:MHD

[Hos12] L. Hospido. Modelling heterogeneity and dynamics in the volatility of individual wages. Journal of Applied Econometrics, 27(3):386–414, April–May 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hendry:2001:SIM

[HP01] David F. Hendry and M. Hashem Pesaran. A special issue in memory of John Denis Sargan: studies in empirical macroeconometrics. Journal of Applied Econometrics, 16(3):197–202, May–June 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Haley:2004:SIR

[HP04a] M. Ryan Haley and Harry J. Paarsch. The stochastic implications of rent maximization: an application to stumpage rates for timber in British Columbia. Journal of Applied Econometrics, 19(1):25–48, January–February 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hu:2004:DFF

[HP04b] Ling Hu and Peter C. B. Phillips. Dynamics of the federal funds target rate: a nonstationary discrete choice approach. Journal of Applied Econometrics, 19(7):851–867, November–December 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Harding:2005:SFC

[HP05] Don Harding and Adrian Pagan. A suggested framework for classifying the modes of cycle research. Journal of Applied Econometrics, 20(2):151–159, 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hendry:2009:MCG

[HP09] David F. Hendry and M. Hashem Pesaran. In memory of Clive Granger: an advisory board member of the Journal. Journal of Applied Econometrics, 24(6):871–873, September–October 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Huber:2021:DST

[HP21] Florian Huber and Michael Pfarrhofer. Dynamic shrinkage in time-varying parameter stochastic volatility in mean models. Journal of Applied Econometrics, 36(2):262–270, March 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hernandez:2007:WCM

[HPH07] Monica Hernandez, Stephen Pudney, and Ruth Hancock. The welfare cost of means-testing: pensioner participation in income support. Journal of Applied Econometrics, 22(3):581– 598, April–May 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Henderson:2008:MWM

[HPR08] Daniel J. Henderson, Christopher F. Parmeter, and R. Robert Russell. Modes, weighted modes, and calibrated modes: evidence of clustering using modality tests. Journal of Applied Econometrics, 23(5):607–638, August 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hall:1997:CCR

[HPS97] Stephen G. Hall, Zacharias Psaradakis, and Martin Sola. Cointegration and changes in regime: the Japanese consumption function. Journal of Applied Econometrics, 12(2):151–168, March 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hall:1999:DPC

[HPS99] Stephen G. Hall, Zacharias Psaradakis, and Martin Sola. Detecting periodically collapsing bubbles: a Markov-switching unit root test. Journal of Applied Econometrics, 14(2):143– 154, March–April 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hevia:2018:RPS

[HPS18] Constantino Hevia, Ivan Petrella, and Martin Sola. Risk premia and seasonality in commodity futures. Journal of Applied Econometrics, 33(6):853–873, September–October 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hayakawa:2023:SDP

[HPS23] Kazuhiko Hayakawa, M. Hashem Pesaran, and L. Vanessa Smith. Short T dynamic panel data models with individual, time and interactive effects. Journal of Applied Econometrics, 38(6):940–967, September–October 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Huang:2019:CAT

[HPY19] Haitao Huang, Liang Peng, and Vincent W. Yao. Comovements and asymmetric tail dependence in state housing prices in the USA: a nonparametric approach. Journal of Applied Econometrics, 34(5):843–849, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hoffman:1996:AFP

[HR96] Dennis L. Hoffman and Robert H. Rasche. Assessing forecast performance in a cointegrated system. Journal of Applied Econometrics, 11(5):495–517, September 1996. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hu:2012:ENM

[HR12] Yingyao Hu and Geert Ridder. Estimation of nonlinear models with mismeasured regressors using marginal information. Journal of Applied Econometrics, 27(3):347–385, April–May 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Herrera:2020:EOS

[HR20] Ana María Herrera and Sandeep Kumar Rangaraju. The effect of oil supply shocks on US economic activity: What have we learned? Journal of Applied Econometrics, 35(2):141–159, March 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hartog:1994:AIJ

[HRV94] J. Hartog, G. Ridder, and M. Visser. Allocation of individuals to job levels under rationing. Journal of Applied Econometrics, 9(4):437–451, October–December 1994. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hall:1990:ASS

[HS90] S. G. Hall and M. J. Stephenson. An algorithm for the solution of stochastic optimal control problems for large nonlinear econometric models. Journal of Applied Econometrics, 5(4): 393–399, September–December 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Horrace:2000:MCB

[HS00] William C. Horrace and Peter Schmidt. Multiple comparisons with the best, with economic applications. Journal of Applied Econometrics, 15(1):1–26, January–February 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Herbst:2014:SMC

[HS14] Edward Herbst and Frank Schorfheide. Sequential Monte Carlo sampling for DSGE models. Journal of Applied Econometrics, 29(7):1073–1098, November–December 2014. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hsu:2021:TMC

[HS21] Yu-Chin Hsu and Shu Shen. Testing monotonicity of conditional treatment effects under regression discontinuity designs. Journal of Applied Econometrics, 36(3):346–366, April–May 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hsiao:2005:AVD

[HSF05] Cheng Hsiao, Yan Shen, and Hiroshi Fujiki. Aggregate vs. disaggregate data analysis — a paradox in the estimation of a money demand function of Japan under the low interest rate policy. Journal of Applied Econometrics, 20(5):579–601, July– August 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Haupt:2010:NEH

[HST10] Harry Haupt, Joachim Schnurbus, and Rolf Tschernig. On nonparametric estimation of a hedonic price function. Journal of Applied Econometrics, 25(5):894–901, August 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hurn:2016:STL

[HST16] A. Stan Hurn, Annastiina Silvennoinen, and Timo Teräsvirta. A smooth transition logit model of the effects of deregulation in the electricity market. Journal of Applied Econometrics, 31(4): 707–733, June–July 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hsiao:2007:EIW

[HSWW07] Cheng Hsiao, Yan Shen, Boqing Wang, and Greg Weeks. Evaluating the impacts of Washington state repeated job search services on the earnings of prime-age female TANF recipients. Journal of Applied Econometrics, 22(2):453–475, March 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hurd:2004:ESS

[HSZ04] Michael D. Hurd, James P. Smith, and Julie M. Zissimopoulos. The effects of subjective survival on retirement and social security claiming. Journal of Applied Econometrics, 19(6):761–775, 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Harvey:2021:CCA

[HT21] Andrew Harvey and Stephen Thiele. Cointegration and control: Assessing the impact of events using time series data. Journal of Applied Econometrics, 36(1):71–85, January–February 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Huang:2009:PEF

[Hua09] Yongfu Huang. The political economy of financial reform: are Abiad and Mody right? Journal of Applied Econometrics, 24 (7):1207–1213, November–December 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See comments [ZDE12].

Huber:2014:ICM

[Hub14] Martin Huber. Identifying causal mechanisms (primarily) based on inverse probability weighting. Journal of Applied Econometrics, 29(6):920–943, September–October 2014. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hughes:1988:SFC

[Hug88] Michael D. Hughes. A stochastic frontier cost function for residential child care provision. Journal of Applied Econometrics, 3(3):203–214, July–September 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hur:2017:MPA

[Hur17] Joonyoung Hur. Monetary policy and asset prices: A Markovswitching DSGE approach. Journal of Applied Econometrics, 32(5):965–982, August 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hussinger:2008:RSF

[Hus08] Katrin Hussinger. R&D and subsidies at the firm level: an application of parametric and semiparametric two-step selection models. Journal of Applied Econometrics, 23(6):729–747, September–October 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Humlum:2013:RYC

[HV13] Maria Knoth Humlum and Rune Majlund Vejlin. The responses of youth to a cash transfer conditional on schooling: a quasi–experimental study. Journal of Applied Econometrics, 28(4):628–649, June–July 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hoderlein:2018:EDW

[HV18] Stefan Hoderlein and Anne Vanhems. Estimating the distribution of welfare effects using quantiles. Journal of Applied Econometrics, 33(1):52–72, January–February 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hurd:2011:SME

[HVW11] Michael Hurd, Maarten Van Rooij, and Joachim Winter. Stock market expectations of Dutch households. Journal of Applied Econometrics, 26(3):416–436, April–May 2011. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hubrich:2010:FES

[HW10] Kirstin Hubrich and Kenneth D. West. Forecast evaluation of small nested model sets. Journal of Applied Econometrics, 25 (4):574–594, June–July 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Horrace:2022:NTT

[HW22] William C. Horrace and Yulong Wang. Nonparametric tests of tail behavior in stochastic frontier models. Journal of Applied Econometrics, 37(3):537–562, April–May 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hsiao:2019:PPS

[HZ19] Cheng Hsiao and Qiankun Zhou. Panel parametric, semiparametric, and nonparametric construction of counterfactuals. Journal of Applied Econometrics, 34(4):463–481, June– July 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

ilKim:2019:RAE

[iKLS19] Kyoo il Kim, Yao Luo, and Yingjun Su. A robust approach to estimating production functions: Replication of the ACF procedure. Journal of Applied Econometrics, 34(4):612–619, June–July 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ilbas:2012:RPU

[Ilb12] Pelin Ilbas. Revealing the preferences of the US Federal reserve. Journal of Applied Econometrics, 27(3):440–473, April– May 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Isiklar:2006:HQD

[ILL06] Gultekin Isiklar, Kajal Lahiri, and Prakash Loungani. How quickly do forecasters incorporate news? Evidence from crosscountry surveys. Journal of Applied Econometrics, 21(6):703– 725, September–October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ivaldi:1996:CFT

[ILOS96] Marc Ivaldi, Norbert Ladoux, Hervé Ossard, and Michel Simioni. Comparing Fourier and translog specifications of multiproduct technology: Evidence from an incomplete panel of French farmers. Journal of Applied Econometrics, 11(6):649– 667, November 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Imbens:1997:BRF

[Imb97] Guido W. Imbens. Book review: The Foundations of Econometric Analysis, David F. Hendry and Mary S. Morgan, Cambridge University Press, Cambridge, ISBN 0-521-38043-X, 558pp, £40 hb. Journal of Applied Econometrics, 12(1):91– 94, January 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ichino:2008:THJ

[IMN08] Andrea Ichino, Fabrizia Mealli, and Tommaso Nannicini. From temporary help jobs to permanent employment: what can we learn from matching estimators and their sensitivity? Journal of Applied Econometrics, 23(3):305–327, April 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Iglesias:2012:IIV

[IP12] Emma M. Iglesias and Garry D. A. Phillips. Improved instrumental variables estimation of simultaneous equations under conditionally heteroskedastic disturbances. Journal of Applied Econometrics, 27(3):474–499, April–May 2012. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Iyer:2008:EET

[IRT08] Krishna G. Iyer, Alicia N. Rambaldi, and Kam Ki Tang. Efficiency externalities of trade and alternative forms of foreign investment in OECD countries. Journal of Applied Econometrics, 23(6):749–766, September–October 2008. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Irvine:2002:SEM

[IS02] Ian J. Irvine and William A. Sims. A simple and efficient method for estimating the magnitude and precision of welfare changes: comment. Journal of Applied Econometrics, 17(1): 81–83, January–February 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Isacsson:2004:EER

[Isa04] Gunnar Isacsson. Estimating the economic return to educational levels using data on twins. Journal of Applied Econometrics, 19(1):99–119, January–February 2004. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Carrion-i-Silvestre:2023:TML

[iSG23] Josep Lluís Carrion i Silvestre and María Dolores Gadea. Testing for multiple level shifts with an integrated or stationary noise component. Journal of Applied Econometrics, 38(6):801– 819, September–October 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ilmakunnas:1994:ECC

[IT94] P. Ilmakunnas and H. Törmä. Energy crises and change of technology. Journal of Applied Econometrics, 9(3):305–320, July–September 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ivaldi:1992:SER

[Iva92] Marc Ivaldi. Survey evidence on the rationality of expectations. Journal of Applied Econometrics, 7(3):225–241, July– September 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ioannides:2003:NEH

[IZ03] Yannis M. Ioannides and Jeffrey E. Zabel. Neighbourhood effects and housing demand. Journal of Applied Econometrics, 18(5):563–584, September–October 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Juselius:2017:RER

[JA17] Katarina Juselius and Katrin Assenmacher. Real exchange rate persistence and the excess return puzzle: The case of Switzerland versus the US. Journal of Applied Econometrics, 32(6): 1145–1155, September–October 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jaggia:1991:STB

[Jag91] Sanjiv Jaggia. Specification tests based on the heterogeneous generalized gamma model of duration: With an application to Kennan’s strike data. Journal of Applied Econometrics, 6(2): 169–180, April–June 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jales:2018:EEM

[Jal18] Hugo Jales. Estimating the effects of the minimum wage in a developing country: A density discontinuity design approach. Journal of Applied Econometrics, 33(1):29–51, January–February 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

James:2017:MAG

[Jam17] Jonathan James. MM algorithm for general mixed multinomial logit models. Journal of Applied Econometrics, 32(4):841–857, June–July 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jarocinski:2010:RMP

[Jar10] Marek Jarociński. Responses to monetary policy shocks in the east and the west of Europe: a comparison. Journal of Applied Econometrics, 25(5):833–868, August 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jacobson:2001:MPA

[JJVW01] Tor Jacobson, Per Jansson, Anders Vredin, and Anders Warne. Monetary policy analysis and inflation targeting in a small open economy: a VAR approach*. Journal of Applied Econometrics, 16(4):487–520, July–August 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jaeger:1990:SAM

[JK90] Albert Jaeger and Robert M. Kunst. Seasonal adjustment and measuring persistence in output. Journal of Applied Econometrics, 5(1):47–58, January–March 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jarrow:2021:IFB

[JK21] Robert A. Jarrow and Simon S. Kwok. Inferring financial bubbles from option data. Journal of Applied Econometrics, 36 (7):1013–1046, November–December 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jochmann:2013:SSV

[JKLGS13] Markus Jochmann, Gary Koop, Roberto Leon-Gonzalez, and Rodney W. Strachan. Stochastic search variable selection in vector error correction models with an application to a model of the UK macroeconomy. Journal of Applied Econometrics, 28 (1):62–81, January–February 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jones:2022:ISF

[JKR22] Callum Jones, Mariano Kulish, and Daniel M. Rees. International spillovers of forward guidance shocks. Journal of Applied Econometrics, 37(1):131–160, January–February 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jungbacker:2014:SDF

[JKvdW14] Borus Jungbacker, Siem Jan Koopman, and Michel van der Wel. Smooth dynamic factor analysis with application to the us term structure of interest rates. Journal of Applied Econometrics, 29(1):65–90, January–February 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jones:2003:IHC

[JL03] Andrew M. Jones and José M. Labeaga. Individual heterogeneity and censoring in panel data estimates of tobacco expenditure. Journal of Applied Econometrics, 18(2):157–177, March–April 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ji:2008:MRP

[JL08] Lu Ji and Tong Li. Multi-round procurement auctions with secret reserve prices: theory and evidence. Journal of Applied Econometrics, 23(7):897–923, November–December 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Janssens:2024:SSU

[JL24] Eva F. Janssens and Robin L. Lumsdaine. Sectoral slowdowns in the United Kingdom: Evidence from transmission probabilities and economic linkages. Journal of Applied Econometrics, 39(1):22–40, January–February 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:2005:AJT

[jLjL05] Myoung jae Lee and Sang jun Lee. Analysis of job-training effects on Korean women. Journal of Applied Econometrics, 20 (4):549–562, May–June 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jones:2014:ABT

[JLR14] Andrew M. Jones, James Lomas, and Nigel Rice. Applying beta–type size distributions to healthcare cost regressions. Journal of Applied Econometrics, 29(4):649–670, June–July 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Johnson:1998:GDT

[JM98] David Johnson and Robert McClelland. A general dependence test and applications. Journal of Applied Econometrics, 13(6): 627–644, November–December 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jorda:2010:PFE

[JM10] Òscar Jordà and Massimiliano Marcellino. Path forecast evaluation. Journal of Applied Econometrics, 25(4):635–662, June– July 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Johar:2014:DCI

[JM14] Meliyanni Johar and Shiko Maruyama. Does coresidence improve an elderly parent’s health? Journal of Applied Econometrics, 29(6):965–983, September–October 2014. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jimenez-Martin:2007:ELC

[JMM07] Sergi Jiménez-Martín and Alfonso R. Sánchez Martín. An evaluation of the life cycle effects of minimum pensions on retirement behavior. Journal of Applied Econometrics, 22(5):923– 950, August 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jiang:2018:TGA

[JMPW18] Lei Jiang, Esfandiar Maasoumi, Jiening Pan, and Ke Wu. A test of general asymmetric dependence. Journal of Applied Econometrics, 33(7):1026–1043, November–December 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jore:2010:CFD

[JMV10] Anne Sofie Jore, James Mitchell, and Shaun P. Vahey. Combining forecast densities from VARs with uncertain instabilities. Journal of Applied Econometrics, 25(4):621–634, June– July 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jin:2019:BPS

[JMY19] Xin Jin, John M. Maheu, and Qiao Yang. Bayesian parametric and semiparametric factor models for large realized covariance matrices. Journal of Applied Econometrics, 34(5):641–660, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jin:2020:TOS

[JO20] Yanchun Jin and Ryo Okui. Testing for overconfidence statistically: a moment inequality approach. Journal of Applied Econometrics, 35(7):879–892, November–December 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jochmans:2020:TCE

[Joc20] Koen Jochmans. Testing for correlation in error-component models. Journal of Applied Econometrics, 35(7):860–878, November–December 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jochmans:2023:TRA

[Joc23] Koen Jochmans. Testing random assignment to peer groups. Journal of Applied Econometrics, 38(3):321–333, April–May 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jones:1989:DHM

[Jon89] Andrew M. Jones. A double-hurdle model of cigarette consumption. Journal of Applied Econometrics, 4(1):23–39, January–March 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Justiniano:2010:MPU

[JP10] Alejandro Justiniano and Bruce Preston. Monetary policy and uncertainty in an empirical small open-economy model. Journal of Applied Econometrics, 25(1):93–128, January–February 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Juvenal:2015:SOM

[JP15] Luciana Juvenal and Ivan Petrella. Speculation in the oil market. Journal of Applied Econometrics, 30(4):621–649, June– July 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Johnston:2014:CMH

[JPPS14] David Johnston, Carol Propper, Stephen Pudney, and Michael Shields. Child mental health and educational attainment: multiple observers and the measurement error problem. Journal of Applied Econometrics, 29(6):880–900, September–October 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jalan:2002:GPT

[JR02] Jyotsna Jalan and Martin Ravallion. Geographic poverty traps? A micro model of consumption growth in rural China. Journal of Applied Econometrics, 17(4):329–346, July–August 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jenkins:2021:MEE

[JRA21] Stephen P. Jenkins and Fernando Rios-Avila. Measurement error in earnings data: Replication of Meijer, Rohwedder, and Wansbeek’s mixture model approach to combining survey and register data. Journal of Applied Econometrics, 36(4):474–483, June–July 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Joskow:1987:PCB

[JS87] Paul L. Joskow and Richard Schmalensee. The performance of coal-burning electric generating units in the United States: 1960–1980. Journal of Applied Econometrics, 2(2):85–109, April 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jeon:2004:REF

[JS04] Byung M. Jeon and Robin C. Sickles. The role of environmental factors in growth accounting. Journal of Applied Econometrics, 19(5):567–591, September–October 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Juarez:2010:NGD

[JS10] Miguel A. Juárez and Mark F. J. Steel. Non-Gaussian dynamic Bayesian modelling for panel data. Journal of Applied Econometrics, 25(7):1128–1154, November–December 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jones:2011:HDH

[JS11] Andrew M. Jones and Stefanie Schurer. How does heterogeneity shape the socioeconomic gradient in health satisfaction? Journal of Applied Econometrics, 26(4):549–579, June–July 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Juodis:2022:RAC

[Juo22] Arturas Juodis. A regularization approach to common correlated effects estimation. Journal of Applied Econometrics, 37 (4):788–810, June–July 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Juselius:2001:EIM

[Jus01] Katarina Juselius. European integration and monetary transmission mechanisms: the case of Italy. Journal of Applied Econometrics, 16(3):341–358, May–June 2001. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jochmans:2022:IVE

[JV22] Koen Jochmans and Vincenzo Verardi. Instrumental-variable estimation of exponential-regression models with two-way fixed effects with an application to gravity equations. Journal of Applied Econometrics, 37(6):1121–1137, September–October 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jacobs:2005:CSS

[JW05] Jan P. A. M. Jacobs and Kenneth F. Wallis. Comparing SVARs and SEMs: two models of the UK economy. Journal of Applied Econometrics, 20(2):209–228, 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Jones:2016:SSR

[JZ16] Lauren E. Jones and Nicolas R. Ziebarth. Successful scientific replication and extension of Levitt (2008): Child seats are still no safer than seat belts. Journal of Applied Econometrics, 31 (5):920–928, August 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kalwij:2007:PTW

[KA07] Adriaan S. Kalwij and Rob Alessie. Permanent and transitory wages of British men, 1975–2001: year, age and cohort effects. Journal of Applied Econometrics, 22(6):1063–1093, September–October 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kalirajan:1990:MEE

[Kal90] K. P. Kalirajan. On measuring economic efficiency. Journal of Applied Econometrics, 5(1):75–85, January–March 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kalouptsidi:2012:MSC

[Kal12] Myrto Kalouptsidi. From market shares to consumer types: Duality in differentiated product demand estimation. Journal of Applied Econometrics, 27(2):333–342, March 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kapteyn:1993:ICM

[Kap93] Arie Kapteyn. International conference on: The microeconometrics of dynamic decision making. Journal of Applied Econometrics, 8(4):425, October–December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kapetanios:2007:DFE

[Kap07] George Kapetanios. Dynamic factor extraction of crosssectional dependence in panel unit root tests. Journal of Applied Econometrics, 22(2):313–338, March 2007. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kasa:1998:ISD

[Kas98] Kenneth Kasa. Identifying the source of dynamics in disaggregated import data. Journal of Applied Econometrics, 13(3): 305–320, May–June 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kaufmann:2010:DFT

[Kau10] Sylvia Kaufmann. Dating and forecasting turning points by Bayesian clustering with dynamic structure: a suggestion with an application to Austrian data. Journal of Applied Econometrics, 25(2):309–344, March 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kaufmann:2020:DCA

[Kau20] Daniel Kaufmann. Is deflation costly after all? the perils of erroneous historical classifications. Journal of Applied Econometrics, 35(5):614–628, August 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Keijsers:2018:CLB

[KDK18] Bart Keijsers, Bart Diris, and Erik Kole. Cyclicality in losses on bank loans. Journal of Applied Econometrics, 33(4):533–552, June–July 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Keen:1986:ZEE

[Kee86] Michael Keen. Zero expenditures and the estimation of Engel curves. Journal of Applied Econometrics, 1(3):277–286, June 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kennet:1994:SMA

[Ken94] D. Mark Kennet. A structural model of aircraft engine maintenance. Journal of Applied Econometrics, 9(4):351–368, October–December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kerssenfischer:2019:PEM

[Ker19] Mark Kerssenfischer. The puzzling effects of monetary policy in VARs: Invalid identification or missing information? Journal of Applied Econometrics, 34(1):18–25, January–February 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kim:2015:PPP

[KFHO15] Hyeongwoo Kim, Ippei Fujiwara, Bruce E. Hansen, and Masao Ogaki. Purchasing power parity and the Taylor rule. Journal of Applied Econometrics, 30(6):874–903, September–October 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Klette:1996:ICS

[KG96] Tor Jakob Klette and Zvi Griliches. The inconsistency of common scale estimators when output prices are unobserved and endogenous. Journal of Applied Econometrics, 11(4):343–361, July 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kumbhakar:1995:LUE

[KH95] Subal C. Kumbhakar and Lennart Hjalmarsson. Labour-use efficiency in Swedish social insurance offices. Journal of Applied Econometrics, 10(1):33–47, January–March 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ke:2022:EIM

[KH22a] Xiao Ke and Cheng Hsiao. Economic impact of the most drastic lockdown during COVID-19 pandemic — the experience of Hubei, China. Journal of Applied Econometrics, 37(1):187–209, January–February 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ke:2022:FC

[KH22b] Xiao Ke and Cheng Hsiao. Featured cover. Journal of Applied Econometrics, 37(1):i, January–February 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ho:2016:SLS

[kHcL16] Tai kuang Ho and Cheng chung Lai. A silver lifeboat, not silver fetters: Why and how the silver standard insulated China from the 1929 Great Depression. Journal of Applied Econometrics, 31(2):403–419, March 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kaufmann:2014:DRE

[KHS14] Hendrik Kaufmann, Florian Heinen, and Philipp Sibbertsen. The dynamics of real exchange rates: a reconsideration. Journal of Applied Econometrics, 29(5):758–773, August 2014. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kidd:1992:RET

[Kid92] Michael P. Kidd. A review of the econometrics toolkit (ET). Journal of Applied Econometrics, 7(3):323–328, July– September 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kiefer:1988:ECJ

[Kie88] Nicholas M. Kiefer. Employment contracts, job search theory, and labour turnover: Preliminary empirical: Results. Journal of Applied Econometrics, 3(3):169–186, July–September 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kiefer:2011:DEC

[Kie11] Nicholas M. Kiefer. Default estimation, correlated defaults, and expert information. Journal of Applied Econometrics, 26 (2):173–192, March 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kilduff:1993:PNR

[Kil93] A. P. Kilduff. Pc-naive 6.01: A review. Journal of Applied Econometrics, 8(2):221–225, April–June 1993. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kilian:1999:ERM

[Kil99] Lutz Kilian. Exchange rates and monetary fundamentals: what do we learn from long-horizon regressions? Journal of Applied Econometrics, 14(5):491–510, September–October 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kim:1996:TFC

[KK96] Chang-Jin Kim and Myung-Jig Kim. Transient fads and the crash of ’87. Journal of Applied Econometrics, 11(1):41–58, January 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kadiyala:1997:NME

[KK97] K. Rao Kadiyala and Sune Karlsson. Numerical methods for estimation and inference in Bayesian VAR-models. Journal of Applied Econometrics, 12(2):99–132, March 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kerkhofs:2003:IEC

[KK03] Marcel Kerkhofs and Peter Kooreman. Identification and estimation of a class of household production models. Journal of Applied Econometrics, 18(3):337–369, May–June 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Knuppel:2022:FUD

[KK22] Malte Knüppel and Fabian Krüger. Forecast uncertainty, disagreement, and the linear pool. Journal of Applied Econometrics, 37(1):23–41, January–February 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kapetanios:2016:FBI

[KKM16] Georges Kapetanios, Lynda Khalaf, and Massimiliano Marcellino. Factor-based identification-robust interference in IV regressions. Journal of Applied Econometrics, 31(5):821–842, August 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kapteyn:1995:MDD

[KKR95] Arie Kapteyn, Nicholas Kieffer, and John Rust. The microeconometrics of dynamic decision making. Journal of Applied Econometrics, 10(S1):S1–S7, December 1995. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kliem:2016:LFR

[KKS16] Martin Kliem, Alexander Kriwoluzky, and Samad Sarferaz. On the low-frequency relationship between public deficits and inflation. Journal of Applied Econometrics, 31(3):566–583, April– May 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kinal:1993:ESE

[KL93] T. Kinal and K. Lahiri. On the estimation of simultaneousequations error-components models with an application to a model of developing country foreign trade. Journal of Applied Econometrics, 8(1):81–92, January–March 1993. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kuan:1995:FER

[KL95] Chung-Ming Kuan and Tung Liu. Forecasting exchange rates using feedforward and recurrent neural networks. Journal of Applied Econometrics, 10(4):347–364, October–December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kim:2001:QFI

[KL01] H. Youn Kim and Junsoo Lee. Quasi-fixed inputs and longrun equilibrium in production: a cointegration analysis. Journal of Applied Econometrics, 16(1):41–57, January–February 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koopman:2005:BDC

[KL05] Siem Jan Koopman and André Lucas. Business and default cycles for credit risk. Journal of Applied Econometrics, 20(2): 311–323, 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kan:2012:LWR

[KL12] Kamhon Kan and Myoung-Jae Lee. Lose weight for a raise only if overweight: Marginal integration for semi-linear panel models. Journal of Applied Econometrics, 27(4):666–685, June–July 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Krolikowski:2024:ALN

[KL24] Pawel M. Krolikowski and Kurt G. Lunsford. Advance layoff notices and aggregate job loss. Journal of Applied Econometrics, 39(3):462–480, April–May 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koopman:2018:DDC

[KLLO18] Siem Jan Koopman, Rutger Lit, André Lucas, and Anne Opschoor. Dynamic discrete copula models for high-frequency stock price changes. Journal of Applied Econometrics, 33(7): 966–985, November–December 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kejriwal:2024:EAP

[KLNT24] Mohitosh Kejriwal, Xiaoxiao Li, Linh Nguyen, and Evan Totty. The efficacy of ability proxies for estimating the returns to schooling: a factor model-based evaluation. Journal of Applied Econometrics, 39(1):3–21, January–February 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kloek:1992:BRM

[Klo92] Teun Kloek. Book review: Misspecification tests in econometrics. L. G. Godfrey, Cambridge University Press, Cambridge, 1988, pp. xii + 252. ISBN 0-521-26616-5. £35 hardcover. Journal of Applied Econometrics, 7(4):433–434, October–December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kloek:1993:BRS

[Klo93] Teun Kloek. Book review: Statistique et modèles econométriques, (statistics and econometric models), in 2 volumes, C. Gourieroux and A. Monfort, Economica, Paris, 1989, ISBN 2-7178-1667-4 and 2-7178-1668-2, paperback, FF 200 each, pp. 565 and 587. Journal of Applied Econometrics, 8(1):109–111, January– March 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kotchoni:2019:MFA

[KLS19] Rachidi Kotchoni, Maxime Leroux, and Dalibor Stevanovic. Macroeconomic forecast accuracy in a data rich environment. Journal of Applied Econometrics, 34(7):1050–1072, November– December 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kondo:2023:HTZ

[KLS23] Illenin O. Kondo, Logan T. Lewis, and Andrea Stella. Heavy tailed but not Zipf: Firm and establishment size in the United States. Journal of Applied Econometrics, 38(5):767–785, August 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kejriwal:2020:MSR

[KLT20] Mohitosh Kejriwal, Xiaoxiao Li, and Evan Totty. Multidimensional skills and the returns to schooling: Evidence from an interactive fixed-effects approach and a linked surveyadministrative data set. Journal of Applied Econometrics, 35 (5):548–566, August 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kim:1991:FPV

[KM91] In-Moo Kim and G. S. Maddala. Flat priors vs. ignorance priors in the analysis of the AR(1) model. Journal of Applied Econometrics, 6(4):375–380, October–December 1991. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kilian:2014:RIS

[KM14a] Lutz Kilian and Daniel P. Murphy. The role of inventories and speculative trading in the global market for crude oil. Journal of Applied Econometrics, 29(3):454–478, April–May 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kleibergen:2014:IIL

[KM14b] Frank Kleibergen and Sophocles Mavroeidis. Identification issues in limited–information Bayesian analysis of structural macroeconomic models. Journal of Applied Econometrics, 29 (7):1183–1209, November–December 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kliem:2022:EMP

[KMG22] Martin Kliem and Alexander Meyer-Gohde. (un)expected monetary policy shocks and term premia. Journal of Applied Econometrics, 37(3):477–499, April–May 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koop:2020:ROG

[KMMP20] Gary Koop, Stuart McIntyre, James Mitchell, and Aubrey Poon. Regional output growth in the United Kingdom: More timely and higher frequency estimates from 1970. Journal of Applied Econometrics, 35(2):176–197, March 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kim:2005:NPE

[KMP05] Chang-Jin Kim, James Morley, and Jeremy Piger. Nonlinearity and the permanent effects of recessions. Journal of Applied Econometrics, 20(2):291–309, 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kim:2008:BCA

[KMP08] Chang-Jin Kim, James Morley, and Jeremy Piger. Bayesian counterfactual analysis of the sources of the great moderation. Journal of Applied Econometrics, 23(2):173–191, March 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kuzin:2013:PVM

[KMS13] Vladimir Kuzin, Massimiliano Marcellino, and Christian Schumacher. Pooling versus model selection for nowcasting GDP with many predictors: empirical evidence for six industrialized countries. Journal of Applied Econometrics, 28(3):392– 411, April–May 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kapetanios:2019:LTP

[KMV19] George Kapetanios, Massimiliano Marcellino, and Fabrizio Venditti. Large time varying parameter VARs: A nonparametric approach. Journal of Applied Econometrics, 34(7):1027– 1049, November–December 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kunst:1990:CMS

[KN90] Robert Kunst and Klaus Neusser. Cointegration in a macroeconomic system. Journal of Applied Econometrics, 5(4):351–365, September–December 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kugler:1993:IRI

[KN93] P. Kugler and K. Neusser. International real interest rate equalization. A multivariate time-series approach. Journal of Applied Econometrics, 8(2):163–174, April–June 1993. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kim:2014:DLR

[KO14] Dukpa Kim and Tatsushi Oka. Divorce law reforms and divorce rates in the USA: an interactive fixed–effects approach. Journal of Applied Econometrics, 29(2):231–245, March 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koenker:1988:ATE

[Koe88] Roger Koenker. Asymptotic theory and econometric practice. Journal of Applied Econometrics, 3(2):139–147, April 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kohli:1989:CEW

[Koh89] U. Kohli. Consistent estimation when the left-hand variable is exogenous over part of the sample period. Journal of Applied Econometrics, 4(3):283–293, July–September 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kolenikov:2001:SRR

[Kol01] Stanislav Kolenikov. Software review: Review of Stata 7. Journal of Applied Econometrics, 16(5):637–646, September– October 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Komunjer:2007:APD

[Kom07] Ivana Komunjer. Asymmetric power distribution: Theory and applications to risk measurement. Journal of Applied Econometrics, 22(5):891–921, August 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koning:2001:CDL

[Kon01] Ruud H. Koning. A comparison of different LaTEX programs. Journal of Applied Econometrics, 16(1):81–92, January– February 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koning:2004:FAF

[Kon04] Ruud H. Koning. FinMetrics: analysis of financial data in S-PLUS. Journal of Applied Econometrics, 19(2):283–290, March–April 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koop:1992:ASM

[Koo92a] G. Koop. Aggregate shocks and macroeconomic fluctuations: A Bayesian approach. Journal of Applied Econometrics, 7(4): 395–411, October–December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koop:1992:OBU

[Koo92b] G. Koop. ‘objective’ Bayesian unit root tests. Journal of Applied Econometrics, 7(1):65–82, January–March 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koop:1992:RP

[Koo92c] Gary Koop. Review of PCBRAP. Journal of Applied Econometrics, 7(1):105–107, January–March 1992. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kooreman:1994:EEM

[Koo94] Peter Kooreman. Estimation of econometric models of some discrete games. Journal of Applied Econometrics, 9(3):255– 268, July–September 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koop:1999:BAC

[Koo99] Gary Koop. Bayesian analysis, computation and communication software. Journal of Applied Econometrics, 14(6):677–689, November–December 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koop:2013:FML

[Koo13] Gary M. Koop. Forecasting with medium and large Bayesian VARS. Journal of Applied Econometrics, 28(2):177–203, March 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kuehnle:2021:RGI

[KOO21] Daniel Kuehnle, Michael Oberfichtner, and Kerstin Ostermann. Revisiting gender identity and relative income within households: a cautionary tale on the potential pitfalls of density estimators. Journal of Applied Econometrics, 36(7):1065–1073, November–December 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kuehnle:2022:C

[KOO22] Daniel Kuehnle, Michael Oberfichtner, and Kerstin Ostermann. Correction. Journal of Applied Econometrics, 36(7):831–840, June–July 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kordas:2006:SBR

[Kor06] Gregory Kordas. Smoothed binary regression quantiles. Journal of Applied Econometrics, 21(3):387–407, April 2006. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Korobilis:2013:VFU

[Kor13] Dimitris Korobilis. Var forecasting using Bayesian variable selection. Journal of Applied Econometrics, 28(2):204–230, March 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kim:2005:NFM

[KOS05] Dong Heon Kim, Denise R. Osborn, and Marianne Sensier. Nonlinearity in the Fed’s monetary policy rule. Journal of Applied Econometrics, 20(5):621–639, July–August 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kapoor:2021:PFA

[KOW21] Sacha Kapoor, Matthijs Oosterveen, and Dinand Webbink. The price of forced attendance. Journal of Applied Econometrics, 36(2):209–227, March 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kodde:1987:PTN

[KP87] David A. Kodde and Franz C. Palm. A parametric test of the negativity of the substitution matrix. Journal of Applied Econometrics, 2(3):227–235, July 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koop:1994:ROL

[KP94] G. Koop and D. J. Poirier. Rank-ordered logit models: An empirical analysis of Ontario voter preferences. Journal of Applied Econometrics, 9(4):369–388, October–December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kreider:2008:IDS

[KP08] Brent Kreider and John Pepper. Inferring disability status from corrupt data. Journal of Applied Econometrics, 23(3):329–349, April 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kelejian:2016:ETS

[KP16] Harry H. Kelejian and Gianfranco Piras. An extension of the J-test to a spatial panel data framework. Journal of Applied Econometrics, 31(2):387–402, March 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kulish:2017:ESM

[KP17] Mariano Kulish and Adrian Pagan. Estimation and solution of models with expectations and structural changes. Journal of Applied Econometrics, 32(2):255–274, March 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kynigakis:2022:DMC

[KP22] Iason Kynigakis and Ekaterini Panopoulou. Does model complexity add value to asset allocation? Evidence from machine learning forecasting models. Journal of Applied Econometrics, 37(3):603–639, April–May 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kodde:1990:ALS

[KPP90] D. A. Kodde, F. C. Plam, and G. A. Pfann. Asymptotic least-squares estimation efficiency considerations and applications. Journal of Applied Econometrics, 5(3):229–243, July– September 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Karlstrom:2004:DPA

[KPS04] Anders Karlstrom, Marten Palme, and Ingemar Svensson. A dynamic programming approach to model the retirement behaviour of blue-collar workers in Sweden. Journal of Applied Econometrics, 19(6):795–807, 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Khwaja:2011:CTE

[KPST11] Ahmed Khwaja, Gabriel Picone, Martin Salm, and Justin G. Trogdon. A comparison of treatment effects estimators using a structural model of AMI treatment choices and severity of illness information from hospital charts. Journal of Applied Econometrics, 26(5):825–853, August 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koop:2005:SBI

[KPT05] Gary Koop, Dale J. Poirier, and Justin Tobias. Semiparametric Bayesian inference in multiple equation models. Journal of Applied Econometrics, 20(6):723–747, September–October 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Keil:1990:CAP

[KR90] M. W. Keil and W. Richardson. A comparison among partial adjustment, rational expectations and error correction estimates of the Canadian demand for money. Journal of Applied Econometrics, 5(3):273–291, July–September 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kraay:2012:IVR

[Kra12] Aart Kraay. Instrumental variables regressions with uncertain exclusion restrictions: a Bayesian approach. Journal of Applied Econometrics, 27(1):108–128, January–February 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Krause:2017:MPC

[Kra17] Melanie Krause. The Millennium peak in club convergence: A new look at distributional changes in the wealth of nations. Journal of Applied Econometrics, 32(3):621–642, April–May 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Krippner:2015:TFN

[Kri15] Leo Krippner. A theoretical foundation for the Nelson–Siegel class of yield curve models. Journal of Applied Econometrics, 30(1):97–118, January–February 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Korenok:2010:IEE

[KRS10] Oleg Korenok, Stanislav Radchenko, and Norman R. Swanson. International evidence on the efficacy of new-Keynesian models of inflation persistence. Journal of Applied Econometrics, 25 (1):31–54, January–February 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koning:1995:SFU

[KRV95] Pierre Koning, Geert Ridder, and Gerard J. Van Den Berg. Structural and frictional unemployment in an equilibrium search model with heterogeneous agents. Journal of Applied Econometrics, 10(S1):S133–S151, December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koop:1991:CCC

[KS91] Gary Koop and Mark F. J. Steel. A comment on: ‘To criticize the critics: An objective Bayesian analysis of stochastic trends’, By Peter C. B. Phillips. Journal of Applied Econometrics, 6(4): 365–370, October–December 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [Phi91b].

Kamhofer:2016:RZR

[KS16] Daniel A. Kamhöfer and Hendrik Schmitz. Reanalyzing zero returns to education in Germany. Journal of Applied Econometrics, 31(5):912–919, August 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kaufmann:2017:IRI

[KS17] Sylvia Kaufmann and Christian Schumacher. Identifying relevant and irrelevant variables in sparse factor models. Journal of Applied Econometrics, 32(6):1123–1144, September–October 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kripfganz:2019:ELD

[KS19] Sebastian Kripfganz and Claudia Schwarz. Estimation of linear dynamic panel data models with time-invariant regressors. Journal of Applied Econometrics, 34(4):526–546, June–July 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koedijk:1998:ETZ

[KSD98] Kees G. Koedijk, Philip A. Stork, and Casper G. De Vries. An EMS target zone model in discrete time. Journal of Applied Econometrics, 13(1):31–48, January–February 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Klein:2023:DDT

[KSN23] Nadja Klein, Michael Stanley Smith, and David J. Nott. Deep distributional time series models and the probabilistic forecasting of intraday electricity prices. Journal of Applied Econometrics, 38(4):493–511, June–July 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kenkel:2001:EPA

[KT01] Donald S. Kenkel and Joseph V. Terza. The effect of physician advice on alcohol consumption: count regression with an endogenous treatment effect. Journal of Applied Econometrics, 16(2):165–184, March–April 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koop:2004:LAH

[KT04] Gary Koop and Justin L. Tobias. Learning about heterogeneity in returns to schooling. Journal of Applied Econometrics, 19 (7):827–849, November–December 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kline:2008:WBB

[KT08] Brendan Kline and Justin L. Tobias. The wages of BMI: Bayesian analysis of a skewed treatment-response model with nonparametric endogeneity. Journal of Applied Econometrics, 23(6):767–793, September–October 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kumbhakar:2011:SES

[KT11] Subal C. Kumbhakar and Efthymios G. Tsionas. Stochastic error specification in primal and dual production systems. Journal of Applied Econometrics, 26(2):270–297, March 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kascha:2015:SIS

[KT15] Christian Kascha and Carsten Trenkler. Simple identification and specification of cointegrated VARMA models. Journal of Applied Econometrics, 30(4):675–702, June–July 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kalamara:2022:MTC

[KTR^{+}22] Eleni Kalamara, Arthur Turrell, Chris Redl, George Kapetanios, and Sujit Kapadia. Making text count: Economic forecasting using newspaper text. Journal of Applied Econometrics, 37 (5):896–919, August 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koopman:2002:SVM

[KU02] Siem Jan Koopman and Eugenie Hol Uspensky. The stochastic volatility in mean model: empirical evidence from international stock markets. Journal of Applied Econometrics, 17(6): 667–689, December 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kreitmeir:2024:DDP

[KÜ24] David Kreitmeir and Thomas Überfuhr. Disease and development-the predicted mortality instrument revisited. Journal of Applied Econometrics, 39(2):327–337, March 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kuklys:2002:BRS

[Kuk02] Wiebke Kuklys. Book review: Stated choice methods: analysis and application, Jordan J. Louviere, David A. Hensher and Joffre D. Swait, Cambridge University Press, ISBN: 0-521-78830-7. Journal of Applied Econometrics, 17(6):701–704, December 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kulaksizoglu:2015:LOC

[Kul15] Tamer Kulaksizoglu. Lag order and critical values of the augmented Dickey–Fuller test: A replication. Journal of Applied Econometrics, 30(6):1010, September–October 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kumar:2012:NEI

[Kum12] Anil Kumar. Nonparametric estimation of the impact of taxes on female labor supply. Journal of Applied Econometrics, 27 (3):415–439, April–May 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kuo:2012:ECU

[Kuo12] Tzu-Chun Kuo. Evaluating Californian under-age drunk driving laws: endogenous policy lags. Journal of Applied Econometrics, 27(7):1100–1115, November–December 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kusters:1991:BRE

[Kus91] Ulrich Kusters. Book review: Elements of statistical computingnumerical computation, R. A. Thisted. Chapman and Hall, New York and London, 1988. ISBN 0-412-01371-1, cloth £28.50, pp. xx + 427. Journal of Applied Econometrics, 6(2): 220–224, April–June 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Karabiyik:2019:CEF

[KUW19] Hande Karabiyik, Jean-Pierre Urbain, and Joakim Westerlund. CCE estimation of factor-augmented regression models with more factors than observables. Journal of Applied Econometrics, 34(2):268–284, March 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Klein:2009:SMB

[KV09] Roger Klein and Francis Vella. A semiparametric model for binary response and continuous outcomes under index heteroscedasticity. Journal of Applied Econometrics, 24(5):735– 762, August 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kelchtermans:2010:PSD

[KV10] Stijn Kelchtermans and Frank Verboven. Participation and study decisions in a public system of higher education. Journal of Applied Econometrics, 25(3):355–391, April–May 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kleinjans:2014:RFP

[KV14] Kristin J. Kleinjans and Arthur Van Soest. Rounding, focal point answers and nonresponse to subjective probability questions. Journal of Applied Econometrics, 29(4):567–585, June– July 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kleibergen:1993:NSG

[KvD93] F. Kleibergen and H. K. van Dijk. Non-stationarity in GARCH models: A Bayesian analysis. Journal of Applied Econometrics, 8(S1):S41–S61, December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kleijn:2006:BMA

[KvD06] Richard Kleijn and Herman K. van Dijk. Bayes model averaging of cyclical decompositions in economic time series. Journal of Applied Econometrics, 21(2):191–212, March 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kole:2017:HIF

[KvD17] Erik Kole and Dick van Dijk. How to identify and forecast bull and bear markets? Journal of Applied Econometrics, 32(1): 120–139, January–February 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kapteyn:1997:IPE

[KvdGVW97] Arie Kapteyn, Sara van de Geer, Huib Van de Stadt, and Tom Wansbeek. Interdependent preferences: an econometric analysis. Journal of Applied Econometrics, 12(6):665–686, November–December 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kastoryano:2022:DEJ

[KvdK22] Stephen Kastoryano and Bas van der Klaauw. Dynamic evaluation of job search assistance. Journal of Applied Econometrics, 37(2):227–241, March 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kennan:1989:SBM

[KW89] John Kennan and Robert Wilson. Strategic bargaining models and interpretation of strike data. Journal of Applied Econometrics, 4(S1):S87–S130, December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Keane:2013:CAM

[KW13] Michael Keane and Nada Wasi. Comparing alternative models of heterogeneity in consumer choice behavior. Journal of Applied Econometrics, 28(6):1018–1045, September–October 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Klossner:2014:EAV

[KW14] Stefan Klößner and Sven Wagner. Exploring all VAR orderings for calculating spillovers? Yes, we can! — a note on Diebold and Yilmaz (2009). Journal of Applied Econometrics, 29(1): 172–179, January–February 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kim:2019:HGP

[KW19] Jaeho Kim and Le Wang. Hidden group patterns in democracy developments: Bayesian inference for grouped heterogeneity. Journal of Applied Econometrics, 34(6):1016–1028, September–October 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Klein:2021:GSM

[KW21] Mathias Klein and Roland Winkler. The government spending multiplier at the zero lower bound: International evidence from historical data. Journal of Applied Econometrics, 36(6):744– 759, September–October 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kapetanios:2010:ETV

[KY10] George Kapetanios and Tony Yates. Estimating time variation in measurement error from data revisions: an application to backcasting and forecasting in dynamic models. Journal of Applied Econometrics, 25(5):869–893, August 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kilian:2002:QUA

[KZ02] Lutz Kilian and Tao Zha. Quantifying the uncertainty about the half-life of deviations from PPP. Journal of Applied Econometrics, 17(2):107–125, March–April 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Koenker:2009:RER

[KZ09] Roger Koenker and Achim Zeileis. On reproducible econometric research. Journal of Applied Econometrics, 24(5):833–847, August 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kilian:2020:DDU

[KZ20] Lutz Kilian and Xiaoqing Zhou. Does drawing down the US Strategic Petroleum Reserve help stabilize oil prices? Journal of Applied Econometrics, 35(6):673–691, September–October 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kilian:2022:OPG

[KZ22] Lutz Kilian and Xiaoqing Zhou. Oil prices, gasoline prices, and inflation expectations. Journal of Applied Econometrics, 37(5):867–881, August 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Klooster:2024:ORI

[KZ24] Jens Klooster and Mikhail Zhelonkin. Outlier robust inference in the instrumental variable model with applications to causal effects. Journal of Applied Econometrics, 39(1):86–106, January–February 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lakdawala:2019:DEM

[Lak19] Aeimit Lakdawala. Decomposing the effects of monetary policy using an external instruments SVAR. Journal of Applied Econometrics, 34(6):934–950, September–October 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lanne:2000:NUR

[Lan00] Markku Lanne. Near unit roots, cointegration, and the term structure of interest rates. Journal of Applied Econometrics, 15(5):513–529, September–October 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lanne:2006:NDI

[Lan06] Markku Lanne. Nonlinear dynamics of interest rate and inflation. Journal of Applied Econometrics, 21(8):1157–1168, December 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Larsen:1990:SRR

[Lar90] Hans Larsen. Software reviews. A review of RATS 3.0: Software for econometric modelling, forecasting, and simulation. Journal of Applied Econometrics, 5(1):87–91, January–March 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Liu:1995:MLE

[LB95] Shi-Miin Liu and B. Wade Brorsen. Maximum likelihood estimation of a GARCH-stable model. Journal of Applied Econometrics, 10(3):273–285, July–September 1995. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Luvsannyam:2019:FAV

[LB19] Davaajargal Luvsannyam and Khuslen Batmunkh. A factoraugmented vector autoregressive (FAVAR) approach for monetary policy: Replication of the empirical results in “measuring the effects of monetary policy”. Journal of Applied Econometrics, 34(5):820–821, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Laudicella:2022:DID

[LD22] Mauro Laudicella and Paolo Li Donni. The dynamic interdependence in the demand of primary and emergency secondary care: a hidden Markov approach. Journal of Applied Econometrics, 37(3):521–536, April–May 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Leamer:1989:PCR

[Lea89] Edward E. Leamer. Planning, criticism, and revision. Journal of Applied Econometrics, 4(S1):S5–S27, December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Leamer:1991:CCC

[Lea91] Edward E. Leamer. Comment on ‘to criticize the critics’. Journal of Applied Econometrics, 6(4):371–373, October–December 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [Phi91b].

Leamer:1997:RTS

[Lea97] Edward E. Leamer. Revisiting Tobin’s 1950 study of food expenditure. Journal of Applied Econometrics, 12(5):533–561, September–October 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lebaron:1992:FIU

[Leb92] B. Lebaron. Forecast improvements using a volatility index. Journal of Applied Econometrics, 7(S1):S137–S149, December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:1995:STE

[Lee95a] Byung-Joo Lee. Separability test for the electricity supply industry. Journal of Applied Econometrics, 10(1):49–60, January–March 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:1995:SPE

[Lee95b] Myoung-Jae Lee. Semi-parametric estimation of simultaneous equations with limited dependent variables: A case study of female labour supply. Journal of Applied Econometrics, 10(2): 187–200, April–June 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:2004:SCS

[Lee04] Myoung-Jae Lee. Selection correction and sensitivity analysis for ordered treatment effect on count response. Journal of Applied Econometrics, 19(3):323–337, May–June 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:2016:EOD

[Lee16] Soohyung Lee. Effect of online dating on assortative mating: Evidence from south Korea. Journal of Applied Econometrics, 31(6):1120–1139, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:2024:PII

[Lee24] Sungwon Lee. Partial identification and inference for conditional distributions of treatment effects. Journal of Applied Econometrics, 39(1):107–127, January–February 2024. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Leong:1998:BRP

[Leo98] Kenneth Leong. Book review: Periodicity and Stochastic Trends in Economic Time Series, Philip Hans Franses, Oxford University Press, Oxford, 1996, £35.00 (hardback), ISBN 0-19-877453-2 (hardback), pp. xii + 230. Journal of Applied Econometrics, 13(1):91–95, January–February 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Leung:2022:DRI

[Leu22] Michael P. Leung. Dependence-robust inference using resampled statistics. Journal of Applied Econometrics, 37(2):270– 285, March 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lewbel:2003:RRF

[Lew03] Arthur Lewbel. A rational rank four demand system. Journal of Applied Econometrics, 18(2):127–135, March–April 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Li:2009:BCU

[LfL09] Ji Li and Lung fei Lee. Binary choice under social interactions: an empirical study with and without subjective data on expectations. Journal of Applied Econometrics, 24(2):257–281, March 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lin:2023:RTM

[LFZC23] Jianhao Lin, Jiacheng Fan, Yifan Zhang, and Liangyuan Chen. Real-time macroeconomic projection using narrative central bank communication. Journal of Applied Econometrics, 38(2): 202–221, March 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lima:2023:EPP

[LG23] Luiz Renato Lima and Lucas Lúcio Godeiro. Equity-premium prediction: Attention is all you need. Journal of Applied Econometrics, 38(1):105–122, January–February 2023. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Liu:1992:UCE

[LGH92] T. Liu, C. W. J. Granger, and W. P. Heller. Using the correlation exponent to decide whether an economic series is chaotic. Journal of Applied Econometrics, 7(S1):S25–S39, December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lin:2019:ERI

[LGR19] Wei Lin and Gloria González-Rivera. Extreme returns and intensity of trading. Journal of Applied Econometrics, 34(7): 1121–1140, November–December 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lundberg:2017:IDS

[LHJC17] Alexander L. Lundberg, Kim P. Huynh, and David T. Jacho-Chávez. Income and democracy: A smooth varying coefficient redux. Journal of Applied Econometrics, 32(3):719–724, April– May 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lien:2017:SWB

[LHL17] Donald Lien, Yue Hu, and Long Liu. Subjective well-being and income: A re-examination of satiation using the regression kink model with an unknown threshold. Journal of Applied Econometrics, 32(2):463–469, March 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Li:1999:ERT

[Li99] Kai Li. Exchange rate target zone models: a Bayesian evaluation. Journal of Applied Econometrics, 14(5):461–490, September–October 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Li:2006:HSC

[Li06] Mingliang Li. High school completion and future youth unemployment: new evidence from high school and beyond. Journal of Applied Econometrics, 21(1):23–53, January–February 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Li:2012:TSV

[Li12] Shanjun Li. Traffic safety and vehicle choice: quantifying the effects of the ‘arms race’ on American roads. Journal of Applied Econometrics, 27(1):34–62, January–February 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lilien:2000:ESR

[Lil00] David M. Lilien. Econometric software reliability and nonlinear estimation in EViews: comment. Journal of Applied Econometrics, 15(1):107–110, January–February 2000. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lin:1992:AEF

[Lin92] Wen-Ling Lin. Alternative estimators for factor GARCH models — a Monte Carlo comparison. Journal of Applied Econometrics, 7(3):259–279, July–September 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lin:1995:JFD

[Lin95] Wen-Ling Lin. Japan’s financial deregulation and linkage of the Gensaki and Euroyen deposit markets. Journal of Applied Econometrics, 10(4):447–467, October–December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Liu:2009:POH

[Liu09] Qianqiu Liu. On portfolio optimization: How and when do we benefit from high-frequency data? Journal of Applied Econometrics, 24(4):560–582, June–July 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Liesenfeld:2000:SVM

[LJ00] Roman Liesenfeld and Robert C. Jung. Stochastic volatility models: conditional normality versus heavy-tailed distributions. Journal of Applied Econometrics, 15(2):137–160, March– April 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lancaster:2010:BQR

[LJ10] Tony Lancaster and Sung Jae Jun. Bayesian quantile regression methods. Journal of Applied Econometrics, 25(2):287–307, March 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Luginbuhl:2004:CEG

[LK04] Rob Luginbuhl and Siem Jan Koopman. Convergence in European GDP series: a multivariate common converging trendcycle decomposition. Journal of Applied Econometrics, 19(5): 611–636, September–October 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lindeboom:2009:HWE

[LK09] Maarten Lindeboom and Marcel Kerkhofs. Health and work of the elderly: subjective health measures, reporting errors and endogeneity in the relationship between health and work. Journal of Applied Econometrics, 24(6):1024–1046, September– October 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Li:2021:UCS

[LK21] Mengheng Li and Siem Jan Koopman. Unobserved components with stochastic volatility: Simulation-based estimation and signal extraction. Journal of Applied Econometrics, 36(5): 614–627, August 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Liu:2012:LPS

[LKL12] Xiaodong Liu, John H. Kagel, and Lung-Fei Lee. Learning from peers in signaling game experiments. Journal of Applied Econometrics, 27(7):1037–1058, November–December 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Li:1996:DTA

[LL96] C. W. Li and W. K. Li. On a double-threshold autoregressive heteroscedastic time series model. Journal of Applied Econometrics, 11(3):253–274, May 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lahiri:2006:MMP

[LL06] Kajal Lahiri and Fushang Liu. Modelling multi-period inflation uncertainty using a panel of density forecasts. Journal of Applied Econometrics, 21(8):1199–1219, December 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lanne:2016:NBV

[LL16] Markku Lanne and Jani Luoto. Noncausal Bayesian vector autoregression. Journal of Applied Econometrics, 31(7):1392– 1406, November–December 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Li:2013:HIE

[LLD13] Shanjun Li, Yanyan Liu, and Klaus Deininger. How important are endogenous peer effects in group lending? Estimating a static game of incomplete information. Journal of Applied Econometrics, 28(5):864–882, August 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lanne:2012:BMS

[LLL12] Markku Lanne, Arto Luoma, and Jani Luoto. Bayesian model selection and forecasting in noncausal autoregressive models. Journal of Applied Econometrics, 27(5):812–830, August 2012.

CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lahaye:2011:JCM

[LLN11] Jérôme Lahaye, Sébastien Laurent, and Christopher J. Neely. Jumps, cojumps and macro announcements. Journal of Applied Econometrics, 26(6):893–921, September–October 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:2023:DNN

[LLP23] Yoonseok Lee, Mary E. Lovely, and Hoang Pham. Dynamic and non-neutral productivity effects of foreign ownership: a nonparametric approach. Journal of Applied Econometrics, 38 (1):24–48, January–February 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Liu:2009:FRV

[LM09] Chun Liu and John M. Maheu. Forecasting realized volatility: a Bayesian model-averaging approach. Journal of Applied Econometrics, 24(5):709–733, August 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lima:2017:SRP

[LM17] Luiz Renato Lima and Fanning Meng. Out-of-sample return predictability: A quantile combination approach. Journal of Applied Econometrics, 32(4):877–895, June–July 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Liu:2018:IMS

[LM18] Jia Liu and John M. Maheu. Improving Markov switching models using realized variance. Journal of Applied Econometrics, 33(3):297–318, April–May 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lim:2005:PPH

[LMM05] G. C. Lim, G. M. Martin, and V. L. Martin. Parametric pricing of higher order moments in S&P500 options. Journal of Applied Econometrics, 20(3):377–404, March–April 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Loaiza-Maya:2021:FBP

[LMMF21] Ruben Loaiza-Maya, Gael M. Martin, and David T. Frazier. Focused Bayesian prediction. Journal of Applied Econometrics, 36(5):517–543, August 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lewis:2023:AGF

[LMPTR23] Daniel J. Lewis, Davide Melcangi, Laura Pilossoph, and Aidan Toner-Rodgers. Approximating grouped fixed effects estimation via fuzzy clustering regression. Journal of Applied Econometrics, 38(7):1077–1084, November–December 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Loaiza-Maya:2018:TSC

[LMSM18] Rubén Loaiza-Maya, Michael S. Smith, and Worapree Maneesoonthorn. Time series copulas for heteroskedastic data. Journal of Applied Econometrics, 33(3):332–354, April–May 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lewis:2022:MRA

[LMST22] Daniel J. Lewis, Karel Mertens, James H. Stock, and Mihir Trivedi. Measuring real activity using a weekly economic index. Journal of Applied Econometrics, 37(4):667–687, June– July 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lahiri:2016:FCR

[LMZ16] Kajal Lahiri, George Monokroussos, and Yongchen Zhao. Forecasting consumption: the role of consumer confidence in real time with many predictors. Journal of Applied Econometrics, 31(7):1254–1275, November–December 2016. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Liao:2023:IER

[LMZ23] Wenting Liao, Jun Ma, and Chengsi Zhang. Identifying exchange rate effects and spillovers of US monetary policy shocks in the presence of time-varying instrument relevance. Journal of Applied Econometrics, 38(7):989–1006, November–December 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:2007:EHP

[LN07] Jaejoon Lee and Charles R. Nelson. Expectation horizon and the Phillips curve: the solution to an empirical puzzle. Journal of Applied Econometrics, 22(1):161–178, January–February 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lutkepohl:2014:DDS

[LN14] Helmut Lütkepohl and Aleksei NetSunajev. Disentangling demand and supply shocks in the crude oil market: how to check sign restrictions in structural vars. Journal of Applied Econometrics, 29(3):479–496, April–May 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Leuven:2008:AAE

[LO08] Edwin Leuven and Hessel Oosterbeek. An alternative approach to estimate the wage returns to private-sector training. Journal of Applied Econometrics, 23(4):423–434, June 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:2018:GBE

[LO18] Soohyung Lee and Chiara Orsini. Girls and boys: Economic crisis, fertility, and birth outcomes. Journal of Applied Econometrics, 33(7):1044–1063, November–December 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lof:2019:EMR

[Lof19] Matthijs Lof. Expected market returns: SVIX, realized volatility, and the role of dividends. Journal of Applied Econometrics, 34(5):858–864, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lord:1991:PFC

[Lor91] M. J. Lord. Price formation in commodity markets. Journal of Applied Econometrics, 6(3):239–254, July–September 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Li:2013:CSV

[LOR13] Qi Li, Desheng Ouyang, and Jeffrey S. Racine. Categorical semiparametric varying-coefficient models. Journal of Applied Econometrics, 28(4):551–579, June–July 2013. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Louzis:2019:SSM

[Lou19] Dimitrios P. Louzis. Steady-state modeling and macroeconomic forecasting quality. Journal of Applied Econometrics, 34(2): 285–314, March 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:2017:DRU

[LOW17] Sokbae Lee, Ryo Okui, and Yoon-Jae Whang. Doubly robust uniform confidence band for the conditional average treatment effect function. Journal of Applied Econometrics, 32(7):1207– 1225, November–December 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lu:2008:ERA

[LP08] Jingfeng Lu and Isabelle Perrigne. Estimating risk aversion from ascending and sealed-bid auctions: the case of timber auction data. Journal of Applied Econometrics, 23(7):871–896, November–December 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lenza:2022:HEV

[LP22] Michele Lenza and Giorgio E. Primiceri. How to estimate a vector autoregression after March 2020. Journal of Applied Econometrics, 37(4):688–699, June–July 2022. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Laudati:2023:IES

[LP23] Dario Laudati and M. Hashem Pesaran. Identifying the effects of sanctions on the Iranian economy using newspaper coverage. Journal of Applied Econometrics, 38(3):271–294, April– May 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:1997:GCM

[LPS97] Kevin Lee, M. Hashem Pesaran, and Ron Smith. Growth and convergence in a multi-country empirical stochastic Solow model. Journal of Applied Econometrics, 12(4):357–392, July 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Li:2004:DDS

[LPT04] Mingliang Li, Dale J. Poirier, and Justin L. Tobias. Do dropouts suffer from dropping out? Estimation and prediction of outcome gains in generalized selection models. Journal of Applied Econometrics, 19(2):203–225, March–April 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:2022:OFU

[LPU22] Tae-Hwy Lee, Shahnaz Parsaeian, and Aman Ullah. Optimal forecast under structural breaks. Journal of Applied Econometrics, 37(5):965–987, August 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Li:2004:PRE

[LR04] Qi Li and Jeff Racine. Predictor relevance and extramarital affairs. Journal of Applied Econometrics, 19(4):533–535, July– August 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Li:2008:MOS

[LR08] Jinhu Li and Jeffrey S. Racine. Maxima: An open source computer algebra system. Journal of Applied Econometrics, 23 (4):515–523, June 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:2017:EPT

[LRS17] Nayoung Lee, Geert Ridder, and John Strauss. Estimation of poverty transition matrices with noisy data. Journal of Applied Econometrics, 32(1):37–55, January–February 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Laurent:2012:FAM

[LRV12] Sébastien Laurent, Jeroen V. K. Rombouts, and Francesco Violante. On the forecasting accuracy of multivariate GARCH models. Journal of Applied Econometrics, 27(6):934–955, September–October 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Liesenfeld:2017:LBI

[LRV17] Roman Liesenfeld, Jean-François Richard, and Jan Vogler. Likelihood-based inference and prediction in spatio-temporal panel count models for urban crimes. Journal of Applied Econometrics, 32(3):600–620, April–May 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Landesmann:1993:SSM

[LS93a] M. Landesmann and A. Snell. Structural shifts in the manufacturing export performance of OECD economies. Journal of Applied Econometrics, 8(2):149–162, April–June 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Laroque:1993:SBE

[LS93b] G. Laroque and B. Salanié. Simulation-based estimation of models with lagged latent variables. Journal of Applied Econometrics, 8(S1):S119–S133, December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Locke:1993:IDF

[LS93c] P. R. Locke and C. L. Sayers. Intra-day futures price volatility: Information effects and variance persistence. Journal of Applied Econometrics, 8(1):15–30, January–March 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Liu:1999:NLC

[LS99] Zhenjuan Liu and Thanasis Stengos. Non-linearities in cross-country growth regressions: a semiparametric approach. Journal of Applied Econometrics, 14(5):527–538, September– October 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Laroque:2002:LMI

[LS02] Guy Laroque and Bernard Salanié. Labour market institutions and employment in France. Journal of Applied Econometrics, 17(1):25–48, January–February 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ley:2009:CJG

[LS09a] Eduardo Ley and Mark F. J. Steel. Comments on ‘jointness of growth determinants’. Journal of Applied Econometrics, 24 (2):248–251, March 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [DW09a].

Ley:2009:EPA

[LS09b] Eduardo Ley and Mark F. J. Steel. On the effect of prior assumptions in Bayesian model averaging with applications to growth regression. Journal of Applied Econometrics, 24(4): 651–674, June–July 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lahiri:2010:MFU

[LS10a] Kajal Lahiri and Xuguang Sheng. Measuring forecast uncertainty by disagreement: The missing link. Journal of Applied Econometrics, 25(4):514–538, June–July 2010. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lubik:2010:LCS

[LS10b] Thomas A. Lubik and Paolo Surico. The Lucas critique and the stability of empirical models. Journal of Applied Econometrics, 25(1):177–194, January–February 2010. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Laroque:2014:IRF

[LS14a] Guy Laroque and Bernard Salanié. Identifying the response of fertility to financial incentives. Journal of Applied Econometrics, 29(2):314–332, March 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:2014:MTH

[LS14b] Soohyung Lee and Azeem M. Shaikh. Multiple testing and heterogeneous treatment effects: re–evaluating the effect of progresa on school enrollment. Journal of Applied Econometrics, 29(4):612–626, June–July 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Laffers:2021:ECD

[LS21] Lukás Lafférs and Bernhard Schmidpeter. Early child development and parents’ labor supply. Journal of Applied Econometrics, 36(2):190–208, March 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

List:2023:MTC

[LSV23] John A. List, Azeem M. Shaikh, and Atom Vayalinkal. Multiple testing with covariate adjustment in experimental economics.

Journal of Applied Econometrics, 38(6):920–939, September– October 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lucas:2017:MFS

[LSZ17] André Lucas, Bernd Schwaab, and Xin Zhang. Modeling financial sector joint tail risk in the Euro area. Journal of Applied Econometrics, 32(1):171–191, January–February 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lee:1991:TSI

[LT91] Tom K. Y. Lee and Y. K. Tse. Term structure of interest rates in the Singapore Asian dollar market. Journal of Applied Econometrics, 6(2):143–152, April–June 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lin:1996:CIC

[LT96] Jin-Lung Lin and Ruey S. Tsay. Co-integration constraint and forecasting: An empirical examination. Journal of Applied Econometrics, 11(5):519–538, September 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Li:2006:CAG

[LT06] Mingliang Li and Justin L. Tobias. Calculus attainment and grades received in intermediate economic theory. Journal of Applied Econometrics, 21(6):893–896, September–October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Linde:2018:SWU

[LT18] Jesper Lindé and Mathias Trabandt. Should we use linearized models to calculate fiscal multipliers? Journal of Applied Econometrics, 33(7):937–965, November–December 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lutkepohl:1999:ISL

[LTW99] Helmut Lütkepohl, Timo Teräsvirta, and Jürgen Wolters. Investigating stability and linearity of a German M1 money demand function. Journal of Applied Econometrics, 14(5):511– 525, September–October 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Laurent:2005:BGB

[LU05] Sébastien Laurent and Jean-Pierre Urbain. Bridging the gap between Ox and Gauss using OxGauss. Journal of Applied Econometrics, 20(1):131–139, January–February 2005. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Luciani:2015:MPH

[Luc15] Matteo Luciani. Monetary policy and the housing market: A structural factor analysis. Journal of Applied Econometrics, 30 (2):199–218, March 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Luintel:2000:RER

[Lui00] Kul B. Luintel. Real exchange rate behaviour: evidence from black markets. Journal of Applied Econometrics, 15(2):161– 185, March–April 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lewis:2019:MNR

[LVG19] Kurt F. Lewis and Francisco Vazquez-Grande. Measuring the natural rate of interest: A note on transitory shocks. Journal of Applied Econometrics, 34(3):425–436, April–May 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Loudon:2000:EAA

[LWY00] Geoffrey F. Loudon, Wing H. Watt, and Pradeep K. Yadav. An empirical analysis of alternative parametric ARCH models. Journal of Applied Econometrics, 15(2):117–136, March–April 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Li:2008:SBI

[LZ08] Tong Li and Xiaoyong Zheng. Semiparametric Bayesian inference for dynamic tobit panel data models with unobserved heterogeneity. Journal of Applied Econometrics, 23(6):699–728, September–October 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mackinnon:1991:A

[Mac91] James Mackinnon. Announcement. Journal of Applied Econometrics, 6(3):315, July–September 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

MacKinnon:1996:NDF

[Mac96] James G. MacKinnon. Numerical distribution functions for unit root and cointegration tests. Journal of Applied Econometrics, 11(6):601–618, November 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

MacKinnon:1999:LOS

[Mac99] James G. MacKinnon. The Linux operating system: Debian GNU/Linux. Journal of Applied Econometrics, 14(4):443–452, July–August 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Machado:2017:USH

[Mac17] Cecilia Machado. Unobserved selection heterogeneity and the gender wage gap. Journal of Applied Econometrics, 32(7):1348– 1366, November–December 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Madsen:1997:BRQ

[Mad97] Jakob B. Madsen. Book review: Quantitative Financial Economics: Stocks, Bonds and Foreign Exchange. Keith Cuthbertson. John Wiley, New York, 1996. ISBN 0-471-95359-8 (cased), 0-471-95360-1 pp470. Price: US$95 (cased), US$60 (paperback). Journal of Applied Econometrics, 12(4):455–457, July 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Madsen:1998:BRB

[Mad98] Jakob B. Madsen. Book review: Basic Econometrics, Damodar N. Gujarati, McGraw-Hill, New York, 1995, ISBN 0-07-025214-9 (paperback), pp. 838. Price US$74.95, £23.50 (paperback). Journal of Applied Econometrics, 13(2):209–212, March–April 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Madeira:2018:TRE

[Mad18] Carlos Madeira. Testing the rationality of expectations of qualitative outcomes. Journal of Applied Econometrics, 33(6):837– 852, September–October 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Malgrange:1988:BRS

[Mal88] Pierre Malgrange. Book review: Structural Sensitivity in Econometric Models, By E. Kuh, J. Neese and P. Hollinger. (John Wiley, New York, 1985, pp. ix + 324, ISBN 0-471-81930-1, cloth $49.50). Journal of Applied Econometrics, 3(3):249– 254, July–September 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mastrogiacomo:2004:RBD

[MAL04] Mauro Mastrogiacomo, Rob Alessie, and Maarten Lindeboom. Retirement behaviour of Dutch elderly households. Journal of Applied Econometrics, 19(6):777–793, 2004. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Marquez:1998:RPP

[Mar98] Jaime Marquez. Review of PcGive professional 9.0 for Windows. Journal of Applied Econometrics, 13(4):411–420, July– August 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Martin:2000:UDS

[Mar00] Gael M. Martin. US deficit sustainability: a new approach based on multiple endogenous breaks. Journal of Applied Econometrics, 15(1):83–105, January–February 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Martins:2001:PSE

[Mar01] Maria Fraga O. Martins. Parametric and semiparametric estimation of sample selection models: an empirical application to the female labour force in Portugal. Journal of Applied Econometrics, 16(1):23–39, January–February 2001. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

March:2002:RSW

[Mar02] Richard A. March. A review of SORITEC for Windows. Journal of Applied Econometrics, 17(1):85–90, January–February 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Moosa:2002:MTS

[MB02] I. A. Moosa and J. L. Baxter. Modelling the trend and seasonals within an AIDS model of the demand for alcoholic beverages in the United Kingdom. Journal of Applied Econometrics, 17(2): 95–106, March–April 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Moral-Benito:2016:GEP

[MB16] Enrique Moral-Benito. Growth empirics in panel data under model uncertainty and weak exogeneity. Journal of Applied Econometrics, 31(3):584–602, April–May 2016. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Musalem:2009:BER

[MBR09] Andrés Musalem, Eric T. Bradlow, and Jagmohan S. Raju. Bayesian estimation of random-coefficients choice models using aggregate data. Journal of Applied Econometrics, 24(3):490– 516, April–May 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Majumder:1990:DPI

[MC90] Amita Majumder and Satya Ranjan Chakravarty. Distribution of personal income: Development of a new model and its application to U.S. income data. Journal of Applied Econometrics, 5(2):189–196, April–June 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McAleer:1989:I

[McA89] Michael McAleer. Introduction. Journal of Applied Econometrics, 4(S1):S1–S3, December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mcaleer:1997:BR

[Mca97] Michael Mcaleer. Book review. Journal of Applied Econometrics, 12(2):191–193, March 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McCall:1994:TPH

[McC94] B. P. McCall. Testing the proportional hazards assumption in the presence of unmeasured heterogeneity. Journal of Applied Econometrics, 9(3):321–334, July–September 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mccullough:1997:RRV

[Mcc97] B. D. Mccullough. A review of RATS v4.2: benchmarking numerical accuracy. Journal of Applied Econometrics, 12(2): 181–190, March 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McCullough:1999:BRM

[McC99a] B. D. McCullough. Book review: Mathematical Statistics for Economics and Business, Ron C. Mittelhammer, Springer-Verlag, New York, 1996, xviii + 723 pp. $57.95, hardback only; solutions manual free to instructors who adopt the text. Journal of Applied Econometrics, 14(1):89–93, January–February 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McCullough:1999:ESR

[McC99b] B. D. McCullough. Econometric software reliability: EViews, LIMDEP, SHAZAM and TSP. Journal of Applied Econometrics, 14(2):191–202, March–April 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McCullough:2000:R

[McC00] B. D. McCullough. Reply. Journal of Applied Econometrics, 15(1):111, January–February 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McCullough:2002:BRP

[McC02] B. D. McCullough. Book review: Probability theory and statistical inference: econometric modelling with observational data, Aris Spanos, Cambridge University Press, 1999, pp. 815 + xxviii, price (pb) $49.95, (hb) $105.00. Journal of Applied Econometrics, 17(4):415–418, July–August 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McCullough:2006:RT

[McC06] B. D. McCullough. A review of TESTU01. Journal of Applied Econometrics, 21(5):677–682, July–August 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McKenzie:1993:MR

[McK93] C. R. McKenzie. Microfit 3.0: A review. Journal of Applied Econometrics, 8(4):413–419, October–December 1993. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McKenzie:1994:BRM

[McK94] Colin McKenzie. Book review: Modelling nonlinear economic relationships, C. W. J. Granger and T. Teräsvirta. Oxford University Press: 1993, ISBN 0-19-877320x, £14.95 paperback, pp. x + 187. Journal of Applied Econometrics, 9(4):479–480, October–December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McKenzie:1997:BRU

[McK97] Colin McKenzie. Book reviews: Undergraduate econometrics, R. C. Hill, W. E. Griffiths, and G. G. Judge, John Wiley, New York, 1997, A$57.95, ISBN 0-471-13993-9, pp. xiv + 366. Instructor’s resource guide to accompany undergraduate econometrics, W. E. Griffiths, R. C. Hill, and L. C. Marsh, John Wiley, New York, 1997, ISBN 0-471-19189-2, pp. 510 + floppy disk. Journal of Applied Econometrics, 12(6):745–749, November–December 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McKenzie:1998:M

[McK98a] C. R. McKenzie. Microfit 4.0. Journal of Applied Econometrics, 13(1):77–89, January–February 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McKenzie:1998:BRT

[McK98b] Colin McKenzie. Book review: Time Series and Dynamic Models, Christian Gourieroux and Alain Monfort, Cambridge University Press, Cambridge, 1997, ISBN 0-521-41146-7 (hardbound), ISBN 0-521-42308-2 (paperback), pp. xv + 668. Price £65.00 (hardbound), £22.95 (paperback). Journal of Applied Econometrics, 13(6):681–684, November–December 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Meddahi:2002:TCB

[Med02] Nour Meddahi. A theoretical comparison between integrated and realized volatility. Journal of Applied Econometrics, 17(5): 479–508, September–October 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Meen:1990:MRT

[Mee90] Geoffrey P. Meen. The measurement of rationing and the treatment of structural change in the UK mortgage market. Journal of Applied Econometrics, 5(2):167–187, April–June 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Melfi:1987:LR

[Mel87] C. A. Melfi. Limdep: A review. Journal of Applied Econometrics, 2(3):251–255, July 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Melnick:1988:PWI

[Mel88] Rafi Melnick. Prices, wages, and import prices in Israel: 1970– 1983. Journal of Applied Econometrics, 3(1):53–67, January 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mellano:1992:BRE

[Mel92] Mauro Mellano. Book review: Econometric modelling of agricultural commodity markets, D. Hallam, Routledge, London and New York, 1990. ISBN 0-415-00405-5, cloth, $35. Pp. iii + 191. Journal of Applied Econometrics, 7(2):223–224, April– June 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Maneesoonthorn:2017:ISE

[MFM17] Worapree Maneesoonthorn, Catherine S. Forbes, and Gael M. Martin. Inference on self-exciting jumps in prices and volatility using high-frequency measures. Journal of Applied Econometrics, 32(3):504–532, April–May 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Maheu:2008:LFS

[MG08] John M. Maheu and Stephen Gordon. Learning, forecasting and structural breaks. Journal of Applied Econometrics, 23(5):

553–583, August 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Moraga-Gonzalez:2013:SNE

[MGSW13] José Luis Moraga-González, Zsolt Sándor, and Matthijs R. Wildenbeest. Semi-nonparametric estimation of consumer search costs. Journal of Applied Econometrics, 28(7):1205– 1223, November–December 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Moser:2014:MPR

[MH14] Mathias Moser and Paul Hofmarcher. Model priors revisited: interaction terms in BMA growth applications. Journal of Applied Econometrics, 29(2):344–347, March 2014. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mhalla:2022:ECH

[MHL22] Linda Mhalla, Julien Hambuckers, and Marie Lambert. Extremal connectedness of hedge funds. Journal of Applied Econometrics, 37(5):988–1009, August 2022. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

MacKinnon:1999:NDF

[MHM99] James G. MacKinnon, Alfred A. Haug, and Leo Michelis. Numerical distribution functions of likelihood ratio tests for cointegration. Journal of Applied Econometrics, 14(5):563–577, September–October 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Milana:1989:BRA

[Mil89a] Carlo Milana. Book review: Applied demand analysis, by R. L. Thomas. (Longman, London: 1987, pp. viii + 104, ISBN 0-582-29723-0. £5.95 (paper)). Journal of Applied Econometrics, 4(1):100–102, January–March 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Milana:1989:BRC

[Mil89b] Carlo Milana. Book review: The Cambridge multisectoral dynamic model of the British economy, edited by T. Barker and W. Peterson, Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-33004-1, £35, pp. xxiv + 507. Journal of Applied Econometrics, 4(3):311–313, July–September 1989.

CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Milana:1991:BRF

[Mil91] Carlo Milana. Book review: Functional form and utility. A Review of Consumer Demand Theory, Arthur S. Goldberger. Westview Press, Boulder, Colorado, 1987, ISBN 0-8133-7489-8, paper $24.95, pp. xii + 121. Journal of Applied Econometrics, 6(3):327–329, July–September 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Milana:1992:BRD

[Mil92] Carlo Milana. Book review: Disaggregation in Econometric Modelling, edited by Terry Barker and M. Hashem Pesaran. Routledge, London, 1990, ISBN 0-415-00918-9, hardbound £45. pp. xiv + 348. Journal of Applied Econometrics, 7(4):431–432, October–December 1992. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Millo:2015:NRS

[Mil15] Giovanni Millo. Narrow replication of ‘A Spatio–Temporal Model of House Prices in the USA’ using R. Journal of Applied Econometrics, 30(4):703–704, June–July 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [BL14a].

Millo:2019:PRR

[Mil19] Giovanni Millo. Private returns to R&D in the presence of spillovers, revisited. Journal of Applied Econometrics, 34(1): 155–159, January–February 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Minicozzi:2003:ESI

[Min03] Alexandra L. Minicozzi. Estimation of sons’ intergenerational earnings mobility in the presence of censoring. Journal of Applied Econometrics, 18(3):291–314, May–June 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mishkin:1988:ITS

[Mis88] Frederic S. Mishkin. The information in the term structure: Some further results. Journal of Applied Econometrics, 3(4): 307–314, October–December 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mizrach:1992:MNN

[Miz92] B. Mizrach. Multivariate nearest-neighbour forecasts of EMS exchange rates. Journal of Applied Econometrics, 7(S1):S151– S163, December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Minesso:2022:DWH

[MKP22] Massimo Ferrari Minesso, Frederik Kurcz, and Maria Sole Pagliari. Do words hurt more than actions? The impact of trade tensions on financial markets. Journal of Applied Econometrics, 37(6):1138–1159, September–October 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Malikov:2016:CSA

[MKT16] Emir Malikov, Subal C. Kumbhakar, and Mike G. Tsionas. A cost system approach to the stochastic directional technology distance function with undesirable outputs: The case of US banks in 2001–2010. Journal of Applied Econometrics, 31 (7):1407–1429, November–December 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Martens:1998:TEC

[MKV98] Martin Martens, Paul Kofman, and Ton C. F. Vorst. A threshold error-correction model for intraday futures and index returns. Journal of Applied Econometrics, 13(3):245–263, May– June 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Montmarquette:1987:EMH

[MM87] Claude Montmarquette and Luc Monty. An empirical model of a household’s choice of activities. Journal of Applied Econometrics, 2(2):145–158, April 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mackinnon:1988:PER

[MM88a] James G. Mackinnon and Ross D. Milbourne. Are price equations really money demand equations on their heads? Journal of Applied Econometrics, 3(4):295–305, October–December 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McCurdy:1988:TMH

[MM88b] Thomas H. McCurdy and Ieuan G. Morgan. Testing the martingale hypothesis in deutsche mark futures with models specifying the form of heteroscedasticity. Journal of Applied Econometrics, 3(3):187–202, July–September 1988. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Montmarquette:1989:DSM

[MM89] Claude Montmarquette and Sophie Mahseredjian. Does school matter for educational achievement? A two-way nested-error components analysis. Journal of Applied Econometrics, 4(2): 181–193, April–June 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Magnus:1995:EAE

[MM95a] Jan R. Magnus and Mary S. Morgan. An experiment in applied econometrics — call for participants. Journal of Applied Econometrics, 10(2):213–216, April–June 1995. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McDonald:1995:DPI

[MM95b] James B. McDonald and Anand Mantrala. The distribution of personal income: Revisited. Journal of Applied Econometrics, 10(2):201–204, April–June 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Magnus:1997:DBD

[MM97a] Jan R. Magnus and Mary S. Morgan. The data: a brief description. Journal of Applied Econometrics, 12(5):651–661, September–October 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Magnus:1997:DE

[MM97b] Jan R. Magnus and Mary S. Morgan. Design of the experiment. Journal of Applied Econometrics, 12(5):459–465, September– October 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Magnus:1997:OE

[MM97c] Jan R. Magnus and Mary S. Morgan. Organization of the experiment. Journal of Applied Econometrics, 12(5):467–476, September–October 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Machado:2000:BCQ

[MM00] José A. F. Machado and José Mata. Box–Cox quantile regression and the distribution of firm sizes. Journal of Applied Econometrics, 15(3):253–274, May–June 2000. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Marcellino:2001:SSM

[MM01] Massimiliano Marcellino and Grayham E. Mizon. Small-system modelling of real wages, inflation, unemployment and output per capita in Italy 1970–1994. Journal of Applied Econometrics, 16(3):359–370, May–June 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mariano:2003:NCI

[MM03] Roberto S. Mariano and Yasutomo Murasawa. A new coincident index of business cycles based on monthly and quarterly series. Journal of Applied Econometrics, 18(4):427–443, July–August 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Maasoumi:2005:RIC

[MM05a] Esfandiar Maasoumi and Daniel L. Millimet. Robust inference concerning recent trends in US environmental quality. Journal of Applied Econometrics, 20(1):55–77, January–February 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Machado:2005:CDC

[MM05b] José A. F. Machado and José Mata. Counterfactual decomposition of changes in wage distributions using quantile regression. Journal of Applied Econometrics, 20(4):445–465, May– June 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mohtadi:2009:RCT

[MM09] Hamid Mohtadi and Antu Panini Murshid. Risk of catastrophic terrorism: an extreme value approach. Journal of Applied Econometrics, 24(4):537–559, June–July 2009. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Millimet:2017:DPD

[MM17] Daniel L. Millimet and Ian K. McDonough. Dynamic panel data models with irregular spacing: With an application to early childhood development. Journal of Applied Econometrics, 32(4):725–743, June–July 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McCracken:2019:EID

[MM19] Michael W. McCracken and Joseph T. McGillicuddy. An empirical investigation of direct and iterated multistep conditional forecasts. Journal of Applied Econometrics, 34(2):181–204, March 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Matas-Mir:2008:ESA

[MMOL08] Antonio Matas-Mir, Denise R. Osborn, and Marco J. Lombardi. The effect of seasonal adjustment on the properties of business cycle regimes. Journal of Applied Econometrics, 23 (2):257–278, March 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

MacKinnon:2014:NDF

[MN14] James G. MacKinnon and Morten Ørregaard Nielsen. Numerical distribution functions of fractional unit root and cointegration tests. Journal of Applied Econometrics, 29(1):161–171, January–February 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

MacKinnon:2023:FRJ

[MNW23a] James G. MacKinnon, Morten Ørregaard Nielsen, and Matthew D. Webb. Fast and reliable jackknife and bootstrap methods for cluster-robust inference. Journal of Applied Econometrics, 38(5):671–694, August 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

MacKinnon:2023:FC

[MNW23b] James G. MacKinnon, Morten Ørregaard Nielsen, and Matthew D. Webb. Featured cover. Journal of Applied Econometrics, 38(5), August 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mavromaras:2004:TLS

[MO04] Kostas G. Mavromaras and Chris D. Orme. Temporary layoffs and split population models. Journal of Applied Econometrics, 19(1):49–67, January–February 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Moench:2012:TSS

[Moe12] Emanuel Moench. Term structure surprises: the predictive content of curvature, level, and slope. Journal of Applied Econometrics, 27(4):574–602, June–July 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Moffitt:1993:BRM

[Mof93] Robert Moffitt. Book review: Modelling and empirical evaluation of labour supply behavior, Isolde Woittiez. Springer-Verlag, 1991, ISBN 3-540-54054-7, pp. 232. Journal of Applied Econometrics, 8(3):330–331, July–September 1993. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Moon:1989:MCC

[Moo89] Choon-Geol Moon. A Monte Carlo comparison of semiparametric tobit estimators. Journal of Applied Econometrics, 4(4): 361–382, October–December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Morrison:1994:CNM

[Mor94] C. J. Morrison. The cyclical nature of markups in Canadian manufacturing: A production theory approach. Journal of Applied Econometrics, 9(3):269–282, July–September 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mora:2008:NDM

[Mor08] Ricardo Mora. A nonparametric decomposition of the Mexican American average wage gap. Journal of Applied Econometrics, 23(4):463–485, June 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Morin:2023:WHW

[Mor23] Annaïg Morin. Workplace heterogeneity and wage inequality in Denmark. Journal of Applied Econometrics, 38(1):123–133, January–February 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mealli:1996:OPJ

[MP96] Fabrizia Mealli and Stephen Pudney. Occupational pensions and job mobility in Britain: Estimation of a random-effects competing risks model. Journal of Applied Econometrics, 11 (3):293–320, May 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Maynard:2001:ROE

[MP01] Alex Maynard and Peter C. B. Phillips. Rethinking an old empirical puzzle: econometric evidence on the forward discount anomaly. Journal of Applied Econometrics, 16(6):671–708, November–December 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Masanjala:2004:SMC

[MP04] Winford H. Masanjala and Chris Papageorgiou. The Solow model with CES technology: nonlinearities and parameter heterogeneity. Journal of Applied Econometrics, 19(2):171–201, March–April 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Moon:2007:EAN

[MP07] Hyungsik Roger Moon and Benoit Perron. An empirical analysis of nonstationarity in a panel of interest rates with factors. Journal of Applied Econometrics, 22(2):383–400, March 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Masanjala:2008:RLR

[MP08] Winford H. Masanjala and Chris Papageorgiou. Rough and lonely road to prosperity: a reexamination of the sources of growth in Africa using Bayesian model averaging. Journal of Applied Econometrics, 23(5):671–682, August 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See comments [Cua11, Pap11].

Miniaci:2010:MVE

[MP10] Raffaele Miniaci and Sergio Pastorello. Mean-variance econometric analysis of household portfolios. Journal of Applied Econometrics, 25(3):481–504, April–May 2010. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McCluskey:2004:MIC

[MQ04] Jill J. McCluskey and Kwamena K. Quagrainie. Measurement of industry conduct with a latent structure. Journal of Applied Econometrics, 19(7):887–897, November–December 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Maynard:2009:PIP

[MQ09] Alex Maynard and Jiaping Qiu. Public insurance and private savings: who is affected and by how much? Journal of Applied Econometrics, 24(2):282–308, March 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Melfi:1988:TEA

[MR88] C. A. Melfi and A. J. Rogers. A test for the existence of allocative inefficiency in firms. Journal of Applied Econometrics, 3 (1):69–80, January 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Monfort:1990:VMS

[MR90] A. Monfort and R. Rabemananjara. From a var model to a structural model, with an application to the wage-price spiral. Journal of Applied Econometrics, 5(3):203–227, July– September 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Marinucci:2001:FSI

[MR01] D. Marinucci and P. M. Robinson. Finite sample improvements in statistical inference with I(1) processes. Journal of Applied Econometrics, 16(3):431–444, May–June 2001. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Meijer:2006:MWE

[MR06] Erik Meijer and Jan Rouwendal. Measuring welfare effects in models with random coefficients. Journal of Applied Econometrics, 21(2):227–244, March 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Meredith:2009:TRE

[MR09] Evan Meredith and Jeffrey S. Racine. Towards reproducible econometric research: the Sweave framework. Journal of Applied Econometrics, 24(2):366–374, March 2009. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Millimet:2016:ETP

[MR16] Daniel L. Millimet and Jayjit Roy. Empirical tests of the pollution haven hypothesis when environmental regulation is endogenous. Journal of Applied Econometrics, 31(4):652–677, June–July 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mukherji:2016:EHD

[MRGB16] Arnab Mukherji, Satrajit Roychoudhury, Pulak Ghosh, and Sarah Brown. Estimating health demand for an aging population: A flexible and robust Bayesian joint model. Journal of Applied Econometrics, 31(6):1140–1158, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mrkaic:2001:SEP

[Mrk01] Mico Mrkaic. Scilab as an econometric programming system. Journal of Applied Econometrics, 16(4):553–559, July–August 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mroz:2012:SFE

[Mro12] Thomas A. Mroz. A simple, flexible estimator for count and other ordered discrete data. Journal of Applied Econometrics, 27(4):646–665, June–July 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mahmud:1987:EDF

[MRS87] S. F. Mahmud, A. L. Robb, and W. M. Scarth. On estimating dynamic factor demands. Journal of Applied Econometrics, 2 (1):69–75, January 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mehta:2010:IAP

[MRS10] Aditi Mehta, Marc Rysman, and Tim Simcoe. Identifying the age profile of patent citations: new estimates of knowledge diffusion. Journal of Applied Econometrics, 25(7):1179–1204, November–December 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Magnac:1995:AII

[MRV95] Thierry Magnac, Jean-Marc Robin, and Michael Visser. Analysing incomplete individual employment histories using indirect inference. Journal of Applied Econometrics, 10(S1): S153–S169, December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Martin:2009:DOM

[MRW09] Gael M. Martin, Andrew Reidy, and Jill Wright. Does the option market produce superior forecasts of noise-corrected volatility measures? Journal of Applied Econometrics, 24(1): 77–104, January–February 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ma:2015:SRP

[MRY15] Shujie Ma, Jeffrey S. Racine, and Lijian Yang. Spline regression in the presence of categorical predictors. Journal of Applied Econometrics, 30(5):705–717, August 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mercenier:1988:MST

[MS88] Jean Mercenier and Khalid Sekkat. Money stock targeting and money supply: An intertemporal optimization approach (with an application to Canada). Journal of Applied Econometrics, 3(3):215–228, July–September 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Machin:1990:UFP

[MS90] Stephen J. Machin and Mark B. Stewart. Unions and the financial performance of British private sector establishments. Journal of Applied Econometrics, 5(4):327–350, September– December 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Moon:1993:TDB

[MS93] C.-G Moon and J. G. Stotsky. Testing the differences between the determinants of Moody’s and Standard and Poor’s ratings: an application of smooth simulated maximum likelihood estimation. Journal of Applied Econometrics, 8(1):51– 69, January–March 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Miranda:1995:EMA

[MS95] Mario J. Miranda and Gary D. Schnitkey. An empirical model of asset replacement in dairy production. Journal of Applied Econometrics, 10(S1):S41–S55, December 1995. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mahieu:1998:EAS

[MS98] Ronald J. Mahieu and Peter C. Schotman. An empirical application of stochastic volatility models. Journal of Applied Econometrics, 13(4):333–360, July–August 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mixon:2006:TUE

[MS06] J. Wilson Mixon, Jr. and Ryan J. Smith. Teaching undergraduate econometrics with GRETL. Journal of Applied Econometrics, 21(7):1103–1107, November 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Metaxoglou:2007:ECE

[MS07] Konstantinos Metaxoglou and Aaron Smith. Efficiency of the California electricity reserves market. Journal of Applied Econometrics, 22(6):1127–1144, September–October 2007. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mumtaz:2009:TVY

[MS09] Haroon Mumtaz and Paolo Surico. Time-varying yield curve dynamics and monetary policy. Journal of Applied Econometrics, 24(6):895–913, September–October 2009. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mastromarco:2015:EFT

[MS15] Camilla Mastromarco and Léopold Simar. Effect of FDI and time on catching up: New insights from a conditional nonparametric frontier analysis. Journal of Applied Econometrics, 30 (5):826–847, August 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Metaxoglou:2017:SPC

[MS17] Konstantinos Metaxoglou and Aaron Smith. State prices of conditional quantiles: New evidence on time variation in the pricing kernel. Journal of Applied Econometrics, 32(1):192– 217, January–February 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Maheu:2018:EBA

[MS18a] John M. Maheu and Yong Song. An efficient Bayesian approach to multiple structural change in multivariate time series. Journal of Applied Econometrics, 33(2):251–270, March 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mumtaz:2018:PUA

[MS18b] Haroon Mumtaz and Paolo Surico. Policy uncertainty and aggregate fluctuations. Journal of Applied Econometrics, 33(3): 319–331, April–May 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Malikov:2019:UML

[MSH19] Emir Malikov, Yiguo Sun, and Diane Hite. (under)mining local residential property values: A semiparametric spatial quantile autoregression. Journal of Applied Econometrics, 34(1):82–109, January–February 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McIntosh:1989:QRA

[MSL89] James McIntosh, Fabio Schiantarelli, and William Low. A qualitative response analysis of UK firms’ employment and output decisions. Journal of Applied Econometrics, 4(3):251–264, July–September 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McIntosh:1986:DMR

[MSN86] James McIntosh, Stephen Satchell, and Anjum Nasim. Differential mortality in rural Bangladesh. Journal of Applied Econometrics, 1(4):345–353, October 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mumtaz:2013:TVD

[MSP13] Haroon Mumtaz and Laura Sunder-Plassmann. Time-varying dynamics of the real exchange rate: an empirical analysis. Journal of Applied Econometrics, 28(3):498–525, April–May 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mumtaz:2021:NEG

[MSP21] Haroon Mumtaz and Laura Sunder-Plassmann. Nonlinear effects of government spending shocks in the USA: Evidence from state-level data. Journal of Applied Econometrics, 36(1):86–97, January–February 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mamuneas:2006:EDR

[MSS06] Theofanis P. Mamuneas, Andreas Savvides, and Thanasis Stengos. Economic development and the return to human capital: a smooth coefficient semiparametric approach. Journal of Applied Econometrics, 21(1):111–132, January–February 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mastromarco:2016:MTE

[MSS16] Camilla Mastromarco, Laura Serlenga, and Yongcheol Shin. Modelling Technical Efficiency in Cross Sectionally Dependent Stochastic Frontier Panels. Journal of Applied Econometrics, 31(1):281–297, January–February 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McAndrew:2012:IRP

[MST12] Clare McAndrew, James L. Smith, and Rex Thompson. The impact of reserve prices on the perceived bias of expert appraisals of fine art. Journal of Applied Econometrics, 27(2): 235–252, March 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mitchell:2013:EAP

[MSW13] James Mitchell, Richard J. Smith, and Martin R. Weale. Efficient aggregation of panel qualitative survey data. Journal of Applied Econometrics, 28(4):580–603, June–July 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McFadden:2000:MMM

[MT00] Daniel McFadden and Kenneth Train. Mixed MNL models for discrete response. Journal of Applied Econometrics, 15(5): 447–470, September–October 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McKenzie:2003:LO

[MT03] C. R. McKenzie and Sumiko Takaoka. 2002: A LIMDEP odyssey. Journal of Applied Econometrics, 18(2):241–247, March–April 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McKenzie:2007:E

[MT07] C. R. McKenzie and Sumiko Takaoka. EViews 5.1. Journal of Applied Econometrics, 22(6):1145–1152, September–October 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Michaud:2011:FFE

[MT11] Pierre-Carl Michaud and Konstantinos Tatsiramos. Fertility and female employment dynamics in Europe: the effect of using alternative econometric modeling assumptions. Journal of Applied Econometrics, 26(4):641–668, June–July 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McKenzie:2012:E

[MT12] C. R. McKenzie and Sumiko Takaoka. EViews 7.2. Journal of Applied Econometrics, 27(7):1205–1210, November–December 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McKenzie:2013:M

[MT13a] C. R. McKenzie and Sumiko Takaoka. Microfit 5.0. Journal of Applied Econometrics, 28(1):170–175, January–February 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Millimet:2013:ETE

[MT13b] Daniel L. Millimet and Rusty Tchernis. Estimation of treatment effects without an exclusion restriction: with an application to the analysis of the school breakfast program. Journal of Applied Econometrics, 28(6):982–1017, September–October 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mirestean:2016:GDR

[MT16] Alin Mirestean and Charalambos G. Tsangarides. Growth Determinants Revisited Using Limited-Information Bayesian Model Averaging. Journal of Applied Econometrics, 31(1): 106–132, January–February 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Merlevede:2021:PEI

[MT21] Bruno Merlevede and Angelos Theodorakopoulos. Productivity effects of internationalisation through the domestic supply chain. Journal of Applied Econometrics, 36(6):808–832, September–October 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mountford:2009:WEF

[MU09] Andrew Mountford and Harald Uhlig. What are the effects of fiscal policy shocks? Journal of Applied Econometrics, 24(6): 960–992, September–October 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mullahy:1997:HEZ

[Mul97] John Mullahy. Heterogeneity, excess zeros, and the structure of count data models. Journal of Applied Econometrics, 12 (3):337–350, May 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Munkin:2022:CRM

[Mun22] Murat K. Munkin. Count Roy model with finite mixtures. Journal of Applied Econometrics, 37(6):1160–1181, September– October 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Murtazashvili:2012:AMI

[Mur12] Irina Murtazashvili. An alternative measure of intergenerational income mobility based on a random coefficient model. Journal of Applied Econometrics, 27(7):1161–1173, November– December 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Melenberg:1996:PSP

[MV96] Bertrand Melenberg and Arthur Van Soest. Parametric and semi-parametric modelling of vacation expenditures. Journal of Applied Econometrics, 11(1):59–76, January 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Maitra:2006:EHC

[MV06] Pushkar Maitra and Farshid Vahid. The effect of household characteristics on living standards in South Africa 1993–1998: a quantile regression analysis with sample attrition. Journal of Applied Econometrics, 21(7):999–1018, November 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mouchart:2007:BPM

[MV07] Michel Mouchart and Marie Vandresse. Bargaining powers and market segmentation in freight transport. Journal of Applied Econometrics, 22(7):1295–1313, December 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Muller:2020:CEM

[MvdKH20] Paul Muller, Bas van der Klaauw, and Arjan Heyma. Comparing econometric methods to empirically evaluate activation programs for job seekers. Journal of Applied Econometrics, 35 (5):526–547, August 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mellander:1992:STE

[MVW92] E. Mellander, A. Vredin, and A. Warne. Stochastic trends and economic fluctuations in a small open economy. Journal of Applied Econometrics, 7(4):369–394, October–December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mitchell:2011:EDF

[MW11] James Mitchell and Kenneth F. Wallis. Evaluating density forecasts: forecast combinations, model mixtures, calibration and sharpness. Journal of Applied Econometrics, 26(6):1023–1040, September–October 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mackinnon:2017:WBI

[MW17] James G. Mackinnon and Matthew D. Webb. Wild bootstrap inference for wildly different cluster sizes. Journal of Applied Econometrics, 32(2):233–254, March 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Morley:2020:I

[MW20] James Morley and Benjamin Wong. Estimating and accounting for the output gap with large Bayesian vector autoregression. Journal of Applied Econometrics, 35(1):1–18, January– February 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mitchell:2023:CDF

[MW23] James Mitchell and Martin Weale. Censored density forecasts: Production and evaluation. Journal of Applied Econometrics, 38(5):714–734, August 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Maasoumi:2024:IFA

[MWWW24] Esfandiar Maasoumi, Jianqiu Wang, Zhuo Wang, and Ke Wu. Identifying factors via automatic debiased machine learning. Journal of Applied Econometrics, 39(3):438–461, April–May 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

McCaig:2007:IWC

[MY07] Brian McCaig and Adonis Yatchew. International welfare comparisons and nonparametric testing of multivariate stochastic dominance. Journal of Applied Econometrics, 22(5):951–969, August 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Mills:1997:SIB

[MZ97] Jeffrey A. Mills and Sourushe Zandvakili. Statistical inference via bootstrapping for measures of inequality. Journal of Applied Econometrics, 12(2):133–150, March 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Malikov:2017:EDU

[MZK17] Emir Malikov, Shunan Zhao, and Subal C. Kumbhakar. Economies of diversification in the US credit union sector.

Journal of Applied Econometrics, 32(7):1329–1347, November– December 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Malikov:2020:EFL

[MZK20] Emir Malikov, Shunan Zhao, and Subal C. Kumbhakar. Estimation of firm-level productivity in the presence of exports: Evidence from China’s manufacturing. Journal of Applied Econometrics, 35(4):457–480, June–July 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Naguib:2022:DEM

[Nag22] Costanza Naguib. Did earnings mobility change after minimum wage introduction? Evidence from parametric and seminonparametric methods in Germany. Journal of Applied Econometrics, 37(7):1379–1402, November–December 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nakamura:1986:FDM

[Nak86] Shinichiro Nakamura. A flexible dynamic model of multiproduct technology for the West German economy. Journal of Applied Econometrics, 1(4):333–344, October 1986. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nakamura:1993:ACM

[Nak93] S. Nakamura. An adjustment cost model of long-term employment in Japan. Journal of Applied Econometrics, 8(2):175–194, April–June 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nakov:2010:JIV

[Nak10] Anton Nakov. Jackknife instrumental variables estimation: replication and extension of Angrist, Imbens and Krueger (1999). Journal of Applied Econometrics, 25(6):1063–1066, September–October 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [AIK99].

Nason:1994:TIL

[NC94] James M. Nason and Timothy Cogley. Testing the implications of long-run neutrality for monetary business cycle models. Journal of Applied Econometrics, 9(S1):S37–S70, December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Neild:1992:BRM

[Nei92] R. R. Neild. Book review: Macroeconomic policy in Britain 1974-1987, A. J. C. Britton. Cambridge University Press, 1991, pp. 365. Price: £30. Journal of Applied Econometrics, 7(3): 329–332, July–September 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Newey:1990:SEB

[New90] Whitney K. Newey. Semiparametric efficiency bounds. Journal of Applied Econometrics, 5(2):99–135, April–June 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ng:1995:RC

[Ng95a] Serena Ng. Review of Coint 2.0. Journal of Applied Econometrics, 10(2):205–210, April–June 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ng:1995:THD

[Ng95b] Serena Ng. Testing for homogeneity in demand systems when the regressors are nonstationary. Journal of Applied Econometrics, 10(2):147–163, April–June 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nyblom:2001:TAS

[NH01] Jukka Nyblom and Andrew Harvey. Testing against smooth stochastic trends. Journal of Applied Econometrics, 16(3): 415–429, May–June 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nibbering:2024:HDM

[Nib24] Didier Nibbering. A high-dimensional multinomial logit model. Journal of Applied Econometrics, 39(3):481–497, April–May 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nishiyama:2005:CEE

[Nis05] Shin-Ichi Nishiyama. The cross-Euler equation approach to intertemporal substitution in import demand. Journal of Applied Econometrics, 20(7):841–872, December 2005. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nunez:2021:OCT

[NO21] Héctor M. Núñez and Jesús Otero. A one covariate at a time, multiple testing approach to variable selection in highdimensional linear regression models: a replication in a narrow sense. Journal of Applied Econometrics, 36(6):833–841, September–October 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nonejad:2016:RRN

[Non16] Nima Nonejad. Replicating the results in ‘A New Model of Trend Inflation’ Using Particle Markov Chain Monte Carlo. Journal of Applied Econometrics, 31(7):1478–1483, November– December 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nelson:2018:DCM

[NPT18] Benjamin Nelson, Gabor Pinter, and Konstantinos Theodoridis. Do contractionary monetary policy shocks expand shadow banking? Journal of Applied Econometrics, 33(2):198–211, March 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nicoletti:2019:SSE

[NR19] Cheti Nicoletti and Birgitta Rabe. Sibling spillover effects in school achievement. Journal of Applied Econometrics, 34(4): 482–501, June–July 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ng:1990:XIW

[NS90] Pin T. Ng and Robin C. Sickles. ‘xplore’-ing the world of nonparametric analysis. Journal of Applied Econometrics, 5(3): 293–297, July–September 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Narendranathan:1993:HDB

[NS93] W. Narendranathan and M. B. Stewart. How does the benefit effect vary as unemployment spells lengthen? Journal of Applied Econometrics, 8(4):361–381, October–December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nason:2008:INK

[NS08] James M. Nason and Gregor W. Smith. Identifying the new Keynesian Phillips curve. Journal of Applied Econometrics, 23 (5):525–551, August 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nowak:2017:TAR

[NS17] Adam Nowak and Patrick Smith. Textual analysis in real estate. Journal of Applied Econometrics, 32(4):896–918, June– July 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nason:2021:MSE

[NS21] James M. Nason and Gregor W. Smith. Measuring the slowly evolving trend in US inflation with professional forecasts. Journal of Applied Econometrics, 36(1):1–17, January–February 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Normandin:1998:SRA

[NSA98] Michel Normandin and Pascal St-Amour. Substitution, risk aversion, taste shocks and equity premia. Journal of Applied Econometrics, 13(3):265–281, May–June 1998. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Noureldin:2012:MHF

[NSS12] Diaa Noureldin, Neil Shephard, and Kevin Sheppard. Multivariate high-frequency-based volatility (HEAVY) models. Journal of Applied Econometrics, 27(6):907–933, September– October 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Natsiopoulos:2022:ABT

[NT22] Kleanthis Natsiopoulos and Nickolaos G. Tzeremes. ARDL bounds test for cointegration: Replicating the pesaran et al. (2001) results for the UK earnings equation using R. Journal of Applied Econometrics, 37(5):1079–1090, August 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [PSS01].

Nguyen:2022:RCH

[NTK22] Trong-Nghia Nguyen, Minh-Ngoc Tran, and Robert Kohn. Recurrent conditional heteroskedasticity. Journal of Applied Econometrics, 37(5):1031–1054, August 2022. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nuesch:1991:BRO

[Nüe91] P. E. Nüesch. Book review: Order restricted statistical inference, T. Robertson, F. T. Wright, R. L. Dykstra Wiley Series in Probability and Mathematical Statistics, John Wiley and Sons, Chichester, 1988. ISBN 0-471-91787-7 cloth, £37.50, pp.

488. Journal of Applied Econometrics, 6(1):105–107, January– March 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nijman:1992:NPD

[NV92] Theo Nijman and Marno Verbeek. Nonresponse in panel data: The impact on estimates of a life cycle consumption function. Journal of Applied Econometrics, 7(3):243–257, July– September 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nyholm:2003:IPI

[Nyh03] Ken Nyholm. Inferring the private information content of trades: a regime-switching approach. Journal of Applied Econometrics, 18(4):457–470, July–August 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Nymoen:1991:SLM

[Nym91] R. Nymoen. A small linear model of wage- and price-inflation in the Norwegian economy. Journal of Applied Econometrics, 6(3):255–269, July–September 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Offstein:2007:EGM

[Off07] Norman Offstein. An extortionary guerrilla movement. Journal of Applied Econometrics, 22(6):995–1011, September–October 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Okunade:1993:PCS

[Oku93] A. A. Okunade. Production cost structure of us hospital pharmacies: Time-series, cross-sectional bed size evidence. Journal of Applied Econometrics, 8(3):277–294, July–September 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Olson:1998:CPS

[Ols98] Craig A. Olson. A comparison of parametric and semiparametric estimates of the effect of spousal health insurance coverage on weekly hours worked by wives. Journal of Applied Econometrics, 13(5):543–565, September–October 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ocal:2000:BCN

[ÖO00] Nadir Öcal and Denise R. Osborn. Business cycle nonlinearities in UK consumption and production. Journal of Applied Econometrics, 15(1):27–43, January–February 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Olea:2019:SCB

[OPM19] José Luis Montiel Olea and Mikkel Plagborg-Møller. Simultaneous confidence bands: Theory, implementation, and an application to SVARs. Journal of Applied Econometrics, 34(1):1–17, January–February 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Owyang:2022:CS

[OPS22] Michael T. Owyang, Jeremy Piger, and Daniel Soques. Contagious switching. Journal of Applied Econometrics, 37(2):415– 432, March 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

ODonnell:2001:EER

[ORD01] Christopher J. O’Donnell, Alicia N. Rambaldi, and Howard E. Doran. Estimating economic relationships subject to firmand time-varying equality and inequality constraints. Journal of Applied Econometrics, 16(6):709–726, November–December 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Onatski:2013:FAL

[ORM13] Alexei Onatski and Francisco Ruge-Murcia. Factor analysis of a large DSGE model. Journal of Applied Econometrics, 28(6): 903–928, September–October 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ormosi:2014:TIP

[Orm14] Peter L. Ormosi. A tip of the iceberg? The probability of catching cartels. Journal of Applied Econometrics, 29(4):549–566, June–July 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ornaghi:2006:AEM

[Orn06] Carmine Ornaghi. Assessing the effects of measurement errors on the estimation of production functions. Journal of Applied Econometrics, 21(6):879–891, September–October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Otrok:2002:EAP

[ORW02] Christopher Otrok, B. Ravikumar, and Charles H. Whiteman. Evaluating asset-pricing models using the Hansen– Jagannathan bound: a Monte Carlo investigation. Journal of Applied Econometrics, 17(2):149–174, March–April 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Osborn:1988:SHP

[Osb88] Denise R. Osborn. Seasonality and habit persistence in a life cycle model of consumption. Journal of Applied Econometrics, 3(4):255–266, October–December 1988. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Opitz:2018:MCR

[OSS18] Sebastian Opitz, Henry Seidel, and Alexander Szimayer. Measuring crisis risk using conditional copulas: An empirical analysis of the 2008 shipping crisis. Journal of Applied Econometrics, 33(2):271–289, March 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Opschoor:2017:CDF

[OvDvdW17] Anne Opschoor, Dick van Dijk, and Michel van der Wel. Combining density forecasts using focused scoring rules. Journal of Applied Econometrics, 32(7):1298–1313, November–December 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Onatski:2010:EPP

[OW10] Alexei Onatski and Noah Williams. Empirical and policy performance of a forward-looking monetary model. Journal of Applied Econometrics, 25(1):145–176, January–February 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Oxley:1996:BRC

[Oxl96] Les Oxley. Book review: Cointegration analysis in econometric modelling, R. I. D. Harris, Harvester, Wheatsheaf, London, 1995, ISBN 0-13-35582-4, pp. ix + 176, £15.95. Journal of Applied Econometrics, 11(2):211–215, March 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Paarsch:1994:LVR

[Paa94] H. J. Paarsch. Limdep, version 6.0: A review. Journal of Applied Econometrics, 9(1):91–98, January–March 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pacini:2019:TSL

[Pac19] David Pacini. The two-sample linear regression model with interval-censored covariates. Journal of Applied Econometrics, 34(1):66–81, January–February 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pagan:1994:ICE

[Pag94] Adrian Pagan. Introduction calibration and econometric research: An overview. Journal of Applied Econometrics, 9(S1): S1–S10, December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pakes:2000:ZG

[Pak00] Ariel Pakes. Zvi Griliches 1930–1999. Journal of Applied Econometrics, 15(5):443–444, September–October 2000. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pak:2018:PCO

[Pak18] Anton Pak. Predicting crude oil prices: Replication of the empirical results in “what do we learn from the price of crude oil?”. Journal of Applied Econometrics, 33(1):160–163, January–February 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Papageorgiou:2011:HUI

[Pap11] Chris Papageorgiou. How to use interaction terms in BMA: Reply to Crespo Cuaresma’s comment on Masanjala and Papageorgiou (2008). Journal of Applied Econometrics, 26(6): 1048–1050, September–October 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [MP08, Cua11].

Parke:1987:MMC

[Par87] William R. Parke. Macroeconometric model comparison and evaluation techniques: A practical appraisal. Journal of Applied Econometrics, 2(2):133–144, April 1987. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Park:2010:HDC

[Par10] Cheolbeom Park. How does changing age distribution impact stock prices? A nonparametric approach. Journal of Applied Econometrics, 25(7):1155–1178, November–December 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Park:2011:AR

[Par11a] Cheolbeom Park. Acknowledgement to referees. Journal of Applied Econometrics, 26(5):888–891, August 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Park:2011:HDC

[Par11b] Cheolbeom Park. How does changing age distribution impact stock prices? A nonparametric approach. Journal of Applied Econometrics, 26(5):886–887, August 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Patel:1990:PPP

[Pat90] Jayendu Patel. Purchasing power parity as a long-run relation. Journal of Applied Econometrics, 5(4):367–379, September– December 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Patton:2006:EMM

[Pat06] Andrew J. Patton. Estimation of multivariate models for time series of possibly different lengths. Journal of Applied Econometrics, 21(2):147–173, March 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pelagatti:2015:EFP

[PC15] Matteo Pelagatti and Emilio Colombo. On the empirical failure of purchasing power parity tests. Journal of Applied Econometrics, 30(6):904–923, September–October 2015. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pedroni:2007:SCB

[Ped07] Peter Pedroni. Social capital, barriers to production and capital shares: implications for the importance of parameter heterogeneity from a nonstationary panel approach. Journal of Applied Econometrics, 22(2):429–451, March 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Peersman:2005:WCE

[Pee05] Gert Peersman. What caused the early millennium slowdown? Evidence based on vector autoregressions. Journal of Applied Econometrics, 20(2):185–207, 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Prieto:2016:TVM

[PEM16] Esteban Prieto, Sandra Eickmeier, and Massimiliano Marcellino. Time variation in macro-financial linkages. Journal of Applied Econometrics, 31(7):1215–1233, November–December 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pendakur:2005:SEL

[Pen05] Krishna Pendakur. Semiparametric estimation of lifetime equivalence scales. Journal of Applied Econometrics, 20(4):

487–507, May–June 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pepper:2005:BAG

[Pep05] John V. Pepper. The Bias Against Guns: why almost everything you’ve heard about gun control is wrong. John Lott, Jr., Regnery Publishing, Inc. 2003, pp. 349. Journal of Applied Econometrics, 20(7):931–942, December 2005. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Persson:1991:BRD

[Per91] Karl Gunnar Persson. Book review: Domesday economy: A new approach to Anglo–Norman history, J. McDonald and G. D. Snooks. Clarendon Press, Oxford, 1986. ISBN 0-19-82854-8, cloth £27.50, pp. xv + 240. Journal of Applied Econometrics, 6(2):219–220, April–June 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Perez:1998:COB

[Per98] Stephen J. Perez. Causal ordering and ‘the bank lending channel’. Journal of Applied Econometrics, 13(6):613–626, November–December 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Perrigne:2010:BRH

[Per10] Isabelle Perrigne. Book review: Harry J. Paarsch and Han Hong, An Introduction to the Structural Econometrics of Auction Data (The MIT Press 2006). Journal of Applied Econometrics, 25(7):1215–1222, November–December 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:1986:ES

[Pes86] M. Hashem Pesaran. Editorial statement. Journal of Applied Econometrics, 1(1):1–4, January 1986. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:1992:Aa

[Pes92a] M. Hashem Pesaran. Announcement. Journal of Applied Econometrics, 7(1):1–2, January–March 1992. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:1992:Ab

[Pes92b] M. Hashem Pesaran. Announcement. Journal of Applied Econometrics, 7(4):335, October–December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:1993:A

[Pes93] M Hashem Pesaran. Announcement. Journal of Applied Econometrics, 8(4):337, October–December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:1994:A

[Pes94] M Hashem Pesaran. Announcement. Journal of Applied Econometrics, 9(4):347, October–December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:1995:A

[Pes95] M. Hashem Pesaran. Announcement. Journal of Applied Econometrics, 10(4):345, October–December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2000:EC

[Pes00] M. Hashem Pesaran. Editorial clarification. Journal of Applied Econometrics, 15(5):543, September–October 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2001:JAEa

[Pes01a] M. Hashem Pesaran. Journal of applied econometrics conference sponsorship grants. Journal of Applied Econometrics, 16 (4):561, July–August 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2001:JAEb

[Pes01b] M. Hashem Pesaran. Journal of applied econometrics distinguished authors. Journal of Applied Econometrics, 16(5):653– 654, September–October 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2001:RSP

[Pes01c] M Hashem Pesaran. The Richard stone prize in applied econometrics. Journal of Applied Econometrics, 16(1):93, January– February 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2003:IRS

[Pes03a] Hashem Pesaran. Introducing a replication section. Journal of Applied Econometrics, 18(1):111, January–February 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2003:JAE

[Pes03b] M. Hashem Pesaran. Journal of applied econometrics scholars programme. Journal of Applied Econometrics, 18(5):619, September–October 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2007:JAE

[Pes07a] M. Hashem Pesaran. Journal of applied econometrics dissertation prize. Journal of Applied Econometrics, 22(7):1395, December 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2007:SPU

[Pes07b] M. Hashem Pesaran. A simple panel unit root test in the presence of cross-section dependence. Journal of Applied Econometrics, 22(2):265–312, March 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2008:MAJ

[Pes08] M. Hashem Pesaran. March 2008 announcement: Journal of Applied Econometrics distinguished authors. Journal of Applied Econometrics, 23(3):391–393, April 2008. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2009:A

[Pes09a] M. Hashem Pesaran. Announcement. Journal of Applied Econometrics, 24(5):865, August 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2009:JAE

[Pes09b] M. Hashem Pesaran. Journal of applied econometrics dissertation prize. Journal of Applied Econometrics, 24(1):207, January–February 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2009:RSP

[Pes09c] M. Hashem Pesaran. The Richard stone prize in applied econometrics. Journal of Applied Econometrics, 24(5):863, August 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2010:JAE

[Pes10a] M. Hashem Pesaran. Journal of applied econometrics distinguished authors. Journal of Applied Econometrics, 25(1):195, January–February 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2010:RSP

[Pes10b] M Hashem Pesaran. The Richard stone prize in applied econometrics. Journal of Applied Econometrics, 25(6):1067–1068, September–October 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2011:JAE

[Pes11] M. Hashem Pesaran. Journal of applied econometrics distinguished authors. Journal of Applied Econometrics, 26(3):545– 546, April–May 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2013:DA

[Pes13] M. Hashem Pesaran. Distinguished authors. Journal of Applied Econometrics, 28(5):900–901, August 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Petit:1989:FMP

[Pet89] Maria Luisa Petit. Fiscal and monetary policy co-ordination: A differential game approach. Journal of Applied Econometrics, 4 (2):161–179, April–June 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Peterson:1991:SRR

[Pet91] Steven P. Peterson. Software review: A review of Stata 2.1. Journal of Applied Econometrics, 6(2):207–212, April–June 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Paap:2000:DMP

[PF00] Richard Paap and Philip Hans Franses. A dynamic multinomial probit model for brand choice with different long-run and short-run effects of marketing-mix variables. Journal of Applied Econometrics, 15(6):717–744, November–December 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pruser:2024:NMT

[PH24] Jan Prüser and Florian Huber. Nonlinearities in macroeconomic tail risk through the lens of big data quantile regressions. Journal of Applied Econometrics, 39(2):269–291, March 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Phillips:1989:BRL

[Phi89] Peter C. B. Phillips. Book review: Lectures on Advanced Econometric Theory, Denis Sargan (edited by Meghnad Desai). Basil Blackwell, Oxford, 1988. Price (cloth): £25, pp. xi + 176. Journal of Applied Econometrics, 4(4):415–418, October– December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Phillips:1991:BRU

[Phi91a] P. C. B. Phillips. Bayesian routes and unit roots: De rebus prioribus semper est disputandum. Journal of Applied Econometrics, 6(4):435–473, October–December 1991. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Phillips:1991:CCO

[Phi91b] P. C. B. Phillips. To criticize the critics: An objective Bayesian analysis of stochastic trends. Journal of Applied Econometrics, 6(4):333–364, October–December 1991. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic). See comments [KS91, Lea91, Poi91, Sim91].

Phillips:1995:NTS

[Phi95] Peter C. B. Phillips. Nonstationary time series and cointegration. Journal of Applied Econometrics, 10(1):87–94, January– March 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Phillips:2001:DEN

[Phi01] Peter C. B. Phillips. Descriptive econometrics for nonstationary time series with empirical illustrations. Journal of Applied Econometrics, 16(3):389–413, May–June 2001. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Parmeter:2007:NEH

[PHK07] Christopher F. Parmeter, Daniel J. Henderson, and Subal C. Kumbhakar. Nonparametric estimation of a hedonic price function. Journal of Applied Econometrics, 22(3):695–699, April– May 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Porro:2009:RRP

[PI09] Giuseppe Porro and Stefano Maria Iacus. Random recursive partitioning: a matching method for the estimation of the average treatment effect. Journal of Applied Econometrics, 24 (1):163–185, January–February 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pierse:1996:GV

[Pie96] Richard G. Pierse. GAUSSX: Version 3.4. Journal of Applied Econometrics, 11(6):687–693, November 1996. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pissarides:1992:BRC

[Pis92] Christopher A. Pissarides. Book review: Challenges for Macroeconomic Modelling, W. Driehuis, M. M. G. Fase, and H. Den Hartog (editors). North-Holland, Amsterdam, 1988. ISBN 0-444-70529-5, cloth, $94.75. pp. xii + 487. Journal of Applied Econometrics, 7(2):221–223, April–June 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Paluch:2012:IVA

[PKH12] Michal Paluch, Alois Kneip, and Werner Hildenbrand. Individual versus aggregate income elasticities for heterogeneous populations. Journal of Applied Econometrics, 27(5):847–869, August 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Parent:2008:UVS

[PL08] Olivier Parent and James P. LeSage. Using the variance structure of the conditional autoregressive spatial specification to model knowledge spillovers. Journal of Applied Econometrics, 23(2):235–256, March 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Lai:2020:EDS

[pLK20] Hung pin Lai and Subal C. Kumbhakar. Estimation of a dynamic stochastic frontier model using likelihood-based approaches. Journal of Applied Econometrics, 35(2):217–247, March 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Phillips:1996:RTF

[PMM96] Peter C. B. Phillips, James W. McFarland, and Patrick C. McMahon. Robust tests of forward exchange market efficiency with empirical evidence from the 1920s. Journal of Applied Econometrics, 11(1):1–22, January 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Proietti:2017:EDE

[PMM17] Tommaso Proietti, Martyna Marczak, and Gianluigi Mazzi. Euromind-D: A density estimate of monthly gross domestic product for the Euro area. Journal of Applied Econometrics, 32 (3):683–703, April–May 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Poirier:1991:CCC

[Poi91] Dale J. Poirier. A comment on ‘to criticize the critics: An objective Bayesian analysis of stochastic trends’. Journal of Applied Econometrics, 6(4):381–386, October–December 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [Phi91b].

Pollock:2002:RTW

[Pol02] D. S. G. Pollock. A review of TSW: the Windows version of the TRAMO–SEATS program. Journal of Applied Econometrics, 17(3):291–299, May–June 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pollak:2013:UHC

[Pol13] Andreas Pollak. Unemployment, human capital depreciation, and unemployment insurance policy. Journal of Applied Econometrics, 28(5):840–863, August 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Postepska:2019:ECI

[Pos19] Agnieszka Postepska. Ethnic capital and intergenerational transmission of educational attainment. Journal of Applied Econometrics, 34(4):606–611, June–July 2019. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Potter:1995:NAU

[Pot95] Simon M. Potter. A nonlinear approach to US GNP. Journal of Applied Econometrics, 10(2):109–125, April–June 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:1992:NDE

[PP92] M. Hashem Pesaran and Simon M. Potter. Nonlinear dynamics and econometrics: an introduction. Journal of Applied Econometrics, 7(S1):S1–S7, December 1992. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Paya:2006:TAE

[PP06] Ivan Paya and David A. Peel. Temporal aggregation of an ESTAR process: some implications for purchasing power parity adjustment. Journal of Applied Econometrics, 21(5):655–668, July–August 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Piatek:2016:MLC

[PP16] Rémi Piatek and Pia Pinger. Maintaining (locus of) control? Data combination for the identification and inference of factor structure models. Journal of Applied Econometrics, 31(4): 734–755, June–July 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Paredes:2023:FTG

[PPQ23] Joan Paredes, Javier J. Pérez, and Gabriel Perez Quiros. Fiscal targets. a guide to forecasters? Journal of Applied Econometrics, 38(4):472–492, June–July 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Picchio:2021:INW

[PPSV21] Matteo Picchio, Claudia Pigini, Stefano Staffolani, and Alina Verashchagina. If not now, when? The timing of childbirth and labor market outcomes. Journal of Applied Econometrics, 36(6):663–685, September–October 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Planas:2004:CID

[PR04] Christophe Planas and Alessandro Rossi. Can inflation data improve the real-time reliability of output gap estimates? Journal of Applied Econometrics, 19(1):121–133, January–February 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesavento:2006:SSC

[PR06] Elena Pesavento and Barbara Rossi. Small-sample confidence intervals for multivariate impulse response functions at long horizons. Journal of Applied Econometrics, 21(8):1135–1155, December 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pettenuzzo:2016:OPC

[PR16] Davide Pettenuzzo and Francesco Ravazzolo. Optimal portfolio choice under decision-based model combinations. Journal of Applied Econometrics, 31(7):1312–1332, November–December 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Prat:2010:RLD

[Pra10] Julien Prat. The rate of learning-by-doing: estimates from a search-matching model. Journal of Applied Econometrics, 25(6):929–962, September–October 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Prieger:2002:FPS

[Pri02] James E. Prieger. A flexible parametric selection model for non-normal data with application to health care usage. Journal of Applied Econometrics, 17(4):367–392, July–August 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Provencher:1995:IHD

[Pro95] Bill Provencher. An investigation of the harvest decision of timber firms in the south-east United States. Journal of Applied Econometrics, 10(S1):S57–S74, December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Proietti:1996:PSS

[Pro96] Tommaso Proietti. Persistence of shocks on seasonal processes. Journal of Applied Econometrics, 11(4):383–398, July 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Prono:2014:RCH

[Pro14] Todd Prono. The role of conditional heteroskedasticity in identifying and estimating linear triangular systems, with applications to asset pricing models that include a mismeasured factor. Journal of Applied Econometrics, 29(5):800–824, August 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Psaradakis:2005:MSC

[PRS05] Zacharias Psaradakis, Morten O. Ravn, and Martin Sola. Markov switching causality and the money-output relationship. Journal of Applied Econometrics, 20(5):665–683, July–August 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pudney:2000:GRP

[PS00] Stephen Pudney and Michael Shields. Gender, race, pay and promotion in the British nursing profession: estimation of a generalized ordered probit model. Journal of Applied Econometrics, 15(4):367–399, July–August 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pagan:2003:SFA

[PS03a] Adrian R. Pagan and Kirill A. Sossounov. A simple framework for analysing bull and bear markets. Journal of Applied Econometrics, 18(1):23–46, January–February 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Page:2003:CBS

[PS03b] Marianne E. Page and Gary Solon. Correlations between sisters and neighbouring girls in their subsequent income as adults. Journal of Applied Econometrics, 18(5):545–562, September– October 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Psaradakis:2003:DCA

[PS03c] Zacharias Psaradakis and Martin Sola. On detrending and cyclical asymmetry. Journal of Applied Econometrics, 18(3): 271–289, May–June 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pinkse:2007:SSM

[PS07] Joris Pinkse and Margaret Slade. Semi-structural models of advertising competition. Journal of Applied Econometrics, 22 (7):1227–1246, December 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Phillips:2009:ETG

[PS09] Peter C. B. Phillips and Donggyu Sul. Economic transition and growth. Journal of Applied Econometrics, 24(7):1153– 1185, November–December 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pendakur:2010:SEC

[PS10] Krishna Pendakur and Stefan Sperlich. Semiparametric estimation of consumer demand systems in real expenditure. Journal of Applied Econometrics, 25(3):420–457, April–May 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Polito:2016:OCH

[PS16] Vito Polito and Peter Spencer. Optimal control of heteroscedastic macroeconomic models. Journal of Applied Econometrics, 31(7):1430–1444, November–December 2016. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Park:2023:RCN

[PS23] Seonyoung Park and Donggyun Shin. Recent changes in the nature of the distribution dynamics of the US county incomes.

Journal of Applied Econometrics, 38(7):1048–1067, November– December 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2001:BTA

[PSS01] M. Hashem Pesaran, Yongcheol Shin, and Richard J. Smith. Bounds testing approaches to the analysis of level relationships. Journal of Applied Econometrics, 16(3):289–326, May– June 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See replication work [NT22].

Psaradakis:2004:MEC

[PSS04] Zacharias Psaradakis, Martin Sola, and Fabio Spagnolo. On Markov error-correction models, with an application to stock prices and dividends. Journal of Applied Econometrics, 19(1): 69–88, January–February 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Polanski:2019:TTT

[PSW19] Arnold Polanski, Evarist Stoja, and Frank Windmeijer. Telling tales from the tails: High-dimensional tail interdependence. Journal of Applied Econometrics, 34(5):779–794, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Preve:2013:ETV

[PT13] Daniel Preve and Yiu-Kuen Tse. Estimation of time–varying adjusted probability of informed trading and probability of symmetric order–flow shock. Journal of Applied Econometrics, 28(7):1138–1152, November–December 2013. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pagan:1988:EAM

[PU88] Adrian Pagan and Aman Ullah. The econometric analysis of models with risk terms. Journal of Applied Econometrics, 3 (2):87–105, April 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pudney:1993:IWI

[Pud93] S. Pudney. Income and wealth inequality and the life cycle. A non-parametric analysis for China. Journal of Applied Econometrics, 8(3):249–276, July–September 1993. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pudney:2004:KGE

[Pud04] Stephen Pudney. Keeping off the grass? An econometric model of cannabis consumption in Britain. Journal of Applied Econometrics, 19(4):435–453, July–August 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pagan:1989:DTM

[PV89] Adrian Pagan and Frank Vella. Diagnostic tests for models based on individual data: A survey. Journal of Applied Econometrics, 4(S1):S29–S59, December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Papke:1996:EMF

[PW96] Leslie E. Papke and Jeffrey M. Wooldridge. Econometric methods for fractional response variables with an application to 401(k) plan participation rates. Journal of Applied Econometrics, 11(6):619–632, November 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Polito:2012:OMP

[PW12] Vito Polito and Mike Wickens. Optimal monetary policy using an unrestricted VAR. Journal of Applied Econometrics, 27(4): 525–553, June–July 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Petrova:2020:FED

[PW20] Yana Petrova and Joakim Westerlund. Fixed effects demeaning in the presence of interactive effects in treatment effects regressions and elsewhere. Journal of Applied Econometrics, 35(7): 960–964, November–December 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Polachek:1996:PET

[PY96] Solomon W. Polachek and Bong Joon Yoon. Panel estimates of a two-tiered earnings frontier. Journal of Applied Econometrics, 11(2):169–178, March 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Perron:2015:UOE

[PY15] Pierre Perron and Yohei Yamamoto. Using OLS to estimate and test for structural changes in models with endogenous regressors. Journal of Applied Econometrics, 30(1):119–144, January–February 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pesaran:2022:MTE

[PY22] M. Hashem Pesaran and Cynthia Fan Yang. Matching theory and evidence on Covid-19 using a stochastic network SIR model. Journal of Applied Econometrics, 37(6):1204–1229, September–October 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Pittau:2006:EEI

[PZ06] Maria Grazia Pittau and Roberto Zelli. Empirical evidence of income dynamics across EU regions. Journal of Applied Econometrics, 21(5):605–628, July–August 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Quandt:1986:UDS

[QR86] Richard E. Quandt and Harvey S. Rosen. Unemployment, disequilibrium and the short run Phillips curve: An econometric approach. Journal of Applied Econometrics, 1(3):235–253, June 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See correction and extension [QR87].

Quandt:1987:UDS

[QR87] Richard E. Quandt and Harvey S. Rosen. Unemployment, disequilibrium and the short-run Phillips curve: Correction and extension. Journal of Applied Econometrics, 2(3):247–249, July 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [QR86].

Qu:2023:UAP

[QT23] Zhongjun Qu and Denis Tkachenko. Using arbitrary precision arithmetic to sharpen identification analysis for DSGE models. Journal of Applied Econometrics, 38(4):644–667, June–July 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Quast:2023:FRO

[QW23] Josefine Quast and Maik H. Wolters. The Federal Reserve’s output gap: The unreliability of real-time reliability tests. Journal of Applied Econometrics, 38(7):1101–1111, November– December 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Racine:1997:FCV

[Rac97] Jeff Racine. Feasible cross–validatory model selection for general stationary processes. Journal of Applied Econometrics, 12 (2):169–179, March 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Racine:2000:CTG

[Rac00] J. Racine. The Cygwin tools: a GNU toolkit for Windows. Journal of Applied Econometrics, 15(3):331–341, May–June 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Racine:2006:SRG

[Rac06] Jeff Racine. Software review: gnuplot 4.0: a portable interactive plotting utility. Journal of Applied Econometrics, 21(1): 133–141, January–February 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Racine:2012:RPI

[Rac12] Jeffrey S. Racine. RStudio: A platform-independent IDE for R and Sweave. Journal of Applied Econometrics, 27(1):167–172, January–February 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Raj:1992:IEP

[Raj92] Baldev Raj. International evidence on persistence in output in the presence of an episodic change. Journal of Applied Econometrics, 7(3):281–293, July–September 1992. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Raj:1995:IHL

[Raj95] Baldev Raj. Institutional hypothesis of the long-run income velocity of money and parameter stability of the equilibrium relationship. Journal of Applied Econometrics, 10(3):233–253, July–September 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rask:1995:STB

[Ras95] Kevin Rask. The structure of technology in Brazilian sugarcane production, 1975-87: An application of a modified symmetric generalized McFadden cost function. Journal of Applied Econometrics, 10(3):221–232, July–September 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ravn:1997:PTS

[Rav97] Morten O. Ravn. Permanent and transitory shocks, and the UK business cycle. Journal of Applied Econometrics, 12(1):27– 48, January 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rossi:2018:IIT

[RdM18] Eduardo Rossi and Paolo Santucci de Magistris. Indirect inference with time series observed with error. Journal of Applied Econometrics, 33(6):874–897, September–October 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rezende:2008:EAL

[Rez08] Leonardo Rezende. Econometrics of auctions by least squares. Journal of Applied Econometrics, 23(7):925–948, November– December 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Racine:2002:URT

[RH02] Jeff Racine and Rob Hyndman. Using R to teach econometrics. Journal of Applied Econometrics, 17(2):175–189, March–April 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ramirez-Hassan:2021:BEE

[RH21] Andrés Ramírez-Hassan. Bayesian estimation of the exact affine Stone index demand system: Replicating the Lewbel and Pendakur (2009) results. Journal of Applied Econometrics, 36 (4):484–491, June–July 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ridder:1986:EHA

[Rid86] G. Ridder. An event history approach to the evaluation of training, recruitment and employment programmes. Journal of Applied Econometrics, 1(2):109–126, April 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rizzi:1991:BRL

[Riz91] Alfredo Rizzi. Book review: Linear structures, J. R. Magnus, Oxford, Oxford University Press, 1988. ISBN 0-19-520655-X, cloth, $49.95, pp. xii + 205. Journal of Applied Econometrics, 6(1):103–105, January–March 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Raper:2000:DMP

[RLS00] Kellie Curry Raper, H. Alan Love, and C. Richard Shumway. Determining market power exertion between buyers and sellers. Journal of Applied Econometrics, 15(3):225–252, May– June 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ruge-Murcia:2000:UFM

[RM00] Francisco J. Ruge-Murcia. Uncovering financial markets’ beliefs about inflation targets. Journal of Applied Econometrics, 15(5):483–512, September–October 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ruge-Murcia:2017:SRB

[RM17] Francisco Ruge-Murcia. Skewness risk and bond prices. Journal of Applied Econometrics, 32(2):379–400, March 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Robinson:1988:SES

[Rob88] P. M. Robinson. Semiparametric econometrics: A survey. Journal of Applied Econometrics, 3(1):35–51, January 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Romeo:1999:CIS

[Rom99] Charles J. Romeo. Conducting inference in semiparametric duration models under inequality restrictions on the shape of the hazard implied by job search theory. Journal of Applied Econometrics, 14(6):587–605, November–December 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Romeu:2004:EGC

[Rom04] Andrés Romeu. ExpEnd: GAUSS code for panel count-data models. Journal of Applied Econometrics, 19(3):429–434, May– June 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ropponen:2011:REC

[Rop11] Olli Ropponen. Reconciling the evidence of Card and Krueger (1994) and Neumark and Wascher (2000). Journal of Applied Econometrics, 26(6):1051–1057, September–October 2011. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). The research referred to appears in Card and Krueger (American Economic Review 1994; 84: 772–793) and Neumark and Wascher (American Economic Review 2000; 90: 1362– 1396).

Rossana:1993:LRI

[Ros93] R. J. Rossana. The long-run implications of the production smoothing model of inventories: An empirical test. Journal of Applied Econometrics, 8(3):295–306, July–September 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rose:2022:SWN

[Ros22] Michael E. Rose. Small world: Narrow, wide, and long replication of Goyal, van der Leij and Moraga–Gonzélez (JPE 2006) and a comparison of EconLit and Scopus. Journal of Applied Econometrics, 37(4):820–828, June–July 2022. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rothman:1992:CPT

[Rot92] P. Rothman. The comparative power of the TR test against simple threshold models. Journal of Applied Econometrics, 7 (S1):S187–S195, December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rodriguez-Poo:2005:STS

[RPSF05] Juan M. Rodríguez-Póo, Stefan Sperlich, and Ana I. Fernández. Semiparametric three-step estimation methods for simultaneous equation systems. Journal of Applied Econometrics, 20(6): 699–721, September–October 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rust:1995:ORS

[RR95] John Rust and Geoffrey Rothwell. Optimal response to a shift in regulatory regime: The case of the US nuclear power industry. Journal of Applied Econometrics, 10(S1):S75–S118, December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Richey:2018:DEM

[RR18] Jeremiah Richey and Alicia Rosburg. Decomposing economic mobility transition matrices. Journal of Applied Econometrics, 33(1):91–108, January–February 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rich:1992:RBF

[RRB92] R. W. Rich, J. E. Raymond, and J. S. Butler. The relationship between forecast dispersion and forecast uncertainty: Evidence from a survey data-arch model. Journal of Applied Econometrics, 7(2):131–148, April–June 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Raj:1986:RFP

[RS86a] Baldev Raj and Pierre L. Siklos. The role of fiscal policy in the St. Louis model: An evaluation and some new evidence. Journal of Applied Econometrics, 1(3):287–294, June 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rees:1986:EAS

[RS86b] Hedley Rees and Anup Shah. An empirical analysis of selfemployment in the U.K. Journal of Applied Econometrics, 1 (1):95–108, January 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Robertson:1992:SSP

[RS92] D. Robertson and J. Symons. Some strange properties of panel data estimators. Journal of Applied Econometrics, 7(2):175– 189, April–June 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Romeo:1999:LDC

[RS99] Charles Romeo and Barry Sopher. Learning and decision costs in one-person games. Journal of Applied Econometrics, 14(4): 335–357, July–August 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rapach:2008:SBG

[RS08a] David E. Rapach and Jack K. Strauss. Structural breaks and GARCH models of exchange rate volatility. Journal of Applied Econometrics, 23(1):65–90, January–February 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ray:2008:PHA

[RS08b] Surajit Ray and N. E. Savin. The performance of heteroskedasticity and autocorrelation robust tests: a Monte Carlo study with an application to the three-factor FAMA–French assetpricing model. Journal of Applied Econometrics, 23(1):91–109, January–February 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rosholm:2009:LMT

[RS09] Michael Rosholm and Lars Skipper. Is labour market training a curse for the unemployed? Evidence from a social experiment. Journal of Applied Econometrics, 24(2):338–365, March 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rossi:2016:FRT

[RS16] Barbara Rossi and Tatevik Sekhposyan. Forecast rationality tests in the presence of instabilities, with applications to Federal reserve and survey forecasts. Journal of Applied Econometrics, 31(3):507–532, April–May 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rebei:2021:TPS

[RS21] Nooman Rebei and Rashid Sbia. Transitory and permanent shocks in the global market for crude oil. Journal of Applied Econometrics, 36(7):1047–1064, November–December 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rodriguez:2022:DTE

[RSU22] Jorge Rodríguez, Fernando Saltiel, and Sergio Urzúa. Dynamic treatment effects of job training. Journal of Applied Econometrics, 37(2):242–269, March 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rahiala:1988:FFP

[RT88] Markku Rahiala and Timo Teräsvirta. Formation of firms’ production decisions in Finnish manufacturing industries. Journal of Applied Econometrics, 3(2):125–137, April 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ranjan:2007:BIG

[RT07] Priya Ranjan and Justin L. Tobias. Bayesian inference for the gravity model. Journal of Applied Econometrics, 22(4): 817–838, June–July 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ryden:1998:SFD

[RTÅ98] Tobias Rydén, Timo Teräsvirta, and Stefan Åsbrink. Stylized facts of daily return series and the hidden Markov model. Journal of Applied Econometrics, 13(3):217–244, May–June 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Restrepo-Tobon:2014:EQL

[RTK14] Diego Restrepo-Tobón and Subal C. Kumbhakar. Enjoying the quiet life under deregulation? Not quite. Journal of Applied Econometrics, 29(2):333–343, March 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ruiz-Tagle:2016:BCC

[RTV16] Jaime Ruiz-Tagle and Francis Vella. Borrowing constraints and credit demand in a developing economy. Journal of Applied Econometrics, 31(5):865–891, August 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rust:1993:GMC

[Rus93] J. Rust. Gauss and MATLAB: A comparison. Journal of Applied Econometrics, 8(3):307–324, July–September 1993. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Runstler:2018:BHC

[RV18] Gerhard Rünstler and Marente Vlekke. Business, housing, and credit cycles. Journal of Applied Econometrics, 33(2):212–226, March 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Reynaerts:2022:ESF

[RV22] Jo Reynaerts and Jakob Vanschoonbeek. The economics of state fragmentation: Assessing the economic impact of secession. Journal of Applied Econometrics, 37(1):82–115, January– February 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ruth:2023:MPE

[RV23] Sebastian K. Rüth and Wouter Van der Veken. Monetary policy and exchange rate anomalies in set-identified SVARs: Revisited. Journal of Applied Econometrics, 38(7):1085–1092, November–December 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Richardson:2013:DDV

[RvdB13] Katarina Richardson and Gerard J. van den Berg. Duration dependence versus unobserved heterogeneity in treatment effects: Swedish labor market training and the transition rate to employment. Journal of Applied Econometrics, 28(2):325– 351, March 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rapach:2005:VRL

[RW05] David E. Rapach and Mark E. Wohar. Valuation ratios and long-horizon stock price predictability. Journal of Applied Econometrics, 20(3):327–344, March–April 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Reese:2016:PPC

[RW16] Simon Reese and Joakim Westerlund. Panicca: Panic on crosssection averages. Journal of Applied Econometrics, 31(6):961– 981, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Riphahn:2003:IED

[RWM03] Regina T. Riphahn, Achim Wambach, and Andreas Million. Incentive effects in the demand for health care: a bivariate panel count data estimation. Journal of Applied Econometrics, 18(4):387–405, July–August 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rabemananjara:1993:TAM

[RZ93] R. Rabemananjara and J. M. Zakoian. Threshold ARCH models and asymmetries in volatility. Journal of Applied Econometrics, 8(1):31–49, January–March 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rees:2003:IWP

[RZH03] Daniel I. Rees, Jeffrey S. Zax, and Joshua Herries. Interdependence in worker productivity. Journal of Applied Econometrics, 18(5):585–604, September–October 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sabani:1991:BRF

[Sab91] Laura Sabani. Book review: Forecasting structural time series models and the Kalman filter, A. C. Harvey. Cambridge University Press, Cambridge, 1989. ISBN 0-521-32196-4, cloth, £55.00 Pp. xvi + 554. Journal of Applied Econometrics, 6 (3):329–331, July–September 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sakaguchi:2020:EAT

[Sak20] Shosei Sakaguchi. Estimation of average treatment effects using panel data when treatment effect heterogeneity depends on unobserved fixed effects. Journal of Applied Econometrics, 35(3): 315–327, April–May 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sakaguchi:2024:PII

[Sak24] Shosei Sakaguchi. Partial identification and inference in duration models with endogenous censoring. Journal of Applied Econometrics, 39(2):308–326, March 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Salanie:1991:WPA

[Sal91] Bernard Salanie. Wage and price adjustment in a multimarket disequilibrium model. Journal of Applied Econometrics, 6(1): 1–15, January–March 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Salemi:1999:ENR

[Sal99] Michael K. Salemi. Estimating the natural rate of unemployment and testing the natural rate hypothesis. Journal of Applied Econometrics, 14(1):1–25, January–February 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sala:2015:DMF

[Sal15] Luca Sala. DSGE models in the frequency domains. Journal of Applied Econometrics, 30(2):219–240, March 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sampson:1991:EPU

[Sam91] Michael Sampson. The effect of parameter uncertainty on forecast variances and confidence intervals for unit root and trend stationary time-series models. Journal of Applied Econometrics, 6(1):67–76, January–March 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sanroman:2015:CPH

[San15] Graciela Sanroman. Cost and preference heterogeneity in risky financial markets. Journal of Applied Econometrics, 30(2):313– 332, March 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sareen:1999:POC

[Sar99] Samita Sareen. Posterior odds comparison of a symmetric low-price, sealed-bid auction within the common-value and the independent-private-values paradigms. Journal of Applied Econometrics, 14(6):651–676, November–December 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Savva:2016:RGH

[Sav16] Christos S. Savva. Replication of Grier, Henry, Olekalns and Shields (2004): the asymmetric effects of uncertainty on inflation and output growth. Journal of Applied Econometrics, 31 (6):1192–1196, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [GHOS04].

Sterbenz:1990:ASE

[SC90] Frederic P. Sterbenz and Giorgio Calzolari. Alternative specifications of the error process in the stochastic simulation of econometric models. Journal of Applied Econometrics, 5(2): 137–150, April–June 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Schaller:1990:RET

[Sch90] Huntley Schaller. A re-examination of the q theory of investment using U.S. firm data. Journal of Applied Econometrics, 5(4):309–325, September–December 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Schafgans:1998:EWD

[Sch98] Marcia M. A. Schafgans. Ethnic wage differences in Malaysia: parametric and semiparametric estimation of the Chinese– Malay wage gap. Journal of Applied Econometrics, 13(5):481– 504, September–October 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Schorfheide:2000:LFB

[Sch00] Frank Schorfheide. Loss function-based evaluation of DSGE models. Journal of Applied Econometrics, 15(6):645–670, November–December 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Schuermann:2005:RRB

[Sch05] Til Schuermann. A review of recent books on credit risk. Journal of Applied Econometrics, 20(1):123–130, January–February 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Schleicher:2007:CCA

[Sch07] Christoph Schleicher. Codependence in cointegrated autoregressive models. Journal of Applied Econometrics, 22(1):137– 159, January–February 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Scheffel:2016:APU

[Sch16] Eric M. Scheffel. Accounting for the political uncertainty factor. Journal of Applied Econometrics, 31(6):1048–1064, September–October 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Scida:2023:SVF

[Sci23] Daniela Scidá. Structural VAR and financial networks: a minimum distance approach to spatial modeling. Journal of Applied Econometrics, 38(1):49–68, January–February 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Scrimgeour:2022:RES

[Scr22] Dean Scrimgeour. Reevaluating the evidence on seasonality in housing market match quality: Replication of Ngai and Tenreyro (2014). Journal of Applied Econometrics, 37(7):1403– 1409, November–December 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Schmidt-Dengler:2003:SMA

[SD03] Philipp Schmidt-Dengler. A structural model of aggregate US job flows: another look. Journal of Applied Econometrics, 18 (1):113–118, January–February 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sun:2018:ETS

[SdV18] Pengfei Sun and Casper G. de Vries. Exploiting tail shape biases to discriminate between stable and student t alternatives. Journal of Applied Econometrics, 33(5):708–726, August 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Shachar:2007:CBR

[SE07] Ron Shachar and Zvi Eckstein. Correcting for bias in retrospective data. Journal of Applied Econometrics, 22(3):657–675, April–May 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Seale:1990:ESF

[Sea90] James L. Seale, Jr. Estimating stochastic frontier systems with unbalanced panel data: The case of floor tile manufactories in Egypt. Journal of Applied Econometrics, 5(1):59–74, January– March 1990. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Semykina:2018:SEA

[Sem18] Anastasia Semykina. Self-employment among women: Do children matter more than we previously thought? Journal of Applied Econometrics, 33(3):416–434, April–May 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sena:1999:SFE

[Sen99] Vania Sena. Stochastic frontier estimation: a review of the software options. Journal of Applied Econometrics, 14(5):579– 586, September–October 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Senyuz:2011:FAP

[Sen11] Zeynep Senyuz. Factor analysis of permanent and transitory dynamics of the US economy and the stock market. Journal of Applied Econometrics, 26(6):975–998, September–October 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sephton:1998:EV

[Sep98] Peter S. Sephton. Easyreg: version 1.12. Journal of Applied Econometrics, 13(2):203–207, March–April 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sephton:2000:FAP

[Sep00] P. S. Sephton. Financial analysis package for GAUSS. Journal of Applied Econometrics, 15(4):433–438, July–August 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sandor:2009:CPE

[SF09] Zsolt Sándor and Philip Hans Franses. Consumer price evaluations through choice experiments. Journal of Applied Econometrics, 24(3):517–535, April–May 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smith:2012:MDU

[SGK12] Michael S. Smith, Quan Gan, and Robert J. Kohn. Modelling dependence using skew t copulas: Bayesian inference and applications. Journal of Applied Econometrics, 27(3):500–522, April–May 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Shapiro:1986:CUC

[Sha86] Matthew D. Shapiro. Capital utilization and capital accumulation: Theory and evidence. Journal of Applied Econometrics, 1(3):211–234, June 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Shachar:2003:PLH

[Sha03] Ron Shachar. Party loyalty as habit formation. Journal of Applied Econometrics, 18(3):251–269, May–June 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Shea:1989:EPR

[She89] Gary S. Shea. Ex-post rational price approximations and the empirical reliability of the present-value relation. Journal of Applied Econometrics, 4(2):139–159, April–June 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Shephard:1993:FNT

[She93] N. Shephard. Fitting nonlinear time-series models with applications to stochastic variance models. Journal of Applied Econometrics, 8(S1):S135–S152, December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sheng:2013:DCD

[She13] Liugang Sheng. Did China diversify its foreign reserves? Journal of Applied Econometrics, 28(1):102–125, January–February 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Shiba:1992:MEL

[Shi92] Tsunemasa Shiba. MICRO-EBA: Leamer’s extreme bounds analysis on GAUSS. Journal of Applied Econometrics, 7(1): 101–103, January–March 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Shively:2001:TSG

[Shi01] Philip A. Shively. Trend-stationary GNP: evidence from a new exact pointwise most powerful invariant unit root test. Journal of Applied Econometrics, 16(4):537–551, July–August 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Shintani:2006:NMC

[Shi06] Mototsugu Shintani. A nonparametric measure of convergence towards purchasing power parity. Journal of Applied Econometrics, 21(5):589–604, July–August 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sun:2011:BAL

[SHK11] Kai Sun, Daniel J. Henderson, and Subal C. Kumbhakar. Biases in approximating log production. Journal of Applied Econometrics, 26(4):708–714, June–July 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Schnurbus:2017:ETG

[SHM17] Joachim Schnurbus, Harry Haupt, and Verena Meier. Economic transition and growth: A replication. Journal of Applied Econometrics, 32(5):1039–1042, August 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Silk:1996:SEC

[Sil96] Julian Silk. Systems estimation: A comparison of SAS, SHAZAM and TSP. Journal of Applied Econometrics, 11 (4):437–450, July 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Silk:1997:TR

[Sil97] Julian Silk. TSP 4.4: a review. Journal of Applied Econometrics, 12(4):445–453, July 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Silva:2001:STN

[Sil01] J. M. C. Santos Silva. A score test for non-nested hypotheses with applications to discrete data models. Journal of Applied Econometrics, 16(5):577–597, September–October 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sims:1991:CCS

[Sim91] Christopher A. Sims. Comment by Christopher A. Sims on ‘To criticize the critics’, by Peter C. B. Phillips. Journal of Applied Econometrics, 6(4):423–434, October–December 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [Phi91b].

Soetevent:2007:DCM

[SK07] Adriaan R. Soetevent and Peter Kooreman. A discrete-choice model with social interactions: with an application to high school teen behavior. Journal of Applied Econometrics, 22(3):

599–624, April–May 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Schmidt:2021:IPF

[SKG21] Patrick Schmidt, Matthias Katzfuss, and Tilmann Gneiting. Interpretation of point forecasts with unknown directive. Journal of Applied Econometrics, 36(6):728–743, September– October 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Schwaab:2017:GCR

[SKL17] Bernd Schwaab, Siem Jan Koopman, and André Lucas. Global credit risk: World, country and industry factors. Journal of Applied Econometrics, 32(2):296–317, March 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Siikamaki:2015:FSS

[SL15] Juha Siikamäki and Douglas M. Larson. Finding sensitivity to scope in nonmarket valuation. Journal of Applied Econometrics, 30(2):333–349, March 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Slichter:2023:EEM

[Sli23] David Slichter. The employment effects of the minimum wage: a selection ratio approach to measuring treatment effects. Journal of Applied Econometrics, 38(3):334–357, April–May 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smith:2004:MPP

[SLKN04] L. Vanessa Smith, Stephen Leybourne, Tae-Hwan Kim, and Paul Newbold. More powerful panel data unit root tests with an application to mean reversion in real exchange rates. Journal of Applied Econometrics, 19(2):147–170, March–April 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Song:1997:CSM

[SLR97] Haiyan Song, Xiaming Liu, and Peter Romilly. A comparative study of modelling the demand for food in the United States and The Netherlands. Journal of Applied Econometrics, 12(5): 593–613, September–October 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smith:1995:APC

[SM95] Jeremy Smith and Michael McAleer. Alternative procedures for converting qualitative response data to quantitative expectations: An application to Australian manufacturing. Journal of Applied Econometrics, 10(2):165–185, April–June 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smith:1987:TEL

[Smi87] Richard J. Smith. Testing for exogeneity in limited dependent variable models using a simplified likelihood ratio statistic. Journal of Applied Econometrics, 2(3):237–245, July 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smith:1989:MME

[Smi89] R. P. Smith. Models of military expenditure. Journal of Applied Econometrics, 4(4):345–359, October–December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smith:1993:ENT

[Smi93] A. A. Smith, Jr. Estimating nonlinear time-series models using simulated vector autoregressions. Journal of Applied Econometrics, 8(S1):S63–S84, December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smith:1995:BRE

[Smi95] Richard J. Smith. Book review: Estimation and inference in econometrics, R. Davidson and J. G. Mackinnon. Oxford University, New York, 1993, pp. 871, ISBN 0-19-506011-3. Price £25.00 hardbound. Journal of Applied Econometrics, 10(3): 339–341, July–September 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smith:2005:POT

[Smi05] Aaron Smith. Partially overlapping time series: a new model for volatility dynamics in commodity futures. Journal of Applied Econometrics, 20(3):405–422, March–April 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smith:2012:DES

[Smi12] Loren K. Smith. Dynamics and equilibrium in a structural model of wide-body commercial aircraft markets. Journal of Applied Econometrics, 27(1):1–33, January–February 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smallwood:2006:GLM

[SN06] Aaron D. Smallwood and Stefan C. Norrbin. Generalized long memory processes, failure of cointegration tests and exchange rate dynamics. Journal of Applied Econometrics, 21(4):409– 417, May–June 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sanso-Navarro:2017:HCS

[SNVCXDE17] Marcos Sanso-Navarro, Maria Vera-Cabello, and Domingo P. Ximénez-De-Embún. Human capital spillovers and regional development. Journal of Applied Econometrics, 32(4):923–930, June–July 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Soderlind:1994:CPR

[Söd94] Paul Söderlind. Cyclical properties of a real business cycle model. Journal of Applied Econometrics, 9(S1):S113–S122, December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sokullu:2016:SPA

[Sok16] Senay Sokullu. A semi-parametric analysis of two-sided markets: An application to the local daily newspapers in the USA. Journal of Applied Econometrics, 31(5):843–864, August 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sollis:2005:EPP

[Sol05] Robert Sollis. Evidence on purchasing power parity from univariate models: the case of smooth transition trendstationarity. Journal of Applied Econometrics, 20(1):79–98, January–February 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sonnberger:1989:BRRa

[Son89a] Harald Sonnberger. Book review: Regression diagnostics: Identifying influential data and sources of collinearity, by D. A. Belsley, K. Kuh and R. E. Welsch. (John Wiley and Sons, New York, 1980, pp. xv + 292, ISBN 0-471-05856-4, cloth $39.95). Journal of Applied Econometrics, 4(1):97–99, January–March 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sonnberger:1989:BRRb

[Son89b] Harald Sonnberger. Book review: Robust regression and outlier detection, P. J. Rousseeuw and A. M. Leroy. John Wiley and Sons, New York, 1987. ISBN 0-471-85233-3. cloth £34.50, pp. xiv + 329. Journal of Applied Econometrics, 4(3):309–311, July–September 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Song:2014:MRS

[Son14] Yong Song. Modelling regime switching and structural breaks with an infinite hidden Markov model. Journal of Applied Econometrics, 29(5):825–842, August 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sordi:1989:BRC

[Sor89] Serena Sordi. Book review: Contributions of business cycle surveys to empirical economics (papers presented at the 18th CIRET Conference, Zurich 1987), edited by K. H. Oppenländer and G. Poser. (Avebury, Aldershot, England: 1988, ISBN 0-566-05629-1, £45.00 hardbound), pp. xxxv + 669. Journal of Applied Econometrics, 4(4):418–421, October– December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sun:2004:UFE

[SP04] Yixiao Sun and Peter C. B. Phillips. Understanding the Fisher equation. Journal of Applied Econometrics, 19(7):869–886, November–December 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Shen:2018:MDH

[SP18] Shulin Shen and Jindong Pang. Measuring the diffusion of housing prices across space and over time: Replication and further evidence. Journal of Applied Econometrics, 33(3):479– 484, April–May 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Spindler:2016:LIV

[Spi16] Martin Spindler. Lasso for instrumental variable selection: A replication study. Journal of Applied Econometrics, 31(2):450– 454, March 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Spagnolo:2005:TUF

[SPS05] Fabio Spagnolo, Zacharias Psaradakis, and Martin Sola. Testing the unbiased forward exchange rate hypothesis using a Markov switching model and instrumental variables. Journal of Applied Econometrics, 20(3):423–437, March–April 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Schoder:2013:CAI

[SPS13] Christian Schoder, Christian R. Proaño, and Willi Semmler. Are the current account imbalances between EMU countries sustainable? Evidence from parametric and non-parametric tests. Journal of Applied Econometrics, 28(7):1179–1204, November–December 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smith:1991:MPN

[SS91] Marlene A. Smith and David J. Smyth. Multiple and pairwise non-nested tests of the influence of taxes on money demand. Journal of Applied Econometrics, 6(1):17–30, January–March 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sickles:1998:ATP

[SS98] Robin C. Sickles and Mary L. Streitwieser. An analysis of technology, productivity, and regulatory distortion in the interstate natural gas transmission industry: 1977–1985. Journal of Applied Econometrics, 13(4):377–395, July–August 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smith:2005:HWD

[SS05] Penelope A. Smith and Peter M. Summers. How well do Markov switching models describe actual business cycles? The case of synchronization. Journal of Applied Econometrics, 20(2):253– 274, 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Simonsen:2006:CMA

[SS06] Marianne Simonsen and Lars Skipper. The costs of motherhood: an analysis using matching estimators. Journal of Applied Econometrics, 21(7):919–934, November 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Serlenga:2007:GMI

[SS07] Laura Serlenga and Yongcheol Shin. Gravity models of intra-EU trade: application of the CCEP–HT estimation in heterogeneous panels with unobserved common time-specific factors. Journal of Applied Econometrics, 22(2):361–381, March 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See replication report [Bal10].

Shephard:2010:RFF

[SS10] Neil Shephard and Kevin Sheppard. Realising the future: forecasting with high-frequency-based volatility (HEAVY) models. Journal of Applied Econometrics, 25(2):197–231, March 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Schaffer:2022:DCR

[SS22] Matthew Schaffer and Nimrod Segev. The deposits channel revisited. Journal of Applied Econometrics, 37(2):450–458, March 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Schmidheiny:2023:ESD

[SS23a] Kurt Schmidheiny and Sebastian Siegloch. On event studies and distributed-lags in two-way fixed effects models: Identification, equivalence, and generalization. Journal of Applied Econometrics, 38(5):695–713, August 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Shu:2023:OPR

[SS23b] Haicheng Shu and Peter Spencer. Oil prices in the real economy. Journal of Applied Econometrics, 38(6):878–897, September– October 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Simonsen:2016:PSD

[SSS16] Marianne Simonsen, Lars Skipper, and Niels Skipper. Price sensitivity of demand for prescription drugs: Exploiting a regression kink design. Journal of Applied Econometrics, 31(2): 320–337, March 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Stengos:2006:ESE

[SSW06] Thanasis Stengos, Yiguo Sun, and Dianqin Wang. Estimates of semiparametric equivalence scales. Journal of Applied Econometrics, 21(5):629–639, July–August 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

SantAnna:2022:CDB

[SSX22] Pedro H. C. Sant’Anna, Xiaojun Song, and Qi Xu. Covariate distribution balance via propensity scores. Journal of Applied Econometrics, 37(6):1093–1120, September–October 2022. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Skalin:1999:ALS

[ST99] Joakim Skalin and Timo Teräsvirta. Another look at Swedish business cycles, 1861–1988. Journal of Applied Econometrics, 14(4):359–378, July–August 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sauer:2021:UWW

[ST21] Robert M. Sauer and Christopher Taber. Understanding women’s wage growth using indirect inference with importance sampling. Journal of Applied Econometrics, 36(4):453–473, June–July 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Startz:2020:I

[Sta20] Richard Startz. The next hundred years of growth and convergence. Journal of Applied Econometrics, 35(1):99–113, January–February 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Stewart:2007:IDU

[Ste07] Mark B. Stewart. The interrelated dynamics of unemployment and low-wage employment. Journal of Applied Econometrics, 22(3):511–531, April–May 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Stewart:2018:NCS

[Ste18] Kenneth G. Stewart. Normalized CES supply systems: Replication of Klump, McAdam, and Willman (2007). Journal of Applied Econometrics, 33(2):290–296, March 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Stinebrickner:2000:SCV

[Sti00] Todd R. Stinebrickner. Serially correlated variables in dynamic, discrete choice models. Journal of Applied Econometrics, 15(6): 595–624, November–December 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Stone:1986:NML

[Sto86] Richard Stone. Nobel memorial lecture 1984. The accounts of society. Journal of Applied Econometrics, 1(1):5–28, January 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Stock:1991:BAU

[Sto91] James H. Stock. Bayesian approaches to the ‘unit root’ problem: A comment. Journal of Applied Econometrics, 6(4):403– 411, October–December 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Storer:1994:UDL

[Sto94] P. Storer. Unemployment dynamics and labour market tightness: An empirical evaluation of matching function models. Journal of Applied Econometrics, 9(4):389–419, October– December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Stoye:2009:BRC

[Sto09] J. Stoye. Book review: Charles F. Manski, Identification for Prediction and Decision (Harvard University Press 2007). Journal of Applied Econometrics, 24(5):857–862, August 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Strachan:2009:CJG

[Str09] Rodney W. Strachan. Comment on ‘Jointness of growth determinants’ by Gernot Doppelhofer and Melvyn Weeks. Journal of Applied Econometrics, 24(2):245–247, March 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [DW09a].

Sueyoshi:1995:CBR

[Sue95] Glenn T. Sueyoshi. A class of binary response models for grouped duration data. Journal of Applied Econometrics, 10(4):411–431, October–December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Summers:2017:CBG

[Sum17] Peter M. Summers. Credit booms gone bust: Replication of Schularick and Taylor (AER 2012). Journal of Applied Econometrics, 32(5):1033–1038, August 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sutherland:2023:FGE

[Sut23] Christopher S. Sutherland. Forward guidance and expectation formation: A narrative approach. Journal of Applied Econometrics, 38(2):222–241, March 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Schotman:1991:BRU

[SV91] Peter C. Schotman and Herman K. Van Dijk. On Bayesian routes to unit roots. Journal of Applied Econometrics, 6(4): 387–401, October–December 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Soderlind:1996:ACA

[SV96] Paul Söderlind and Anders Vredin. Applied cointegration analysis in the mirror of macroeconomic theory. Journal of Applied Econometrics, 11(4):363–381, July 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sarno:2005:MFS

[SV05] Lucio Sarno and Giorgio Valente. Modelling and forecasting stock returns: exploiting the futures market, regime shifts and international spillovers. Journal of Applied Econometrics, 20 (3):345–376, March–April 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sorensen:2014:REI

[SV14] Kenneth L. Sørensen and Rune Vejlin. Return to experience and initial wage level: do low wage workers catch up? Journal of Applied Econometrics, 29(6):984–1006, September–October 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sokullu:2022:ICC

[SV22] Senay Sokullu and Christine Valente. Individual consumption in collective households: Identification using repeated observations with an application to PROGRESA. Journal of Applied Econometrics, 37(2):286–304, March 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Straetmans:2008:EUS

[SVW08] S. T. M. Straetmans, W. F. C. Verschoor, and C. C. P. Wolff. Extreme US stock market fluctuations in the wake of 9/11. Journal of Applied Econometrics, 23(1):17–42, January– February 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smith:1986:EIC

[SW86] P. N. Smith and M. R. Wickens. An empirical investigation into the causes of failure of the monetary model of the exchange rate. Journal of Applied Econometrics, 1(2):143–162, April 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smets:2005:CSF

[SW05] Frank Smets and Raf Wouters. Comparing shocks and frictions in US and euro area business cycles: a Bayesian DSGE approach. Journal of Applied Econometrics, 20(2):161–183, 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Savage:2008:LFD

[SW08] Scott J. Savage and Donald M. Waldman. Learning and fatigue during choice experiments: a comparison of online and mail survey modes. Journal of Applied Econometrics, 23(3):351– 371, April 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Semykina:2013:EDP

[SW13] Anastasia Semykina and Jeffrey M. Wooldridge. Estimation of dynamic panel data models with sample selection. Journal of Applied Econometrics, 28(1):47–61, January–February 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Semykina:2018:BRP

[SW18a] Anastasia Semykina and Jeffrey M. Wooldridge. Binary response panel data models with sample selection and selfselection. Journal of Applied Econometrics, 33(2):179–197, March 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Spolaore:2018:ADN

[SW18b] Enrico Spolaore and Romain Wacziarg. Ancestry and development: New evidence. Journal of Applied Econometrics, 33 (5):748–762, August 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Sy:2020:SRS

[SW20] Malick Sy and Liuren Wu. The shale revolution and shifting crude dynamics. Journal of Applied Econometrics, 35(2):160– 175, March 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Stevens:2021:EAL

[SW21] Arnoud Stevens and Joris Wauters. Is euro area lowflation here to stay? Insights from a time-varying parameter model with survey data. Journal of Applied Econometrics, 36(5):566– 586, August 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Swann:2001:SPC

[Swa01] Christopher A. Swann. Software for parallel computing: the LAM implementation of MPI. Journal of Applied Econometrics, 16(2):185–194, March–April 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Swanson:2020:PIR

[SXY20] Norman R. Swanson, Weiqi Xiong, and Xiye Yang. Predicting interest rates using shrinkage methods, real-time diffusion indexes, and model combinations. Journal of Applied Econometrics, 35(5):587–613, August 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Smith:1991:PDM

[SZ91] Gregor W. Smith and Stanley E. Zin. Persistent deficits and the market value of government debt. Journal of Applied Econometrics, 6(1):31–44, January–March 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Stengos:2006:IPP

[SZ06] T. Stengos and E. Zacharias. Intertemporal pricing and price discrimination: a semiparametric hedonic analysis of the personal computer market. Journal of Applied Econometrics, 21 (3):371–386, April 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Simar:2007:SIA

[SZ07] Léopold Simar and Valentin Zelenyuk. Statistical inference for aggregates of Farrell-type efficiencies. Journal of Applied Econometrics, 22(7):1367–1394, December 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Shi:2018:SEB

[SZ18] Zhentao Shi and Huanhuan Zheng. Structural estimation of behavioral heterogeneity. Journal of Applied Econometrics, 33 (5):690–707, August 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Szroeter:1994:BRL

[Szr94] Jerzy Szroeter. Book review: Learning and practicing econometrics, William E. Griffiths, R. Carter Hill and George G. Judge. John Wiley, New York, 1993, ISBN 0-471-59951-4, pp.

866. price £19.95. Journal of Applied Econometrics, 9(2):225– 226, April–June 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Szydlowski:2019:ECM

[Szy19] Arkadiusz Szydłowski. Endogenous censoring in the mixed proportional hazard model with an application to optimal unemployment insurance. Journal of Applied Econometrics, 34 (7):1086–1101, November–December 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Terasvirta:1992:CNB

[TA92a] T. Terasvirta and H. M. Anderson. Characterizing nonlinearities in business cycles using smooth transition autoregressive models. Journal of Applied Econometrics, 7(S1):S119–S136, December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Tunali:1992:MSS

[TA92b] Insan Tunali and Ragui Assaad. Market structure and spells of employment and unemployment: Evidence from the construction sector in Egypt. Journal of Applied Econometrics, 7(4): 339–367, October–December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Tan:2010:NOT

[Tan10] Chih Ming Tan. No one true path: uncovering the interplay between geography, institutions, and fractionalization in economic development. Journal of Applied Econometrics, 25 (7):1100–1127, November–December 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Taylor:1989:MPS

[Tay89] John B. Taylor. Monetary policy and the stability of macroeconomic relationships. Journal of Applied Econometrics, 4(S1): S161–S178, December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Terasvirta:2001:NLE

[TE01] Timo Teräsvirta and Ann-Charlotte Eliasson. Non-linear error correction and the UK demand for broad money, 1878– 1993. Journal of Applied Econometrics, 16(3):277–288, May– June 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Temple:1998:RTA

[Tem98] Jonathan R. W. Temple. Robustness tests of the augmented Solow model. Journal of Applied Econometrics, 13(4):361–375, July–August 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Terasvirta:1988:RPG

[Ter88] Timo Teräsvirta. A review of PC-give: A statistical package for econometric modelling. Journal of Applied Econometrics, 3(4): 333–340, October–December 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Terrell:1996:IMC

[Ter96] Dek Terrell. Incorporating monotonicity and concavity conditions in flexible functional forms. Journal of Applied Econometrics, 11(2):179–194, March 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Terza:2002:AAE

[Ter02] Joseph V. Terza. Alcohol abuse and employment: a second look. Journal of Applied Econometrics, 17(4):393–404, July– August 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Teyssiere:1998:XIS

[Tey98] Gilles Teyssière. XploRe 4.0, an interactive statistical computing environment. Journal of Applied Econometrics, 13(6): 673–679, November–December 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Teyssiere:2005:STS

[Tey05] Gilles Teyssière. Structural time series modelling with STAMP 6.02. Journal of Applied Econometrics, 20(4):571–577, May– June 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Tsui:2004:CHE

[TH04] Albert K. Tsui and Kin-Yip Ho. Conditional heteroscedasticity of exchange rates: further results based on the fractionally integrated approach. Journal of Applied Econometrics, 19(5): 637–642, September–October 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Thiele:2020:I

[Thi20] Stephen Thiele. Modeling the conditional distribution of financial returns with asymmetric tails. Journal of Applied Econometrics, 35(1):46–60, January–February 2020. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Thoursie:2004:RSS

[Tho04] Peter Skogman Thoursie. Reporting sick: are sporting events contagious? Journal of Applied Econometrics, 19(6):809–823, 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Thomas:2019:SQG

[Tho19] James Thomas. The signal quality of grades across academic fields. Journal of Applied Econometrics, 34(4):566–587, June– July 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Timmermann:1995:CTP

[Tim95] Allan Timmermann. Cointegration tests of present value models with a time-varying discount factor. Journal of Applied Econometrics, 10(1):17–31, January–March 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Tsionas:2014:FHP

[TK14] Efthymios G. Tsionas and Subal C. Kumbhakar. Firm heterogeneity, persistent and transient technical inefficiency: a generalized true random–effects model. Journal of Applied Econometrics, 29(1):110–132, January–February 2014. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Tobias:2003:FSH

[TL03] Justin L. Tobias and Mingliang Li. A finite-sample hierarchical analysis of wage variation across public high schools: evidence from the NLSY and high school and beyond. Journal of Applied Econometrics, 18(3):315–336, May–June 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Tanizaki:1994:PFS

[TM94] H. Tanizaki and R. S. Mariano. Prediction, filtering and smoothing in non-linear and non-normal cases using Monte Carlo integration. Journal of Applied Econometrics, 9(2):163– 179, April–June 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Tobin:1997:CPJ

[Tob97] James Tobin. Comments by Professor James tobin. Journal of Applied Econometrics, 12(5):647–650, September–October 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Town:1992:MWS

[Tow92] R. J. Town. Merger waves and the structure of merger and acquisition time-series. Journal of Applied Econometrics, 7(S1): S83–S100, December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Tunali:1997:CRA

[TP97] Insan Tunali and Jonathan B. Pritchett. Cox regression with alternative concepts of waiting time: the New Orleans yellow fever epidemic of 1853. Journal of Applied Econometrics, 12 (1):1–25, January 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Trede:1998:APM

[Tre98] Mark M. Trede. The age profile of mobility measures: an application to earnings in West Germany. Journal of Applied Econometrics, 13(4):397–409, July–August 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Trivedi:1987:SR

[Tri87] P. K. Trivedi. Software reviews. Journal of Applied Econometrics, 2(1):77, January 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Trivedi:1988:SRS

[Tri88] Pravin K. Trivedi. Software reviews: Statistical software tools (SST) version 1.1: A review. Journal of Applied Econometrics, 3(3):235–239, July–September 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Trivedi:1991:SST

[Tri91] Pravin K. Trivedi. Statistical software tools (SST) version 2.0: An update. Journal of Applied Econometrics, 6(1):97– 101, January–March 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Trivedi:1993:APL

[Tri93] P. K. Trivedi. An analysis of publication lags in econometrics. Journal of Applied Econometrics, 8(1):93–100, January–March 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Trivedi:1997:EME

[Tri97] Pravin K. Trivedi. Econometric models of event counts. Journal of Applied Econometrics, 12(3):199–201, May 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Troccoli:2024:DPP

[Tro24] Claudia Troccoli. Does paid parental leave affect children’s schooling outcomes? Replicating Danzer and Lavy (2018). Journal of Applied Econometrics, 39(2):356–362, March 2024. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Tse:1998:CHY

[Tse98] Y. K. Tse. The conditional heteroscedasticity of the yen-dollar exchange rate. Journal of Applied Econometrics, 13(1):49–55, January–February 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Tsionas:2002:SFM

[Tsi02] Efthymios G. Tsionas. Stochastic frontier models with random coefficients. Journal of Applied Econometrics, 17(2):127–147, March–April 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Tsionas:2006:IDS

[Tsi06] Efthymios G. Tsionas. Inference in dynamic stochastic frontier models. Journal of Applied Econometrics, 21(5):669–676, July– August 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Tsiaplias:2021:CIE

[Tsi21] Sarantis Tsiaplias. Consumer inflation expectations, income changes and economic downturns. Journal of Applied Econometrics, 36(6):784–807, September–October 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Tufte:1998:CRC

[Tuf98] David Tufte. CATS in RATS: cointegration analysis of time series: version 1.01. Journal of Applied Econometrics, 13(3): 321–330, May–June 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Turkington:1998:BRH

[Tur98] Darrell Turkington. Book review: Handbook of Matrices, Helmut Lütkepohl, Wiley, New York, 1996, ISBN 0-471-96688-6 (hardback), pp. xvi + 304. Price A$195.95 (hardback), A$89.95 (paperback). Journal of Applied Econometrics, 13(4):421–423, July–August 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Turner:2009:TCU

[Tur09] Paul Turner. Testing for cointegration using the Johansen approach: are we using the correct critical values? Journal of Applied Econometrics, 24(5):825–831, August 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Turnovsky:1987:AMD

[TW87] Stephen J. Turnovsky and Mark E. Wohar. Alternative modes of deficit financing and endogenous monetary and fiscal policy in the U.S.A. 1923–1982. Journal of Applied Econometrics, 2 (1):1–25, January 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Toda:2017:FTS

[TW17] Alexis Akira Toda and Kieran James Walsh. Fat tails and spurious estimation of consumption-based asset pricing models. Journal of Applied Econometrics, 32(6):1156–1177, September–October 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Toda:2021:EMD

[TW21] Alexis Akira Toda and Yulong Wang. Efficient minimum distance estimation of Pareto exponent from top income shares. Journal of Applied Econometrics, 36(2):228–243, March 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Turner:1989:DPL

[TWW89] David S. Turner, Kenneth F. Wallis, and John D. Whitley. Differences in the properties of large-scale macroeconometric models: The role of labour market specifications. Journal of Applied Econometrics, 4(4):317–344, October–December 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Traum:2015:WDG

[TY15] Nora Traum and Shu-Chun S. Yang. When does government debt crowd out investment? Journal of Applied Econometrics, 30(1):24–45, January–February 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Teulings:2014:ERM

[TZ14] Coen N. Teulings and Nikolay Zubanov. Is economic recovery a myth? Robust estimation of impulse responses. Journal of Applied Econometrics, 29(3):497–514, April–May 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Todorov:2022:IGU

[TZ22] Viktor Todorov and Yang Zhang. Information gains from using short-dated options for measuring and forecasting volatility. Journal of Applied Econometrics, 37(2):368–391, March 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Unayama:2006:ECA

[Una06] Takashi Unayama. The Engel curve for alcohol and the rank of demand systems. Journal of Applied Econometrics, 21(7): 1019–1038, November 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Uysal:2015:DRE

[Uys15] S. Derya Uysal. Doubly robust estimation of causal effects with multivalued treatments: An application to the returns to schooling. Journal of Applied Econometrics, 30(5):763–786, August 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

VanBiesebroeck:2007:CAP

[Van07] Johannes Van Biesebroeck. Complementarities in automobile production. Journal of Applied Econometrics, 22(7):1315–1345, December 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

VanOphem:2011:FVD

[Van11] Hans Van Ophem. The frequency of visiting a doctor: Is the decision to go independent of the frequency? Journal of Applied Econometrics, 26(5):872–879, August 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

VanVuuren:2017:USF

[Van17] Aico Van Vuuren. Using a structural-form model to analyze the impact of home ownership on unemployment duration. Journal of Applied Econometrics, 32(4):858–876, June–July 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Vakratsas:2002:SLH

[VB02] Demetrios Vakratsas and Frank M. Bass. A segment-level hazard approach to studying household purchase timing decisions. Journal of Applied Econometrics, 17(1):49–59, January– February 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Vickner:1998:CES

[VD98] Steven S. Vickner and Stephen P. Davies. Comment: on the estimation of simultaneous-equations error-components models with an application to a model of developing country foreign trade. Journal of Applied Econometrics, 13(6):671, November– December 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

vanDijk:2014:FIR

[vDKvdWW14] Dick van Dijk, Siem Jan Koopman, Michel van der Wel, and Jonathan H. Wright. Forecasting interest rates with shifting endpoints. Journal of Applied Econometrics, 29(5):693–712, August 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

vanDijk:2005:DBC

[vDvDF05] Dick van Dijk, Herman K. van Dijk, and Philip Hans Franses. On the dynamics of business cycle analysis: Editors’ introduction. Journal of Applied Econometrics, 20(2):147–150, 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

vanderWeide:2002:GGM

[vdW02] Roy van der Weide. GO–GARCH: a multivariate generalized orthogonal GARCH model. Journal of Applied Econometrics, 17(5):549–564, September–October 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Veall:1986:EEP

[Vea86] Michael R. Veall. On estimating the effects of peak demand pricing. Journal of Applied Econometrics, 1(1):81–93, January 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Veall:1991:SR

[Vea91] Michael R. Veall. Shazam 6.2: A review. Journal of Applied Econometrics, 6(3):317–320, July–September 1991. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Veall:1992:BPM

[Vea92] Michael R. Veall. Bootstrapping the process of model selection: an econometric example. Journal of Applied Econometrics, 7(1):93–99, January–March 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Vella:1992:STS

[Vel92] F. Vella. Simple tests for sample selection bias in censored and discrete choice models. Journal of Applied Econometrics, 7(4): 413–421, October–December 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Verdier:2016:EDP

[Ver16] Valentin Verdier. Estimation of dynamic panel data models with cross-sectional dependence: Using cluster dependence for efficiency. Journal of Applied Econometrics, 31(1):85–105, January–February 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Verdier:2020:ATE

[Ver20] Valentin Verdier. Average treatment effects for stayers with correlated random coefficient models of panel data. Journal of Applied Econometrics, 35(7):917–939, November–December 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

VanDijk:1999:TAP

[VFL99] Dick Van Dijk, Philip Hans Franses, and André Lucas. Testing for ARCH in the presence of additive outliers. Journal of Applied Econometrics, 14(5):539–562, September–October 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Velasquez-Giraldo:2018:FED

[VGCBHJC18] Mateo Velásquez-Giraldo, Gustavo Canavire-Bacarreza, Kim P. Huynh, and David T. Jacho-Chavez. Flexible estimation of demand systems: A copula approach. Journal of Applied Econometrics, 33(7):1109–1116, November–December 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

vanHasselt:2022:BAA

[vHBB22] Martijn van Hasselt, Christopher R. Bollinger, and Jeremy W. Bray. A Bayesian approach to account for misclassification in prevalence and trend estimation. Journal of Applied Econometrics, 37(2):351–367, March 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Villani:2009:SSP

[Vil09] Mattias Villani. Steady-state priors for vector autoregressions. Journal of Applied Econometrics, 24(4):630–650, June–July 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Vinod:2000:RGW

[Vin00] H. D. Vinod. Review of GAUSS for Windows, including its numerical accuracy. Journal of Applied Econometrics, 15(2): 211–220, March–April 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Vinod:2003:RMV

[Vin03] H. D. Vinod. Review of mathStatica (v.1): an add-on to Mathematica. Journal of Applied Econometrics, 18(4):485–491, July– August 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Visco:1988:BRA

[Vis88] Ignazio Visco. Book review: Advanced Econometric Methods. By Thomas B. Fomby, R. Carter Hill AND Stanley R. Johnson. (Springer Verlag, New York, Berlin, Heidelberg, Tokyo, 1984. ISBN 0-387-90908-7. $54.00. pp. xix + 624). Journal of Applied Econometrics, 3(2):165–168, April 1988. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

VanSoest:1987:MEA

[VK87] Arthur Van Soest and Peter Kooreman. A micro-econometric analysis of vacation behaviour. Journal of Applied Econometrics, 2(3):215–226, July 1987. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

vanNorden:1996:RST

[vN96] Simon van Norden. Regime switching as a test for exchange rate bubbles. Journal of Applied Econometrics, 11(3):219–251, May 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

VanDriel:1997:DFU

[VNZ97] Hans Van Driel, Venuta Nadall, and Kees Zeelenberg. The demand for food in the United States and The Netherlands: a systems approach with the CBS model. Journal of Applied Econometrics, 12(5):509–532, September–October 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

vanOurs:2006:CCJ

[vO06] Jan C. van Ours. Cannabis, cocaine and jobs. Journal of Applied Econometrics, 21(7):897–917, November 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Vogelvang:1992:HTC

[Vog92] E. Vogelvang. Hypotheses testing concerning relationships between spot prices of various types of coffee. Journal of Applied Econometrics, 7(2):191–201, April–June 1992. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

vanOphem:2019:CRI

[vOvGvGB19] Hans van Ophem, Noud van Giersbergen, Kees Jan van Garderen, and Maurice Bun. The cyclicality of R&D investment revisited. Journal of Applied Econometrics, 34(2):315– 324, March 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

vanOurs:2011:CUM

[vOW11] Jan C. van Ours and Jenny Williams. Cannabis use and mental health problems. Journal of Applied Econometrics, 26(7):1137– 1156, November–December 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

vanOordt:2019:SRB

[vOZ19] Maarten van Oordt and Chen Zhou. Systemic risk and bank business models. Journal of Applied Econometrics, 34(3):365– 384, April–May 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

VanPraag:1993:CAM

[VV93] B. M. S. Van Praag and E. M. Vermeulen. A count-amount model with endogenous recording of observations. Journal of Applied Econometrics, 8(4):383–395, October–December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Vella:1998:WWD

[VV98] Francis Vella and Marno Verbeek. Whose wages do unions raise? A dynamic model of unionism and wage rate determination for young men. Journal of Applied Econometrics, 13(2): 163–183, March–April 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

VanderKlaauw:2013:CSH

[VV13] Bas Van der Klaauw and Jan C. Van Ours. Carrot and stick: how re-employment bonuses and benefit sanctions affect exit rates from welfare. Journal of Applied Econometrics, 28(2): 275–296, March 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

VanSoest:2014:HSR

[VV14] Arthur Van Soest and Hana Vonkova. How sensitive are retirement decisions to financial incentives? A stated preference analysis. Journal of Applied Econometrics, 29(2):246–264, March 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

VanVuuren:2000:MEE

[VVR00] Aico Van Vuuren, Gerard J. Van Den Berg, and Geert Ridder. Measuring the equilibrium effects of unemployment benefits dispersion. Journal of Applied Econometrics, 15(6):547–574, November–December 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wagner:2015:EKC

[Wag15] Martin Wagner. The environmental Kuznets curve, cointegration and nonlinearity. Journal of Applied Econometrics, 30(6): 948–967, September–October 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wang:2006:CPC

[Wan06] Zijun Wang. Convergence of productivity: a comment. Journal of Applied Econometrics, 21(7):1101–1102, November 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ward:2017:SDZ

[War17] Felix Ward. Spotting the danger zone: Forecasting financial crises with classification tree ensembles and many predictors. Journal of Applied Econometrics, 32(2):359–378, March 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Watson:1989:MR

[Wat89] Mark W. Watson. MTS: A review. Journal of Applied Econometrics, 4(2):205–206, April–June 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Watson:1994:CPa

[Wat94a] Mark W. Watson. Call for papers. Journal of Applied Econometrics, 9(4):481, October–December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Watson:1994:CPb

[Wat94b] Mark W. Watson. Call for papers. Journal of Applied Econometrics, 9(S1):S161, December 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Watanabe:1999:NLF

[Wat99] Toshiaki Watanabe. A non-linear filtering approach to stochastic volatility models with an application to daily stock returns. Journal of Applied Econometrics, 14(2):101–121, March–April 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Watson:2007:JAE

[Wat07] Mark Watson. Journal of applied econometrics annual lecture series. Journal of Applied Econometrics, 22(3):701, April–May 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Walker:2007:IPN

[WBAB07] Joan L. Walker, Moshe Ben-Akiva, and Denis Bolduc. Identification of parameters in normal error component logit-mixture (NECLM) models. Journal of Applied Econometrics, 22(6): 1095–1125, September–October 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Winkelmann:2016:EMP

[WBL16] Lars Winkelmann, Markus Bibinger, and Tobias Linzert. ECB monetary policy surprises: Identification through cojumps in interest rates. Journal of Applied Econometrics, 31(4):613–629, June–July 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wilson:2004:NAR

[WC04] Paul W. Wilson and Kathleen Carey. Nonparametric analysis of returns to scale in the US hospital industry. Journal of Applied Econometrics, 19(4):505–524, July–August 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Warne:2017:MPL

[WCC17] Anders Warne, Günter Coenen, and Kai Christoffel. Marginalized predictive likelihood comparisons of linear Gaussian statespace models with applications to DSGE, DSGE–VAR, and VAR models. Journal of Applied Econometrics, 32(1):103–119, January–February 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wedel:1993:LCP

[WDBR93] M. Wedel, W. S. Desarbo, J. R. Bult, and V. Ramaswamy. A latent class Poisson regression model for heterogeneous count data. Journal of Applied Econometrics, 8(4):397–411, October– December 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Weale:1992:EDM

[Wea92] Martin Weale. Estimation of data measured with error and subject to linear restrictions. Journal of Applied Econometrics, 7(2):167–174, April–June 1992. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Weber:1995:COC

[Web95] Christian E. Weber. Cyclical output, cyclical unemployment, and Okun’s coefficient: A new approach. Journal of Applied Econometrics, 10(4):433–445, October–December 1995. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Weeks:2001:BRM

[Wee01] Melvyn Weeks. Book review: Measurement error and latent variables in econometrics, Tom Wansbeek and Erik Meijer, advanced textbooks in economics: editors C. Bliss and M. D. Intriligator, North-Holland, Amsterdam. Journal of Applied Econometrics, 16(6):749–753, November–December 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Weeks:2002:BRI

[Wee02] Melvyn Weeks. Book reviews: Introductory Econometrics: A Modern Approach, Jeffrey M. Wooldridge, South-Western College Publishing, 2000, 1-538-85013-2, 824, Price (hb) $98.75. Journal of Applied Econometrics, 17(2):191–193, March–April 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Weeks:2003:BRD

[Wee03] M. Weeks. Book review: Discrete choice methods with simulation, Kenneth E. Train, Cambridge University Press, 2003, ISBN: 0-521-81696-3, pp. 334. Journal of Applied Econometrics, 18(3):379–383, May–June 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Weiss:1996:ETS

[Wei96] Andrew A. Weiss. Estimating time series models using the relevant cost function. Journal of Applied Econometrics, 11 (5):539–560, September 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wells:2003:RFM

[Wel03] Curt Wells. Retesting Fair’s (1978) model on infidelity. Journal of Applied Econometrics, 18(2):237–239, March–April 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Westerlund:2008:PCT

[Wes08] Joakim Westerlund. Panel cointegration tests of the Fisher effect. Journal of Applied Econometrics, 23(2):193–233, March 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Westerlund:2020:CSA

[Wes20] Joakim Westerlund. A cross-section average-based principal components approach for fixed- t panels. Journal of Applied Econometrics, 35(6):776–785, September–October 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wang:2001:FPG

[WFBM01] Kai-Li Wang, Christopher Fawson, Christopher B. Barrett, and James B. McDonald. A flexible parametric GARCH model with an application to exchange rates. Journal of Applied Econometrics, 16(4):521–536, July–August 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wegmuller:2023:UWE

[WG23] Philipp Wegmüller and Christian Glocker. US weekly economic index: Replication and extension. Journal of Applied Econometrics, 38(6):977–985, September–October 2023. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Westerlund:2011:NPT

[WH11] Joakim Westerlund and Wolfgang Hess. A new poolability test for cointegrated panels. Journal of Applied Econometrics, 26 (1):56–88, January–February 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wheeler:2003:EAE

[Whe03] Christopher H. Wheeler. Evidence on agglomeration economies, diseconomies, and growth. Journal of Applied Econometrics, 18 (1):79–104, January–February 2003. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

White:1988:SR

[Whi88] K. J. White. Shazam: A response. Journal of Applied Econometrics, 3(3):245–247, July–September 1988. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

White:1989:NFW

[Whi89] Kenneth J. White. NAG Fortran Workstation Library: A review. Journal of Applied Econometrics, 4(1):93–96, January– March 1989. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wilke:2006:SPE

[Wil06a] Ralf A. Wilke. Semi-parametric estimation of consumptionbased equivalence scales: the case of Germany. Journal of Applied Econometrics, 21(6):781–802, September–October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Willis:2006:MPR

[Wil06b] Jonathan L. Willis. Magazine prices revisited. Journal of Applied Econometrics, 21(3):337–344, April 2006. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Williams:2019:CAU

[Wil19] Benjamin Williams. Controlling for ability using test scores. Journal of Applied Econometrics, 34(4):547–565, June–July 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Winkelmann:2004:HCR

[Win04] Rainer Winkelmann. Health care reform and the number of doctor visits — an econometric analysis. Journal of Applied Econometrics, 19(4):455–472, July–August 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Westerlund:2017:TPP

[WKN17a] Joakim Westerlund, Hande Karabiyik, and Paresh Narayan. Testing for predictability in panels with general predictors. Journal of Applied Econometrics, 32(3):554–574, April–May 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wrenn:2017:CPE

[WKN17b] Douglas H. Wrenn, H. Allen Klaiber, and David A. Newburn. Confronting price endogeneity in a duration model of residential subdivision development. Journal of Applied Econometrics, 32(3):661–682, April–May 2017. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wren-Lewis:1986:EMU

[WL86] Simon Wren-Lewis. An econometric model of U.K. manufacturing employment using survey data on expected output. Journal of Applied Econometrics, 1(4):297–316, October 1986. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wickens:2001:ESI

[WM01] Michael R. Wickens and Roberto Motto. Estimating shocks and impulse response functions. Journal of Applied Econometrics, 16(3):371–387, May–June 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wohar:1996:PPV

[Woh96] Mark E. Wohar. PcGive professional (version 8) and reviews (MicroTSP for Windows version 1.1a): a comparative review. Journal of Applied Econometrics, 11(1):105–115, January 1996. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wolfson:1993:CCA

[Wol93] P. Wolfson. Compositional change, aggregation, and dynamic factor demand. Estimates on a panel of manufacturing firms.

Journal of Applied Econometrics, 8(2):129–148, April–June 1993. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wolak:2007:QSS

[Wol07] Frank A. Wolak. Quantifying the supply-side benefits from forward contracting in wholesale electricity markets. Journal of Applied Econometrics, 22(7):1179–1209, December 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wolters:2015:EPD

[Wol15] Maik H. Wolters. Evaluating point and density forecasts of DSGE models. Journal of Applied Econometrics, 30(1):74–96, January–February 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wolters:2018:HBB

[Wol18] Maik H. Wolters. How the baby boomers’ retirement wave distorts model-based output gap estimates. Journal of Applied Econometrics, 33(5):680–689, August 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wooldridge:2005:SSI

[Woo05] Jeffrey M. Wooldridge. Simple solutions to the initial conditions problem in dynamic, nonlinear panel data models with unobserved heterogeneity. Journal of Applied Econometrics, 20 (1):39–54, January–February 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Westerlund:2019:CFP

[WPN19] Joakim Westerlund, Yana Petrova, and Milda Norkute. CCE in fixed-T panels. Journal of Applied Econometrics, 34(5):746– 761, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wang:2018:HPP

[WPS18] Wuyi Wang, Peter C. B. Phillips, and Liangjun Su. Homogeneity pursuit in panel data models: Theory and application. Journal of Applied Econometrics, 33(6):797–815, September– October 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wright:1999:TUR

[Wri99] Jonathan H. Wright. Testing for a unit root in the volatility of asset returns. Journal of Applied Econometrics, 14(3):309–318, May–June 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wright:2013:ERT

[Wri13] Jonathan H. Wright. Evaluating real–time var forecasts with an informative democratic prior. Journal of Applied Econometrics, 28(5):762–776, August 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Windmeijer:1997:ECD

[WS97] F. A. G. Windmeijer and J. M. C. Santos Silva. Endogeneity in count data models: an application to demand for health care. Journal of Applied Econometrics, 12(3):281–294, May 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Watkins:2014:MEI

[WVG14] John G. T. Watkins, Andrey L. Vasnev, and Richard Gerlach. Multiple event incidence and duration analysis for credit data incorporating non-stochastic loan maturity. Journal of Applied Econometrics, 29(4):627–648, June–July 2014. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wei:2013:RRL

[WW13] Min Wei and Jonathan H. Wright. Reverse regressions and long–horizon forecasting. Journal of Applied Econometrics, 28 (3):353–371, April–May 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wheelock:2018:ESE

[WW18] David C. Wheelock and Paul W. Wilson. The evolution of scale economies in US banking. Journal of Applied Econometrics, 33 (1):16–28, January–February 2018. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wu:2020:NIR

[WX20] Jing Cynthia Wu and Fan Dora Xia. Negative interest rate policy and the yield curve. Journal of Applied Econometrics, 35(6):653–672, September–October 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Watanabe:2023:DMZ

[WY23] Tsutomu Watanabe and Tomoyoshi Yabu. The demand for money at the zero interest rate bound. Journal of Applied Econometrics, 38(6):968–976, September–October 2023. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wagenvoort:2014:UCR

[WZ14] Rien J. L. M. Wagenvoort and Sanne Zwart. Uncovering the common risk–free rate in the European Monetary Union. Journal of Applied Econometrics, 29(3):394–414, April–May 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Wang:2019:PPW

[WZP19] Wendun Wang, Xinyu Zhang, and Richard Paap. To pool or not to pool: What is a good strategy for parameter estimation and forecasting in panel regressions? Journal of Applied Econometrics, 34(5):724–745, August 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Xun:2016:SET

[XL16] Zhou Xun and Michel Lubrano. Simulation estimation of twotiered dynamic panel tobit models with an application to the labour supply of married women: A comment. Journal of Applied Econometrics, 31(4):756–761, June–July 2016. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Xie:2012:MIN

[XQ12] Hui Xie and Yi Qian. Measuring the impact of nonignorability in panel data with non-monotone nonresponse. Journal of Applied Econometrics, 27(1):129–159, January–February 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Xu:2013:NEE

[Xu13] Pai Xu. Nonparametric estimation of entry cost in first–price procurement auctions. Journal of Applied Econometrics, 28 (6):1046–1065, September–October 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Yamamoto:2019:BII

[Yam19] Yohei Yamamoto. Bootstrap inference for impulse response functions in factor-augmented vector autoregressions. Journal of Applied Econometrics, 34(2):247–267, March 2019. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Yatchew:2000:SEE

[Yat00] A. Yatchew. Scale economies in electricity distribution: a semiparametric analysis. Journal of Applied Econometrics, 15(2): 187–210, March–April 2000. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Yamamoto:2022:IFA

[YH22] Yohei Yamamoto and Naoko Hara. Identifying factoraugmented vector autoregression models via changes in shock variances. Journal of Applied Econometrics, 37(4):722–745, June–July 2022. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Yang:2021:SEV

[YHLL21] Bingduo Yang, Christian M. Hafner, Guannan Liu, and Wei Long. Semiparametric estimation and variable selection for single-index copula models. Journal of Applied Econometrics, 36(7):962–988, November–December 2021. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Yang:2006:EMC

[YHLW06] Jian Yang, Cheng Hsiao, Qi Li, and Zijun Wang. The emerging market crisis and stock market linkages: further evidence. Journal of Applied Econometrics, 21(6):727–744, September– October 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Yalta:2008:GLP

[YL08] A. Talha Yalta and Riccardo Lucchetti. The GNU/Linux platform and freedom respecting software for economists. Journal of Applied Econometrics, 23(2):279–286, March 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Yang:2006:MFS

[YT06] Z. L. Yang and Y. K. Tse. Modelling firm-size distribution using Box–Cox heteroscedastic regression. Journal of Applied Econometrics, 21(5):641–653, July–August 2006. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Yuengert:1994:IER

[Yue94] A. M. Yuengert. Immigrant earnings, relative to what? The importance of earnings function specification and comparison points. Journal of Applied Econometrics, 9(1):71–90, January– March 1994. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Yalta:2007:GNA

[YY07] A. Talha Yalta and A. Yasemin Yalta. GRETL 1.6.0 and its numerical accuracy. Journal of Applied Econometrics, 22(4): 849–854, June–July 2007. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Yin:2015:HMM

[YZ15] Xiangkang Yin and Jing Zhao. A hidden Markov model approach to information-based trading: Theory and applications. Journal of Applied Econometrics, 30(7):1210–1234, November– December 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Zafar:2011:CSE

[Zaf11] Basit Zafar. Can subjective expectations data be used in choice models? Evidence on cognitive biases. Journal of Applied Econometrics, 26(3):520–544, April–May 2011. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Zandberg:2012:PEF

[ZDE12] Eelco Zandberg, Jakob De Haan, and J. Paul Elhorst. The political economy of financial reform: How robust are Huang’s findings? Journal of Applied Econometrics, 27(4):695–699, June–July 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [Hua09].

Zheng:2004:PCD

[Zhe04] Buhong Zheng. Poverty comparisons with dependent samples. Journal of Applied Econometrics, 19(3):419–428, May–June 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Zhang:2020:FSR

[ZHJJ20] Hongwei Zhang, Qiang He, Ben Jacobsen, and Fuwei Jiang. Forecasting stock returns with model uncertainty and parameter instability. Journal of Applied Econometrics, 35(5):629– 644, August 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Zhou:2020:RWO

[Zho20] Xiaoqing Zhou. Refining the workhorse oil market model. Journal of Applied Econometrics, 35(1):130–140, January–February 2020. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Zeileis:2005:VMS

[ZK05] Achim Zeileis and Christian Kleiber. Validating multiple structural change models — a case study. Journal of Applied Econometrics, 20(5):685–690, July–August 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ziebarth:2014:EEG

[ZK14] Nicolas R. Ziebarth and Martin Karlsson. The effects of expanding the generosity of the statutory sickness insurance system. Journal of Applied Econometrics, 29(2):208–230, March 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Zeileis:2005:MSC

[ZLKH05] Achim Zeileis, Friedrich Leisch, Christian Kleiber, and Kurt Hornik. Monitoring structural change in dynamic econometric models. Journal of Applied Econometrics, 20(1):99–121, January–February 2005. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Zellner:1998:AFF

[ZR98] Arnold Zellner and Hang Ryu. Alternative functional forms for production, cost and returns to scale functions. Journal of Applied Econometrics, 13(2):101–127, March–April 1998. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).